A Complete Bibliography of Publications in Metrika

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

26 July 2025 Version 1.71

Title word cross-reference

(α, β) [408, 445]. (k_{1}, k_{2}) [2072]. (n − k + 1) [1704, 1597]. {−1, 1} [1324]. 1 [461, 475, 1288]. 1/\sqrt{n} [1354]. 100/\sqrt{n} [1624]. 1135 [461]. 135 [461]. 15 [1324]. 16 [1169]. 175 [475]. 18 [1433]. 2 [759, 239, 246, 1403, 638, 1298]. 20 [1169]. 24 [1229]. 2^{k−p} [1351]. 2^{m−k} [1419]. 2^{n} [1336, 1212]. 2^{n−m} [1854]. 2 × 2 [2309, 99, 1074, 1100]. 32 [1352]. 36 [1397]. 4 [485]. 50 [831]. 75 [475]. < [533]. ^{+} [639]. _{p} [1732]. _{pmk} [1194]. A [1407, 799, 1207, 2173]. α [630, 1624, 1904, 2204]. AR(1) [1610, 1749, 1974]. arcsin [935]. X̄ [881]. β [363]. c [1444, 2277]. C′′_{pk} [1259, 1293]. C_{pm} [1493]. chi^{2} [28, 28]. CM [1121]. D [2216, 1418, 1562, 1798, 847, 1986, 1328, 1770, 2065, 1068, 1581, 1444, 2277, 1765, 1298, 2137]. δ [911]. E [688, 699, 808, 1110, 781, 1163].

E(s^{2}) [1502, 1495, 1557]. ℓ_{2} [2224]. F [746, 306, 1111, 2167, 604, 932]. G [746, 2071, 2349]. h [435]. H_{o} [505]. I^{2} [2009]. I_{L} [2208]. INAR(1) [1417]. K [820, 2197, 1542, 2328, 351, 1898, 2032, 1407, 1381, 2024, 2205, 2191, 1406, 288, 2071, 945, 1671, 505, 239, 246, 2087, 373, 778, 559]. K_{φ} [1359]. k ≤ 10 [1108]. L [1844, 1360, 1938]. L^{1} [1362]. L^{p} [2334]. L_{1} [999, 1250]. L_{2} [1349]. L_{p} [1494]. L_{q} [2156, 1785]. LM [1638]. LR [1638]. M [1337, 1588, 1338, 1461, 1522, 778, 1942, 1979, 1850, 1824, 1243, 332, 1935]. P (X < Y) [2327]. D [1405]. N [346, 353, 90, 48, 2205, 1618, 2071, 1671, 505, 1704, 2087, 1597, 559, 1298]. n^{−1/2} [1250]. n ≤ 2^{n−m−2} [1854]. P [1826, 932, 1645, 1227]. P (X < Y) [1237, 937]. P [Y < X] [1300]. PBIB(3) [1269]. ϕ [1075, 1244]. ϕ_{p} [1555]. Q [817, 2272, 2195, 1955]. q = n/N > 0 [571].

R [2011, 1776, 2128, 1696, 484, 513]. r∗ [1738]. ρ [2331]. r × c [284]. S [1121, 780, 1453]. s^{b} [1156]. s^{k−p} [1156]. s^{m} [1645]. s_{1} × s_{2} × s_{3} [1328]. Σp_{i}f(p_{i}) ≥ p_{i}f(q_{i}) [361]. SLE_{8} [1815]. T [1734, 342, 2188, 931, 1098, 1111, 1233, 167, 1544, 511, 1638]. T_{1} [431]. T_{3} [223]. θ [2335]. U [927, 1253, 1062, 1818, 1625, 642, 654, 1607]. V [2284]. φ [2248, 1055]. D⃗ [2220]. W [1638]. x^{2} [205]. X_{k+s:n} − X_{k:n} [825]. Z [2240, 427].

-aberration [2335]. -Charts [461]. -circulant [2032, 1381]. -computation [2349]. -copulas [1618]. -dependent [332]. -determinant [1324]. -dimensional [1243]. -dimensionaler [48]. -Discrepancy [1349]. -Divergence [1359]. -divergences [1075]. -elementiges [353]. -END [1935]. -estimates [1121, 1337, 1942]. -estimation [1588, 1461, 1696]. -estimators [1522, 1979, 1850, 1453, 1824]. -Felder-Tafeln [485]. -functionals [1938]. -kernel [1338]. -level [2195, 759]. -likelihood [2156, 1785]. -lists [1406]. -mean [484]. -measure [2009]. -method [911]. -mixing [2248, 1055, 1244, 2331]. -Mixtures [2204]. -model [931, 1098]. -Norm [1362, 1250, 1494]. -optimal [1444, 2277, 1555, 699, 2216, 1418, 1562, 1798, 808, 1502, 1405, 847, 1986, 2011, 2208, 1110, 1207, 1328, 1770, 1495, 1557, 1776, 2065, 1068, 1581, 2173, 1765, 2220, 1298, 2137]. -optimality [688, 1407, 781, 2128, 2284, 1163]. -Optimum [642]. -out-of- [2205, 2071, 1671, 1704, 2087, 1597]. -penalized [2224]. -person [346]. -Personen-Spiele [90]. -process [2237]. -quantile [2334]. -quantiles [630]. -record [945]. -records [1542]. -reduced [817]. -run [1433]. -runs [2072]. -sequence [1955]. -shrinking [1354]. -stable [1904]. -Stage [1403, 373]. -statistics [435, 1253, 927, 1844, 1360, 1062, 1818, 1625, 778, 654]. -Stichproben-Fall [239, 246]. -te [513]. -Test [167, 2167, 604, 932]. -Testes [604]. -Tests [511]. -th [2024]. -Transform [427]. -type [1607]. -value [1227, 1826, 932]. -Verfahren [28].

0 [1372, 1371]. 0-387-00135-2 [1371].

0-387-20078-9 [1372].

1 [441, 238, 269]. 19 [420]. 1939 [120]. 1958 [18]. 1968 [323].

2 [1371]. 2012 [1814]. 26 [420]. 27 [1478].

3 [1370]. 3-540-03079-4 [1370]. 3rd [2044].

4 [1370].

5th [7].

9 [1372].

A-optimal [2257]. a-priori [1245]. AAI [2309]. Abelian [1322]. Aberration [1419, 1272, 864, 1170, 2323, 1080, 2335, 1545, 1640, 2227, 2159]. abgegrenzten [241, 268]. ablehnung [505]. ableitbares [277]. Ableitung [106]. Abschätzung [5]. abschnittsweisen [116]. absolute [1616, 2345]. Abstände [265, 501]. Abständen [528]. abstract [11]. abundance [2042]. Abweichung [562]. accelerated [1447, 1538, 2192]. Acceleration [1991]. acceptance [765]. accepted [851, 895]. accepting [2220]. account [407]. accounting [1311]. Accumulation [1793]. Accurate [949, 809]. Acknowledgement [547]. active [292]. activity [108]. Adams [1380]. Adaptive [2276, 2049, 2189, 1581]. addition [2187]. additional [790, 2219]. Additive [2307, 2171, 1774, 1964, 60, 1928, 2041, 2201, 2252, 1700, 470, 1981, 463, 1659]. additive-multiplicative [2041]. adequacy [1748]. adjacent [1448, 984, 1721]. Adjunktionsmatrizen [84]. Adjusted [1364, 2016]. adjustments [2115, 479]. Admissibility [837, 1886, 1044, 1594, 1172, 590]. Admissible [769, 452, 1500, 1024]. admitting [1322]. advances [390]. affine [1152]. after [2344, 915, 925]. Against [1408, 91, 952, 2265, 712, 649]. age [1273]. ageing [2123, 2114, 2091]. aggregate [2309]. aggregated [621, 1720]. agreement [1486]. ahead [1581]. algebra [1319]. algebraic [2035, 1539]. algorithm [1118, 1442, 1653, 1746, 2189, 725, 1986, 1842, 1438, 299, 2025, 2220, 1463]. Algorithms [1121, 859]. aliased [1982, 2073]. Allocation [1407, 577, 1329, 1686, 174, 1277, 1289, 421, 571, 1263, 449]. allocations [912, 907]. allowed [387]. allowing [1777]. Almost [700, 1625, 460, 919, 1892, 1665]. Alternative [1399, 1499, 680, 975, 1690, 537]. alternativen [101]. alternatives [1022, 1190, 952, 680, 1268, 1828]. Alternativtests [537]. always [1100]. am [8]. Amarts [640]. among [1678, 1608, 2317]. amount [363]. analog [2249]. Analogiegerät [185]. Analogy [842]. Analyse [44, 51, 58, 314, 65, 327]. Analyses [305]. Analysis [1313, 826, 720, 1343, 1413, 1394, 1366, 1362, 1371, 44, 51, 58, 486, 1673, 1336, 1119, 2035, 2241, 180, 1035, 2009, 2257, 285, 276, 477, 1238, 1695, 1019, 1835, 1283, 2133, 1928, 2041, 1863, 1031, 554, 250, 624, 1016, 2193, 2234, 32, 327, 1496, 1058, 1011, 2231, 1212, 1106, 1808, 543, 392, 1731, 2144, 100, 295, 728, 1033, 31, 65, 1463, 1480, 1370]. Analytic [2154, 2305]. Analytical [1474, 1049, 2103]. Analyzing [1432, 1746]. ANCOVA [44, 51, 58]. Anderson [50]. Änderungen [323]. Andrea [1387].

Angabefehlern [334]. angekommenen [342]. angewandten [626]. angle [2301]. angle-based [2301]. Ankündigung [517]. Ankunftszeiten [342]. Anlage [110]. Annäherung [205]. Announcement [188, 524, 545, 130]. ANOVA [1653, 2343, 2010, 1139, 1178, 401, 1132]. ANOVA-type [2010]. Antonio [1372]. Antony [1387]. Antwortklassen [353]. Anwendung [398, 43, 153, 140, 93]. Anwendungen [156]. Application [1349, 358, 2105, 141, 2281, 859, 900, 2026, 1570, 707, 514, 1460, 1577, 2345, 2085, 1657, 1047, 2083, 756, 192, 68, 129, 1201]. Applications [583, 1395, 967, 1365, 1891, 927, 1951, 1929, 1756, 1730, 2140, 2336, 1046, 1817, 1531, 1989, 854, 733, 1863, 2002, 236, 1494, 668, 1009, 1949, 1325, 2005, 1666, 1935, 1743, 2350, 1476]. applied [1019, 2219, 626, 1370]. apportionment [1263]. appraisal [486, 102]. Approach [2197, 486, 1997, 898, 1673, 2138, 2275, 1519, 876, 108, 1901, 1902, 840, 2337, 1486, 1794, 2168, 1930, 1467, 1732, 1028, 1931, 783, 2158, 1208, 565, 1983, 1293, 978, 292, 1113, 2184, 716]. approaches [1819]. Approximate [2058, 490, 847, 2224, 1749, 191, 2082]. approximately [419, 1053, 449]. approximating [475, 985]. Approximation [1375, 210, 527, 279, 1738, 692, 1728, 1947, 742, 860, 497, 858, 2305, 901, 1796]. Approximations [793, 2110, 1618]. approximative [996]. Arbeit [59]. arbitrary [243]. arc [1722]. ARCH [2162]. Archimedean [2135, 2013]. arcsine [1767]. Area [1626]. argument [838]. ARIMA [514]. Arising [1395, 1007, 884]. arithmetic [2269]. ARMA [621, 1105, 651, 647]. ARMA-model [651]. ARMAX [534]. armed [610]. arms [1239]. array [924]. arrays [1442, 141, 1645, 1179, 1326, 1352, 1433, 1584, 2232, 1028, 2005, 1765, 2187]. arrival [367, 339]. arrivals [217, 197, 354, 350, 394, 349]. artificial [1913]. aspect [1331]. aspects [844, 250, 390, 2062, 618]. assays [134]. Assessing [2059]. Assessment [1577, 1293, 335]. Associated [1398, 518, 1538, 1919, 2096, 1963, 924, 1901, 1902, 1734, 1946, 2019, 1864, 2343, 1111, 1039]. association [2007, 2309, 1035, 1152, 720, 1461, 2356, 1271]. assumption [562, 1511]. assumptions [1261]. Asymmetric [1349, 1369, 971, 1070, 1309, 2346, 2198, 1015, 1114, 1259]. Asymmetrical [2143, 1410, 1272, 688, 1678]. Asymptotic [2240, 1951, 1246, 1651, 1103, 1890, 1812, 996, 1868, 2235, 976, 2196, 1308, 2299, 811, 2176, 682, 1402, 1124, 974, 1436, 2127, 1261, 1593, 1670, 1711, 315, 1939, 1228, 2077, 1522, 267, 694, 708, 1334, 1862, 1849, 1850, 1453, 1337, 876, 2243, 1253, 1725, 872, 860, 616, 695, 1173, 360, 378, 291, 836, 2312, 2324, 652, 2172]. Asymptotically [1023, 2222, 955, 728, 1655, 1581]. Asymptotics [1697, 1353, 1567, 1979, 2282, 2297]. Asymptotik [98]. asymptotische [616, 139, 652]. Atkinson [1387]. attributes [667, 1017, 1458]. aufgrund [336]. Aufteilung [384]. augmented [2195, 991]. Ausdehnung [264]. Ausgewählte [161, 168, 239, 246]. Ausgleichung [437]. Aussagen [326]. Ausschußanteilen [511]. Ausschußhäufigkeit [116]. Austria [155]. Auswahlsatz [333]. Auswahlverfahren [229]. Auszeichnung [271]. autocorrelated [427, 1298]. autocovariance [1892, 2179]. Automatic [1384]. automatischen [110]. automorphism [1322]. autoregression [1171, 2190]. autoregressive [518, 1651, 2156, 2266, 1770, 2166, 1885, 801, 3, 1979, 1664, 2100, 2286, 1634]. auxiliary [2240, 919, 1099, 2074, 763, 944]. availability [252]. average [518, 918, 2349, 38, 1757, 1282, 1466, 578, 316, 905]. averages [563, 436, 471]. averaging [2003, 2344, 2161, 858, 901, 1836, 2321, 1870]. avoiding [1549]. Award [271]. AX=B [1230]. axiomatic [470]. Axiomensystem [202]. axioms [132, 138, 2061].

B [1824]. B-spline [1824]. Bad

[7, 8, 191, 7, 8]. Bahadur

[587, 2345, 2278, 2046, 418, 754, 1906]. balance [1770]. balanced [1542, 1329, 141, 768, 1549, 2247, 1279, 634, 2166, 1278, 1894, 760, 401, 1766, 2182, 1910, 1909]. Balking [269, 238]. balls [1732]. Ban [1201]. Banach [484, 1817]. Band [20, 2201]. bandit [1239, 610]. bands [738, 1985]. Barnett [2242]. Bartlett [886, 770, 1752]. Barycentric [1986]. base [438]. Based [1412, 1404, 1915, 927, 2329, 664, 2066, 1506, 1755, 1803, 1907, 2040, 1382, 1756, 1797, 1473, 2217, 1868, 2283, 1722, 1994, 876, 1076, 1750, 2192, 937, 2088, 1083, 2230, 1523, 1786, 2048, 1972, 1314, 1697, 1879, 1253, 1025, 1899, 1913, 2188, 750, 1957, 2133, 1636, 2181, 1714, 1785, 1343, 2002, 1534, 624, 2051, 2276, 791, 994, 2062, 1036, 659, 817, 1670, 1042, 2027, 2018, 2198, 2158, 2278, 1075, 2106, 1939, 251, 968, 1001, 1455, 2046, 1777, 1968, 1472, 1293, 1949, 2020, 1561, 2028, 1981, 2305, 255, 2064, 901, 778, 2177, 2301, 2245, 2254, 1736, 2357, 2079, 1037, 2358, 870]. based [1463, 1144, 1638, 2180, 1620, 2350, 1870, 1973, 2095, 1057, 1050]. baseline [2115]. basic [863, 28]. Basis [454, 653]. batch [197, 394]. batches [367]. bathtub [2028]. bathtub-shaped [2028]. Bayes [669, 1238, 1827, 1783, 1881, 1661, 837, 412, 783, 1711, 1575, 1245, 714, 697]. Bayesian [1959, 1182, 1915, 2334, 1598, 1994, 2246, 1487, 1750, 840, 2163, 2226, 1871, 2121, 2323, 2260, 1451, 421, 1194, 1293, 1362, 564, 2182, 2313, 1566, 1607, 1296]. Bayesian-like [1194]. Bayessche [355]. bearing [132, 138]. Becker [912]. Bedeutung [107]. bedingter [437]. Begriffe [166]. Begündung [653].

Behandlung [212, 13, 59]. behavior [1273, 1812, 2098, 455, 1868, 1311, 1559, 1453, 1057, 584]. behaviour [651, 155, 1862, 2312, 2324, 2238]. Behnken [970]. Behrens [182, 70, 2115, 430, 521]. Beitrag [392, 314, 53]. bekannten [191]. bekannter [191, 241]. Bell [2120]. Bellman [567]. Bemerkung [437, 333, 405, 60, 65]. Bemerkungen [275, 526, 392, 622, 59, 337]. bending [2000]. Bennett [420]. Beobachtungen [437, 464, 10]. Beobachtungsmodellen [340]. Beobachtungswerte [122]. Berechnung [149, 191, 17, 84, 15]. Bericht [18]. Berichtigung [89]. Berkson [2218]. Berlin [1372, 1370]. Bern [142]. Bernoulli [330, 2303, 169]. Bernstein [1941, 909]. Berry [1511, 403, 329, 434, 836, 2331]. Berry-Esséen [836]. Berücksichtigung [423]. Bessel [214]. Best [1046, 922, 625, 2216, 1133, 610, 532, 1208, 615, 643, 447, 1427, 932]. bester [240]. Bestimmung [176, 297, 355, 52, 79, 67, 235, 38, 411, 528]. beta [2281, 1641, 306, 646, 146, 1111, 2333, 1749, 1456]. beta- [146]. beta-binomial [2281]. betaverteilten [146]. better [1273, 2155]. betting [370]. Between [1356, 518, 1738, 998, 2007, 2138, 1944, 1490, 1412, 608, 97, 1648, 1042, 1788, 1271, 1896, 2167, 1545, 1936, 255, 926, 705, 1505, 1640, 1563, 697, 1344]. Bevölkerung [324]. Bevölkerungs [273]. Bevölkerungs-verteilung [273]. Beweis [28]. Bewertung [110, 554]. Bhattacharyya [1603]. biadditive [1139]. Bias [1569, 1898, 109, 1308, 97, 1137, 787, 1097, 274, 433, 1768, 1566]. Bias-corrected [1569]. biased [2217, 1952, 1930, 1681, 1554, 1984]. Biases [563, 1263]. Bickel [2249]. big [2163]. biggest [210]. bilinear [1192, 1701]. bimatrix [1641]. binary [2178, 699, 1746, 2229, 2133, 1955, 1623]. bindings [556]. Bindungen [556]. Binomial [1078, 193, 2322, 86, 99, 91, 145, 375, 561, 2281, 831, 1570, 1330, 2081, 682, 2052, 657, 2042, 806, 658, 2282, 993, 1444, 1417, 573, 2136, 822]. Biologie [31]. biology [31]. biometrical [7]. Biometrische [7]. bipartite [2341]. Birnbaum [1919, 2019, 2198, 1999, 2095]. birth [884]. birth-death-with-immigration [884]. bis [342]. Bivariate [2306, 1395, 854, 1912, 658, 1248, 1541, 1919, 2152, 186, 249, 351, 1994, 561, 632, 1320, 976, 1831, 2126, 2039, 1452, 646, 1348, 1501, 1510, 849, 1864, 1302, 2298, 1281, 1467, 657, 1075, 1485, 1857, 1217, 2134, 1949, 1808, 504, 1999, 798, 1866, 1572, 1995, 1492, 2286, 1872]. Blackwell [371]. Block [602, 863, 1893, 2341, 1750, 1322, 1791, 1892, 2247, 2015, 781, 1207, 634, 2166, 834, 1278, 1143, 1291, 1591, 1163, 2159, 199]. Blocked [1982, 1903, 1894, 1011, 2159]. Blöcken [199]. blocking [1179]. Blocks [1405, 1074, 882, 1156, 913, 1108]. Blockwise [1462]. Blomqvist [1456]. BLUEs [2027]. BLUPs [1672, 1922]. Blutdruckwerte [323]. Bonferroni [1319]. Book [755, 821, 1066, 1081, 1101, 1115, 1141, 1150, 1166, 1223, 1231, 1240, 1264, 1294, 1304, 1315, 818, 788, 807, 734, 816, 1214, 693, 805, 810, 687, 1195, 1384, 775, 736, 953, 794, 1373, 1385, 947, 585, 813, 1372, 592, 566, 672, 1370, 753, 902, 633, 928, 1196, 605, 684, 676, 797, 650, 576, 603, 747, 749, 771, 855, 894, 833, 1175, 691, 936, 767, 1478, 601, 732, 773, 800, 880, 1389, 1188, 620, 1197, 703, 777, 792, 689, 613, 965, 745, 715, 548, 1185, 758, 1387, 536, 1186, 786, 920, 1205, 823, 764, 717, 1371, 914, 617, 941, 713, 1386, 660]. Book [724, 611, 873, 850, 1388, 958, 623, 727, 1476, 696, 722, 1477, 1383, 706, 679, 1275, 839, 1187, 710, 34, 206, 62, 480, 216, 516, 40, 131, 459, 409, 473, 137, 248, 260, 87, 95, 159, 56, 432, 9, 465, 443, 76, 81, 178, 165, 296, 317, 377, 393, 49, 19, 171, 499, 143, 69, 525, 27].

Book-reviews [894, 880, 873]. Boolean [1963, 811]. Bootstrap [2344, 1228, 1416, 1750, 1791, 1385, 1892, 1462, 2345, 2290, 1911, 1006, 1464, 2238, 1796, 1906, 1940]. Bootstrapping [1717, 964, 2014, 1584, 1085]. both [2141]. Bound [1349, 2076, 946, 957, 403, 329, 294, 2331]. boundaries [946]. boundary [690]. bounded [386, 2281, 785, 804, 1125, 1852]. Bounds [831, 2354, 1282, 419, 1671, 1564, 822, 1603, 1054, 1883, 1938, 1210, 1016, 1511, 1759, 653, 1319, 1654, 1843, 559]. box [1686, 486, 2085, 970, 1828]. brain [557]. branching [1965, 942, 1796]. BRD [324]. Breakdown [1585, 1739, 1050]. Brownian [1528, 844, 1445]. Buchbesprcchungen [189]. Buchbesprechung [195, 487]. Buchbesprechungen [34, 206, 62, 480, 216, 516, 40, 131, 459, 409, 473, 137, 248, 260, 87, 95, 159, 56, 432, 9, 465, 443, 76, 81, 178, 165, 296, 317, 377, 393, 49, 19, 171, 499, 143, 69, 525, 27]. Bucklew [1372]. Buckley [1620, 2317]. bulk [217]. Bulletin [6]. Bundesrepublik [327]. Burn [1760]. Burn-in [1760]. Burr [1437]. business [24]. bzw [519].

C [1389, 1387, 1114, 1194, 271]. cable [38]. calculation [17, 15]. calculations [1789]. calculus [1532]. calibrating [1930]. Calibration [1457, 1289, 1568, 1682]. cancer [2315]. canonical [304]. capability [923, 1493, 1465, 1114, 1194, 1293]. capacity [864, 1274]. capital [121]. Capture [1777, 2068]. Capture-recapture [1777]. cards [1159]. care [1652]. Carlo [1386, 1473, 1617]. carryover [1882]. Cascaded [507]. case [1200, 1596, 1237, 1890, 2076, 1078, 996, 2125, 2243, 1486, 439, 1036, 204, 237, 985, 66, 815, 925, 133, 1012, 1331, 649, 1664, 1620, 2030, 2297]. case-cohort [1620]. case-control [1331]. cases [500]. catalogue [1327]. catch [1722].

Categorical [1359, 1032, 1536]. Cauchy [865, 1787]. causal [2270, 1282, 31]. celebrate [1333]. cell [563, 1577]. cells [2309, 1168]. Censored [1357, 1915, 2338, 1642, 2239, 976, 1072, 937, 1846, 2003, 1025, 2157, 750, 1878, 1680, 1938, 1091, 890, 1691, 2078, 1900, 1485, 468, 2102, 1916, 861, 1802, 1888, 1887, 2079, 2313, 2055, 1869, 1761, 2095]. censoring [1599, 1890, 1914, 2347, 738, 982, 2026, 2298, 1998, 1670, 2027, 1669, 1131, 1567, 728, 1234]. censorship [888]. census [1288]. cent [1624]. Centered [1349]. Central [1529, 657, 658, 555, 1625, 332]. Cerioli [1387]. Certain [1420, 1286, 1415, 688, 1135, 1326, 250, 624, 852, 2278, 701, 444, 395]. chain [1238]. chains [441, 1800, 1751]. chance [1454]. Change [927, 1343, 1355, 1668, 1190, 2212, 1489, 1953, 1313, 1076, 1923, 2049, 2221, 1062, 2044, 1717, 2237, 2014, 2045, 1106, 1136, 870]. change-over [1136]. change-point [1953, 1313, 1923, 2049, 2221, 2044, 1717]. change-points [2212]. changeover [1766]. changepoint [2293]. changes [2050, 1911, 1940, 1943, 930]. channel [879]. channels [507]. Chao [2076]. character [1107, 1558, 1093, 975]. Characteristic [1343, 962, 1755, 1314, 799, 675, 2278, 1480]. characteristics [1096, 1310, 1027, 1589, 563, 662]. Characterization [1404, 825, 2279, 1721, 2096, 568, 575, 2039, 704, 1090, 1939, 1885, 2046, 1271, 470, 417, 400, 2103, 709, 1500, 481, 512, 1736, 2079, 580]. Characterizations [1219, 669, 1378, 968, 1209, 1455, 1056, 1693, 1866, 1492, 711, 1448, 644, 916, 702, 1286, 992, 1001]. characterized [998, 921, 893]. characterizing [1441, 2170, 877]. charakterisieren [469]. Charakterisierung [422, 575, 580]. Charakteristiken [219]. charakteristischen [176]. Chart [1078].

Charts [461, 1242, 1547, 1652, 1017, 881, 1019, 1943]. Chaudhuri [705]. Chebyshev [1046, 875]. Checking [2202]. checks [1775]. chemical [1770]. Chernoffsches [202]. chessboard [1514]. Chi [1358, 584, 1711, 2016, 2240, 285, 674]. Chi-Square [1358, 2016, 1711, 285, 584, 674]. chi-squared [2240]. Choice [1287, 557, 996, 1992, 2173, 1608, 1781, 2030]. choices [1446]. chosen [2087]. Christofides [1267]. circle [1798, 1735]. circles [1644]. circulant [2032, 1381]. Circular [1892, 1279, 2149, 1009, 1252, 748, 148]. class [1689, 1542, 223, 699, 180, 2160, 1622, 632, 1658, 1260, 979, 612, 2090, 763, 884, 431, 2167, 310, 444, 1808, 1491, 291, 434, 598, 1724, 558]. Classes [1905, 290, 344, 669, 2189, 1095, 701, 2091, 1707, 184]. Classical [624, 1280, 2218, 1980]. Classification [543, 2181, 2314, 2016, 2301]. classifier [1709]. Clear [1410, 2143, 1733, 1924]. climb [1000]. Close [1358]. closeness [1778, 1959]. cluster [1787, 789]. clustered [1968, 1623]. clustering [2116]. clusters [1034, 343]. code [2086]. codes [2031]. codifference [1646]. coding [253]. Coefficient [2174, 1879, 1801, 1680, 2213, 2183, 157, 1674, 1976, 2124, 2069, 1931, 1134, 2308, 796, 491, 1392, 1632, 1811, 1498, 1775, 1748, 1851, 1663, 2206]. Coefficients [1751, 2138, 876, 1892, 477, 723, 1105]. cognition [828]. coherent [1832, 1987, 2199, 2256, 2250, 1806, 1482, 2006, 2037, 2139, 1978, 1934]. cohort [2268, 1620]. cold [1832, 2139]. Collapsibility [1016]. collected [1997]. collection [173]. Colloquium [7]. column [1149, 1797, 1074, 1873, 759, 1982, 2092]. column-orthogonal [1873]. Columns [1379, 780, 642]. COM [2025]. combination [2022, 653]. combinations [2235, 2140]. Combined [817, 2268, 1654, 573, 2227].

Combining [1148, 1288, 2064]. comlexity [956]. comments [418, 622]. common [1237, 2211, 1473, 2014, 2084, 1769]. Communication [2267, 2259]. Communication-efficient [2267, 2259]. commutative [1893]. compact [2111, 938, 911]. compared [24]. Comparing [1407, 1551, 2129, 2315, 1162]. Comparison [1351, 2294, 1063, 502, 2056, 1135, 587, 1742, 705, 642, 729, 449]. Comparisons [1369, 1550, 1927, 2123, 2316, 2332, 2023, 1280, 1954, 2038, 1809, 1042, 2013, 801, 959, 2006, 2037, 1561, 2269]. Competing [1413, 1255, 1998, 2055]. Complete [1104, 1149, 599, 1846, 2039, 2078, 2074, 1271, 1291, 1582]. completeness [701]. complex [1267, 1600, 250, 2074]. component [1832, 2234, 961, 2071, 862, 2187, 1628]. components [1473, 1960, 917, 2199, 2205, 2256, 2296, 2250, 784, 2036, 1809, 760, 251, 1806, 2006, 2167, 2139, 1704, 2087, 1978, 1597, 1811, 1921, 1934, 1660, 1924]. componentwise [1991]. composed [2205, 2250]. Composite [2107, 2089, 2334, 637, 2273, 680, 2144, 2259]. compositional [1662]. Compound [306, 1779, 802, 1820]. compounding [2004]. Comprehensive [1201]. compromise [2073, 697]. Compromised [1051]. Computation [1578, 2349]. Computational [466, 670, 1477]. compute [1121]. computer [1873]. computers [15]. Computing [1718, 1118, 1986, 191]. concave [2140, 1418, 1426, 946, 2036]. concavity [1688]. concentration [523]. concept [1989, 371]. Concepts [1477, 1585]. Concerning [198, 127, 157, 429, 661]. concomitant [1890]. Concomitants [1395, 1895, 1485, 694, 1866]. concordant [649]. concurrent [2287]. condition [1347, 2171, 1834, 1869]. Conditional [711, 1072, 1367, 2156, 2033, 992, 1409, 1975, 1292, 2328, 1951, 921, 583, 952, 2098, 2021, 568, 1441, 2039, 1987, 1617, 2219, 1010, 2205, 2256, 1732, 2191, 630, 1593, 2108, 2252, 1137, 1001, 1455, 1693, 481, 877, 2116, 718, 1705]. conditionals [2306, 1912]. conditioned [112]. conditioning [795, 744]. Conditions [496, 770, 1594, 967, 1055, 2079, 1244]. Confidence [1759, 134, 186, 1716, 903, 738, 1835, 1493, 2022, 2084, 1465, 2201, 2042, 815, 2315, 1707, 1334, 2238, 1781, 2030, 2297, 1985, 2361]. confirmation [92]. conflict [2202]. conflicting [1233, 1638]. confounding [1854, 2258, 2099, 1789, 2279, 2359]. congruential [814, 904]. connected [664]. Connection [1545, 1936]. Connections [1640, 1678, 1505]. considered [648]. considering [1689]. Consistency [596, 730, 872, 2244, 1230, 1342, 444, 500, 2136, 739, 2328, 1338, 1988, 1691, 2191, 531, 278, 899, 1256, 1942, 2053, 2262, 1619, 1833]. Consistent [1534, 1192, 302, 457, 1040, 1306, 1250]. consisting [2256]. constancy [1674]. constant [2048, 893, 1743]. constants [1046]. Constrained [1140, 2202, 1986, 2264, 1443, 2238]. Constraint [1369]. Constraints [1408, 1366, 1686, 2226, 534, 1153, 181]. constructed [2200]. Constructing [1420, 2086, 2138, 2031, 1854, 1028]. Construction [2194, 1322, 1877, 1170, 1795, 804, 2112, 2236, 1328, 1894, 1679, 2232, 1160, 2359, 1442, 829, 1716, 2122, 1502, 759, 1220, 132, 138, 1495, 1557, 2131, 2103, 1024, 1781]. Constructions [1397, 2341]. consulting [2279]. consumer [843]. consumption [38, 274]. Contacts [21, 42, 63, 82, 103, 124, 144, 160, 179, 200, 221, 242, 261, 282, 303, 320, 341, 362, 383, 404, 425, 446, 467, 488, 509, 530, 551, 572, 593, 614, 635, 656, 677, 698, 719, 740, 761, 782, 803, 824, 845, 866, 887, 908, 929, 948, 966, 973, 980, 987, 995, 1002, 1008, 1014, 1021, 1029, 1038, 1045, 1052, 1059, 1067, 1073, 1082, 1088, 1094, 1102, 1109, 1116, 1130, 1142, 1151, 1158, 1167, 1176, 1189, 1198, 1206, 1215, 1224, 1232, 1241, 1249, 1257, 1265, 1276, 1285, 1295, 1305, 1316]. Contacts [1332, 1341, 1350, 1361, 1374, 1390, 1400, 1411, 1421, 1430, 1440, 1450, 1459, 1468, 1479, 1488, 1497, 1507, 1517, 1524, 1540, 1548, 1556, 1565, 1574, 1583, 1592, 1602, 1612, 1621, 1629, 1637, 1647, 1656, 1667, 1675, 1685, 1694, 1703, 1712, 1719, 1727, 1737, 1745, 1753, 1762, 1773, 1782, 1790, 1799, 1807, 1813, 1822, 1830, 1838, 1845, 1853, 1860, 1867, 1874]. Containing [1410, 1733, 1924]. Contaminated [1377, 1399, 1133, 1683]. contamination [2287, 869, 2265, 1344]. content [1122]. content-corrected [1122]. contingency [2329, 99, 284, 366, 952, 1622, 1932, 1713, 1585, 1983, 33, 30, 37, 1026, 420]. continuity [1815, 2282]. continuous [1965, 108, 2126, 569, 483, 451, 1145, 391, 2169, 489, 5, 2075, 1595, 1819, 1464, 1572, 1908, 1664]. contoured [1161]. Contradictions [1031]. contrast [1810, 380, 811, 1555, 403, 278, 329, 500]. contrasts [863]. contribution [290, 392]. contributions [861]. Control [1652, 1078, 1407, 1242, 1313, 1551, 1063, 1943, 1331, 1268, 1819, 1162, 638, 1550, 716]. controlled [2352]. controlling [1905]. conventional [455]. Convergence [942, 484, 2189, 198, 1842, 2141, 2185, 1533, 1329, 2318, 555, 397, 1437, 1055, 482, 1244, 1665]. converging [154]. converting [1435]. convex [233, 283, 2066, 785, 1883, 2036]. convex/concave [2036]. convolution [1780, 1820, 594]. convolutions [1926]. cooperative [346]. Copas [2193]. Copas-like [2193]. Copenhagen [130]. copula [2098, 2135, 2013, 1949]. copula-based [1949]. Copulas [1420, 1541, 1756, 2242, 2295, 1714, 1618, 1290, 1615, 1808]. Corrected [1423, 1569, 1122]. Correcting [2325]. Correction [2038, 2172, 109, 2282]. corrections [1242, 886, 1127]. correlated [385, 1892, 1279, 1570, 791, 354, 310, 394, 1164, 349, 1514, 1882, 2207, 1566]. correlation [2138, 180, 2048, 1412, 298, 1976, 957, 1134, 2308, 1936]. correlations [1636, 15, 2288]. correspondence [2035]. corresponding [1896]. cosine [2263]. cost [577, 1986, 2360, 173, 191, 675]. cost-constrained [1986]. cost-effective [173]. Cost-optimized [675]. Costa [636]. Costa-optimized [636]. costs [579, 173]. count [2322, 2068, 2164, 2002, 2047, 2100]. countably [1834]. Counting [169]. counts [2281]. covariance [44, 51, 58, 1797, 2109, 704, 2299, 2168, 1098, 250, 1726, 1764, 2234, 1529, 1018, 1552, 2057, 2358]. covariances [2024, 563, 1560, 1284]. covariate [2281, 1901, 1902, 2287, 1091, 2224, 2141, 2136, 1606, 2082]. covariate-driven [2281]. covariates [1518, 2115, 1627, 1123, 2129, 1715, 2353, 1950, 1709, 1431, 1861, 1925]. covariation [1817]. coverage [1849, 2238]. Cox [2085, 1606, 1828]. CP [2004]. crack [1657]. Cramér [1375, 198, 1266, 1650, 654]. Cramér-type [1650]. credit [1530]. Criteria [1380, 1312, 452, 1260, 1826, 1678, 1193, 2350]. Criterion [1563, 2003, 515, 655, 1004, 1090, 1063, 2335, 1836, 1165]. Criterion-robust [1563]. critical [1965, 1796]. cross [1149, 1087, 2014, 2124, 1484, 1284, 1995]. cross-covariances [1284]. cross-ratio [1995]. cross-sectional [2124]. cross-sectionally [2014, 1484]. crosses [1591]. crossings [690]. crossover [2175, 1882]. cubic [848]. Cuboidal [1825]. Cucconi [2102]. cum [876, 226]. CUMIN [1547]. Cumulative [2320, 2101, 2329, 1951, 1435, 2062, 1945, 2218, 2134, 1784]. cure [2025]. current [2069]. curtailed [591]. curve [1815, 1023, 1741, 1091, 1956, 1936, 909]. curved [860]. curves [1083, 2129, 2233, 1058]. CUSUM [1078, 969, 1875]. CUSUM-procedure [1875]. CVaR [1812]. cycles [128, 135]. cyclic [1549, 725, 695, 1033].

D [371, 1371]. D. [371]. damage [589]. Dantzig [1758]. daraus [277]. Darling [1875]. Darstellung [338, 273]. Data [1489, 2174, 2283, 1878, 1357, 1657, 1359, 1393, 1915, 2068, 1997, 1653, 1951, 1642, 1755, 2105, 1270, 1119, 91, 1812, 2164, 2217, 2098, 1722, 2304, 876, 1964, 976, 1644, 1631, 1226, 461, 2021, 1846, 2132, 2230, 1523, 644, 2202, 2029, 1025, 2285, 1227, 2097, 1899, 1913, 2126, 2157, 1099, 750, 1152, 2163, 670, 2041, 2181, 2219, 1600, 1123, 1032, 1183, 2043, 1662, 1048, 1680, 1863, 1031, 1938, 1091, 2002, 24, 2051, 1330, 2276, 1668, 514, 1577, 1726, 2141, 2124, 2069, 1649, 1683, 1168, 2268, 2314, 2315, 1496, 1047, 1387, 1699, 2016, 2102, 1968, 2054, 1635, 1911, 1940, 2218, 2327]. data [1011, 1522, 2231, 2028, 1428, 955, 1887, 1720, 1731, 1536, 2210, 2321, 2301, 2245, 491, 1623, 1875, 1572, 1908, 2079, 1256, 2358, 1033, 970, 2055, 1026, 2153, 2100, 2259, 2262, 1620, 1851, 1869, 1761, 1925, 1344, 1888]. data-based [1899]. dating [489]. DDA [1843]. DDR [324, 271]. death [884]. deciding [1608]. Decision [483, 1339, 609, 612, 1905]. decisions [636]. decomposition [1710, 2343, 14, 598]. decompositions [940]. deconvolution [1810]. decreasing [1917, 1426, 1564]. Dedication [151]. deep [2351]. Default [1756]. defects [799]. deficiency [2224, 1711]. Defined [1356, 1955]. definite [986]. Definition [733, 316]. definitive [1894]. degree [1941]. degrees [878]. deliverables [191]. Demographie [123]. Demometrie [123]. denoising [1842]. denoting [243]. densest [2180]. densities [1424, 2039, 804, 1426, 448]. Density [1367, 1377, 748, 1768, 739, 596, 678, 1677, 1907, 2040, 1722, 943, 2249, 2336, 1820, 1747, 2332, 2226, 1684, 1757, 1250, 1261, 1511, 1683, 2074, 889, 444, 1564, 1533, 1925]. departure [66, 339]. departures [394]. Dependence [1147, 1831, 1889, 1074, 1800, 1601, 1531, 1786, 2135, 1725, 1283, 2351, 1436, 1261, 2124, 977, 2061, 1456, 1582]. dependency [770]. Dependent [1416, 2294, 1519, 1446, 1812, 2098, 2148, 2287, 1576, 1048, 2014, 1668, 1460, 2036, 1593, 1462, 972, 1060, 1105, 1291, 1802, 332, 1891, 1484, 2302, 1918, 2262, 2089, 1453]. depending [339]. depth [2345, 1657, 1005, 1057]. deren [43, 30, 37]. Derflinger [1384]. Derivation [1320, 1307, 1922]. derivatives [1466, 1203, 2288, 683]. derived [304]. Design [2272, 1803, 1835, 830, 491, 1037, 1379, 846, 1518, 1730, 863, 141, 1112, 876, 1588, 881, 1221, 939, 859, 1823, 991, 1199, 634, 791, 390, 1894, 1004, 1090, 1613, 913, 438, 251, 1153, 1581, 926, 1682, 1724, 2220, 1160, 1480, 1607, 1165, 1193, 1453, 1085, 1371]. Design-based [1803]. design-cum-model [876]. design-optimality [991]. designed [1600]. Designs [1351, 1368, 1418, 1381, 1405, 1369, 1419, 1213, 1353, 1415, 1132, 1410, 1272, 1312, 1329, 664, 1318, 688, 699, 829, 812, 892, 2272, 1321, 1746, 1929, 1149, 1562, 1798, 2032, 2194, 1074, 2257, 2341, 1323, 1322, 1586, 2122, 109, 808, 1992, 776, 1841, 1877, 2031, 2200, 2247, 1170, 1279, 1429, 2015, 1718, 751, 848, 1502, 1873, 1095, 1795, 1854, 720, 780, 1087, 2143, 847, 1260, 1611, 1986, 2011, 2208, 1110, 2086, 2112, 2195, 1156, 1225, 1555, 2118, 2258, 2154, 781, 1207, 2236, 2292, 1328, 1770, 2166, 1495, 1557, 2150, 834, 1180, 1278, 1274, 1287, 2131, 2323, 852, 1920, 2099, 1679, 1776]. designs [2065, 2128, 1080, 977, 760, 967, 835, 1068, 2284, 1554, 1521, 1654, 1143, 2222, 1291, 2335, 1254, 1444, 2093, 1178, 1331, 1888, 1887, 2277, 926, 401, 1591, 1766, 2175, 2173, 642, 1163, 1108, 1425, 1162, 1563, 2280, 1514, 1136, 1269, 1982, 2220, 2263, 1789, 970, 1882, 1327, 1856, 2227, 1903, 2092, 2180, 2073, 1298, 1566, 2214, 2137, 2359, 1733, 1924, 2159]. desirability [1346]. destructive [2025]. detectability [1698, 2042]. Detecting [2212, 930, 1943, 2254]. detection [1953, 2049, 2293, 2221, 1062, 2014, 2045]. detector [870]. determinant [1042, 1324, 1552]. Determination [1079, 1207, 1330, 401]. determine [667]. determined [795, 744, 1993, 1590]. Determining [38]. deterministic [871, 1527]. deutschen [120]. Deutschland [327]. Deutungen [458]. development [128, 135, 327, 208]. developments [820]. deviancy [2329]. deviation [562, 1133, 918, 2008, 237, 2345, 1650, 490, 2300, 1344]. deviations [663, 654, 2311]. device [2105, 1449]. DFRA [1843]. Diagnostic [1071, 2193, 2322, 1972, 2337, 917, 1016, 1759, 1137]. diagnostics [997, 1069, 1544]. dial [477]. diallel [1149, 1087, 1591]. dice [1825]. Dichtefunktionals [462]. Dichten [528, 454]. Dickey [1981]. dieser [156]. Dieter [1373]. differ [257]. Difference [237, 314, 1118, 608, 97, 1557, 815, 825, 2017, 433, 1851]. differences [44, 51, 58, 729, 44, 51, 58]. different [2205, 2250, 1698, 173, 1859]. differentiability [911]. differential [1730, 2058, 2332, 1527, 2150, 301]. differentiated [637]. diffusion [2058, 2051, 2276, 1553, 1744]. digraph [1722]. dimension [1643, 1747, 1826, 2116, 2206]. dimensional [1651, 1035, 2228, 2132, 2230, 2029, 2179, 1865, 1801, 2352, 974, 2267, 1243, 2016, 2054, 1164, 2144, 208, 2203, 2358, 2207, 2206, 2361]. dimensionaler [48]. dimensionality [723]. dimensions [2357]. diminishing [1424, 1009]. direct [978]. directed [538, 429]. direction [1601, 2012]. directional [1730, 1636]. direkter [437]. Dirichlet [2121]. discontinuous [678, 1126, 396]. discounted [610]. Discrepancies [1356, 1654]. Discrepancy [1349, 1929, 1086, 1996, 2112, 1254, 311]. Discrete [1367, 1254, 3, 2209, 1441, 346, 872, 1486, 1504, 1481, 1611, 2024, 2205, 2256, 2296, 1443, 1478, 994, 951, 358, 1467, 32, 783, 1701, 1319, 133, 884, 1217, 1131, 1079, 2173, 1258, 1427, 1464]. discrete-time [1701]. discretely [2199]. discretization [2251, 793, 2169, 2075]. Discriminant [2133, 1362, 32, 2144, 295]. discriminating [1724]. Discrimination [1382, 1226, 1563]. discriminatory [71]. Discussion [1819]. diskrete [207, 43]. dispersed [2306]. dispersion [180, 1723]. Dispersionsmaßen [422]. Dispersive [1660, 1580]. displacement [1682]. dissimilarities [543]. Distance [1004, 1376, 885, 1988, 2002, 1880, 2150, 1063, 2077, 2222, 535, 1949, 448, 369, 2116, 2180, 550, 2172]. distances [1549, 255]. distinct [1499]. distinction [1344]. distorted [2213, 2006]. Distortion [1344, 2233, 1665]. Distributed [2289, 1571, 2199, 146, 2234, 2267, 889, 748, 597, 2301, 2259]. Distribution [1395, 1307, 1941, 1639, 1358, 1391, 1404, 1726, 204, 287, 1394, 1134, 712, 1376, 2245, 1416, 1200, 39, 1303, 1915, 2068, 1951, 1803, 2120, 544, 2239, 533, 562, 145, 368, 375, 2096, 1133, 2217, 2283, 1722, 1994, 1914, 1320, 903, 976, 1076, 1072, 2021, 865, 1735, 2146, 285, 1846, 935, 1840, 738, 840, 854, 575, 580, 2126, 1947, 2110, 2012, 750, 831, 389, 762, 1077, 1570, 702, 1286, 2226, 2219, 1086, 849, 1996, 742, 2, 211, 860, 1785, 1863, 515, 1302, 1435, 1111, 1534, 182, 250, 624, 886, 890, 197, 616, 595, 2078, 994, 1937, 2104, 1281]. distribution [1300, 439, 1732, 397, 2155, 1772, 2101, 1436, 1577, 2127, 237, 2027, 300, 213, 263, 1145, 1172, 1009, 1301, 1503, 468, 1983, 1777, 1857, 2308, 658, 2282, 503, 520, 1251, 1707, 2001, 1252, 2218, 2134, 187, 374, 2028, 490, 825, 1005, 1568, 255, 625, 256, 504, 291, 1767, 627, 2075, 257, 1439, 709, 173, 191, 482, 2177, 1595, 1039, 1567, 512, 541, 1346, 1933, 2079, 2312, 2324, 1245, 1666, 2047, 313, 1664, 591, 1723, 718, 1721, 550, 944, 1296, 2095, 822, 39]. Distribution-free [1726, 712, 2245, 255, 257, 313]. Distributional [1114]. Distributions [795, 921, 744, 1420, 1848, 198, 245, 1181, 1396, 1689, 1542, 1182, 1844, 711, 2306, 998, 1347, 752, 1147, 1246, 1103, 2209, 2211, 2316, 1219, 1448, 1603, 1963, 1017, 1774, 1140, 561, 632, 937, 1222, 1441, 881, 2332, 1641, 1734, 413, 1831, 306, 2346, 644, 1946, 1827, 1899, 916, 2188, 569, 790, 1883, 1957, 2354, 2299, 646, 1348, 1501, 1688, 885, 2142, 1510, 2023, 1161, 1426, 1000, 1864, 2022, 2024, 475, 979, 1661, 1111, 2360, 1443, 1952, 2084, 992, 2078, 969, 1124, 652, 2052, 951, 1467, 1648, 2072, 657, 1437, 1769, 962, 2108, 2198]. distributions [1711, 1683, 1779, 1155, 2149, 1904, 2085, 686, 1587, 1912, 2006, 2083, 2255, 968, 1001, 1209, 1455, 884, 508, 2327, 1217, 1018, 1578, 1056, 1825, 1491, 256, 434, 1263, 875, 1505, 257, 1573, 1693, 1464, 1736, 1866, 1995, 1666, 1608, 510, 877, 1492, 1705, 582, 1204, 558]. disturbance [2176]. Divergence [1366, 1359, 1907, 2040, 538, 429]. divergences [1944, 1075]. diverging [1811]. divided [237]. divisibility [1689]. divisible [1510, 634]. division [1247]. Dixon [204]. DNA [1371]. DNID [2296]. Do [917, 1100]. Dodge [411]. Doeblin [598]. Dokumentation [280]. Domain [1356, 1412]. Donoho [1057]. doppelten [55]. dose [210]. Dosis [210]. double [812, 892, 2170, 1841, 218, 433]. doubly [2109, 1025, 1871]. down [1442, 2000]. Dr [230]. drei [636]. dreidimensionale [161, 168]. drift [2058]. drifting [1019]. drifts [1550]. driven [1489, 2281]. Drop [1518]. Drop-the-loser [1518]. dropouts [2215, 1431]. DS [1178, 1379]. DS-optimal [1178, 1379]. dual [1146, 365]. duality [725]. duels [385]. dullness [541]. duplicate [55]. Durbin [1617]. durch [528, 205, 25, 65]. Durchschnitte [156]. durchschnittlichen [323, 38]. during [327]. Dutch [102]. dynamic [442, 1917, 766, 616, 1648, 1681, 1819, 1861]. dynamischen [616].

E. [1388, 35]. Early [1394, 2225]. east [2152]. ebener [343]. EBLUPs [1910, 1909]. Eckart [196, 196]. ecology [281]. Ecometrics [281]. Econometric [150, 194, 427, 616, 327]. econometrics [639]. Economic [812, 846, 892, 631, 791, 461, 881, 38, 24, 128, 135, 327, 626, 208]. economics [281, 68, 129]. ed [1383]. EDF [1209]. Edgeworth [1738, 2325]. Editorial [988, 1469, 1763, 1401, 856]. education [1037]. Effect [450, 369, 1792, 1112, 2261, 1970, 882, 2099, 1982, 2073]. effect-number [1982]. Effective [1203, 1442, 879, 173, 10, 1805]. effectiveness [285]. Effects [2174, 1342, 1410, 1242, 688, 1309, 2058, 2015, 2143, 2349, 1207, 2292, 2150, 834, 1282, 327, 2284, 1216, 604, 2182, 2210, 1500, 1765, 2190, 1922, 1623, 1984, 1910, 1909, 1882, 2082, 1733]. Effekten [604]. Efficacies [1402]. Efficiencies [1318, 1402, 2046, 754, 926]. Efficiency [2056, 1042, 1164, 2272, 249, 1135, 726, 1253, 1279, 1260, 1168, 2278, 419, 1939, 1143, 1561, 220, 456, 1334, 1514, 46]. Efficient [1255, 2174, 276, 1992, 2069, 1417, 1423, 2175, 1724, 1628, 1321, 1718, 681, 2183, 784, 2267, 967, 955, 1162, 2259]. efforts [173]. Efron [1009]. EFSBP [1739]. Ehrenmitglied [230]. Ehrung [88]. Eigenschaft [52, 79]. Eigenschaften [336]. eigenvalue [1529]. eigenvalues [950]. Eight [977]. Eight-run [977]. einem [636, 101, 392, 241, 228, 268, 35]. einfache [537, 10, 653]. einfacher [28]. Einheiten [173]. einige [43, 275, 526, 328, 392, 10]. einiger [652, 101]. einparametrigen [582]. einschlägigen [110]. Eisenhart [634]. electronic [15]. elektronischen [15]. elementaren [152, 172]. elementary [900, 494]. elementiges [353]. elements [1271]. elementwise [1230]. elementwise-weighted [1230]. Elfving [2128, 1243]. Elliptic [1382]. elliptical [1876, 1844, 1470, 979, 2085, 1496, 343]. elliptically [1571, 1899, 790]. Elliptische [343]. Empiric [714]. Empirical [1923, 1238, 2029, 1865, 1680, 1343, 1466, 2197, 2254, 1632, 1755, 1314, 2157, 2193, 2260, 2278, 1515, 444, 1428, 1268, 1427, 1975, 1851, 697]. empirischen [219, 493]. END [2005, 1935]. ended [218]. endlich [537]. entire [125]. Entrepreneurial [155]. Entropie [12]. Entropy [1365, 254, 445, 1236, 2332, 931, 1210, 2062, 429, 1145, 1784, 12]. Entscheidungen [636, 115]. Entscheidungskriterien [202]. Entscheidungsproblem [202]. Entwicklung [327, 120]. Entwicklungen [328]. environment [1809]. environmental [192]. EPC [1019]. epps [742]. epps-pulley [742]. equal [584, 1288]. equalities [1614]. equality [2329, 1907, 351, 382, 2299, 1672, 2084, 1858, 1635]. equation [1730, 2130, 567, 451, 301, 932]. equations [2058, 1527, 429]. equi [2306]. equi-dispersed [2306]. equidistant [1679]. Equidistribution [904]. equipped [1832, 2139]. Equivalence [2093, 1312, 2274, 2167]. equivariant [2285, 809]. equivocation [507]. Erdos [1875]. Erdos-type [1875]. Ereignisse [262, 372]. Ereignissen [559]. Ergebnissen [537]. ergodic [2241, 1993, 436, 471]. ergodicity [1751]. Erhebungen [334]. Erhebungskosten [173]. Erkennen [94, 828]. Erklärung [120]. Erlang [969]. Ermittlung [334]. Erneuerungsprozesse [474]. Erratum [1902, 1940, 1888, 1910]. Error [1423, 949, 1673, 2338, 1792, 1023, 1061, 768, 2249, 891, 2156, 616, 1191, 2104, 2319, 784, 2353, 653, 961, 2327, 163, 1097, 1484, 2288, 1427, 1608, 1638, 2107, 1886, 2082, 2137, 1628, 1659]. error-in-variable [2107]. Errors [1364, 1571, 1040, 1627, 2196, 427, 2219, 2011, 2213, 997, 2266, 1770, 619, 1233, 2166, 407, 1880, 2129, 652, 1192, 1230, 2248, 2218, 648, 1696, 457, 1164, 1514, 1606, 1882, 1980, 2089, 2207, 1566, 2214, 2186, 1663, 1453]. Errors-in-Variables [1364, 1040, 619, 1192, 1230, 648, 1980]. Erschöpftheitsbegriff [371]. erster [616]. erwartungstreuer [405]. Erwartungswert [342]. Erwartungswerte [61, 501]. Erweitertes [330]. Erzeugung [146]. Esseen [1511, 329, 434, 403, 836, 2331]. Esseen-type [434]. estimate [2339, 2275, 389, 267, 433, 1805, 2262, 5]. estimated [1812, 1017, 2349]. Estimates [685, 1121, 1337, 976, 1644, 380, 435, 885, 554, 97, 1250, 662, 1764, 1261, 1511, 403, 419, 278, 289, 329, 359, 447, 909, 1623, 1942, 1690, 1768, 2317, 1665, 2095, 1833]. Estimating [876, 1499, 1901, 2340, 982, 772, 2237, 1010, 2050, 1509, 2010, 1137, 498, 1114, 1646, 1933, 944, 2211, 1558, 495, 939, 1099, 1213, 1155, 1211, 1904, 438, 586, 925, 907, 1126, 763, 986, 1457, 627, 2288, 1765, 491, 1902]. Estimation [1242, 1596, 1571, 1107, 1120, 1157, 2174, 1395, 1093, 1716, 943, 1335, 1076, 891, 865, 2097, 2179, 2219, 1377, 1794, 2213, 1340, 1091, 931, 979, 1098, 890, 298, 1300, 1772, 2124, 1771, 300, 1590, 1604, 2149, 673, 915, 468, 2290, 628, 1015, 2001, 311, 1520, 1999, 1097, 2075, 963, 1439, 542, 184, 1258, 1425, 981, 1392, 1796, 697, 1634, 596, 868, 950, 743, 1997, 1677, 1673, 1237, 2060, 1598, 1803, 544, 533, 224, 599, 2217, 2283, 919, 1994, 1627, 2246, 632, 971, 1070, 700, 552, 1061, 741, 1588, 864, 461, 2021, 2336, 1820, 2111, 1747, 1658, 2058, 2287]. estimation [1970, 1523, 2170, 1941, 785, 1810, 766, 1697, 1899, 1948, 2337, 1913, 960, 872, 1990, 750, 1783, 2156, 2226, 1481, 2349, 608, 1123, 849, 1714, 2043, 1839, 2168, 1655, 1881, 2298, 1461, 2183, 2264, 1435, 1661, 1848, 1534, 2360, 1443, 2121, 595, 619, 956, 2113, 407, 2224, 2104, 1192, 29, 1516, 1715, 1274, 1438, 1758, 2201, 2234, 2141, 450, 412, 2267, 2289, 2018, 2069, 1950, 1855, 1339, 783, 1711, 1683, 1701, 938, 1168, 2106, 1145, 2074, 476, 760, 565, 686, 820, 2315, 251, 1777, 1692, 503, 508, 520, 1472, 1251, 302, 2218, 2327, 1018, 1669, 2134, 1949, 748]. estimation [540, 1956, 309, 187, 827, 647, 1106, 2305, 1568, 955, 641, 1696, 1417, 457, 1720, 862, 769, 778, 220, 226, 539, 606, 789, 379, 274, 2210, 867, 1064, 909, 491, 16, 2357, 1113, 1138, 316, 905, 714, 1256, 1744, 1811, 1849, 2326, 30, 37, 1606, 313, 1706, 1977, 2331, 2153, 2107, 2089, 2030, 2297, 1234, 1861, 1628, 582, 1013, 1925]. estimations [652, 1614]. Estimator [1364, 1355, 1342, 739, 1118, 344, 678, 2328, 2068, 1951, 1040, 1306, 2040, 2076, 871, 919, 726, 874, 960, 2208, 2325, 1662, 2330, 1702, 2343, 2161, 2351, 1691, 1230, 2191, 1900, 1593, 2117, 832, 2108, 2268, 479, 2255, 531, 1194, 615, 643, 1055, 190, 1552, 1508, 1097, 460, 2310, 975, 1333, 606, 1595, 227, 2057, 1244, 2311, 718, 1620]. Estimators [1473, 1357, 922, 1353, 1376, 223, 290, 344, 1876, 2105, 1119, 581, 1792, 2160, 1633, 2098, 1820, 1083, 669, 1827, 2189, 645, 2110, 681, 811, 1988, 1589, 1938, 2002, 2051, 1616, 1233, 837, 1191, 695, 563, 132, 138, 1124, 1044, 974, 439, 784, 817, 1670, 1042, 1788, 983, 1155, 1594, 1126, 1434, 763, 889, 721, 431, 1522, 448, 500, 625, 1533, 798, 482, 2177, 590, 1969, 177, 293, 312, 1500, 899, 148, 1979, 2053, 2059, 1850, 1024, 1026, 1638, 697, 1870, 1453, 1619, 1824]. Eur69.95 [1370]. Europäische [150]. European [194]. EV [2034, 1632]. Evaluating [1360]. Evaluation [1229, 553, 554]. Evaluations [1957, 1898]. Event [1372, 1964, 2041, 33, 1201, 25]. events [905, 559, 372, 1373]. evidence [1125]. evolution [1383]. evolving [1471]. EWMA [1943]. Exact [1642, 1914, 1798, 847, 2027, 754, 2017, 1538, 86, 99, 284, 1797, 182, 2078, 2046, 1181, 1298]. exactly [440]. exakten [205]. examination [584]. exceedance [225]. Exceedances [1967, 1559]. exceptional [1643]. excess [690]. exchange [1118]. exchangeability [1576, 1989]. exchangeable [2109, 1639, 1007]. exclusion [1320]. exhaustion [371]. exist [599]. Existence [1364, 1549, 1815, 751, 71, 1623, 2095]. Existenz [388, 57, 71, 240, 405]. exists [1551]. Exogenous [2103]. expansion [1738, 2325, 682, 708]. expansions [2243, 1084, 1725, 1266]. expectation [481]. expectations [1360, 921, 583, 1441, 1883, 1957, 2354, 1210, 1564, 1843, 877]. Expected [1366, 916, 946, 985, 1228, 1001, 401]. Expectile [2231, 2328, 2293]. experience [102]. experimental [1929, 776, 1718, 2150, 2128, 835, 2093]. Experiments [912, 1336, 1149, 141, 1645, 158, 1841, 1873, 1095, 1087, 1986, 1555, 759, 1287, 1009, 1521, 1153, 1011, 1212, 2173, 2187]. explanation [208]. explanatory [648]. explicit [1633]. exploratory [1033]. Exploring [1738, 1032, 1387]. explosive [2176]. exponent [1839]. Exponential [1422, 1297, 482, 1935, 1237, 1538, 1599, 1642, 2251, 2239, 533, 581, 2235, 1994, 1914, 1562, 976, 1490, 937, 522, 766, 1025, 580, 1760, 217, 2226, 860, 1027, 1881, 2022, 1302, 1661, 286, 1210, 197, 2078, 1281, 1300, 439, 496, 2101, 204, 237, 985, 783, 1826, 686, 815, 925, 2315, 825, 1181, 1567, 510, 1721, 1872, 1660, 838]. Exponentialfamilien [348, 240]. exponentiality [1636, 1280, 1939]. exponentially [889]. Exponentialverteilung [580]. Exponentialverteilungen [94]. Exponentiated [1307, 2148]. exponentielle [43]. Exponentieller [164, 122]. extended [1741, 1291, 2203]. extendible [1714]. Extending [991]. Extension [1060, 1804, 1541, 2209, 2242, 1432, 589]. extensions [1147, 849, 1319, 295]. extrapolation [683]. Extrema [265]. extremal [1213, 1767, 510]. Extreme [954, 2055, 2294, 1786, 1785, 1863, 1251, 731, 1754]. extreme-value [1786, 1754]. Extremes [1373, 629, 1246, 2148, 668, 1573, 1872]. extremum [939].

factor [2228, 1309, 276, 1095, 2011, 2182, 1765, 1982, 1924]. Factorial [1368, 1419, 1410, 1272, 1312, 1336, 1645, 1179, 1309, 1586, 1841, 1156, 759, 2292, 1328, 1080, 977, 1011, 1254, 1212, 2263, 1327, 2359, 1733]. Factorials [1349, 1678, 1640]. Factors [1368, 882, 2274, 1095, 2292, 2335, 1575, 1298]. fail [1660]. fail-safe [1660]. failed [1076, 2199, 2296]. Failure [1077, 1642, 2202, 2354, 1426, 2296, 1299, 2271]. Failures [1394]. Falk [1373]. fall [1993, 239, 246]. Falle [33, 41]. Familien [244]. families [1321, 2251, 581, 1893, 1314, 1883, 804, 860, 1443, 1616, 2101, 1826, 289, 721, 1202, 798]. Family [1355, 2209, 2316, 1580, 1722, 943, 2242, 766, 733, 1510, 2023, 1881, 1302, 1661, 2360, 837, 783, 1711, 1181, 1886]. Farlie [1000]. FDR [1905]. feature [2033, 1976, 2016, 2054]. Federal [327]. Fehlerabschätzung [381]. fehlerbehafteten [464]. Fehlerhinweis [549]. Fehlern [652]. Fehlerrechnung [107]. Felder [485]. Fertigung [241, 268]. fewer [1992]. FGM [1541, 1147, 1420]. Fiducial [1980]. field [907, 1521]. fields [1658, 1937, 1476]. filling [2350]. filter [2045]. filtering [766]. finance [1478, 1743]. finding [2187]. Finite [435, 1057, 1447, 1273, 664, 1107, 1347, 2316, 495, 1238, 960, 906, 918, 1481, 1611, 1589, 72, 132, 138, 298, 662, 1516, 412, 2149, 438, 498, 586, 1482, 628, 1581, 618, 830, 564, 769, 963, 606, 590, 1258, 1427, 177, 349, 1608, 1050]. finitely [537]. fire [385]. First [1364, 776, 2150, 1798, 1220, 616, 1885, 801, 1766, 1664]. first-order [1798, 801]. Fisher [1089, 70, 2115, 430, 766, 2012, 1183, 182, 96, 1604, 1009, 1485, 2255, 1173, 1916, 521, 2079]. Fit [1367, 752, 1821, 1755, 2120, 2096, 1972, 1878, 2142, 1380, 917, 2044, 2034, 1218, 994, 1071, 1684, 1757, 1880, 1937, 680, 2278, 968, 1209, 1020, 2047, 1498]. Fitting [1644, 2249, 1793, 1152, 670, 2059]. five [646, 2101]. Fixed [1353, 1588, 738, 1207, 1168, 2315, 806, 1911, 1940, 1500, 1567, 30, 37, 2136]. fixed-design [1588]. fixed-width [738, 2315]. flation [2164]. Flexible [1340, 1631]. fluctuating [4]. fluctuation [1743]. Focused [2003, 1836]. folded [2188, 1916, 1995]. folding [1053]. Foldover [1419, 2092]. Föllmer [1478]. Follow [1841]. Follow-up [1841]. following [1972]. Forced [1449]. forecast [616, 496]. Forecasting [102, 1861]. Foreword [1, 104, 321]. Formaldehyd [25, 65]. formaldehyde [25, 65]. Formale [219]. formation [132, 138, 2036]. forms [236, 1698, 986]. formula [1320]. formulae [435]. formulas [1308]. formulation [346, 680]. forward [1387]. Four [2061, 2086, 2112, 2227]. four-level [2086, 2112, 2227]. Fourier [477, 1948, 2263]. Fourier-cosine [2263]. Fourier-type [1948]. Fractional [1419, 1713, 2263, 234, 305, 1272, 1312, 1528, 345, 2241, 386, 1445, 1179, 1309, 1586, 2196, 1841, 2292, 247, 258, 294, 364, 1327, 1733, 1156]. fractionated [1336, 1212]. Frage [2, 115, 140]. Fragen [326]. frailty [1928, 2114, 1513, 1995, 1958]. frame [2064]. framework [648]. Frankfurt [8, 8]. Frankfurt/M [8]. Fréchet [1579, 1496]. free [759, 1726, 1983, 2046, 255, 712, 257, 2245, 313]. French [44, 51, 58]. Frequencies [1366]. Frequency [1412, 2097, 2245, 2331]. Frequency-domain [1412]. Friedman [475]. Frisch [84]. Frontiers [638]. führen [505]. Full [1414]. Fuller [1981]. fully [1575]. Function [1418, 1626, 1376, 344, 678, 1596, 2339, 1755, 1803, 1236, 86, 284, 214, 2217, 2320, 1562, 971, 1070, 735, 1569, 2111, 1892, 1941, 640, 1314, 1728, 982, 2165, 750, 2039, 1152, 772, 2330, 2002, 1435, 994, 682, 956, 1771, 2278, 1211, 1145, 1172, 1015, 2218, 2134, 1055, 754, 1671, 825, 1646, 1005, 1568, 1682, 2288, 2177, 1595, 274, 1053, 1064, 295, 2245, 16, 1346, 1995, 550, 2297, 944]. Functional [2174, 2273, 1829, 596, 2328, 1951, 1040, 345, 2287, 2285, 2161, 567, 2063, 2141, 2289, 429, 451, 391, 2231, 2151, 2171, 1875, 1744, 2153, 1971, 2089, 2067]. functionals [234, 305, 844, 1938, 574, 1268, 247, 258, 294]. Functions [1343, 1398, 1542, 1292, 2275, 1473, 2320, 943, 2021, 1901, 1902, 2332, 738, 799, 1688, 2219, 2033, 2298, 1443, 397, 2101, 494, 1466, 1788, 893, 962, 1590, 1650, 1126, 1482, 2233, 628, 1905, 417, 444, 255, 694, 408, 1614, 1908, 877]. Funktion [16]. Funktionen [519]. Funktionsleiter [338]. Further [218, 820, 360, 378, 1065]. future [2239, 2271, 2147]. Fuzzy [1677, 1409, 1393, 898, 1174, 1227, 1262, 1228, 1047, 1579, 990, 1256]. FWCI [2315]. FWER [1905].

G [441, 1384, 238, 269, 1371]. Galton [954]. galvanic [1695]. game [579, 212]. game-theoretic [212]. Games [1454, 346]. Gamma [741, 1783, 1391, 2105, 2275, 413, 306, 2142, 2154, 146, 245, 1155, 1926, 1439, 718, 1859, 2096, 2244, 2277]. gamma-distributed [146]. Gamma-Minimax [1391, 741]. gammaverteilten [146]. gap [916]. GARCH [2281, 1948, 1684]. Gastvorlesung [8]. Gauss [875, 575]. Gaussian [2060, 2318, 2211, 1603, 1658, 2135, 575, 872, 389, 211, 1559, 532, 250, 2084, 497, 2094, 779, 793, 1855, 468, 1858, 1491, 1476, 1862, 2059, 1723, 1463]. GD [634]. GD-PBIBD [634]. Geburten [120]. Geburten- [120]. Gedanken [115]. gegebenen [505]. Gelände [215]. Geleitwort [1, 104, 321]. gemeinsam [597]. Gemeinsame [150]. gemischten [604]. Genauigkeit [118, 119, 215]. General [2164, 364, 1689, 1542, 934, 1246, 1040, 2316, 2009, 2221, 1095, 1854, 1501, 1481, 2121, 1495, 1557, 2155, 2117, 1788, 2252, 763, 770, 1595, 1640, 1614, 1572, 1789, 396, 2359, 607, 350]. generalisation [2035, 2309, 658, 504]. Generalised [1710, 1641, 468]. generalization [196, 372]. Generalizations [2072, 408]. Generalized [1368, 1929, 1946, 1420, 1827, 1493, 1897, 1465, 1920, 1342, 1784, 875, 2215, 1245, 1431, 1204, 1272, 1182, 1292, 2212, 518, 1120, 1219, 1448, 1424, 2320, 2291, 386, 1627, 561, 1184, 1222, 1238, 2253, 854, 1170, 2019, 1253, 2337, 1054, 982, 1570, 2330, 1785, 1863, 1210, 567, 992, 837, 1998, 1300, 1266, 1080, 1297, 213, 263, 479, 1912, 1554, 1635, 2091, 1579, 400, 1999, 521, 1767, 445, 627, 709, 177, 1039, 1736, 1811, 2100, 1872, 1234, 2082]. generated [2235, 1643, 1883, 972]. generating [1292, 2320, 2002, 1494, 1172, 2220]. generation [1386, 1384]. generic [1578]. genomic [1577]. Gentle [1386, 1477]. Geodesic [1735]. geographischen [273]. geometric [2068, 1774, 2165, 1777, 2269, 1693]. geordneter [265]. German [12, 34, 39, 105, 152, 172, 90, 1, 106, 6, 142, 150, 230, 271, 272, 322, 636, 176, 206, 297, 235, 426, 83, 562, 262, 107, 355, 422, 342, 474, 519, 62, 275, 319, 162, 437, 323, 424, 60, 480, 216, 353, 347, 516, 40, 110, 131, 280, 459, 388, 469, 492, 513, 575, 580, 196, 556, 7, 8, 308, 352, 398, 409, 473, 526, 111, 506, 210, 277, 137, 161, 168, 47, 584, 2, 324, 248, 38, 371, 828, 464, 84, 370, 48, 554, 146, 260, 199, 616, 87, 95, 139, 244, 325, 485, 159, 326]. German [202, 652, 340, 343, 167, 52, 101, 96, 114, 73, 85, 183, 402, 327, 61, 411, 675, 328, 115, 79, 56, 15, 77, 98, 207, 225, 265, 264, 432, 458, 501, 653, 9, 116, 531, 33, 212, 14, 57, 43, 71, 117, 153, 465, 348, 528, 443, 140, 330, 76, 231, 415, 604, 81, 154, 178, 505, 537, 28, 239, 246, 50, 104, 151, 219, 5, 321, 331, 392, 410, 18, 279, 165, 240, 203, 626, 53, 454, 462, 296, 317, 377, 393, 597, 13, 49, 55, 64, 118, 149, 173, 191, 241, 228, 268]. German [333, 423, 414, 67, 94, 511, 622, 405, 19, 314, 334, 205, 25, 171, 16, 120, 499, 35, 126, 148, 193, 259, 487, 93, 59, 17, 270, 316, 493, 164, 559, 384, 121, 10, 335, 80, 143, 30, 37, 69, 78, 156, 336, 381, 529, 31, 65, 122, 337, 372, 229, 525, 123, 166, 273, 318, 357, 27, 338, 215, 582, 416, 558]. Germany [327, 68, 129]. Gesamtheit [90]. geschichteten [333]. Gesetzes [372]. Gewichtung [437, 164, 122]. gewisser [265]. gewöhnlichen [554, 264]. GI [441, 72]. GI/M/1 [441]. GI/M/I [72]. Gibbs [1991, 1825]. GIG [1505]. Gini [2320, 1742]. Giorgio [318, 318]. given [1290, 449, 1296]. gleichmäßig [240]. gleichverteilten [342]. Gleichwahrscheinlichkeit [584]. gleitenden [156]. global [1470, 2257]. Glücksspirale [426]. GMLE [1620]. gold [1759]. golden [1333]. Good [1385]. Goodness [1821, 1367, 1380, 2044, 2034, 1218, 994, 1757, 2047, 1498, 752, 1755, 2120, 2096, 1972, 1878, 2142, 1684, 1937, 680, 2278, 968, 1209]. Goodness-of-fit [1821, 1380, 2044, 2034, 1218, 994, 1757, 2047, 1498, 1755, 2120, 2096, 1972, 1878, 2142, 1684, 1937, 2278, 968, 1209]. goods [121]. GPSD [1814]. gradual [2237]. graph [1328]. Graphical [1043, 1035, 1800, 1452, 13]. graphischen [338]. graphisches [78]. Gray [2031, 2086]. Green [982]. Grenzwertsatz [77]. Grenzwertsätze [265]. grids [1164]. großem [78]. Größen [176, 119]. größten [210]. group [1322, 2163, 634, 2232, 2106, 1415]. grouped [1653, 29]. grouping [442]. Groups [1368]. Groupwise [2116]. Growing [1471, 2029]. growth [1023, 1741, 1058, 1936]. Grubbs [770]. Grundgesamtheit [101]. Grundgleichungen [308, 352]. Grundzüge [152, 172]. Gruyter [1478]. Guest [8]. guide [2121]. Gumbel [1000, 2242, 504]. Gut [259].

H [392, 420, 271]. H. [1478, 392, 1476]. Haar [1425]. Haar-wavelet [1425]. Haavelmo [274]. Haavelmo-bias [274]. habil [231]. Hadamard [1169, 1229, 1397, 2200, 1220, 1274, 1325]. Hain [105]. half [1915]. half-logistic [1915]. halfspace [1057]. Handbook [1477]. Hans [322, 322]. Härdle [1477, 1370]. harmonic [477, 2303]. Harrington [1346]. Hausdorff [1643]. having [2154, 245, 1749]. Hazard [1392, 998, 2123, 1868, 1348, 1688, 1426, 1466, 893, 2252, 1699, 1669, 1513]. hazards [1964, 1831, 2041, 2352, 2268, 1700, 2307, 2080, 1606, 1234]. hc [230]. health [1652, 192]. heavy [2265, 2198, 1012, 2055]. heavy-tailed [2198, 1012]. Heidelberg [1370]. Heidelberg-New [1370]. Held [1476]. Hellinger [1949]. Help [21, 42, 63, 82, 103, 124, 144, 160, 179, 200, 221, 242, 261, 282, 303, 320, 341, 362, 383, 404, 425, 446, 467, 488, 509, 530, 551, 572, 593, 614, 635, 656, 677, 698, 719, 740, 761, 782, 803, 824, 845, 866, 887, 908, 929, 948, 966, 973, 980, 987, 995, 1002, 1008, 1014, 1021, 1029, 1038, 1045, 1052, 1059, 1067, 1073, 1082, 1088, 1094, 1102, 1109, 1116, 1130, 1142, 1151, 1158, 1167, 1176, 1189, 1198, 1206, 1215, 1224, 1232, 1241, 1249, 1257, 1265, 1276, 1285, 1295, 1305, 1316]. Help [1332, 1341, 1350, 1361, 1374, 1390, 1400, 1411, 1421, 1430, 1440, 1450, 1459, 1468, 1479, 1488, 1497, 1507, 1517, 1524, 1540, 1548, 1556, 1565, 1574, 1583, 1592, 1602, 1612, 1621, 1629, 1637, 1647, 1656, 1667, 1675, 1685, 1694, 1703, 1712, 1719, 1727, 1737, 1745, 1753, 1762, 1773, 1782, 1790, 1799, 1807, 1813, 1822, 1830, 1838, 1845, 1853, 1860, 1867, 1874]. helpful [1006]. Hermite [1818, 2177]. heterogeneity [2009, 2352, 2081, 1777]. heterogeneous [2294, 1490, 2148, 1960, 2265, 2013, 1978, 2302, 1934, 1660, 1859]. heteroscedastic [2029, 2011, 1837, 2214, 1752, 2137]. Heteroscedasticity [1544]. heteroscedatic [996]. heteroscesdastic [2111]. Heuristic [1375, 1336, 1212]. hidden [1810, 2081]. Hierarchical [543, 1034]. High [2203, 2207, 2206, 1035, 2228, 2230, 2179, 1865, 1801, 2352, 2267, 1458, 2245, 2358, 2361]. High-dimensional [2203, 2207, 2206, 1035, 2228, 2230, 1865, 1801, 2352, 2267, 2358, 2361]. high-frequency [2245]. high-quality [1458]. higher [1084, 616, 1769, 1650]. Highest [2052]. highly [1336, 1212]. Hilbert [2171]. Hilfe [84, 616]. hill [1000]. histogram [933, 2181]. histogram-based [2181]. histograms [569, 490]. Historical [68, 129]. history [1539]. hit [553, 25, 65]. hit-theoretical [65]. Hochrechnung [424]. höheren [388]. höherer [616]. Hölder [1528]. Hölderian [1345]. Hommel [653]. HOMMEL’schen [653]. homogene [519]. Homogeneity [2197, 145, 375, 2230, 2126, 2078, 712]. homogeneous [2040, 754]. homogener [519]. Homogenitätsmaß [325]. Hörmann [1384]. Horvitz [223, 290, 431]. Hotelling [1128]. HR [2253]. HT [438]. HT-method [438]. Huang [1147]. human [1311]. Hüsler [1373]. Hybrid [2051, 1413, 1914, 2347, 2026, 1998, 2095]. hyper [217, 197]. hyper-exponential [197]. hyper-Poisson [217]. hyperball [1213]. hyperbolic [1985]. hypercomplex [624]. hypercube [2122, 1877, 1795, 2222, 1816].

Hyperdeterminantal [1526]. hypergeometric [213, 263, 256, 714]. hypergeometrischen [193]. hyperspace [127]. Hypothese [584]. Hypothesen [604, 537]. Hypotheses [1363, 1393, 1174, 671, 1227, 637, 1016, 779, 793, 604, 537, 990, 1385, 1388]. Hypothesis [201, 778, 898, 2228, 2109, 584, 1575, 10, 313, 2067].

i.i.d [2235]. ideal [281]. identifiability [1347, 1834]. identification [222, 232, 1442, 2178, 2119, 427, 534]. Identifikation [78]. Identifying [1702]. Identitätstest [203]. identities [1804]. identity [896, 661]. IFR [746, 1957]. ihre [327, 458, 148]. II [37, 1915, 1599, 1890, 1914, 937, 1025, 352, 1878, 634, 1670, 2027, 1251, 378, 1802, 1520, 37, 2095]. II. [869]. III [1924]. Ill [112]. Ill-conditioned [112]. illustrations [2315]. im [531, 33, 41, 239, 246, 18, 55, 273, 215]. images [2031]. Imbedded [441]. immigration [1965, 942, 884]. imperfect [1994, 1990]. Implications [1259, 2347]. implicit [1590]. imploring [653]. importance [2113]. Important [546, 2292, 1804]. Imprecise [2192, 1740]. Improved [1363, 1251, 479, 1707, 1097]. Improvement [860, 428]. Improving [721, 1910, 1909]. Imputation [1973, 2270, 1713, 1051]. IMS [142, 130]. IMS-Tagung [142]. Inaccuracy [253, 2088, 538, 1648, 1681, 1945, 429, 254, 2020, 463]. inactivation [25, 65]. inactivity [1987, 2091, 1934]. inadmissible [2346]. INAR [1918]. INARCH [2004]. incidence [913]. incident [2268]. incidental [302]. inclusion [1320, 440, 751, 1452, 1288]. inclusion-probabilities [440]. incomplete [2126, 2181, 1713, 634, 1278, 1143, 609, 955]. Inconsistency [1355]. incorporating [2309, 2150]. increase [1100]. Increasing [2140, 1100, 1426, 1584, 2036, 1826]. increment [2252]. increments [2357]. Indefinite [345]. Independence [1402, 1399, 1622, 1041, 790, 1281, 2054, 1561]. independent [222, 232, 2056, 1519, 1123, 2205, 1417]. Index [52, 1951, 2309, 1569, 2029, 923, 1493, 2033, 1680, 2213, 2273, 1668, 1742, 2018, 2108, 1904, 1114, 1194, 1259, 1293, 1020, 1346, 2326, 1824]. indices [39, 1465]. individual [2210, 2190]. individually [1720]. induced [2209, 2332, 1995]. inductive [1389]. industriellen [328]. ineffective [210]. Inequalities [2135, 1046, 1042, 1297, 875]. Inequality [1408, 1360, 583, 507, 361, 1742, 391, 756, 434, 1891, 757, 1935]. Inequivalent [1169, 1229]. Inference [1919, 1893, 2084, 1422, 1742, 514, 2028, 1642, 1651, 1965, 1914, 1072, 1487, 2192, 1791, 2344, 2270, 2163, 1474, 1161, 2276, 1389, 2224, 2353, 2027, 2158, 1009, 1047, 2083, 1699, 2185, 1968, 1259, 830, 1012, 1456, 523, 2151, 1039, 2017, 1632, 1980, 1851, 2207, 1973, 1344]. inferences [1311, 733, 1572]. inferential [1538]. inferiority [1567]. infimum [1849]. infinite [1689, 1747, 1646, 1834]. infinitely [1239, 1510]. inflated [1950]. inflates [1624]. inflation [2076]. Influence [1626, 1137, 1496, 1463, 1473, 2330, 1125]. influential [1702]. Information [1414, 1938, 209, 1089, 2240, 1746, 863, 2320, 919, 776, 2003, 1740, 1099, 1835, 1260, 1183, 2258, 1098, 538, 1148, 1897, 1649, 1208, 1604, 2074, 1485, 763, 254, 1173, 1847, 470, 1916, 400, 955, 2064, 363, 408, 445, 428, 463, 1836, 2079, 944, 838, 506]. information- [538]. information-improvement [428]. informative [982, 1431, 1234]. inhomogene [519]. initial [991, 1287]. Inner [2070]. innovation [1684]. innovations [2318, 1948, 1885, 1974, 1918]. Inset [538]. Inset-inaccuracy [538]. Inspection [1201, 1200, 1096, 1078, 900, 579]. Instantaneous [1394]. Institute [150, 194].

Instituts [230]. instrumental [2107]. insulation [38]. insurance [2337]. Integer [2184, 1610, 2281, 2221, 293, 2100, 2286]. integer-estimators [293]. Integer-valued [2184, 1610, 2281, 2221, 2100, 2286]. Integral [2142, 1193]. integrated [994]. Integration [48]. integrative [1019]. intensity [1530, 695, 2104, 1053]. inter [863, 1486]. inter-block [863]. inter-rater [1486]. interacting [913, 472]. interaction [1043, 634, 956, 1698, 1977, 1924]. interactions [1309, 1765]. Interactive [1035, 1034, 1032, 1031]. intercept [2154]. interdependent [1954, 2038]. interference [2247, 1429, 2015, 1180]. Internationalen [230]. interpenetrating [251, 588]. interpolants [787]. Interpretation [162, 808]. interruptions [269]. interval [1997, 1722, 1716, 785, 960, 1152, 2084, 2042, 2315, 5, 1849, 313, 1144]. interval-valued [1152]. Intervallbreite [5]. intervals [1915, 1835, 1493, 2022, 2052, 1465, 815, 1707, 3, 2238, 2361]. intra [180]. intra-class [180]. Intraclass [325]. intraday [2261]. Introduction [1030, 1478, 663, 1372]. invariance [770, 570, 558]. invariant [1899, 848, 1260, 1145, 1595]. Invarianzeigenschaft [558]. inventory [4, 356]. inverse [2060, 2211, 1603, 389, 211, 2258, 2084, 2094, 1855, 468, 1858, 1491, 2310, 573, 2059, 1723, 1463]. inversion [154]. inversive [904]. invertibility [621]. Investigating [1613, 1530]. investigation [1144]. Investigations [1371]. Investitionsgüter [121]. involving [583, 733, 1172]. IPPS [705]. irregular [455]. Irrtumswahrscheinlichkeitsschranke [653]. ISBN [1372, 1370, 1371]. isolating [2212]. isotonic [1432, 692]. Issue [2044, 1814]. issues [1122]. item [2246]. items [1076]. Iterated [1398]. Iterationsverfahren [154]. Iterative [1377, 1793, 154]. IV [538, 2359, 1924].

J [1384, 8, 1388, 1477]. J. [8]. Jackknife [2197, 2321]. Jacobi [196, 196, 2343]. Jahre [231]. Jahren [323]. James [1372, 1386, 1620, 2317, 1870]. Januar [7]. jedem [199]. Jefferson [1380]. Jenkins [486]. jittered [2146]. Joint [194, 903, 2041, 1559, 2216, 1320, 1964, 1631, 2168, 1587, 1896, 1252, 2312, 2324, 1723]. jointly [597]. Jordan [622]. Joseph [105, 136]. Journal [2342, 2348, 2355, 2362]. jubilee [1333]. Jumps [1530]. Jürg [1373]. Justification [1288, 1272]. justifying [1869].

Kabel [38]. Kabel- [38]. Kalman [766]. kanonischen [264]. Kantorovich [1042, 1049]. Kantorovich-type [1042]. Karhunen [1695]. Kausalanalyse [31]. Kenward [2070]. Kernel [2217, 1747, 750, 2074, 2064, 2288, 739, 1792, 2160, 1600, 1338, 2094, 2141, 560, 1956, 1268, 2321, 2171, 870]. Kernel-based [2064]. kernel-smoothing [1956]. kernel-type [2160]. kernel-weighted [1268, 2321]. Ketten [519, 47, 519, 398]. key [1252]. Keynesian [274]. kill [369]. kinds [1764]. Klasse [381, 558]. Klassifikation [355]. klassifikations [23]. klassifikations-problem [23]. klassifizierten [340]. kleinem [270]. kleinen [61]. Knockoffs [2352]. knot [2183]. knowledge [609]. known [452, 1098, 191]. Koeffizient [325]. Kolmogorov [1313, 885, 133, 396]. kombiniertes [23]. Konferenz [150]. Konfidenzgrenzen [116]. Konfidenzintervall [277]. Konfidenzintervalle [43, 85]. Konsistenz [454]. Konstruktionsbasis [162]. Kontingenzkoeffizient [33, 41]. Kontingenztafeln [41, 30, 37, 33]. kontinuierlichen [411]. Kontrollbehandlungen [199]. konvergierendes [154]. Konzentration [275, 328]. Korn [1371]. Korrektur [487].

Korrekturen [336]. Korrelationen [15]. Korrelationskoeffizienten [185, 264]. korrelativer [113]. Kosten [191, 259]. Kosten-optimale [259]. Kostenoptimale [636, 675, 415]. kostenoptimaler [235, 411]. Kotz [1147]. Kovarianzdichte [402]. Kovarianzen [501]. Koziol [982]. Kugelverteilung [176]. Kullback [1847]. Kunden [342]. Kunststoff [38]. kupierter [78]. kurtosis [2346, 968, 1001, 1202]. kutosis [540].

L [1370, 1732, 531, 1371, 1388, 1476, 35]. L. [531]. Lack [1378, 1880, 1759]. lack-of-fit [1880]. LAD [1761]. Lagrangian [651]. Lahiri [883]. Laplace [1072, 2027, 627, 2310, 2017, 2095]. Large [950, 924, 869, 555, 11, 1808, 641, 2311, 2098, 2008, 1495, 2131, 2267, 1650, 1252, 663, 2005, 2182, 654, 1781, 2030, 2297]. Large-sample [950, 2098, 1252]. large-scale [2267]. larger [2340]. Largest [690, 820, 2302]. Lasso [2248, 2049]. latent [1631, 1282]. later [1864]. Latin [2122, 1877, 1795, 2222]. lattice [814, 569]. lattice-and [569]. Law [924, 522, 555, 11, 1398, 372]. Laws [1380, 1512, 2005, 1373]. leading [1789]. Least [2178, 871, 1364, 1152, 1118, 2189, 2208, 956, 1230, 974, 1788, 787, 961, 1434, 274, 2053, 2311, 1833]. least-squares [1788]. Leben [272]. Lebensdauer [506]. Lebensdauerverteilung [67]. Lebensdauerverteilungen [270]. lecture [8]. Lee [2172, 2112, 2077]. left [1900, 1593, 508, 520]. left-truncated [1900, 1593]. Lehmann [1388, 645, 35]. Lehmann-unbiased [645]. Lehne [505]. Leibler [1847]. Leitungsisolierungen [38]. Lemma [990, 1393, 1689, 853]. length [2217, 1017, 881, 1019, 1681, 339, 1389]. Lepage [2090]. Lepage-type [2090]. Lerntheorie [347]. less [1759, 870]. level [1609, 1598, 1112, 2032, 2194, 1901, 1902, 1841, 1326, 1433, 2200, 1221, 1381, 2086, 2112, 2195, 1589, 759, 2236, 2292, 2323, 852, 977, 1640, 1334, 1765, 1597, 1327, 1160, 2227, 1903, 1298, 2359, 1733, 1924]. levels [2272, 1095, 2258, 1216, 2182]. leverage [1730]. Leydold [1384]. Liefermengen [191]. Life [1413, 1447, 998, 1538, 1537, 214, 943, 2192, 962, 1482, 2091, 1671, 1131, 1144, 1296]. life-testing [943]. Life-tests [1413, 1538]. lifelengths [1806]. lifetime [752, 1919, 1953, 2105, 1729, 1504, 849, 2062, 1648, 2252, 1699, 1704, 1978, 1597]. lifetimes [2239, 2199, 2205, 2256, 2296, 1872]. light [2274]. like [125, 2193, 1194]. Likelihood [1960, 1408, 2042, 1342, 2197, 1914, 735, 976, 1490, 1923, 1658, 2058, 276, 413, 478, 1003, 2029, 1025, 2337, 2157, 2344, 872, 2156, 1865, 2299, 389, 1474, 1048, 2330, 1680, 1785, 515, 2264, 2130, 1534, 2078, 695, 2224, 29, 1438, 2193, 1670, 1769, 2260, 1466, 2018, 2158, 938, 2268, 2255, 1515, 267, 1428, 1508, 570, 709, 2210, 899, 867, 1567, 2254, 1623, 1572, 1908, 1632, 1975, 2203, 2184, 1606, 1608, 718, 1851, 1869, 697, 1234, 2095]. likelihood-based [1534]. Limit [1895, 1687, 555, 951, 1625, 332, 1416, 742, 1460, 1529, 2106, 256, 809]. limited [441, 339]. Limiting [1380, 1578]. Limits [922, 843, 949, 134, 186, 1122, 97, 184]. Lin [2200]. Lindley [2291]. line [2251, 897, 2075]. Linear [1369, 1402, 922, 1342, 1692, 1366, 1353, 222, 232, 234, 305, 223, 290, 344, 743, 2328, 1174, 1673, 2119, 2216, 934, 2309, 2235, 2228, 2140, 365, 386, 1627, 1070, 1588, 2249, 427, 808, 2287, 1238, 2029, 1740, 1741, 869, 2337, 2179, 1728, 1865, 1481, 1611, 1672, 1801, 2043, 2330, 1555, 2118, 1717, 2154, 2022, 2034, 2161, 1837, 2266, 157, 112, 1191, 132, 138, 1937, 587, 1148, 2319, 1764, 1758, 2094, 2353, 1289, 1544, 2117, 2063, 1004, 1090, 2065, 2267, 2289, 66, 1788, 1829, 2069, 1529, 1594, 673, 301, 1047, 1554, 1635, 2248, 1749, 431, 540, 2167, 796, 447, 1106]. linear [604, 1646, 694, 978, 163, 181, 292, 2215, 1333, 482, 1819, 1258, 1427, 247, 258, 294, 1500, 728, 1425, 1614, 2357, 364, 1632, 1811, 1805, 1852, 1942, 2358, 1177, 313, 1850, 870, 1431, 1706, 1977, 2107, 1851, 1971, 2089, 2207, 1566, 1165, 1193, 1065, 2082, 2186, 1973, 1453, 1619, 1659, 1761, 2067, 2361]. lineare [604, 424]. linearen [464, 84]. Linearity [984, 1884, 1219, 1971]. linearization [1742]. linearized [2173]. lines [127]. LINEX [1596, 1661, 2360, 1771, 1211, 1064]. liquidity [2162]. List [322]. lists [1406]. literature [390]. Liu [1702, 1638]. lives [893]. Local [1588, 1791, 534, 587, 2094, 2046, 2328, 1944, 1528, 2196, 1702, 1461, 1508, 1535, 1268]. localization [2125]. Locally [678, 147, 790, 532, 1399, 1444, 2137, 768, 2176, 2154, 938]. Location [645, 1596, 1237, 1337, 865, 2148, 1314, 1827, 2226, 2023, 1616, 2315, 1692, 309, 1201, 2017, 1781, 1050]. location-scale [2148, 1314, 2023, 1616]. locations [2022, 815]. Loéve [1695]. Loewner [1090]. Log [1418, 1688, 2192, 1238, 1426, 1555, 1772, 1769, 1439, 728, 1852, 1886]. Log-concave [1418, 1426]. Log-concavity [1688]. log-gamma [1439]. log-linear [1238, 1852]. log-logistic [1772]. log-normal [1769]. log-rank [2192]. Logarithm [1398]. logarithmic [632]. logarithmischen [338]. logarithmischer [331]. logconcave [1587]. Logistic [1395, 1915, 594, 1715, 1772, 504, 16, 2145]. logistischen [16]. logit [2224, 2333]. Loglinear [1366]. lognormal [1172, 1690]. logratio [1794]. Lohnpolitik [327]. lokal [54]. Lokalisations [422]. Lokalisations-und [422]. Lomax [1251, 2001]. Long [252, 1757, 1880, 1937, 2129, 1943, 2326, 1453]. long-memory [1943]. long-range-dependent [1453]. Long-run [252]. longitudinal [1631, 2021, 1123, 2043, 1522, 1731, 2313, 1861]. look [1626, 832]. Lorenz [231, 774, 231]. loser [1518]. loss [1542, 843, 1596, 2283, 971, 1070, 785, 1899, 931, 1661, 2360, 1616, 1771, 1211, 1145, 1015, 1595, 1064, 1886]. losses [2337]. lot [900, 906, 918]. Lotterie [426]. Lotto [584]. lotto-numbers [584]. low [2228, 1768]. low-dimensional [2228]. Lowe [136]. Lower [1349, 2258, 1654, 2076, 1854, 559, 1789, 2359]. Lower-order [2258, 1789, 2359]. Lq [2330]. Lq-likelihood [2330]. LRT [2167]. LSE [2318]. LTS [1127]. Lukacs [1146].

M [441, 8, 72, 531, 1371, 441, 238, 269]. M. [531]. M.-L [531]. M.L.R. [289]. M.L.R.-families [289]. M/G/1 [441, 238, 269]. machine [387]. main [1112, 1309, 882, 2099, 2182, 1765, 8]. main-effect [1112]. maintenance [2250]. majority [653]. Management [150, 194]. MANOVA [2182, 1752]. many [452, 1239, 537, 349]. map [2031, 2086]. March [18]. Marco [1387]. Marginal [1222, 2329, 366, 2126, 1123, 563, 1757, 1937, 926, 2254, 30, 37, 2047, 420]. Marginalizing [1536]. marginals [1831, 1749]. marked [1452, 695]. market [4]. Markoff [398]. Markoffsche [519]. Markoffscher [519]. Markov [1476, 1821, 441, 1800, 1658, 1238, 380, 872, 1961, 2290, 267, 1751]. Markovian [1487, 598]. Markowsche [47]. Marshall [1756, 1714, 1872]. martingales [690]. marxistischen [93]. März [18]. mass [2052]. massive [2304, 2301]. matches [102]. material [38]. Materialverbrauches [38]. Mathematical [6, 22, 18, 109, 828]. Mathematics [1478, 68, 129]. Mathematische [18, 6]. mathematischen [93]. mathematischer [153, 140]. mathematisches [828]. Matrices [1414, 1397, 1379, 776, 1169, 1229, 2179, 382, 2299, 720, 2168, 1220, 157, 112, 1274, 1726, 1529, 1325]. Matricvariate [1734]. matrix [950, 1876, 1797, 2109, 1734, 704, 931, 979, 1098, 1042, 1594, 154, 1520, 1505, 661, 2057, 2358]. Matrixinversion [154]. Matrixsummierungsverfahren [279]. max [1510, 235]. max-infinitely [1510]. Max-Prinzips [235]. Maxbias [1083]. Maxima [501, 1437, 1692, 1862, 2312, 2324, 1859]. Maximal [1253, 1419, 1729, 1324]. maximale [33, 41, 505]. Maximin [2180, 1718, 967, 2222, 2277]. maximizing [1535]. Maximum [864, 1658, 526, 389, 439, 1342, 2147, 1133, 976, 2058, 276, 413, 478, 1025, 872, 1636, 2156, 1077, 1348, 1474, 1426, 1863, 2264, 695, 29, 1438, 1670, 938, 1587, 2255, 33, 490, 267, 756, 1508, 709, 2210, 899, 867, 1623, 1606, 1608, 1664, 718, 697, 1234, 2095]. maximum-likelihood [276]. MCD [1127]. McShane [1371]. MDS [1034]. Mean [1061, 784, 1482, 2083, 1427, 2358, 2329, 998, 1609, 2060, 2241, 2211, 1157, 1558, 2283, 971, 741, 484, 999, 1947, 2012, 906, 1504, 2033, 1589, 515, 2014, 2050, 1509, 397, 1726, 1769, 2345, 779, 793, 893, 1855, 1211, 565, 820, 925, 763, 1015, 731, 1943, 754, 1671, 1012, 5, 401, 1097, 539, 606, 1064, 1597, 1723, 1796, 1057, 148]. Means [1420, 1907, 224, 351, 1898, 2228, 1070, 2340, 569, 608, 1551, 1771, 801, 2310, 981, 697, 1550]. measurability [278]. Measure [2329, 1778, 2007, 2009, 2088, 2356, 695, 1648, 1681, 1904, 2020, 1582]. Measurement [1423, 209, 949, 2338, 562, 2213, 1191, 2353, 1696, 523, 1606, 2082, 2137, 2186, 1663, 1659, 1104]. measurements [39, 1289, 1162]. Measures [2346, 1725, 1139, 2066, 1944, 2000, 1812, 2261, 125, 733, 2250, 1897, 2061, 1496, 470, 400, 1456, 1202, 363, 445, 463, 1665]. measuring [13, 55, 191]. mechanistic [2082]. Median [83, 289, 1303, 2223, 2146, 1099, 831, 2345, 419, 2282, 359, 521, 1144, 822].

Median-Quartile-Test [83]. medians [186]. meeting [130, 194]. mehrdimensionale [264, 225]. mehrdimensionaler [454]. mehrere [415]. mehreren [384]. mehrerer [501]. mehrfacher [15]. Mehrfachkorrelationen [84]. Mehrheitstestkombination [653]. Mehrstufige [406]. mehrstufigen [416]. Memory [1378, 1757, 1880, 1937, 2129, 1943, 2326]. Mengenfunktionen [60]. Merkmal [675]. Merkmale [5, 101, 415]. Merkmalen [114, 384]. Merkmalwerte [241, 268]. message [1389]. messende [562, 13, 55, 149, 191, 241, 268, 423, 59]. messender [64]. Messung [275, 328, 273]. Messungen [215]. Meßwerte [437]. meta [2009, 2193]. meta-analysis [2009, 2193]. Method [1340, 2119, 440, 2049, 1323, 413, 2200, 196, 2157, 874, 2270, 2163, 1863, 1460, 1494, 1769, 1243, 1290, 438, 961, 1515, 535, 1956, 154, 911, 1428, 226, 191, 16, 870, 2082, 2187, 1705]. Methode [16, 314]. Methoden [153, 140, 93, 335]. methodologies [959]. Methods [1380, 1953, 224, 276, 1718, 670, 2142, 1475, 619, 187, 1457, 600, 311, 2064, 1263, 2075, 100, 449, 1383, 335, 313, 1477, 1386]. metric [2336, 1496]. Metrika [420, 1814]. metrischen [219]. Michael [1373]. Micro [24, 1720]. micro-aggregated [1720]. Micro-economic [24]. Microarray [1371]. microstructure [2245]. Midzuno [883, 827, 190]. mikro [122]. mikro-ökonomische [122]. mildly [2176]. MIMSQE [768]. Mini [235]. minification [1248]. Minima [236, 2312, 2324]. Minimal [1260, 634, 1729, 704]. Minimally [1418]. Minimally-supported [1418]. Minimax [785, 483, 1391, 1435, 1443, 139, 983, 1145, 1521, 2134, 1595, 2214, 1673, 455, 452, 741, 1827, 1899, 790, 804, 1020, 708, 883]. minimax-tests [804]. Minimierung [528]. minimization [1810]. Minimum [544, 632, 1405, 885, 1419, 2002, 595, 1880, 503, 508, 520, 1949, 2335, 2159, 1272, 678, 2272, 2040, 1890, 1470, 1812, 599, 864, 1170, 380, 1854, 1077, 1348, 2023, 811, 1988, 1426, 2360, 1389, 2323, 403, 760, 1587, 278, 329, 535, 1545, 448, 500, 1552, 311, 798, 665, 1113, 1789, 2227, 2359]. minimum-concomitant [1890]. minimum-discrepancy [311]. minres [276]. Mischverteilungen [78]. misclassification [1330, 1608]. Mises [2012, 2255, 1375, 1084, 1947, 555, 1266, 1009, 1532]. missing [1335, 2132, 2285, 1715, 2353, 2063, 2260, 1950, 1709, 2314, 1428, 1731, 1836, 981, 1113, 2313, 1026, 1431, 1775, 1748, 2153, 1851, 1973, 1925]. missingness [1536]. misspecification [2280, 1984]. misspecified [1434, 1065]. Mitteilung [170]. Mitteilungsblatt [6]. mittels [235]. Mittelwert [5]. Mittelwerte [423, 148]. Mittelwerten [316]. Mixed [1729, 2174, 1225, 1342, 1415, 2227, 1733, 1924, 688, 2032, 2194, 976, 1970, 1221, 1429, 2015, 1672, 2043, 2112, 1207, 2236, 538, 2224, 2319, 1764, 2010, 2323, 2065, 2018, 2185, 2167, 2305, 604, 2321, 2313, 1910, 1909, 1882, 1431, 1706, 1160, 2082]. mixed-data [2321]. mixed-effect [1970]. mixed-effects [604, 1910, 1909]. Mixed-level [1733, 1924, 2032, 2194, 1221, 2112, 2236, 2323, 1160]. Mixing [2318, 1362, 2248, 1055, 757, 1244, 2331]. Mixture [1405, 1447, 1927, 1347, 2178, 2239, 1140, 413, 762, 2244, 1952, 1516, 2149, 1134, 2255, 1560, 1020, 2269, 1536, 2220]. Mixtures [2204, 2000, 2123, 1756, 2316, 1555, 897, 2264, 912, 1482, 2006, 2083, 1573, 1834]. MLE [1103, 2244, 1961]. MLE-process [1961]. MLEs [1599, 1741]. MM [1353]. MM-Estimators [1353]. MME [2243]. MNBU [2155]. modal [2304]. mode [766, 1010, 1691, 1593, 2262, 1085]. Model [1364, 634, 1422, 1472, 1669, 1568, 1423, 1984, 1392, 1870, 1915, 1571, 664, 688, 1610, 1919, 2105, 1651, 1120, 99, 214, 1470, 1473, 2228, 1917, 2246, 876, 1964, 2281, 2249, 1631, 891, 1923, 2293, 2111, 2287, 1970, 2003, 1810, 2202, 2247, 1879, 1740, 1429, 1741, 1948, 2337, 2344, 982, 1760, 1449, 900, 1019, 2156, 1928, 2041, 1481, 1123, 2244, 2352, 1555, 1680, 991, 2343, 2161, 931, 1098, 1837, 2266, 616, 1952, 619, 1191, 1277, 1192, 1230, 974, 1465, 2191, 817, 1715, 2193, 2201, 1462, 2063, 832, 2124, 1788, 1829, 2069, 1950, 2198, 2268, 2252, 1604, 1826, 1594, 1451, 498, 586]. model [961, 1009, 1134, 1700, 2307, 1912, 1699, 1216, 2025, 589, 1139, 2248, 557, 1749, 651, 472, 1522, 796, 1131, 1106, 1936, 1696, 1520, 978, 2277, 1974, 2151, 1484, 1154, 2173, 173, 1553, 789, 1427, 1836, 2321, 728, 1614, 1922, 2280, 2162, 1852, 2053, 2184, 1918, 2004, 1606, 1882, 1026, 1977, 2311, 1980, 1638, 1851, 2089, 2207, 1234, 2082, 1453, 1634, 1085, 838, 1824, 1344, 2352]. Model-based [1472, 664]. Model-calibration [1568]. model-restricted [1026]. Model-X [2352]. Modeling [2100, 1951, 1530, 2246, 2087]. Modell [353, 228, 334]. Modelle [355, 424, 604, 219]. Modelling [2322, 878, 1378]. Modells [616, 173]. Models [2174, 1358, 1342, 1366, 1415, 1447, 222, 232, 743, 1927, 1966, 1876, 1174, 2306, 1318, 2178, 1538, 1599, 934, 1270, 1040, 1306, 1119, 1446, 1756, 1255, 1023, 1035, 2164, 2160, 871, 1963, 996, 1730, 1530, 1622, 1627, 1798, 943, 700, 1569, 1588, 768, 2021, 1800, 1043, 2148, 2196, 427, 808, 2230, 2221, 1531, 644, 1238, 1003, 2029, 2019, 2015, 2097, 2189, 2157, 2344, 1452, 1865, 1474, 1611, 1672, 2011, 2208, 811, 1801, 2043, 2033, 2330, 2118, 2213, 2154, 991, 1207, 897, 2161, 2183, 157, 2081, 994, 2113, 128, 135, 1124, 2224, 1180, 2319, 514, 1516, 1900, 1674, 1764, 1758, 2010]. models [1976, 1544, 2117, 655, 1004, 1090, 912, 1698, 1776, 2042, 2065, 1243, 1788, 2018, 1931, 1613, 1213, 2114, 2149, 1060, 1885, 673, 801, 2284, 621, 1554, 1271, 1635, 1949, 540, 2167, 618, 861, 2305, 1444, 604, 1646, 955, 1513, 1999, 1535, 2103, 2215, 2269, 1536, 1258, 2182, 2210, 1500, 1053, 1425, 1614, 1623, 1113, 1563, 1724, 1995, 1632, 1811, 1805, 1942, 1979, 1975, 2220, 2263, 2313, 2059, 1910, 1909, 313, 1850, 1431, 1706, 1723, 1463, 1958, 1498, 1775, 1748, 1856, 2107, 1921, 1872, 1480, 1132, 1566, 2214, 1165, 1193, 1065, 1861, 2082, 2137, 2186, 1663, 1973, 2206, 1619, 1628, 1659, 1761, 1985, 2067, 2361]. models [1833]. Moderate [2300]. modification [2076, 201]. Modifications [1000]. Modified [1863, 293, 214, 478, 2360, 2090, 503, 520, 1620, 1752]. modularity [2127, 2300]. Modularization [2026]. modulus [1815, 904]. Mögliche [326]. möglichen [636]. Möglichkeiten [280]. molecular [1383]. Moment [1364, 1714, 1619, 1292, 2178, 776, 1786, 435, 1172, 1545]. Moment-based [1714, 1786]. Momente [388, 513]. Momentengleichung [348]. Momentenproblem [244]. Moments [1307, 1644, 896, 1054, 2188, 1501, 2024, 2010]. Monge [1049]. Monitoring [1284, 1553, 1652, 1171, 1458, 1201, 1550]. monographs [26]. Monotone [1382, 1800, 2354, 2290, 1536, 2313]. Monotonicity [735, 1348, 1599, 1688]. Monte [1386, 1473, 1617]. Morgenstern [1395, 1994, 1000]. Mori [1477]. Mortara [318]. mosaic [1032]. Most [1399, 2225, 147, 1079, 173, 809]. motion [844]. motions [1528]. motivation [1354]. Moving [436, 471, 518, 1757]. MST [1034]. MTAR [1484]. Multi [2032, 1381, 1107, 1061, 879, 2011, 1961, 2065, 2315, 251, 2016, 1458, 975, 542, 2182, 1640, 649, 208]. multi-attributes [1458]. multi-channel [879]. multi-character [1107, 975]. multi-classification [2016]. multi-dimensional [208]. multi-factor [2011]. Multi-level [2032, 1381, 1640]. multi-response [2065]. multi-sided [649]. multi-stage [1061, 2315, 251, 542]. multi-step [1961]. multi-way [2182]. multicategory [2301]. multichotomous [1823]. Multidimensional [2008, 225]. Multilinear [540]. Multinomial [1391, 2202, 515, 1438, 1577, 1455, 1635, 266, 2305]. multiparameter [1022]. Multiple [1368, 671, 1331, 1190, 1966, 2212, 2119, 1454, 1740, 859, 848, 1099, 2270, 1570, 2033, 2168, 1605, 2130, 2045, 15, 1905, 1696, 1535, 729, 2317, 697, 2159, 1985, 1833]. multiple-index [2033]. multiple-outlier [1966]. multiplicative [700, 2097, 2041, 1958]. multiplier [1086, 651]. multiply [2270]. Multipolishing [1092]. Multiresolution [1695, 933]. multiresponse [1776]. multistage [959]. Multivariate [1927, 998, 180, 1963, 1490, 1364, 999, 2012, 1111, 1378, 893, 959, 1105, 1363, 1615, 577, 1596, 1844, 711, 2123, 1962, 224, 919, 741, 1643, 1800, 937, 730, 1043, 1901, 1902, 1734, 2170, 2048, 1084, 1741, 1725, 1370, 2188, 960, 1077, 1688, 1474, 702, 849, 1426, 2343, 931, 1098, 250, 624, 951, 1192, 1230, 2155, 1726, 962, 2252, 1594, 1904, 1387, 1058, 503, 508, 520, 770, 1251, 1018, 309, 1561, 1520, 1456, 564, 1284, 1693, 930, 899, 867, 512, 1623, 2313, 1033, 1850, 1638, 1057, 1050]. Münzner [322]. mutual [1582]. MV [1407]. MV-optimality [1407].

nach [110, 575, 84, 241, 268]. Nachruf [319, 96, 50, 318]. Näherungsverfahren [191]. näherungsweisen [297]. naive [832]. nat [231]. national [17]. Natur [828]. Natural [1481, 630, 934, 2251, 22, 631, 1302, 2101, 828]. Nauheim [7, 8]. navigation [2342, 2348, 2355, 2362]. near [1890, 1437, 1862]. near-maxima [1437, 1862]. Nearest [2166, 1376]. nearly [1877, 1795, 1028]. nearness [1778].

Necessary [1594, 967, 615, 643, 1869]. need [158]. Negative [1407, 2322, 99, 145, 375, 561, 552, 831, 2022, 2081, 657, 1460, 985, 686, 2315, 806, 1455, 658, 2282, 1444, 256, 1417, 312, 2136, 822]. negative-binomial [2322]. negatively [1576, 237, 1891]. neighbor [1279, 2166]. neighborhoods [1354]. neighbors [2328, 2191]. Neighbour [1376, 2150]. Nested [100, 1074, 2122, 780, 2182, 1563, 1724]. Network [2190, 2300]. networks [2351, 2127, 55]. neue [16]. neuer [437]. neuere [328]. neural [2351]. Newly [851, 895]. Neyman [8, 8, 680, 789, 990, 1393, 1104]. Nichtgültigkeit [584]. Nichtkonsistenz [531]. nichtparametrische [10]. nichtparametrischen [35, 381]. nichtparametrisches [529, 506]. nichtquadratischer [33, 41]. Nielsen [1383]. Niveau [52, 79]. NN [2328, 2191]. no [599, 1644]. noise [2287, 45, 2245]. noisy [1010, 2104, 2218, 2327]. Non [1364, 2243, 1414, 1576, 1412, 1843, 699, 2040, 2241, 224, 368, 1023, 1157, 365, 552, 2257, 2111, 942, 984, 346, 2050, 514, 657, 1711, 938, 1650, 531, 658, 470, 33, 2327, 1749, 754, 500, 963, 2175, 578, 312, 1567, 598, 2079, 35, 1245, 2313, 1721, 1869, 1886]. non-additive [470]. non-adjacent [984, 1721]. non-central [657, 658]. non-cooperative [346]. Non-correlation [1412]. non-deficiency [1711]. non-ergodic [2241]. Non-exchangeability [1576]. Non-Existence [1364]. non-homogeneous [2040, 754]. non-inferiority [1567]. non-linear [365, 1749]. non-M.-L. [531]. non-Markovian [598]. non-monotone [2313]. non-negative [552, 312]. non-normal [1023]. non-normality-Special [514]. non-null [368]. non-overlapping [1650]. non-parametric [224, 2257, 35, 1869].

Non-positive [1843]. non-precise [1245]. non-regular [2175, 1886]. non-regularity [2079]. non-response [963]. non-sequential [578]. non-simultaneous [2050]. non-square [33]. non-standard [2327, 500]. non-stationary [942]. Noncentral [1641, 285, 762]. Nonclassical [1398]. nonconcave [1801]. nonconstant [1318]. nonhomogeneous [1821]. nonidentical [1246, 1704]. nonignorable [2132, 2353, 2215, 1731]. nonisomorphic [1442, 1433]. Nonlinear [860, 1408, 1342, 1925, 1318, 1306, 996, 1923, 2293, 814, 2189, 534, 1124, 2201, 1434, 1153, 1949, 2313, 1850, 2311, 2082]. nonlinearity [1139]. nonnegative [986, 827, 1908]. nonnegativity [639]. nonnormal [949, 768]. Nonparametric [1997, 1226, 2021, 1820, 1970, 2298, 2266, 956, 2113, 2104, 1250, 1683, 938, 2054, 2218, 2327, 1018, 1456, 909, 1775, 2153, 596, 1951, 2338, 2339, 1519, 2285, 681, 1883, 1600, 1938, 2129, 1462, 2108, 288, 1568, 291, 729, 728, 10, 1690, 1607, 1705, 1628, 1085]. Nonregular [1368, 1274, 1312, 837]. nonresponse [2246, 662]. nonsymmetric [1235]. nonuniform [1384, 497]. nonvalidity [584]. Norm [1362, 1250, 1494, 760]. Normal [1358, 1404, 2306, 1907, 562, 1023, 1133, 971, 1070, 741, 1735, 2340, 575, 762, 1077, 1688, 702, 608, 997, 2264, 157, 624, 1509, 1769, 1771, 983, 1211, 820, 1134, 2083, 968, 1001, 1209, 1707, 1097, 1974, 2310, 539, 709, 1064, 512, 910, 1995, 1138, 1921]. normality [1755, 1337, 730, 2170, 669, 872, 742, 514, 1261, 1593, 66, 315, 2046, 2308, 360, 378, 1522, 267, 694, 836]. normalization [1246]. normalized [484]. normalizing [802]. normally [597]. normals [1560]. normalverteilter [355, 597]. Normalverteilung [469, 575, 331]. Normalverteilungen [78, 48, 85]. Normalverteilungsannahme [562]. North [2152]. North-east [2152]. Note [1815, 99, 186, 249, 351, 380, 133, 247, 396, 233, 223, 1270, 2251, 1337, 1133, 1171, 1239, 935, 1041, 1523, 781, 1207, 1220, 2130, 1438, 1776, 1080, 2268, 476, 815, 251, 45, 1678, 1896, 615, 643, 374, 945, 911, 68, 129, 373, 257, 539, 456, 312, 665, 481, 512, 757, 1256, 1942, 2238, 970, 2297, 420, 1050, 765]. Notes [1988, 407]. notice [546]. notion [1739]. Notiz [215]. novel [2275]. Nuclear [1201]. nuisance [1218, 1837]. null [368]. Nullhypothese [106]. number [2328, 1890, 1076, 2199, 2296, 2183, 2191, 1437, 1577, 1809, 1168, 1216, 1386, 578, 173, 2182, 905, 1811, 1805, 1982, 2073]. Numbers [1373, 814, 904, 584, 146, 2005, 924, 555, 11]. Numerical [1049]. numerischen [48, 52, 79].

OAVS [913]. oben [241, 268]. obere [559]. Oberwolfach [18]. Obituary [319, 96, 50, 318]. object [543]. objectives [1454]. objects [2, 1068, 748]. Objekten [2]. observation [1797, 1551, 1624, 1203]. Observations [1382, 1377, 2040, 1890, 2243, 1993, 413, 1279, 872, 790, 1010, 483, 1702, 2104, 1593, 237, 2260, 1060, 889, 1291, 266, 1533, 981, 1113, 2312, 2324, 10, 1298]. observed [1072, 563, 1701, 955]. observers [649]. obtain [1290]. Obtained [1397, 1220]. obtaining [276]. occupancy [1320]. occurrence [648]. odds [2069, 1912, 1271]. off [382]. Officers [1159]. offset [369]. ohne [106]. ökonometrische [327]. ökonometrischen [616]. Ökonomie [93]. ökonomische [122]. ökonomischen [337]. ökonomischer [410]. Old [187, 430]. Olkin [1756, 1714, 1872]. OLSE [1164]. Olympia [426]. Olympia-Lotterie [426]. One [1915, 1382, 1624, 1363, 1581, 1579, 649, 1985, 2339, 2105, 1651, 1103, 671, 952, 2076, 891, 610, 1551, 791, 837, 1465, 779, 793, 1290, 1216, 309, 1615, 1178, 728, 1721, 1752, 582]. one- [728]. one-dimensional [1651]. one-inflation [2076]. one-parameter [1103, 582]. One-sample [1915, 1579]. one-shot [2105]. One-sided [649, 1985, 671, 952, 791, 779, 793]. One-step [1581]. one-way [891, 1465, 1216, 1752]. ones [2212, 1955]. online [1036]. only [1499, 257]. operating [799]. Operationscharakteristiken [399]. operator [1053]. Optimal [1312, 664, 581, 1898, 442, 1730, 2304, 1074, 1407, 1221, 1429, 1686, 848, 1783, 1883, 1369, 1611, 2118, 2161, 2292, 1180, 662, 1289, 912, 913, 1125, 1009, 1554, 1153, 1291, 266, 1591, 683, 1201, 1403, 1162, 1575, 1882, 1856, 2030, 699, 1538, 2216, 1023, 1812, 1418, 1562, 1798, 2257, 1179, 808, 2247, 939, 2015, 859, 1502, 1405, 1087, 847, 1986, 2011, 2208, 1110, 2086, 1655, 2154, 1207, 1328, 1770, 1277, 1495, 1557, 1287, 1894, 1776, 2065, 907, 1068, 2222, 1581, 1444, 1178, 1888, 1887, 2277, 1766, 2173, 191, 1819, 1765, 728, 449, 1425, 1563, 2280, 2220, 1706, 2073, 1298]. optimal [1165, 1193, 1379, 1781, 2137, 716, 1555]. optimale [54, 340, 331, 384, 259]. optimalen [297, 333]. optimaler [191]. Optimalität [604]. Optimality [829, 1645, 2341, 1279, 834, 760, 1143, 1415, 1514, 688, 1407, 808, 637, 991, 781, 496, 1004, 1090, 2128, 1063, 2284, 604, 2093, 373, 1163]. optimally [772]. optimization [826, 2219, 1129, 600]. optimized [636, 675]. Optimum [577, 174, 586, 835, 642, 667, 1318, 461, 438, 421, 1216, 571]. optional [1624, 1630]. options [2225]. optisches [185]. oracle [2310]. Order [1397, 1778, 1182, 1292, 1966, 2294, 2066, 1895, 2216, 795, 744, 2209, 1219, 1448, 1424, 1868, 2140, 1798, 1932, 1222, 984, 1884, 776, 841, 896, 2253, 1229, 1960, 2019, 1084, 916, 1054, 2165, 1883, 1957, 2354, 1854, 772, 2023, 2208, 2258, 2130, 1210, 616, 992, 2150, 1250, 1516, 1954, 2038, 2036, 1769, 1771, 1297, 2013, 1650, 1155, 1885, 801, 2315, 1485, 1047, 1482, 1324, 954, 1173, 1847, 2001, 731, 615, 643, 647, 1802, 825, 1564, 1843, 641, 694, 574, 2302, 1439, 1334, 1736, 1666, 1744, 2184, 1789, 1664, 877, 2279, 1781, 2030, 1861, 2359, 2187]. order-of-addition [2187]. order-restricted [1771]. ordered [1190, 1483, 1438, 237, 476, 2006, 1934]. Ordering [2148, 2019, 2302, 774, 712, 1958, 1660]. orderings [1780, 1580, 2140, 1490, 2253, 1486, 2091]. Orders [1351, 1169, 1896, 651]. ordinal [644, 543]. ordinary [554, 961]. Ordnung [616]. orientation [1657]. oriented [327]. Ornstein [2241, 1530, 1970]. Orthogonal [141, 882, 1266, 1442, 1893, 1645, 1323, 1179, 2122, 1877, 1326, 1352, 1433, 1873, 1795, 2099, 2232, 1028, 1657, 1258, 1427, 1765, 1903, 2187]. orthogonally [1260]. Oskar [50, 50]. Other [1399, 363, 445, 449]. outcome [1715]. outcomes [102]. outgoing [846]. outlier [1966, 869]. outliers [2119, 840, 286, 450, 1981, 648]. overdispersed [2100]. overdispersion [1652]. overhead [577]. overlapping [1884, 307, 1650]. overshoot [946]. overview [862].

P [1388]. Paar [183]. paarigen [10]. packings [2180]. pair [795]. Paired [1369, 2056, 1992, 2126, 1726, 10]. pairs [671, 1992]. pairwise [1460, 959]. Paley [1397]. panel [1653, 108, 2230, 2124, 1911, 1940, 1484, 2210]. panels [2014]. Panjer [2209]. paper [418]. papers [851, 895, 2044, 1930]. parabola [1009]. parabolic [2357]. paradox [92]. Parallel [2163, 1960, 1990, 2296, 358, 2036]. parallel/series [2036]. paralleled [252]. parallelism [1571]. Parallelität [117]. Parameter [1395, 849, 411, 67, 270, 1596, 1237, 1103, 533, 1914, 1644, 865, 785, 1783, 646, 2244, 2352, 1785, 2237, 2360, 1616, 2078, 837, 1044, 1422, 1772, 1758, 2117, 1701, 468, 993, 889, 1153, 2028, 2305, 625, 1417, 778, 257, 1933, 1138, 30, 37, 591, 718, 1296, 582].

Parameter- [270]. parameterfreie [239, 246]. Parameterfunktionen [414]. Parameters [1357, 1571, 1599, 2275, 544, 1633, 871, 1171, 903, 461, 1846, 645, 1990, 389, 885, 811, 1837, 890, 2084, 2078, 791, 1191, 450, 300, 1590, 1155, 476, 2315, 479, 1858, 1692, 2001, 302, 309, 625, 1696, 5, 1751, 1439, 729, 789, 312, 258, 867, 2017, 1572, 2238, 1859, 2095]. Parameterschätzproblemen [582]. Parameterschätzung [30, 37]. Parametric [234, 1810, 1699, 1464, 2357, 1396, 1821, 224, 452, 2257, 2021, 1385, 2226, 1435, 1443, 1788, 1614, 35, 364, 1632, 1869, 1705]. parametrized [885]. Paretian [1948]. Pareto [544, 903, 595, 494, 2108, 287, 300, 1503, 1251, 187, 374, 1056, 625, 641, 1039, 541, 1736, 1248, 1933, 2079]. Pareto-type [2108]. parities [121]. parsimonious [670]. Part [844]. Partial [1543, 2073, 1313, 1798, 484, 1280, 1559, 1976, 2268, 1692, 1345, 2093, 2215, 1971, 2089, 2067]. Partially [1586, 141, 1639, 2029, 982, 1801, 2213, 2161, 634, 1758, 2201, 2063, 1829, 2069, 1632, 1811, 1234, 1973, 1619, 1659]. particles [1471]. particular [283, 1881, 1661]. partielle [167]. partitioned [1788]. partly [1701]. passing [387]. Past [1365, 1546, 2088, 2165, 1504, 2062, 1648, 1681, 1945, 1704, 1597]. paternity [335]. patients [557]. Pattern [1368, 2099, 1536, 1982, 2073, 2350]. pattern-mixture [1536]. patterns [553, 1864]. Paul [231]. pay [382]. pay-off [382]. Payley [1220]. PBIB [720]. PBIBD [634]. PDE [2170]. Pearson [2016, 1520, 990, 1393]. peeling [1842]. Penalized [1003, 2330, 2255, 2351, 2224, 2184]. penalty [1801]. percentage [831]. Percentile [1616, 2039, 2090, 2238]. percolation [1816]. performance [1792, 1812, 1530, 1842, 961]. period [1584, 327]. Periodic [1200, 2241, 1584]. periodically [1892]. periodicity [2261]. permutation [618, 1385]. permutations [2236, 2092]. Permutationstest [277]. permutons [2295]. person [346]. Personen [90]. perturbed [836, 1553]. Petersburger [330]. Pfadkoeffizienten [167]. phase [2249, 354, 421, 350]. phase-type [354]. phases [996]. Phillip [1385]. PHR [1966]. physiological [1695]. Pitman [1778, 1959]. places [2]. plan [1112]. planar [343, 1708, 789]. Pläne [149, 191, 2125, 241, 268, 423, 74]. Plänen [13, 59]. planes [127]. planning [4, 1296]. Plans [1419, 1200, 636, 667, 1179, 1309, 1549, 882, 799, 675, 212, 13, 55, 191]. plastics [38]. Plätze [2]. play [1199, 191]. play-the-winner-type [1199]. playing [1159]. plot [2194, 2143, 2292, 2323, 2227, 1733]. plots [1032, 1092]. plug [1665]. plug-in [1665]. plus [1112, 1645, 2360, 2232]. pmk [1114]. Point [1343, 1355, 1376, 927, 1489, 1953, 1313, 1076, 1923, 2049, 2221, 1526, 939, 831, 1452, 2044, 1559, 1717, 2360, 1668, 343, 299, 1213, 2045, 1739, 1006, 1064, 1138, 1256, 870]. points [1190, 2212, 2009, 935, 1793, 1494, 1050]. Pointwise [2336]. Poisson [262, 1303, 2322, 2339, 1610, 802, 1120, 1963, 2291, 632, 2125, 2243, 2146, 1820, 2221, 854, 217, 211, 1027, 1655, 1871, 197, 695, 2052, 2104, 2018, 1950, 1779, 476, 2284, 1857, 2185, 2025, 754, 1444, 1508, 2277, 899, 867, 1053, 2059, 1918, 1463, 2100, 2286, 697, 1634]. Poisson-inverse [211]. Polar [1445]. policies [2250]. policy [327]. politischen [93]. Pólya [458]. Pólya-Verteilung [458]. polygon [2331]. Polygonal [1108]. Polynomial [1418, 1423, 1811, 1562, 1941, 859, 847, 787, 1724, 2137, 1985]. polynomials [2343]. population [664, 1107, 1716, 495, 874, 960, 1589, 286, 662, 204, 412, 1855, 287, 1208, 1211, 438, 498, 586, 763, 628, 1015, 618, 830, 1005, 769, 963, 539, 606, 590, 981]. Populations [1382, 2275, 1803, 351, 762, 608, 819, 1771, 820, 2315, 806, 1858, 2001, 1916, 1564, 1843, 564, 177, 910]. portfolio [1470, 1129, 2057]. portfolios [1812]. Positive [1407, 1110, 1843, 312]. Positively [1398]. positivity [227]. possessing [2354]. possible [636, 610, 1566]. possibly [2126, 2314, 1434]. posterior [766, 2052]. posteriori [1923]. postponable [269]. Potentialtheorie [398]. power [1907, 2040, 86, 284, 735, 1820, 904, 1253, 1100, 1948, 1728, 2325, 682, 1301, 503, 508, 520, 71, 417, 1535]. Powerful [1399, 147, 1079, 2225]. powers [1533, 1138]. pp [1370]. PPSWR [705]. practical [1630]. practice [1537]. praktische [107]. preassigned [751]. Precedence [1506]. Precedence-type [1506]. precise [1245]. precision [1520]. Predicting [2271, 731]. Prediction [1542, 2239, 1391, 2233, 978, 1915, 664, 2216, 1487, 840, 1025, 1611, 1871, 1148, 2052, 2319, 809, 2147, 1144, 1607]. predictions [616]. Predictive [606, 1182, 2226]. predictor [1427]. predictors [2154, 662, 2254]. Preface [1117, 1317, 1814, 1525, 857]. Preferred [1340]. Preisparitäten [121]. Preliminary [130, 1233, 1191, 1638]. prescribed [843, 846]. Presence [1351, 1413, 1518, 180, 840, 556, 1091, 1330, 1998, 2129, 662, 2074, 302, 2218, 1669, 1428, 2254, 944]. present [1925]. Preservation [2036]. pretest [1804]. prevalence [2224]. prevalent [2268]. previous [1609, 1299]. price [514, 121]. principal [1473, 2234]. principle [570, 2300]. principles [1037]. Prinzip [330]. Prinzips [235]. prior [2202, 1835]. prior-data [2202]. priori [1245]. priority [395]. privacy [1823]. privacy-protecting [1823]. probabilistic [2103]. probabilities [2329, 440, 751, 1086, 2303, 1297, 1708, 225, 666, 1288, 46]. Probability [526, 1262, 1720, 209, 737, 1107, 1944, 1335, 826, 2332, 125, 22, 1835, 804, 584, 2002, 439, 653, 1472, 756, 369, 55, 888, 559, 714, 1849, 1935, 1492, 1013, 370]. probability-densities [804]. probability-like [125]. probable [327]. Probeentnahmen [116]. Probenahme [20, 118]. probit [2305]. Problem [636, 79, 410, 279, 283, 1489, 387, 70, 1313, 455, 2115, 430, 1432, 356, 707, 1435, 29, 1172, 367, 354, 531, 23, 600, 1049, 192, 521, 571, 395, 394, 2310, 258, 294, 757, 1247, 339, 349, 1690, 1065, 531]. Probleme [280, 17]. Problems [1359, 1022, 147, 452, 1239, 579, 762, 1570, 2044, 612, 421, 309, 175, 310, 350, 491, 582, 17]. procedure [846, 1755, 1793, 1655, 1881, 806, 993, 901, 573, 1334, 1875, 1608, 1104]. Procedures [667, 812, 892, 1598, 1558, 943, 1072, 1972, 1996, 1848, 791, 496, 659, 985, 686, 756, 705]. Proceedings [466]. Process [1293, 802, 2241, 1792, 2291, 881, 1787, 923, 1493, 1559, 1027, 2237, 1655, 1010, 1871, 2104, 1465, 1961, 1289, 779, 793, 2252, 45, 884, 1114, 1194, 1943, 754, 28, 1006, 930, 899, 867, 1053, 208, 1664, 1796, 716]. Processes [1376, 518, 1821, 1965, 2318, 1313, 1530, 1588, 1820, 1817, 640, 1639, 1526, 869, 2179, 380, 872, 1019, 1452, 607, 2176, 629, 532, 2051, 2121, 2276, 695, 1462, 66, 783, 1203, 2290, 1458, 1105, 2169, 267, 1268, 789, 1969, 598, 1248, 1744, 2326, 2100, 2286, 1550, 471, 558]. processes- [869]. product [1096, 896, 2024, 893, 287, 1503, 926, 460, 220, 226, 606, 1563, 2147]. product-type [460]. Production [4, 1458, 146]. Productivity [327]. Productivity-oriented [327]. products [1436, 1625, 597]. Produkten [597]. Produktionsplanung [140]. Produktivitätsorientierte [327]. Prof [271]. Prof. [230, 271, 231]. Profile [1369, 1496]. Prognosefehlers [616]. Prognosemodelle [122]. Prognosen [616, 164]. program [365, 386]. programming [233, 234, 305, 283, 345, 442, 1718, 112, 163, 181, 292, 1819, 247, 258, 294, 364]. Progressive [1447, 1915, 1890, 2026, 1998, 1670, 2027]. progressively [1915, 1642, 1914, 1072, 1802, 570]. Projection [1272, 1326, 1397, 2200, 2272, 1929, 1352, 725, 1220, 2345, 2350, 1050]. projections [1169, 1229]. projektive [338]. proof [518, 28]. proper [2317]. Properties [1397, 1928, 1365, 2204, 737, 1889, 2339, 2123, 1528, 1246, 1103, 1756, 1890, 2291, 976, 1831, 669, 904, 1326, 1352, 2200, 2019, 733, 811, 917, 2176, 1662, 1426, 629, 1605, 1220, 2205, 695, 974, 2155, 852, 1670, 962, 245, 1587, 2083, 2077, 2172, 1114, 1345, 945, 1802, 825, 1808, 1926, 2080, 256, 482, 2177, 598, 1850, 2092, 1934]. Property [1378, 1673, 2000, 1424, 1774, 1179, 1007, 1767, 883, 541, 558]. Proportion [1362, 868, 1598, 1093, 1716, 1794, 2106]. Proportional [1513, 1392, 2123, 440, 1831, 1793, 2352, 2069, 2268, 1912, 1699, 1560, 2080, 1263, 449, 1606, 1234]. proportions [1499, 1993, 1849]. protecting [1823]. Prozentpunktschätzung [270]. Prozesse [558]. Prüfaufwands [173, 228]. Prüfpläne [636, 235, 675, 259]. Prüfplänen [212]. Prüfung [562, 584, 55, 191, 241, 268, 423, 59, 10, 13, 64, 149, 205, 259]. Prüfverfahren [239, 246]. Prüfverteilungen [161, 168]. PrY [533]. Pseudo [1428]. pseudorandom [814, 904]. psycho [1695]. psycho-physiological [1695]. psychologischen [52, 79]. Publications [322]. Publisher [2038]. pulley [742]. Punkthaufen [343]. Punktprozesse [402]. Punktprozessen [388, 513]. purely [2060, 2360]. Pyrrho [853].

QMLE [2221]. quadrant [1460]. Quadrat [584]. Quadrat-Tests [584]. Quadratic [1532, 345, 939, 236, 1213, 760, 986, 1615, 2151, 1024]. quadratically [768]. Quadrupling [2131]. Qualitative [1969, 1093, 2335, 415]. qualitativen [114]. quality [846, 1313, 1458, 638]. Quantile [1119, 960, 2062, 1968, 2334, 2049, 2132, 1543, 2003, 2157, 2273, 2113, 2191, 1976, 2353, 2260, 1829, 2307, 2233, 2271, 1720, 2151, 2144, 2321, 2254, 1975, 2153, 2107, 2259, 2089, 1973, 2206]. Quantile-based [2062]. quantiles [1993, 630, 1251, 1457, 641, 836, 1906]. quantitative [1270, 1558, 1449, 1589, 675, 2335]. quantitatives [675]. quantities [1590]. quantized [1168]. quarter [2223]. Quartile [83, 1118]. Quasi [752, 692, 969, 1423, 2210, 867, 233, 1474, 2018, 899, 2184]. Quasi-isotonic [692]. quasi-likelihood [2018, 2184]. Quasi-maximum [2210, 867, 899]. Quasi-stationarity [969]. quaternary [2031]. question [2]. queue [1487, 218, 339]. Queueing [36, 350, 441, 367, 354, 310, 192, 395, 394, 339, 349]. Queues [217, 197]. Queuing [238, 72]. Quick [899]. quotient [510]. Quotientenmethode [314].

R [1371]. Radmacher [1371]. raking [563]. Random [1367, 1839, 972, 1409, 2169, 1398, 577, 2294, 2007, 2138, 1246, 2318, 1962, 1446, 1512, 2235, 1490, 891, 1643, 1384, 826, 158, 2148, 1658, 2058, 1993, 924, 738, 1576, 1941, 1483, 2019, 2135, 2285, 594, 1617, 1996, 607, 1007, 1225, 629, 2296, 146, 940, 1937, 1897, 1945, 834, 1465, 497, 1584, 1460, 1954, 2038, 1809, 2063, 1529, 2013, 1228, 907, 11, 1047, 1203, 1587, 2284, 1216, 2185, 1271, 2091, 1579, 618, 1926, 1891, 2005, 2302, 963, 597, 1288, 184, 1476, 888, 2182, 1500, 1922, 1623, 148, 2312, 2324, 316, 1935, 1775, 1748, 2153, 2297, 1453, 1085, 1925, 1833, 1386]. random-effects [1922]. randomization [682, 560]. Randomized [1310, 868, 1270, 1598, 1135, 1558, 1716, 1267, 1499, 1184, 1449, 1277, 1630, 769, 1154]. randomly [1091, 748, 2087, 2311, 1761]. randomness [796]. Randsummen [30, 37]. Rang [54]. range [756, 1453]. ranges [756]. Ranggrößen [61, 126]. Rangtests [357, 354]. Rangvariable [556]. Rank [2230, 1414, 1419, 1837, 1402, 1399, 796, 147, 1489, 368, 1255, 2192, 2224, 587, 819, 602, 964, 649, 1982, 556]. Rank-based [2230]. rank-deficiency [2224]. Ranked [1558, 1755, 1157, 1994, 865, 1011]. ranking [1994, 915]. rankings [1446]. Ranks [2295, 659, 1818, 972, 1981]. rapidly [154]. Rare [1373, 1372, 372]. Rasmus [1383]. Rate [555, 998, 2123, 1868, 1077, 1348, 1426, 2141, 2252, 1912, 1699, 2025, 1669, 836]. rater [1486]. Rates [1665, 2354, 893, 1533, 339]. Ratio [1408, 226, 134, 186, 224, 919, 726, 735, 1490, 1960, 874, 960, 2299, 1785, 515, 2078, 563, 97, 2042, 957, 1271, 190, 570, 220, 888, 1567, 981, 1995, 2203]. Raum [273]. Re [1744]. Re-weighted [1744]. Real [2293, 2251, 2049, 1913, 1863, 2075]. Real-time [2293, 2049]. Realisation [559]. realization [559]. realizations [243]. realized [2261]. really [1006]. rearrangement [600]. recapture [2068, 1777]. Rechenmaschinen [15]. rechteckverteilter [501]. Rechtsvorschriften [110]. reciprocal [2094]. Recognition [989]. Reconstruction [1546, 1299]. Record [1395, 1039, 1089, 1778, 1609, 1506, 1898, 1441, 1523, 916, 1183, 1649, 1339, 2233, 1784, 945, 2028, 1866, 2147, 1492, 1721]. records [1542, 1609, 1546, 2152, 1962, 1774, 1967, 1687, 2024, 1299, 957]. Recovered [997]. recovery [863]. rectangular [829]. Recurrence [1292, 2188]. recurrent [1964, 2041]. recursion [942]. Recursive [181, 1323, 1261, 647, 955, 1244]. redescending [1338]. reduced [1199, 2051, 817]. reduction [2116, 2206]. redundant [358]. reduzierter [506]. reexamination [274]. Referees [989]. reference [863]. refined [2282]. Refining [2305]. regarding [1122]. region [903, 2030, 2297]. regional [2224]. Regionalstatistik [111]. regions [1781]. Regression [1418, 2111, 1283, 1357, 1404, 531, 1353, 1392, 1174, 1673, 2334, 1121, 2178, 2338, 2339, 2119, 1919, 1306, 1792, 1219, 1448, 1313, 1146, 2160, 871, 996, 1562, 1798, 2304, 876, 2249, 2196, 984, 808, 2132, 1543, 1884, 1432, 2003, 692, 644, 766, 939, 1740, 2285, 2189, 1913, 859, 848, 2157, 2156, 1865, 723, 2181, 1481, 847, 1611, 1600, 2208, 1110, 860, 1717, 554, 2343, 997, 1069, 2034, 2273, 1218, 157, 2265, 1233, 2081, 956, 2129, 1148, 1016, 1715, 1900, 1674, 2094, 2353, 1462, 1544, 2063, 2141, 832, 1698, 2260, 2267, 2289, 1466, 1829, 2018, 1950, 299, 2198, 2333, 1213, 2314]. regression [1451, 673, 1434, 1657, 479, 2233, 2271, 2284, 1515, 2025, 2248, 1949, 1055, 1345, 2231, 796, 861, 1106, 1444, 1020, 1568, 1696, 1508, 1535, 457, 2151, 1333, 2288, 1258, 1427, 1836, 2254, 491, 2171, 1113, 1724, 1572, 1244, 1805, 1852, 1975, 2053, 2136, 1690, 1850, 2311, 1463, 1856, 2153, 1492, 1721, 1638, 2107, 2259, 2145, 1971, 2089, 2214, 932, 2186, 1973, 1628, 1985, 2067, 1085, 1833, 31]. regressions [1164, 2321, 31]. Regressions- [31]. Regressionsanalyse [554, 229]. Regressionskoeffizienten [106, 167, 203]. regressors [1164, 1833]. regular [2292, 2175, 1982, 1327, 1886]. regularity [1792, 452, 2079]. regularization [956]. Regularized [2132]. Reihen [14, 454]. Reiss [1373]. Reiz [353]. Related [1404, 664, 2306, 629, 475, 2072, 1180, 1056, 1456]. relation [1738, 1725]. Relations [2024, 1292, 2188, 1788]. Relationship [1356, 518, 998, 926, 1177]. relationships [1896]. Relative [1402, 2114, 408, 1673, 685, 249, 1722, 288, 1847, 463]. Relevance [1048, 639]. relevation [2088, 2020, 2080].

Reliability [1990, 1422, 2139, 2080, 927, 2000, 1832, 2274, 849, 2205, 2250, 2062, 358, 925, 945, 2087, 674]. Remaining [1537, 893, 1806]. remark [645, 447, 647]. Remarks [209, 924, 1527, 392, 578, 1065, 65]. reneging [36, 238, 269]. renewal [930]. repair [878, 1990]. repairable [878]. Repeated [874, 1162, 2186]. replacement [440, 174, 46]. replicated [699, 1586, 109]. replicates [1321, 1269]. replications [2034]. Report [18]. repräsentativer [336]. representation [560, 1263]. Representations [1934, 304, 2345, 1906]. representing [943]. reproducing [1471, 2171]. Republic [327]. rer [231]. research [26, 626]. residual [998, 1609, 1236, 2283, 2048, 1827, 2256, 1945, 962, 1806, 1482, 2091, 1784, 2020, 1671, 1978, 1872]. residual-based [2048]. Residuals [1416, 1780, 1345, 1766]. resilience [2114]. resolution [413, 1920, 2359, 1924]. Resolvable [1278, 1321, 1269]. resource [1036]. resources [631]. resp [645]. Response [2174, 433, 868, 1318, 1746, 1270, 1598, 1135, 1558, 1716, 2246, 1267, 1184, 1310, 644, 1695, 1449, 1277, 407, 1282, 2141, 2065, 2314, 1554, 1630, 769, 963, 1154, 1836, 1563, 1775, 1851, 1607]. responses [1157, 1499, 2132, 2285, 2353, 2063, 2289, 1731, 1748, 2153, 1973]. restricted [1022, 452, 1658, 477, 1509, 1771, 1155, 2001, 1614, 1026, 1065]. restricting [1499]. restriction [157]. Restrictions [1379, 1932, 1438, 1635, 1933]. result [403]. Results [1704, 453, 664, 1951, 1538, 1895, 1236, 996, 2235, 1645, 2031, 610, 1854, 1501, 1988, 2008, 218, 858, 2106, 1709, 1700, 360, 378, 537, 1049, 2302, 1736]. Reverse [1007]. reversed [2088, 1831, 2354, 1348, 1688, 1426, 1513]. Review [1294, 1304, 1315, 1384, 1373, 1385, 1372, 1370, 1478, 1389, 1387, 1371, 1386, 1388, 100, 1476, 1477, 1383, 755, 821, 818, 807, 734, 805, 687, 736, 794, 813, 1280, 749, 771, 691, 773, 800, 390, 703, 777, 792, 689, 745, 715, 758, 786, 823, 764, 717, 724, 696, 722, 679]. reviewers [1930]. Reviews [34, 1066, 1081, 1101, 1115, 1141, 1150, 1166, 1223, 1231, 1240, 1264, 206, 1214, 62, 1195, 480, 216, 516, 40, 131, 459, 1196, 409, 473, 137, 248, 1175, 260, 87, 95, 159, 1188, 1197, 56, 432, 1185, 9, 1186, 1205, 465, 443, 76, 81, 178, 165, 296, 317, 377, 393, 49, 19, 171, 499, 143, 69, 1275, 525, 27, 1187, 788, 816, 693, 810, 775, 953, 947, 585, 592, 566, 672, 753, 902, 633, 928, 605, 684, 676, 797, 650, 576, 603, 747, 855, 894, 833, 936, 767, 601, 732, 880, 620, 613, 965, 548, 536, 920, 914]. reviews [617, 941, 713, 660, 611, 873, 850, 958, 623, 727, 706, 839, 710]. revisited [430]. reward [985]. reweighting [1473]. rho [1561]. RHR [2253]. Riani [1387]. ridge [1233, 2319, 479]. Riesz [1840]. Right [2338, 1131, 1878, 1680, 2298, 1900]. Right-censored [2338, 1680, 1900]. Risikoaversion [162]. Risikos [139]. Risk [681, 685, 1255, 1530, 2360, 1847, 1553, 1113, 1665, 2297]. Risk-efficient [681]. Risks [1413, 1901, 1902, 1998, 2055]. Robust [1959, 2105, 1907, 2040, 1149, 1633, 1879, 2285, 1899, 2337, 2265, 1129, 659, 1516, 1976, 2234, 2333, 1126, 1981, 1731, 2144, 2301, 2280, 1396, 1572, 1120, 1083, 1846, 1003, 848, 2110, 2270, 1783, 1655, 1881, 2121, 619, 2198, 1521, 1563, 1128, 1856, 1566, 2214]. Robustness [1122, 1136, 514, 447, 1969, 2059, 970]. ROC [1956]. Roger [2070]. Rolf [1373]. Rolf-Dieter [1373]. Romano [1388]. Ronald [96]. room [441]. Root [1416, 2196, 1484]. roots [2130]. Rosenblatt [2249]. Rosenthal [1891]. Rosenthal-type [1891]. rotatability [776, 841]. rotatable [2150]. Rounding [1380, 1086, 1996]. row [1149, 1074, 759, 642]. row-column [1149, 1074, 759]. row-columns [642]. Rows [1379, 780]. rowwise [2005]. RRSB [1754]. Rue [1476]. ruin [1553]. rule [934, 971, 879, 612]. rules [1783, 1086, 266]. Run [1351, 1017, 881, 1019, 1639, 1433, 2131, 977, 252, 98]. Run-Testes [98]. Runs [1483, 1645, 2229, 1946, 1352, 2072, 1955].

S [1389]. s.o [2315]. Saddlepoint [2110, 2012]. safe [1660]. Sagoroff [272, 319]. Saisonbereinigung [410, 337]. same [1310]. Sample [1078, 1330, 1273, 950, 1915, 147, 2098, 1994, 1267, 865, 108, 1993, 1253, 1100, 1686, 1947, 1161, 682, 132, 138, 407, 390, 1262, 29, 819, 1529, 806, 968, 1001, 2308, 1252, 731, 1127, 309, 1579, 1808, 641, 1079, 836, 578, 2312, 2324, 10, 1851, 1057, 1050]. sampled [1499, 2276, 621]. Samples [1382, 1404, 1953, 2056, 1519, 145, 375, 1072, 937, 2148, 1878, 890, 2078, 288, 251, 1485, 468, 2327, 1916, 1949, 1417, 588, 778, 910, 1984, 313, 2331, 1781, 2030, 2297, 1619, 2095]. Sampling [70, 91, 307, 666, 1403, 737, 577, 1107, 667, 1755, 1157, 1558, 1716, 440, 1061, 1335, 1549, 751, 174, 799, 900, 1794, 1225, 1589, 1991, 2360, 1475, 2113, 132, 138, 1732, 1137, 438, 251, 421, 701, 1472, 502, 190, 618, 830, 564, 769, 220, 963, 1051, 542, 708, 883, 1288, 177, 573, 433, 910, 981, 1480, 46, 1013, 765]. satisfying [1912]. saturated [2099]. Sätzen [622]. Saunders [1919, 2019, 2198, 1999, 2095]. scale [1756, 1133, 1083, 2148, 1314, 645, 1783, 2023, 931, 979, 2360, 1218, 1616, 2078, 1124, 1044, 1772, 2267, 1155, 2046, 1692, 2001, 721, 729, 1138, 1781, 1859]. scale-contaminated [1133]. scale-free [2046]. scaled [1098, 1954, 2038]. scales [304]. scaling [2318]. scanning [879]. scatter [1050]. scattered [1644, 1494]. scatterers [343]. scattering [12]. Scharfe [559]. Schätzbare [414]. Schätzen [506, 402]. Schätzer [531, 526, 405]. Schätzfolgen [381]. Schätzfunktionen [148, 126].

Schätzproblem [529]. Schätzproblemen [492]. Schätzung [464, 454, 462, 16, 316]. Schätzungen [554, 101]. Schätzverfahren [652]. Schätzwerte [64]. scheduling [387]. scheme [737, 190, 883]. schemes [1146, 2347, 2026, 969, 972, 3, 1550]. Schichtabgrenzung [331]. Schied [1478]. Schlecht [259]. Schlechtanteile [191]. Schlechtstücke [116]. schnell [154]. Schnittkreise [176]. Schranken [559]. Schriftenverzeichnis [322]. Schur [1926]. Schützenberger [757]. sciences [22, 150, 194]. scientific [1930]. Score [1423, 1569]. scores [1930]. scoring [766]. Scott [789]. scrambled [1630]. Screening [852, 812, 846, 892, 1095, 2033, 791, 1976, 1894, 2016, 2054, 2144]. search [1112, 1095, 1991, 1387]. Seat [1263, 1308]. Second [2315, 1861, 776, 2208, 1047, 2302, 173, 1334, 1744]. Second-order [2315, 2208, 1334, 1744]. second-stage [173]. Secretary [579]. sectional [2124]. sectionally [2014, 1484]. sections [1615]. seemingly [1628]. seen [299]. seismic [1201]. seismology [2336]. selected [1997, 772, 1211, 1015, 641]. Selecting [1535, 932]. Selection [1208, 934, 1879, 2344, 1801, 2352, 1991, 2183, 659, 1758, 2193, 1894, 655, 1931, 1604, 1826, 1709, 1451, 915, 806, 1669, 2305, 1836, 2184, 1723, 2073, 2145, 2317, 46, 1663, 1659, 1761]. selections [2057]. selector [1758]. self [1997, 1979, 1882]. self-selected [1997]. self-weighted [1979]. Seltene [262]. seltener [372]. Semi [2063, 2161, 1278, 1248, 1575, 1632, 2067]. semi-balanced [1278]. Semi-functional [2063, 2161, 2067]. semi-parametric [1632]. semi-Pareto [1248]. semidefinite [1718]. semifoldover [1841]. semifoldovers [1903]. Semiotische [152, 172]. Semiparametric [1715, 1950, 1606, 1255, 2270, 1123, 1900, 1522, 1532, 2053]. Sen [827, 190, 883]. sense [587]. sensitive [1558, 1499, 1310, 1823, 870, 1480]. sensitivity [1598, 1093, 2257, 1589, 1759]. separable [484]. Separating [1560]. sequence [1483, 1559, 1864, 169, 1955]. sequences [1962, 2229, 629, 1961, 2225, 1862]. Sequential [743, 1627, 971, 1070, 738, 608, 565, 1268, 1064, 1138, 1805, 2060, 2275, 1868, 1222, 681, 1062, 1717, 1655, 1881, 1235, 1605, 2130, 1661, 1848, 2360, 612, 1287, 783, 1711, 801, 915, 993, 578, 888]. sequentially [925]. Sequentielle [74]. serial [1936]. serially [2098, 1918]. serie [2047]. Serienfertigung [140]. serienweisem [116]. Series [1412, 2322, 453, 486, 1951, 632, 1226, 1043, 2221, 1791, 1892, 2133, 2296, 2356, 1691, 2036, 2141, 655, 2045, 503, 508, 520, 14, 1749, 1949, 1981, 1646, 392, 1284, 1268, 2177, 1269, 2184]. series-parallel [2296]. server [1487]. servers [349]. service [217, 197, 367, 354, 350]. Session [18]. set [1755, 1157, 1558, 1994, 865, 1901, 1902, 640]. sets [2294, 1445, 1643, 1993, 1557, 622, 2279]. Setting [843, 938]. settings [2175, 1921]. setup [2071]. Several [1396, 147, 1096, 2211, 2084, 1551, 2078, 1769, 1858, 2064]. sFDA [2231]. Shannon [1649]. Shape [2066, 1165, 1710, 2219, 2244, 1785, 2084, 1858, 1933, 1995, 591, 718, 1296, 1859]. shaped [897, 1534, 2028]. Sharp [1054, 1210, 559]. sheep [1159]. sheets [1445]. Shewhart [1242, 1019]. shift [436]. shock [2103, 313]. shocks [2294, 2019]. short [263, 1539, 2210]. shortest [1835]. shot [2105]. Shrinkage [868, 1237, 2043, 1855, 1977, 1876, 971, 2168, 1804, 2057, 2184, 1026]. shrinkage-pretest [1804]. shrinking [1354]. Sided [1363, 892, 671, 952, 580, 1835, 791, 779, 793, 1301, 649, 1985]. sign [249, 351]. signed [368]. significance [906, 918]. significant [222, 232, 2254]. signifikanzniveau [505]. Simar [1370]. SIMEX [1697]. SIMEX-based [1697]. Simon [1371].

Simple [1307, 1828, 1599, 2035, 440, 879, 1864, 1191, 2061, 1657, 537, 28, 10, 653]. simplest [2101]. simplicially [1161]. simplifying [1789]. simulation [1988, 1372]. Simulations [1036]. Simultan [75]. Simultaneous [1096, 1706, 1651, 2050, 2201, 565, 959, 1985, 2361]. Single [1953, 1415, 2212, 1951, 812, 892, 744, 1797, 1569, 1487, 2029, 1680, 2213, 2273, 2018, 1824]. Single-group [1415]. single-index [1569, 2029, 1680, 2213, 2273, 2018, 1824]. singular [1065]. sinusoidal [1164]. Sir [96, 96]. Site [1201]. Six [1321, 1068]. size [1078, 440, 1074, 1686, 1986, 1330, 1952, 806, 367, 1911, 1940, 1108, 910, 2312, 2324, 2297]. size- [1986]. size-biased [1952]. sizes [1100, 834, 2131, 588]. skew [1689, 1844, 1688, 2264, 2149, 1974, 1921]. skew-elliptical [1844]. skew-normal [1688, 1974, 1921]. skew-symmetric [2149]. skewness [1966, 968, 1001, 540]. skin [1695]. Slawtscho [319, 272, 319]. Slope [134, 2230]. Slow [1437]. Small [1309, 1373, 1127, 1953, 882, 1947, 2051, 2276, 1788, 1911, 1940, 1564, 1843, 1514, 313]. smallest [2019, 2302]. Smooth [1155, 2125, 1972, 750, 917, 1071, 680]. smoothed [1569]. smoothers [1338]. Smoothing [1913, 569, 1003, 496, 1956, 870, 1132, 1705]. smoothness [2171]. soccer [102]. social [2140]. Society [150, 194]. Solidarity [598]. Solution [600, 339, 305, 430]. solutions [276, 1527, 2154, 483, 451, 391, 163]. Some [453, 1236, 2009, 844, 1395, 644, 1352, 610, 1854, 1501, 2205, 250, 2250, 858, 2106, 1618, 1709, 1700, 1365, 1587, 1363, 175, 618, 400, 256, 2204, 578, 295, 1736, 313, 1026, 2092, 1065, 1329, 2339, 829, 2035, 1023, 2160, 871, 735, 730, 1531, 854, 2189, 2110, 762, 1988, 946, 1848, 940, 992, 1124, 390, 652, 1261, 429, 421, 1804, 470, 1896, 1325, 2269, 1765, 981, 1934, 844, 392, 10]. sooner [1864]. source [2125]. Sozialismus [93]. sozialwissenschaftlichen [107]. space [1471, 1747, 484, 1817, 1003, 72, 1581, 2357, 339, 349, 2350]. space-filling [2350]. spaces [2336, 1262, 1732, 1079, 2171]. spacings [2235, 1490, 1643, 1883, 1957, 1687, 1863, 1650, 2147]. Span [55]. spanning [2341]. Spannweitenpläne [55]. Sparse [2351, 2118, 2248, 2259, 2145]. Sparse-penalized [2351]. Sparsified [2361]. Spatial [1461, 1588, 1747, 1515, 2231, 2357, 1977]. spatially [1164, 2262]. spatio [2356]. spatio-temporal [2356]. SPC [1019]. SPC/ [1019]. SPDEs [2357]. Spearman [1561]. Special [2044, 514, 18, 1814, 844, 863, 707, 560, 434]. species [472]. specific [1759]. specification [1041, 2245]. specified [1273, 1309, 1053, 1765]. spectral [1725, 1904]. spectrum [1481, 1611, 1258, 1427]. speed [1329]. Spezialtagung [18]. Sphäre [316]. sphere [1494, 316, 550]. spheres [1732]. spherical [2283]. spherically [1827, 1899]. sphericity [2203]. Spiel [330]. Spiele [90]. spieltheoretische [212]. Spline [2018, 1879, 1913, 2183, 2321, 1811, 1132, 1824]. Spline-based [2018, 1879]. splines [956]. split [2194, 2143, 2292, 1278, 2323, 1143, 2227, 1733]. split-block [1278, 1143]. split-plot [2194, 2143, 2292, 2323, 2227, 1733]. spot [2325]. SPR [637]. spread [288]. Springer [1372, 1370, 1371]. Square [1358, 2016, 1136, 2329, 952, 1061, 285, 918, 584, 1711, 33, 674]. squared [2240, 1673, 784, 1427, 1886]. squared-log [1886]. Squares [1364, 2178, 871, 2189, 1152, 2208, 956, 1230, 974, 2319, 1788, 787, 961, 1434, 401, 274, 2053, 2311, 1833]. Stabilität [115]. Stability [1001, 163]. stabilizing [802]. stable [1948, 1839, 1904, 1105]. stable-Paretian [1948]. Stage [1403, 1997, 667, 846, 1061, 1335, 1881, 2022, 2360, 1277, 2201, 985, 686, 815, 2315, 251, 701, 721, 373, 542, 173, 1334, 910]. stages [743, 173]. standard [237, 1759, 2327, 500, 1097]. standardisierte [167]. Standardized [2277]. standby [1832, 2071, 2139]. Stange [59]. star [897, 1534]. star-shaped [897, 1534]. stärkste [75]. starting [558]. Startverteilungen [558]. state [1889, 1965, 1003, 339]. stationärer [402]. stationarity [969, 1578]. Stationary [1412, 2014, 1376, 942, 1238, 1086, 2176, 1010, 655, 2169, 1263, 2162]. Statistic [1359, 1416, 744, 742, 1785, 182, 2319, 1060, 2102, 133, 1567, 396]. Statistical [26, 1965, 2241, 1311, 477, 1161, 1662, 2353, 668, 2158, 472, 2151, 789, 455, 452, 1622, 1370, 553, 733, 900, 102, 624, 1930, 1047, 14, 600, 523, 1388, 184, 1247, 638, 1177, 1973, 1344, 1389, 1383]. statistician [299]. statisticians [1333]. Statistics [6, 466, 1375, 1253, 607, 18, 1477, 1778, 1182, 1292, 927, 1966, 2294, 1844, 1360, 1895, 2216, 795, 1219, 1448, 1424, 599, 1868, 2140, 1222, 826, 984, 1884, 896, 1639, 2253, 2019, 916, 1054, 2165, 1062, 1883, 1957, 2354, 435, 772, 2023, 555, 475, 24, 2130, 236, 1210, 634, 992, 1233, 1036, 1818, 1954, 2038, 204, 1266, 1529, 1339, 287, 1297, 2013, 1485, 1482, 954, 1173, 1847, 1625, 1252, 731, 945, 1802, 1539, 825, 602, 1564, 1843, 255, 641, 694, 291, 626, 570, 574, 778, 2302, 1006, 1439, 964, 1736, 654, 1037, 1666, 2055]. statistics [877, 1781, 2030, 1050]. Statistik [18, 105, 152, 172, 6, 107, 626]. Statistiker [136]. Statistische [355, 110]. Statistischen [230, 275, 347, 280, 326, 67, 334, 336]. statistischer [115, 14]. status [2069]. Steffensen [1360]. Stein [1689, 1899, 1460, 2057, 1870]. Stein-type [2057]. step [1442, 1538, 1599, 2183, 1961, 1581]. step-down [1442]. step-stress [1538, 1599]. stepwise [299]. Sterbewahrscheinlichkeiten [324, 338].

Stereologie [308, 352]. Stetige [527, 205]. stetigem [353]. stetiger [5]. Stichproben [107, 353, 239, 246, 61, 265, 333]. Stichproben-Theorie [353]. Stichprobenplan [411]. Stichprobenpläne [415]. Stichprobenplanes [416]. Stichprobentheorie [414]. Stichprobenumfang [139, 270, 78]. Stichprobenumfanges [297]. Stichprobenumfangs [384]. Stichprobenverfahren [406, 111]. stigmatized [1803]. Stochastic [1780, 1599, 2316, 385, 2332, 1954, 2038, 1809, 2013, 2006, 2037, 1927, 1889, 2123, 2209, 2058, 692, 1810, 1527, 1486, 2023, 1991, 1010, 1871, 858, 1090, 66, 1701, 301, 1896, 472, 901, 163, 181, 292, 712, 2269, 1154, 1819, 1969, 598, 208, 1193, 558, 210, 1478]. stochastically [1917, 2290]. stochastische [527, 210]. stochastischen [279]. stochastischer [558]. stock [514]. stopping [971, 1848, 612, 266, 1819, 2225]. Straight [897]. Straight-line [897]. strata [1686]. Strategies [1403, 950, 1135, 1239, 586, 2315, 1804, 701, 502, 1288]. strategy [489, 827]. Stratification [827, 442, 571, 449, 2350]. stratified [577, 174, 1277, 390, 564]. Stream [350]. Strecker [392]. Strength [1369, 1917, 1152, 1422, 2232, 2071]. Stress [1422, 1447, 1915, 1538, 1599, 1917, 2071]. Stress-strength [1422, 1917, 2071]. Streuflächen [343]. Streuung [106, 12]. Streuungszerlegung [183]. Strong [1512, 1342, 1055, 1244, 1833, 739, 1527, 2008, 1691, 2232, 2005, 1942, 924, 555]. strongly [1905]. structural [2244, 619, 1911, 1940]. Structure [2347, 1340, 1147, 1023, 814, 842, 2033, 770, 1936, 2004]. structured [1893]. structures [1035, 1893, 534, 1704, 2358]. Student [2188, 1233, 1638]. Student- [2188, 1638]. Studentized [1069, 654]. Studien [626]. Studies [1478, 70, 91, 826, 1695, 1282, 1036, 2, 626]. study [1473, 1832, 1987, 1027, 472, 1597]. Stufe [173]. Stufen [173]. Sub [463, 251]. sub-samples [251]. Subject [1366, 1627, 534, 1536, 1550]. Subjective [564, 209]. submartingale [1007]. submitted [191]. subplot [2292]. subpopulations [307]. subranges [1083]. Subsampling [1012, 2304]. Subset [1356]. substitution [874, 1617]. Subsystems [1414]. success [2303]. Suche [110]. Sufficiency [379, 417]. sufficient [599, 634, 967, 560, 2116, 1834, 1869]. Suffizienz [492]. sum [1559, 2319, 184]. summation [497]. Summen [597]. sums [1313, 1512, 484, 1996, 1436, 1650, 2185, 1345, 1533, 597]. superadditive [1891]. superadditivity [754]. supercritical [2318]. superiority [1065]. superpopulation [498, 586]. Supersaturated [1381, 1746, 2032, 2194, 2200, 1221, 1502, 1495, 1557, 1679, 1160]. supplementary [1557]. Support [1405, 2111]. supported [1418]. supporting [968]. sure [1625, 1665]. surface [907, 1563, 1607]. surfaces [776]. surrogate [2113, 1908]. Survey [1475, 1333, 2056, 876, 108, 1099, 1823, 1600, 1794, 2298, 1137, 255, 963, 1051, 708, 491, 981, 1480, 1816]. surveys [1107, 1267, 407, 2074, 489, 1630, 2064, 975]. survival [1631, 982, 1928, 1938, 2298, 1091, 893, 962, 1613, 2054, 557, 1808, 1995]. survivor [772]. suspected [2340]. SV [1710]. Switching [531, 2071]. symmetric [1827, 1899, 1683, 2149, 628]. symmetrical [2257, 790]. Symmetrieprüfung [485]. symmetrischen [90]. symmetry [1147, 366, 952, 1075, 420]. synthesis [1538]. system [879, 72, 2199, 2205, 2296, 2250, 358, 2036, 1809, 1806, 1699, 2071, 1671, 2139, 1597]. systematic [1225, 220]. systematische [324]. systems [1889, 878, 1729, 441, 1868, 285, 1832, 1960, 1990, 1987, 2256, 1701, 252, 1482, 2006, 2037, 238, 269, 2087, 1978, 1934, 1660].

table [2309, 99, 284, 263]. tables

[2329, 366, 952, 1622, 1932, 1100, 1713, 1585, 1983, 33, 30, 37, 1026, 420]. Tafeln [485]. Tagung [142]. Tail [1626, 1668, 2108, 1297, 1904, 1456, 2326, 1057]. tailed [2198, 1012]. tails [2055]. taking [407]. target [918]. Tatsachen [166]. te [513]. teaching [1036]. technique [283, 386, 1267, 1184, 1310, 1630]. Techniques [32, 1228]. technology [1037]. technology-based [1037]. Teil [85]. temporal [2356, 2357]. Term [1340, 370, 2280]. terminal [1964, 2041]. terms [2156, 962, 1847, 1693]. Test [83, 2049, 1407, 1412, 167, 1408, 1201, 67, 94, 1396, 35, 1416, 420, 843, 949, 2070, 1489, 636, 562, 86, 99, 91, 145, 249, 284, 368, 375, 1470, 1797, 2228, 1076, 2249, 1923, 2192, 1041, 2230, 2170, 1786, 2048, 1972, 478, 869, 1728, 790, 1636, 584, 742, 1235, 1605, 24, 1218, 2078, 1233, 1191, 1668, 682, 1684, 1937, 2193, 675, 1063, 1759, 2045, 288, 653, 770, 993, 212, 23, 651, 2167, 604, 756, 291, 521, 712, 173, 729, 888, 2245, 2225, 1162, 674, 2203, 2358, 1743, 1128, 510, 591, 1638, 1971, 932, 1296, 1752]. test-statistics [1233]. Testen [511]. Testes [604, 98]. Testing [746, 1174, 952, 2109, 1932, 1227, 2126, 1785, 286, 1674, 779, 1858, 1857, 1635, 1911, 2197, 1393, 2004, 1748, 1705, 1022, 1273, 774, 898, 2105, 671, 1120, 214, 943, 2192, 2097, 723, 1570, 2142, 2084, 2078, 2081, 793, 1168, 2106, 1228, 535, 651, 1578, 2167, 796, 1331, 778, 990, 1567, 1575, 10, 2047, 313, 1498, 1634, 1940, 1388]. Tests [1190, 366, 75, 1314, 1358, 906, 918, 584, 1402, 1281, 66, 1075, 1399, 1363, 54, 505, 240, 10, 1921, 2067, 1022, 1447, 2240, 2322, 927, 147, 752, 1821, 1506, 1538, 2056, 1907, 2040, 1519, 2120, 581, 127, 351, 2096, 1255, 1023, 1313, 996, 735, 730, 1846, 1385, 1253, 1100, 2012, 1878, 2299, 804, 917, 1280, 637, 1717, 532, 2034, 1837, 2266, 157, 994, 1071, 1757, 1880, 680, 587, 1413, 1726, 819, 2278, 1939, 2090, 968, 1209, 1983, 2046, 71, 1561, 754, 796, 1579, 1981, 537, 1079, 373, 1484, 257, 578, 13, 55, 191, 1181, 1464, 2254, 649, 728, 1638]. Tests [52, 79, 57, 43, 71, 511]. Testverfahren [485]. tetrachorischen [185]. text [1036]. th [1898, 2024]. their [1147, 1395, 141, 2031, 1352, 1725, 435, 1380, 132, 138, 1274, 1485, 1365, 1047, 408, 2005, 148, 1666]. Theorem [60, 196, 1146, 1993, 2128, 1529, 1650, 28, 1245, 555, 1625, 28, 332]. Theorems [198, 1687, 951, 1460, 400]. theoretic [1339, 1208, 212, 838]. Theoretical [1561, 1746, 65]. Theorie [353, 526, 53, 156, 357, 354, 156]. theories [125]. theorists [1754]. Theory [1476, 2350, 290, 1537, 158, 2196, 1710, 631, 607, 483, 250, 538, 1328, 1475, 132, 138, 407, 1262, 253, 663, 618, 192, 609, 445]. there [158]. third [841]. Thompson [223, 290, 431]. Three [1146, 2022, 686, 743, 636, 1889, 1321, 882, 1326, 1433, 1095, 646, 2258, 815, 993, 1334, 1327, 2279, 2359]. three-level [1327, 2359]. three-order [2279]. Three-stage [2022, 686, 815, 1334]. three-state [1889]. threshold [2220, 2286]. Time [1351, 1412, 2322, 453, 486, 1951, 1965, 871, 1171, 1964, 1631, 1226, 2021, 1310, 2049, 2293, 1471, 1043, 2221, 1791, 1892, 1946, 1760, 1987, 2133, 2041, 1123, 1864, 197, 2356, 1691, 2141, 655, 793, 783, 1701, 2045, 2071, 2091, 1105, 2169, 1749, 1949, 3, 175, 1981, 1646, 392, 1284, 1268, 1819, 2210, 905, 2184, 2047, 1872, 1934, 1478]. time-discretization [793]. time-independent [1123]. time-intervals [3]. time-transformed [1760]. time-varying [1171, 2021, 2210]. times [1528, 1848, 1299, 2271, 1567]. Todesprozeß [120]. Todesprozesses [464]. tolerance [1122]. tolerances [1114, 1259]. Toleranzbereich [241, 268]. tool [1033]. Top [1406, 1446]. Top- [1406]. total [664, 1107, 1662, 1230, 438]. Totally [1110]. totals [563, 30, 37]. tough [1000]. toxicological [2105]. trace [1098]. traditional [24]. Training [1382]. trajectories [2313]. transform [2066, 2088, 1883, 2142, 2020, 2080, 2310, 427]. Transformation [338, 2309, 196, 2333, 1144]. transformation-based [1144]. Transformationen [469]. transformations [802, 1878, 436, 1828]. Transformationsverfahrens [196]. transformed [1760, 1711, 1872]. transforming [1318]. transforms [358]. transient [339]. transienter [398]. translation [669]. Transportation [283, 1049]. trapezoidal [1204]. treatment [2349, 1551, 212, 1217, 13, 1247]. Treatments [1407, 1329, 1551, 1063, 642, 1162]. Treaty [1201]. Tree [1972, 2246]. trees [954]. Trefferereignis [25]. treffertheoretischen [65]. Trend [1351, 652, 1519, 86, 91, 759, 1578, 649]. Trendextrapolation [80]. trending [1791]. Trennanalyse [73]. trennscharfer [57, 71]. Trennung [78]. Trennverfahren [114]. trials [86, 2229, 2303, 1864, 169, 1955]. triangular [924]. trigonometric [1798, 1856]. trigonometrischer [454]. Trimmed [1900, 1898, 1947, 1057]. Truncated [1357, 1610, 1103, 581, 632, 2188, 389, 1897, 1945, 1900, 1593, 819, 2308, 508, 520, 1217, 570, 1163, 1995]. truncation [707]. Tsallis [2062]. tubes [1319]. Tukey [1057]. tumor [2105]. tumour [557]. Turing [2068]. Two [387, 1382, 1112, 1095, 2360, 1404, 1422, 1764, 2201, 819, 985, 418, 1202, 1480, 737, 2294, 1997, 1329, 667, 1944, 846, 892, 2275, 1907, 533, 2316, 2098, 996, 1622, 1914, 1335, 2249, 1310, 1309, 413, 2340, 882, 1841, 904, 2200, 1253, 580, 1835, 762, 702, 356, 1672, 608, 2244, 1881, 2022, 2292, 2183, 1713, 2078, 1277, 682, 97, 2129, 1648, 974, 1180, 237, 852, 1042, 2232, 1771, 977, 1155, 1092, 476, 2315, 1301, 421, 2046, 1271, 701, 2001, 721, 615, 643, 472, 309, 2028, 1560, 2139, 537, 625, 1164, 257, 642, 173, 1765, 910, 728, 2357, 1269, 10]. two [2227, 1721, 1903, 1298, 1851, 1924]. two- [2227]. two-dimensional [1164]. two-factor [1309, 1765, 1924]. Two-level [1112, 1841, 2200, 2292, 852, 977, 1765, 1903, 1298]. Two-parameter [1422, 533, 1914, 2244, 2078, 2028, 625]. two-phases [996]. Two-sample [819, 1253, 682, 10, 1851]. two-sided [892, 580, 1835, 1301]. Two-stage [2360, 2201, 985, 1997, 667, 846, 1335, 1881, 1277, 721, 910]. two-step [2183]. two-way [1622, 1713, 1092, 728]. Type [1375, 1395, 1025, 799, 780, 1542, 1506, 2040, 1313, 2160, 1994, 942, 1948, 1873, 1199, 1207, 2010, 1042, 2108, 1650, 2090, 354, 350, 1491, 1520, 434, 395, 363, 408, 445, 1891, 460, 2057, 1875, 654, 1607, 1599, 1642, 1914, 937, 1878, 1670, 2027, 2102, 1802, 2095, 1915, 1890]. type- [363]. Type-I [1599, 1642, 2102, 2095]. Type-II [1599, 1914, 937, 1878, 1670, 1802, 2095, 1915, 1890]. types [1893, 2205, 2250]. Typisierung [113].

U [1873]. U-type [1873]. überall [57, 71, 10]. überlappenden [85]. Überschreitungswahrscheinlichkeiten [77, 225]. Übervölkerung [166]. Uhlenbeck [2241, 1530, 1970]. Uhlmann [1676]. ultrahigh [2132, 2016, 2054, 2144]. ultrahigh-dimensional [2132, 2054]. UMP [1022]. UMPU [1100]. UMPU-tests [1100]. UMVUE [1103, 937]. unabhängiger [501]. unavailability [970]. unbekannten [355]. unbekannter [268]. Unbiased [1589, 1690, 1022, 678, 2216, 544, 599, 919, 632, 700, 645, 595, 817, 419, 760, 503, 508, 520, 289, 359, 827, 625, 798, 769, 460, 379, 1427]. unbiasedness [735, 1256]. uncertain [1751, 1805]. uncertainty [733, 400]. underdispersed [2100]. Underlying [1396]. Unequal [1407, 737, 1335, 2303, 1472, 588, 778, 666, 910, 1013]. unequally [699]. unexpected [1673]. Unfavorable [582]. unfolded [1916]. Unfolding [1053]. Ungewißheit [202]. Ungünstige [582]. ungünstigster [528]. unified [1844, 2138, 1519, 158, 2026, 407, 1467, 1217, 1736, 716]. Uniform [2195, 2191, 798, 739, 2328, 2272, 1337, 1424, 1722, 935, 2031, 1286, 1996, 2236, 1691, 1732, 2131, 1055, 1936, 482, 1666, 1244]. uniformity [1545, 1640]. Uniformly [809, 599, 1494, 1079]. unimodal [448, 875]. unique [290]. Uniquely [922]. uniqueness [304, 1527, 2095]. Unit [1484, 1416, 1494, 905]. units [1335, 1499, 2150, 173, 46]. unity [2196]. Univariable [122]. Univariate [1420, 1689, 486, 1688, 2298, 1896]. Universal [2160, 1908, 1044, 71]. Universally [2247, 2015, 1766, 10]. universally-effective [10]. universe [298]. universes [132, 138]. Unknown [1396, 881, 2226, 791, 2104, 1590, 2288, 910]. Unlacing [1816]. unordering [431]. unpaired [288]. unrelated [1628]. Unreplicated [1351]. unsere [105]. unten [241, 268]. untere [559]. Unterschiede [324]. unterschiedlichen [173]. unterschiedlicher [173]. Untersuchung [426]. Untersuchungen [506, 2]. Untervölkerung [166]. unverfälschte [75]. unverfälschter [240]. unverzerrte [64]. unvollständigen [176, 199]. unwirksamen [210]. up-dating [489]. upon [2296, 1777]. Upper [294, 386, 97, 957, 1843, 559]. urn [1329, 1750, 1199]. urn-based [1750]. Ursachenforschung [326]. Use [521, 1113, 109, 1069, 2186]. used [1273, 2155]. useful [1537, 619, 332]. Using [1367, 1362, 2318, 99, 919, 1716, 865, 1820, 568, 2200, 1718, 960, 1099, 2039, 720, 772, 1794, 1098, 1328, 2121, 616, 1034, 1277, 956, 2113, 1339, 1063, 1604, 763, 1203, 2233, 2271, 1692, 1568, 955, 641, 963, 1682, 1439, 606, 184, 1866, 1346, 2280, 2107, 1705, 1824]. Utilitätskurve [162]. utilities [1446, 733]. utility [1446].

Vague [1393]. validity [838]. value [1786, 1227, 1785, 1863, 2084, 1826, 1228, 708, 1754, 2055, 1721, 932]. valued [1610, 2281, 1817, 2221, 1152, 2184, 2100, 2286]. Values [1395, 1089, 1778, 1609, 1506, 1898, 1441, 916, 2233, 1001, 993, 1784, 1039, 1866, 148, 316, 2147, 2313, 1492, 1721]. VaR [1812]. varentropy [2332]. Variability [1958, 1199]. Variable [1801, 1758, 316, 1723, 2145, 1663, 1659, 1879, 556, 2352, 1991, 791, 834, 1282, 1920, 1931, 1228, 367, 1579, 1615, 648, 369, 2312, 2324, 396, 2107, 1761]. Variablen [73, 652, 148]. Variables [1364, 1367, 1409, 1398, 222, 232, 2294, 2007, 2138, 1040, 1512, 2235, 1490, 924, 1576, 1483, 2135, 594, 900, 1823, 1996, 1007, 2296, 940, 619, 652, 1192, 1230, 1897, 1945, 1460, 1954, 2038, 32, 494, 972, 2013, 2314, 11, 1047, 1587, 648, 1926, 457, 332, 1891, 2005, 2302, 597, 184, 148, 316, 1805, 1935, 1980, 2107, 1859, 2159, 765]. Variance [802, 919, 2349, 1409, 1106, 1013, 344, 678, 1318, 544, 1470, 599, 632, 495, 552, 768, 891, 568, 881, 1697, 874, 1996, 2033, 1589, 515, 250, 595, 784, 1764, 2117, 412, 983, 760, 498, 565, 586, 251, 2308, 503, 508, 520, 1472, 1707, 2167, 827, 1579, 190, 1646, 5, 862, 798, 769, 627, 963, 542, 173, 100, 590, 227, 665, 728, 1921]. variance-estimation [552]. variance-estimator [227]. variance-truncated [2308]. variance/covariance [1764]. Variances [2197, 1396, 518, 1729, 1481, 563, 778, 1258, 910, 1743]. Variante [314, 493]. variants [2035]. Varianz [101, 5, 173, 241, 268, 164]. Varianzen [61, 501]. Variate [314, 1876, 1384, 1641]. Variation [1424, 2325, 1662, 1009]. variogram [2110].

Various [1402, 878, 943, 1046, 1654]. Varying [2174, 1392, 1107, 871, 1171, 2021, 1879, 1801, 1680, 2213, 2183, 1674, 1976, 2124, 2069, 1931, 3, 2210, 258, 1632, 1811, 1498, 1775, 1748, 46, 1663, 2206]. Varying-Coefficient [2174, 1392, 1680, 2213, 2183, 1674, 2069, 1632, 1498, 2206]. Vaterschaftsbegutachtung [335]. vector [2283, 1827, 236, 659, 1836, 2190]. vectors [1827, 2050, 1726]. Vektoren [355]. verallgemeinerten [511]. verallgemeinerter [120]. Verallgemeinerung [196, 458, 372]. verbandstheoretische [152, 172]. Verbesserungen [437]. Verfahren [355, 28, 118, 78, 166]. Verfahrenskontrollen [336]. Vergleich [399, 485, 118, 193]. Vergleiche [183]. Verhalten [437, 139, 584]. Verkehrszählungen [118]. Verlag [1370]. vermutlichen [327]. verschiebungsinvarianten [492]. verschiedene [156]. verschiedenen [458]. versions [770, 1561, 1456]. Versuchs [74]. Versuchsanordnungen [199]. Versuchspläne [340]. Verteilung [176, 2, 616, 207, 458, 173, 191, 228, 423, 205, 273, 262, 205, 193]. Verteilungen [652, 43, 582, 43, 528]. verteilungsfreien [106]. Verteilungsfunktion [493]. Verteilungsindizes [39]. Verteilungsmaße [39]. Vertical [128, 135]. Vertrauensgrenzen [167]. Verzerrte [64]. Verzweigungsprozesse [53]. via [1120, 1219, 1320, 1441, 1710, 1941, 1841, 725, 2325, 1801, 2086, 1991, 1557, 1758, 1976, 2184, 2116, 2279, 1492, 1721, 932, 2359]. vielen [537]. view [1603, 939, 299]. virus [65, 25]. Virusinaktivierung [25, 65]. visualization [1034, 1033]. vitality [2165]. volatility [2261, 1474, 2245]. Volkseinkommens [120]. Volksvermögens [17]. völliger [202]. vom [20]. von-Mises [1009]. Voraussetzung [106]. vorgegeben [30, 37]. vorgelegten [191]. Vorkenntnissen [423]. Vorkommen [116]. Vorliegen [556]. Vorschlag [437]. vs [1063].

W [1370, 1371, 1477]. W. [1384]. wachsendem [139]. wage [327]. Wahl [73, 416]. Wahlergebnissen [424]. während [327]. Wahrscheinlichkeit [559]. Wahrscheinlichkeits [370]. Wahrscheinlichkeits- [370]. Wahrscheinlichkeitslernen [347]. Wahrscheinlichkeitsnetz [55]. waiting [441, 1946, 1864, 72, 175, 349]. Wald [2040, 1060]. Wald-type [2040]. walks [826]. Wallace [1389]. Walter [376]. warehouse [356]. Warner [1499, 1277]. warranted [1334]. Warteschlange [342]. Warteschlangenproblem [207]. Wartungsstrategien [474]. Watson [954]. wavelet [1842, 1462, 1425, 1925]. way [1622, 891, 853, 1713, 1465, 1092, 1216, 1178, 2182, 728, 1752]. ways [1499]. Weak [2229, 1531, 198, 397, 2005, 1774, 1441, 1989, 1687, 2351, 1679]. weaken [1622]. weakly [1462]. wealth [17]. weapon [285]. web [1036]. web-based [1036]. Wegmüller [376]. Weibull [1200, 1633, 1307, 1846, 840, 1960, 2130, 886, 890, 1394, 1754, 591, 1144, 1296]. weighing [1770, 1068]. Weight [1418, 1562]. Weighted [1434, 1399, 2160, 1512, 1716, 2337, 1048, 1230, 2345, 1788, 2158, 1779, 1552, 2093, 2087, 2144, 1268, 2321, 1744, 1979, 2053, 2311]. Weighting [1407, 1839]. weights [1470, 1812, 1125, 1682]. Weinman [937]. Welch [182]. welfare [2140]. well [452]. wenn [505]. Werner [1676]. Wettbegriff [370]. whether [746, 1273, 257]. Which [1340]. whose [30, 37]. Widmung [151]. width [738, 2315, 5]. Wigner [1529]. Wilcoxon [2056, 368]. Wilhelm [230, 271]. William [271]. Williams [1136]. Winding [1708]. Winkler [230, 271]. winner [1199].

Winsorized [1900]. wire [38]. Wirken [272]. wirksame [10]. Wirksamkeit [101]. Wirkungen [327]. wirtschaftliche [327]. wirtschaftlichen [38]. wirtschaftlicher [149]. wirtschaftlichste [173, 228]. Wirtschaftsforschung [626]. Wirtschaftspolitik [93]. Wirtschaftsprozessen [47]. wirtschaftsstatistischer [119]. Wirtschaftswissenschaften [153]. Wishart [950, 1505, 661]. Wissenschaften [219]. without [440, 2015, 2154, 682, 1759]. Wordtype [1368]. working [2199, 1809]. workings [2070]. workshop [142, 2044]. Wright [1371]. Wynn [2189].

X [533, 2352]. XII [1437]. XVIII [1370]. XVIII/486 [1370]. XX [639].

Yates [2258]. Yates-order [2258]. year [1333]. Years [231]. York [1370, 1371]. Young [196, 196].

Zahl [173]. Zahlen [584]. Zahlenfolgen [279]. Zählungen [402]. zeichnerische [13, 59]. Zeit [342]. zeitgenössische [105]. Zeitparameter [353]. Zeitraumes [327]. Zeitreihen [314, 117, 410, 164, 80, 337]. Zeitreihenanalyse [392]. Zenga [1742]. Zentrale [77]. Zerlegung [14, 228]. Zero [1610, 632, 2112, 1950]. zero-inflated [1950]. Zhao [1371]. zirkulären [148]. zufällige [316, 324]. zufälligen [148]. zufälliger [355]. Zufallskomponenten [228]. Zufallsvariablen [244, 597]. Zufallsvariabler [501]. Zufallszahlen [146]. zugrundeliegende [28]. zum [333, 94, 371, 79, 410]. Zur [119, 338, 355, 342, 275, 437, 110, 526, 2, 485, 52, 115, 354, 154, 505, 392, 626, 53, 191, 67, 405, 314, 334, 120, 59, 10, 78, 65, 337, 123, 90, 106, 297, 422, 388, 48, 554, 73, 98, 140, 604, 240, 462, 270, 316, 357, 416]. Zusammenhänge [113]. zusätzlichen [199]. zwei [353, 161, 168, 537, 118, 173]. zwei- [161, 168]. zweier [485, 85, 399]. zweifach [340]. zweifache [458, 228]. zweiseitigen [580]. Zweistichprobenhypothese [10]. zweiter [173].

References

Anderson:1958:GGF

[1] Prof. Oskar Anderson, Prof. Wilhelm Winkler, Prof. Hans Kellerer, and Prof. Slawtscho Sagoroff. Geleitwort. (German) [Foreword]. Metrika, 1(1):1– 2, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02613388.

Hetz:1958:UFV

[2] Dipl.-Math. Dr. W. Hetz and Dipl.- Math. Dr. H. Klinger. Untersuchungen zur Frage der Verteilung von Objekten auf Plätze. (German) [Studies on the question of the distribution of objects in places]. Metrika, 1(1):3–20, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613389.

Quenouille:1958:DAS

[3] Prof. M. H. Quenouille. Discrete autoregressive schemes with varying timeintervals. Metrika, 1(1):21–27, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613390.

Dano:1958:PIP

[4] S. Danö and E. L. Jensen. Production and inventory planning in a fluctuating market. Metrika, 1(1):28–40, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613391.

Schaffer:1958:AVS

[5] Dr. K.-A. Schäffer and Dipl. Math. M. Nourney. Abschätzung der Varianz stetiger Merkmale aus Intervallbreite und Mittelwert. (German) [Estimate of the variance of continuous parameters of interval width and mean]. Metrika, 1(1):41–56, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613392.

Anonymous:1958:MMS

[6] Anonymous. Mitteilungsblatt für mathematische Statistik. (German) [Bulletin of mathematical statistics]. Metrika, 1(1):57–68, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF02613393.

Geppert:1958:BCB

[7] M. P. Geppert. Das 5. Biometrische Colloquium in Bad Nauheim (24.–26. Januar 1958). (German) [The 5th biometrical colloquium in Bad Nauheim (24–26 Januar 1958)]. Metrika, 1 (1):69–71, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613394.

Geppert:1958:GJN

[8] M. P. Geppert. Gastvorlesung von J. Neyman in Frankfurt/M. und Bad Nauheim. (German) [Guest lecture by J. Neyman in Frankfurt am Main and Bad Nauheim]. Metrika, 1(1):72–73, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02613395.

Munzner:1958:BGB

[9] H. Münzner, S. Koller, M. Nourney, and K.-A. Schäffer. Buchbesprechungen. (German) [Book reviews]. Metrika, 1(1):73–80, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613396.

Walter:1958:EEN

[10] E. Walter. Einige einfache nichtparametrische überall wirksame Tests zur Prüfung der Zweistichprobenhypothese mit paarigen Beobachtungen. (German) [Some simple nonparametric universally-effective tests for testing the two-sample hypothesis with paired observations]. Metrika, 1(1):81–88, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613397.

Nasr:1958:LLN

[11] Dr. S. K. Nasr. A Law of Large Numbers for abstract random variables. Metrika, 1(1):89–98, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613398.

Adam:1958:ESG

[12] Dr. A. Adam. Entropie und Streuung. (German) [Entropy and scattering]. Metrika, 1(1):99–110, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613399.

Stange:1958:ZBP

[13] Prof. Dr. K. Stange. Die zeichnerische Behandlung von Plänen für messende Prüfung. (German) [The graphical treatment plans for measuring tests]. Metrika, 1(1):111–129, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF02613400. See

remark [59].

Pfanzagl:1958:ZSR

[14] Dozent Dr. J. Pfanzagl. Über die Zerlegung statistischer Reihen. (German) [On the decomposition of statistical series]. Metrika, 1(1):130–147, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613401.

Moos:1958:BMK

[15] Dipl.-Ing. H. Moos. Berechnung mehrfacher Korrelationen auf elektronischen Rechenmaschinen. (German) [Calculation of multiple correlations on electronic computers]. Metrika, 1 (1):148–153, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613402.

Tintner:1958:NMS

[16] Prof. Dr. G. Tintner. Eine neue Methode für die Schätzung der logistischen Funktion. (German) [A new method for the estimation of the logistic function]. Metrika, 1(1):154–157, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613403.

Vinski:1958:PBV

[17] Dr. I. Vinski. Probleme der Berechnung des Volksvermögens. (German) [Problems of calculation of national wealth]. Metrika, 1(1):158–170, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613404.

Schmetterer:1958:BSM

[18] L. Schmetterer. Bericht über die Spezialtagung über Mathematische Statistik im März 1958 in Oberwolfach. (German) [Report on the special session on mathematical statistics in March 1958 in Oberwolfach]. Metrika, 1(1):173–174, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613405.

Strecker:1958:BGB

[19] H. Strecker. Buchbesprechungen. (German) [Book reviews]. Metrika, 1 (1):175–176, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613406.

Stange:1958:PB

[20] Prof. Dr. K. Stange. Probenahme vom band. Metrika, 1(1):177–222, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613407.

Anonymous:1958:HC

[21] Anonymous. Help & contacts. Metrika, 1(1):??, ???? 1958. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Fisher:1959:MPN

[22] Prof. Sir Ronald A. Fisher. Mathematical probability in the natural sciences. Metrika, 2(1):1–10, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613720.

Pfanzagl:1959:KTK

[23] Dozent Dr. Johann Pfanzagl. Ein kombiniertes test & klassifikationsproblem. Metrika, 2(1):11–45, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613721.

Jochems:1959:MEB

[24] D. B. Jochems and H. Neudecker. Micro-economic business test data compared with traditional statistics. Metrika, 2(1):46–61, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF02613722.

Stur:1959:VDF

[25] Dr. med. Dietrich Stur. Die Virusinaktivierung durch Formaldehyd als Trefferereignis. (German) [Virus inactivation by formaldehyde as a hit event]. Metrika, 2(1):62–77, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613723.

Anonymous:1959:SRM

[26] Anonymous. Statistical research monographs. Metrika, 2(1):78, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02613724.

Wolff:1959:BGB

[27] Wolff, Franz Ferschl, H. Kregeloh, and P. Korndörfer. Buchbesprechungen. (German) [Book reviews]. Metrika, 2(1):79–88, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613725.

Rybarz:1959:EBV

[28] Prof. Dr. J. Rybarz. Ein einfacher Beweis für das dem χ^{2}-Verfahren zugrundeliegende Theorem. (German) [A simple proof of the chi^{2} process basic theorem]. Metrika, 2(1):89–93, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613726.

Kulldorff:1959:PML

[29] Gunnar Kulldorff. A problem of maximum likelihood estimation from a grouped sample. Metrika, 2(1):94–99, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613727.

Weichselberger:1959:PKDa

[30] Dr. Kurt Weichselberger. Über die Parameterschätzung bei Kontingenztafeln, deren Randsummen vorgegeben sind (I). (German) [On the parameter estimation in contingency tables whose marginal totals are fixed (I)]. Metrika, 2(1):100–130, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613728.

Wette:1959:RKB

[31] Dr. Reimut Wette. Regressions- und Kausalanalyse in der Biologie. (German) [Regression and causal analysis in biology]. Metrika, 2(1):131–137, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613729.

Linhart:1959:TDA

[32] H. Linhart. Techniques for discriminant analysis with discrete variables. Metrika, 2(1):138–149, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613730.

Pawlik:1959:MKIa

[33] Kurt Pawlik. Der maximale Kontingenzkoeffizient im Falle nichtquadratischer Kontingenztafeln. (German) [The maximum contingency in the event of nonsquare contingency tables]. Metrika, 2(1):150–166, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613731.

Adam:1959:BGB

[34] A. Adam. Buchbesprechungen. (German) [Book reviews]. Metrika, 2 (1):167–168, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613732.

Uhlmann:1959:ENT

[35] Dr. Werner Uhlmann. Zu einem nichtparametrischen Test von E. L. Lehmann. (German) [On a nonparametric test of E. L. Lehmann]. Metrika, 2(1):169–185, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF02613733.

Haight:1959:QR

[36] Frank A. Haight. Queueing with reneging. Metrika, 2(1):186–197, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613734.

Weichselberger:1959:PKDb

[37] Dr. Kurt Weichselberger. Über die Parameterschätzung bei Kontingenztafeln, deren Randsummen vorgegeben sind (II). (German) [On the parameter estimation in contingency tables whose marginal totals are fixed (II)]. Metrika, 2(1):198–229, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613735.

Heumann:1959:BDW

[38] H. Heumann and E. Nobis. Bestimmung des durchschnittlichen und des wirtschaftlichen Materialverbrauches von Kunststoff bei Kabel- und Leitungsisolierungen. (German) [Determining the average and the economic material consumption of plastics in cable and wire insulation]. Metrika, 2(1):230–238, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613736.

Adam:1959:VVG

[39] Dr. Adolf Adam. Verteilungsmaße und Verteilungsindizes. (German) [Distribution measurements and distribution indices]. Metrika, 2(1):239–244, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613737.

Fels:1959:BGB

[40] E. Fels, E. Walter, O. Schürz, H. Kregeloh, A. Klingst, and A. Adam. Buchbesprechungen. (German) [Book reviews]. Metrika, 2(1):245–255, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613738.

Pawlik:1959:MKIb

[41] K. Pawlik. Der maximale kontingenzkoeffizient im falle nichtquadratischer kontingenztafeln. Metrika, 2(1):256, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02613739.

Anonymous:1959:HC

[42] Anonymous. Help & contacts. Metrika, 2(1):??, ???? 1959. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Pfanzagl:1960:TKE

[43] Professor Dr. Johann Pfanzagl. Tests und Konfidenzintervalle für exponentielle Verteilungen und deren Anwendung auf einige diskrete Verteilungen. (German) []. Metrika, 3(1):1–25, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613435.

Abt:1960:ACAa

[44] Dr. Klaus Abt. Analyse de covariance et analyse par différences. (French) [ANCOVA analysis and differences]. Metrika, 3(1):26–45, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613436.

Nasr:1960:NNP

[45] Dr. Saad K. Nasr. A note on a noise process. Metrika, 3(1):46–52, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613437.

Zarkovic:1960:ESV

[46] S. S. Zarkovic. On the efficiency of sampling with varying probabilities and the selection of units with replacement. Metrika, 3(1):53–60, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613438.

Henn:1960:MKW

[47] Prof. Dr. Rudolf Henn. Markowsche Ketten bei Wirtschaftsprozessen. (German) []. Metrika, 3(1):61–73, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613439.

Ihm:1960:NID

[48] Dr. Peter Ihm. Zur numerischen Integration n-dimensionaler Normalverteilungen. (German) []. Metrika, 3(1):74–78, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613440.

Stange:1960:BGB

[49] K. Stange, S. Sagoroff, Dr. Franz Ferschl, and Dr. Schwarz. Buchbesprechungen. (German) [Book reviews]. Metrika, 3(1):79–88, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613441.

Sagoroff:1960:NOA

[50] S. Sagoroff. Nachruf für Oskar Anderson. (German) [Obituary for Oskar Anderson]. Metrika, 3(1):89–94, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613442.

Abt:1960:ACAb

[51] Dr. Klaus Abt. Analyse de covariance et analyse par différences. (French) [ANCOVA analysis and differences]. Metrika, 3(1):95–116, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613443.

Lienert:1960:INB

[52] Dr. G. A. Lienert and O. Ebel. Ein Index zur numerischen Bestimmung der Niveau–Eigenschaft eines psychologischen Tests. (German) []. Metrika, 3(1):117–127, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613444.

Schneider:1960:BTV

[53] B. Schneider. Beitrag zur Theorie der Verzweigungsprozesse. (German) []. Metrika, 3(1):128–142, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613445.

Pfanzagl:1960:LOR

[54] Professor Dr. J. Pfanzagl. Über lokal optimale rang–tests. Metrika, 3(1):143–150, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613446.

Stange:1960:SMP

[55] Prof. Dr. K. Stange. Spannweitenpläne für messende Prüfung im doppelten Wahrscheinlichkeitsnetz. (German) [Span of plans for measuring tests in duplicate probability networks]. Metrika, 3(1):151–165, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613447.

Mittenecker:1960:BGB

[56] Erich Mittenecker and J. Pfanzagl. Buchbesprechungen. (German) [Book reviews]. Metrika, 3(1):166–168, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613448.

Pfanzagl:1960:EUT

[57] Prof. Dr. J. Pfanzagl. Über die Existenz überall trennscharfer Tests. (German) []. Metrika, 3(1):169–176, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613449.

Abt:1960:ACAc

[58] Dr. Klaus Abt. Analyse de covariance et analyse par différences. (French) [ANCOVA analysis and differences]. Metrika, 3(1):177–211, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613450.

Vas:1960:BAK

[59] Éva Vas. Bemerkungen zur Arbeit von K. Stange über “Die zeichnerische Behandlung von Plänen für messende Prüfung”. (German) []. Metrika, 3 (1):212–214, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613451. See [13].

Eberl:1960:BTA

[60] Dr. Walther Eberl. Bemerkung über ein Theorem für additive Mengenfunktionen. (German) []. Metrika, 3 (1):215–217, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613452.

Ludwig:1960:EVR

[61] Dr. Otto Ludwig. Über Erwartungswerte und Varianzen von Ranggrößen in kleinen Stichproben. (German) []. Metrika, 3(1):218–222, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613453.

Brandau:1960:BGB

[62] Eberhard Brandau, P. Korndörfer, Erich Mittenecker, and G. Reißig. Buchbesprechungen. (German) [Book reviews]. Metrika, 3(1):234–240, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613454.

Anonymous:1960:HC

[63] Anonymous. Help & contacts. Metrika, 3(1):??, ???? 1960. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Stange:1961:VUS

[64] K. Stange. Verzerrte und unverzerrte Schätzwerte bei messender Prüfung. (German) []. Metrika, 4(1):1–29, ???? 1961. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613864.

Wette:1961:BTA

[65] Dr. R. Wette. Bemerkung zur treffertheoretischen Analyse der Virusinaktivierung durch Formaldehyd. (German) [Remarks on the hit-theoretical analysis of virus inactivation by formaldehyde]. Metrika, 4(1):30–36, ???? 1961. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613865.

Lomnicki:1961:TDN

[66] Z. A. Lomnicki. Tests for departure from normality in the case of linear stochastic processes. Metrika, 4(1):37–62, ???? 1961. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613866.

Stormer:1961:STB

[67] H. Störmer. Über einen statistischen Test zur Bestimmung der Parameter einer Lebensdauerverteilung. (German) []. Metrika, 4(1):63–77, ???? 1961. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613867.

Sagoroff:1961:HNA

[68] Prof. Dr. S. Sagoroff. Historical note on the application of mathematics in economics in Germany. Metrika, 4(1):78–81, ???? 1961. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613868.

Weichselberger:1961:BGB

[69] K. Weichselberger, A. Adam, F. Ferschl, S. Koller, and A. Linder. Buchbesprechungen. (German) [Book reviews]. Metrika, 4(1):82–88, ???? 1961. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613869.

Bennett:1961:SSB

[70] B. M. Bennett and P. Hsu. Sampling studies on the Behrens–Fisher problem. Metrika, 4(1):89–104, ???? 1961. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613870.

Pfanzagl:1961:EUT

[71] Prof. Dr. J. Pfanzagl. Über die Existenz überall trennscharfer Tests. (German) [On the universal existence of discriminatory power tests]. Metrika, 4(1):105–106, ???? 1961. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613871.

Jaiswal:1961:QSG

[72] N. K. Jaiswal. The queuing system GI/M/I with finite waiting space.

Metrika, 4(1):107–125, ???? 1961. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF02613872.

Linhart:1961:WVT

[73] H. Linhart. Zur Wahl von Variablen in der Trennanalyse. (German) []. Metrika, 4(1):126–139, ???? 1961. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613873.

Vogel:1961:SVP

[74] W. Vogel. Sequentielle Versuchs– Pläne. Metrika, 4(1):140–157, ???? 1961. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613874.

Bierlein:1961:SSU

[75] Dr. D. Bierlein. Simultan stärkste unverfälschte tests. Metrika, 4(1): 158–168, ???? 1961. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613875.

Reissig:1961:BGB

[76] G. Reißig and O. Ludwig. Buchbesprechungen. (German) [Book reviews]. Metrika, 4(1):169–172, ???? 1961. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613876.

Morgenstern:1961:ZGU

[77] Prof. Dr. D. Morgenstern. Der Zentrale Grenzwertsatz für Überschreitungswahrscheinlichkeiten.

(German) []. Metrika, 4(1):173–175, ???? 1961. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613877.

Weichselberger:1961:GVT

[78] Dr. K. Weichselberger. Über ein graphisches Verfahren zur Trennung von Mischverteilungen und zur Identifikation kupierter Normalverteilungen bei großem Stichprobenumfang. (German) []. Metrika, 4(1):178–229, ???? 1961. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613878.

Michel:1961:PNB

[79] Dr. L. Michel. Zum Problem der numerischen Bestimmung der Niveau–Eigenschaft eines psychologischen Tests. (German) []. Metrika, 4 (1):230–236, ???? 1961. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613879.

Wartmann:1961:TZG

[80] Dr. R. Wartmann. Trendextrapolation bei Zeitreihen. (German) []. Metrika, 4(1):237–246, ???? 1961. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613880.

Roppert:1961:BGB

[81] J. Roppert, S. Sagoroff, G. Bruckmann, F. Ferschl, and G. Fasching. Buchbesprechungen. (German) [Book reviews]. Metrika, 4(1):247–252, ????

1961. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613881.

Anonymous:1961:HC

[82] Anonymous. Help & contacts. Metrika, 4(1):??, ???? 1961. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Bauer:1962:MQT

[83] Dipl.-Math. R. K. Bauer. Der “Median-Quartile-Test”. (German) []. Metrika, 5(1):1–16, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02616180.

Hoch:1962:BLM

[84] Dr. J. Hoch. Berechnung von linearen Mehrfachkorrelationen mit Hilfe von Adjunktionsmatrizen nach Frisch. (German) []. Metrika, 5(1):17–30, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02616181.

Linhart:1962:KUT

[85] H. Linhart. Konfidenzintervalle für den überlappenden Teil zweier Normalverteilungen. (German) []. Metrika, 5(1):31–48, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02616182.

Bennett:1962:ETT

[86] Professor B. M. Bennett. On an exact test for trend in binomial trials and its power function. Metrika, 5(1):49–53, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02616183.

Kellerer:1962:BGBa

[87] H. Kellerer, H. Schneeweiß, O. Ludwig, R. E. Quandt, F. Ferschl, J. Pfanzagl, G. Schmid, W. Winkler, G. Reißig, and H. Kremser. Buchbesprechungen. (German) [Book reviews]. Metrika, 5(1):54–79, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02616184.

Linder:1962:E

[88] Professor Dr. Arthur Linder. Ehrung. Metrika, 5(1):79, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02616185.

Anonymous:1962:B

[89] Anonymous. Berichtigung. Metrika, 5 (1):80, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF02616186.

Ahrens:1962:GSP

[90] J. Ahrens. Zur Gesamtheit der symmetrischen n-Personen-Spiele. (German) []. Metrika, 5(1):81–95, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02616187.

Bennett:1962:SST

[91] Prof. B. M. Bennett and Prof. P. Hsu. Sampling studies on a test against trend in binomial data. Metrika, 5 (1):96–104, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02616188.

Canfield:1962:PC

[92] J. Canfield. On the paradox of confirmation. Metrika, 5(1):105–118, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02616189.

Varga:1962:AMM

[93] Prof. Dr. S. Varga. Die Anwendung von mathematischen Methoden in der marxistischen politischen Ökonomie und in der Wirtschaftspolitik des Sozialismus. (German) []. Metrika, 5(1):119–127, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02616190.

Stormer:1962:TEE

[94] Dr. H. Störmer. Ein Test zum Erkennen von Exponentialverteilungen. (German) []. Metrika, 5(1): 128–137, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02616191.

Kellerer:1962:BGBb

[95] H. G. Kellerer and J. Pfanzagl. Buchbesprechungen. (German) [Book reviews]. Metrika, 5(1):138–140, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02616192.

Linder:1962:NSR

[96] A. Linder. Nachruf für Sir Ronald Fisher. (German) [Obituary for Sir Ronald Fisher]. Metrika, 5(1):141–144, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02616193.

Koop:1962:ULD

[97] J. C. Koop. On upper limits to the difference in bias between two ratio estimates. Metrika, 5(1):145–149, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02616194.

Morgenstern:1962:ART

[98] Professor Dr. D. Morgenstern. Zur Asymptotik des Run-Testes. (German) []. Metrika, 5(1):150–153, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02616195.

Bennett:1962:NET

[99] Professor B. M. Bennett. Note on an exact test for the 2 × 2 contingency table using the negative binomial model. Metrika, 5(1):154–157, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02616196.

Storm:1962:NAV

[100] L. E. Storm. Nested analysis of variance: Review of methods. Metrika, 5(1):158–183, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02616197.

Likes:1962:WES

[101] J. Likes. Wirksamkeit einiger Schätzungen der Varianz einer Grundgesamtheit mit einem alternativen Merkmale. (German) []. Metrika, 5(1): 184–193, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02616198.

Jochems:1962:FOS

[102] Dr. Senior Research Officer D. B. Jochems. Forecasting the outcomes of soccer matches: A statistical appraisal of the Dutch experience. Metrika, 5 (1):194–207, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02616199.

Anonymous:1962:HC

[103] Anonymous. Help & contacts. Metrika, 5(1):??, ???? 1962. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Sagoroff:1963:GGF

[104] S. Sagoroff. Geleitwort. (German) [Foreword]. Metrika, 6(1):i–ii, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.

com/accesspage/article/10.1007/
BF02613348.

Adam:1963:JHU

[105] Prof. Dr. A. Adam. Joseph Hain und unsere zeitgenössische Statistik. (German) []. Metrika, 6(1):1–9, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613349.

Anderson:1963:VAS

[106] Prof. Dr. O. Anderson and Dipl.- Math. R. K. Bauer. Zur verteilungsfreien Ableitung der Streuung von Regressionskoeffizienten ohne Voraussetzung der Nullhypothese. (German) []. Metrika, 6(1):10–17, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613350.

Blind:1963:PBF

[107] Prof. Dr. A. Blind. Die praktische Bedeutung der Fehlerrechnung bei Stichproben in der sozialwissenschaftlichen Statistik. (German) []. Metrika, 6 (1):18–29, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613351.

Dalenius:1963:PAC

[108] T. Dalenius. The panel approach to continuous sample survey activity. Metrika, 6(1):30–36, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF02613352.

Deming:1963:CMB

[109] W. E. Deming, LL.D., Ph.D. On the correction of mathematical bias by use of replicated designs. Metrika, 6(1):37–42, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613353.

Fels:1963:SBA

[110] Prof. Dr. E. Fels. Statistische Bewertung einer automatischen Anlage zur Suche nach einschlägigen Rechtsvorschriften. (German) []. Metrika, 6(1):43–49, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613354.

Gunzert:1963:SRG

[111] Prof. Dr. R. Gunzert. Stichprobenverfahren und Regionalstatistik. (German) []. Metrika, 6(1):50–59, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613355.

Kendall:1963:ICM

[112] Prof. Dr. M. G. Kendall. Illconditioned matrices in linear programming. Metrika, 6(1):60–64, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613356.

Koller:1963:TKZ

[113] Prof. Dr. Dr. S. Koller. Typisierung korrelativer Zusammenhänge. Metrika, 6(1):65–75, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613357.

Linder:1963:TQM

[114] Prof. Dr. A. Linder. Trennverfahren bei qualitativen Merkmalen. (German) []. Metrika, 6(1):76–83, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613358.

Menges:1963:GFS

[115] Prof. Dr. G. Menges. Gedanken zur Frage der Stabilität statistischer Entscheidungen. (German) []. Metrika, 6(1):84–94, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613359.

Munzner:1963:KAA

[116] Prof. Dr. H. Münzner. Konfidenzgrenzen für die Ausschußhäufigkeit bei abschnittsweisen Probeentnahmen und serienweisem Vorkommen der Schlechtstücke. (German) []. Metrika, 6(1):95–99, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613360.

Pfanzagl:1963:PZG

[117] Prof. Dr. J. Pfanzagl. Über die Parallelität von Zeitreihen. (German) []. Metrika, 6(1):100–113, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613361.

Stange:1963:VGZ

[118] Prof. Dr. K. Stange. Vergleich der Genauigkeit von zwei Verfahren der Probenahme bei Verkehrszählungen. (German) []. Metrika, 6(1):114–132, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613362.

Strecker:1963:GWG

[119] Prof. Dr. H. Strecker. Zur genauigkeit wirtschaftsstatistischer Größen. Metrika, 6(1):133–142, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613363.

Tintner:1963:VGT

[120] G. Tintner and J. K. Sengupta. Ein verallgemeinerter Geburten- und Todesprozeß zur Erklärung der Entwicklung des deutschen Volkseinkommens, 1851–1939. (German) []. Metrika, 6 (1):143–147, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613364.

Wagenfuhr:1963:PIG

[121] Prof. Dr. R. Wagenführ. Preisparitäten für Investitionsgüter. (German) [Price parities for capital goods]. Metrika, 6(1):148–154, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613365.

Wetzel:1963:UMO

[122] Prof. Dr. W. Wetzel. Univariable mikro-ökonomische Prognosemodelle mit exponentieller Gewichtung der Beobachtungswerte. (German) []. Metrika, 6(1):155–186, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613366.

Winkler:1963:DDG

[123] Hofrat Prof. Dr. Dr. h. c. W. Winkler. Von der Demographie zur Demometrie. (German) []. Metrika, 6(1):187–198, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613367.

Anonymous:1963:HCa

[124] Anonymous. Help & contacts. Metrika, 6(1):??, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Fels:1963:APL

[125] E. M. Fels. About probability-like measures for entire theories. Metrika, 7(1):1–22, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613958.

Uhlmann:1963:RSG

[126] W. Uhlmann. Ranggrößen als Schätzfunktionen. (German) []. Metrika, 7(1):23–40, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613959.

Bennett:1963:TCL

[127] B. M. Bennett. On tests concerning lines and planes in hyperspace. Metrika, 7(1):41–46, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613960.

Korenjak:1963:VMEa

[128] Dipl.-Ing. F. Korenjak. Vertical models for economic development and cycles. Metrika, 7(1):47–70, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613961.

Sagoroff:1963:HNA

[129] S. Sagoroff. Historical note on the application of mathematics in economics in Germany. Metrika, 7(1):71–75, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613962.

Anonymous:1963:PAI

[130] Anonymous. Preliminary announcement of the IMS Copenhagen meeting. Metrika, 7(1):76, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02613963.

Ferschl:1963:BGB

[131] F. Ferschl, W. Oehmisch, M. Behara, and O. Heinisch. Buchbesprechungen. (German) [Book reviews]. Metrika, 7(1):76–80, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613964.

Koop:1963:ASFa

[132] Dr. J. C. Koop. On the axioms of sample formation and their bearing on the construction of linear estimators in sampling theory for finite universes. Metrika, 7(1):81–114, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613965.

Noether:1963:NKS

[133] G. E. Noether. Note on the Kolmogorov statistic in the discrete case. Metrika, 7(1):115–116, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02613966.

Bennett:1963:SRA

[134] B. M. Bennett. Slope ratio assays and confidence limits. Metrika, 7 (1):117–120, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613967.

Korenjak:1963:VMEb

[135] Dipl.-Ing. F. Korenjak. Vertical models for economic development and cycles. Metrika, 7(1):121–142, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613968.

Hampe:1963:JLS

[136] Prof. Dr. A. Hampe. Joseph Lowe als statistiker. Metrika, 7(1):143–158, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613969.

Heinisch:1963:BGB

[137] O. Heinisch, E. P. Billeter, Saad K. Nasr, K. Krickeberg, Professor Dr. W. Wetzel, H. S. Hoffmann, G. Reissig, and Josef Roppert. Buchbesprechungen. (German) [Book reviews]. Metrika, 7(1):159–164, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613970.

Koop:1963:ASFb

[138] J. C. Koop. On the axioms of sample formation and their bearing on the construction of linear estimators in sampling theory for finite universes. Metrika, 7(1):165–204, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613971.

Kellerer:1963:AVM

[139] H. G. Kellerer. Das asymptotische Verhalten des Minimax–Risikos bei wachsendem Stichprobenumfang. (German) []. Metrika, 7(1):205–212, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613972.

Rabar:1963:FPA

[140] F. Rabár and M. Krajcsovits. Zur Frage der Produktionsplanung mit der Anwendung mathematischer Methoden in der Serienfertigung. (German) []. Metrika, 7(1):213–230, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613973.

Chakravarti:1963:OPB

[141] I. M. Chakravarti. Orthogonal and partially balanced arrays and their application in design of experiments. Metrika, 7(1):231–243, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF02613974.

Anonymous:1963:ITB

[142] Anonymous. IMS-Tagung 1964 in Bern. (German) [IMS workshop 1964 in Bern]. Metrika, 7(1):243, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02613975.

Wegmuller:1963:BGB

[143] W. Wegmüller and W. Oehmisch. Buchbesprechungen. (German) [Book reviews]. Metrika, 7(1):243–244, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02613976.

Anonymous:1963:HCb

[144] Anonymous. Help & contacts. Metrika, 7(1):??, ???? 1963. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Bennett:1964:THS

[145] B. M. Bennett. On a test of homogeneity for samples from a negative binomial distribution. Metrika, 8(1):1–4, ???? 1964. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613705.

Johnk:1964:EBG

[146] Dipl.-Volkswirt M. D. Jöhnk. Erzeugung von betaverteilten und gammaverteilten Zufallszahlen. (German) [Production of beta- and gammadistributed random numbers]. Metrika, 8(1):5–15, ???? 1964. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613706.

Andrews:1964:LMP

[147] F. C. Andrews and D. R. Truax. Locally most powerful rank tests for several sample problems. Metrika, 8(1):16–24, ???? 1964. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613707.

Uhlmann:1964:MIS

[148] Dozent Dr. W. Uhlmann. Mittelwerte und ihre Schätzfunktionen bei zirkulären zufälligen Variablen. (German) [Mean values and their estimators in circular random variables]. Metrika, 8(1):25–47, ???? 1964. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613708.

Stange:1964:BWP

[149] Prof. Dr. K. Stange. Die Berechnung wirtschaftlicher Pläne für messende Prüfung. (German) []. Metrika, 8 (1):48–82, ???? 1964. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613709.

Anonymous:1964:GEK

[150] Anonymous. Gemeinsame Europäische Konferenz von Econometric Society und The Institute of Management Sciences. (German) []. Metrika, 8(1):83– 84, ???? 1964. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF02613710.

Sagoroff:1964:WGD

[151] S. Sagoroff. Widmung. (German) [Dedication]. Metrika, 8(1):85, ???? 1964. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02613711.

Adam:1964:SVG

[152] Prof. Dr. A. Adam. Semiotische und verbandstheoretische Grundzüge einer elementaren Statistik. (German) []. Metrika, 8(1):87–128, ???? 1964. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613712.

Pfanzagl:1964:AMM

[153] Prof. Dr. J. Pfanzagl. Über die Anwendung mathematischer Methoden in den Wirtschaftswissenschaften. (German) []. Metrika, 8(1):129–151, ???? 1964. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613713.

Roppert:1964:SKI

[154] Dr. J. Roppert. Ein schnell konvergierendes Iterationsverfahren zur Matrixinversion. (German) [A rapidly converging iterative method for matrix inversion]. Metrika, 8(1):152–154, ???? 1964. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613714.

Streissler:1964:EBA

[155] E. Streissler and P. Hoschka. Entrepreneurial behaviour in Austria. Metrika, 8(1):155–184, ???? 1964. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF02613715.

Weichselberger:1964:TGD

[156] Prof. Dr. K. Weichselberger. Über eine Theorie der gleitenden Durchschnitte und verschiedene Anwendungen dieser Theorie. (German) []. Metrika, 8 (1):185–230, ???? 1964. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613716.

Kabe:1964:TCC

[157] D. G. Kabe. On tests concerning coefficient matrices of linear restriction with normal regression models. Metrika, 8(1):231–234, ???? 1964. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613717.

Dalenius:1964:TNU

[158] Docent T. Dalenius and Professor B. Matérn. Is there a need for a unified theory of random experiments? Metrika, 8(1):235–247, ???? 1964. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613718.

Koller:1964:BGB

[159] S. Koller, Professor W. Wegmüller, and A. Panzera. Buchbesprechungen. (German) [Book reviews]. Metrika, 8(1):248–250, ???? 1964. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613719.

Anonymous:1964:HC

[160] Anonymous. Help & contacts. Metrika, 8(1):??, ???? 1964. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Hengartner:1965:AZDa

[161] O. Hengartner. Ausgewählte zwei- und dreidimensionale Prüfverteilungen. (German) []. Metrika, 9(1):1–37, ???? 1965. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614168.

Buhlmann:1965:RIK

[162] H. Bühlmann. Die Risikoaversion als Interpretation und Konstruktionsbasis der Utilitätskurve. (German) []. Metrika, 9(1):38–46, ???? 1965. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF02614169.

Sengupta:1965:SSU

[163] J. K. Sengupta, C. Millham, and G. Tintner. On the stability of solutions under error in stochastic linear programming. Metrika, 9(1):47–60, ???? 1965. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614170.

VonPrillwitz:1965:VPE

[164] H. Von Prillwitz. Die Varianz der Prognosen bei Exponentieller Gewichtung der Zeitreihen. (German) []. Metrika, 9(1):61–68, ???? 1965. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02614171.

Schmid:1965:BGB

[165] G. Schmid, G. Bruckmann, H. Müller-Merbach, G. Fischer, A. Adam, G. Bruckmann, E. M. Meixner, Max-Detlev Jöhnk, M. P. Geppert, and O. Pichler. Buchbesprechungen. (German) [Book reviews]. Metrika, 9 (1):69–84, ???? 1965. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02614172.

Winkler:1965:UUB

[166] Wilhelm Winkler. Übervölkerung, Untervölkerung Begriffe, Verfahren, Tatsachen. (German) []. Metrika, 9 (1):85–102, ???? 1965. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02614173.

LeRoy:1965:TVP

[167] H. L. Le Roy. t-Test und Vertrauensgrenzen für Pfadkoeffizienten (standardisierte partielle Regressionskoeffizienten). (German) []. Metrika, 9(1):103–104, ???? 1965. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02614174.

Hengartner:1965:AZDb

[168] O. Hengartner. Ausgewählte zwei- und dreidimensionale Prüfverteilungen. (German) []. Metrika, 9(1):105–148, ???? 1965. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614175.

Kurth:1965:CSB

[169] R. Kurth. Counting as a sequence of Bernoulli trials. Metrika, 9(1):149–152, ???? 1965. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614176.

Anonymous:1965:M

[170] Anonymous. Mitteilung. Metrika, 9(1): 153, ???? 1965. CODEN MTRKA8.

ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF02614177.

Tichy:1965:BGB

[171] Gunther Tichy, Dr. Gerhart Bruckmann, F. Ferschl, M. Kricke, and Hans Prillwitz. Buchbesprechungen. (German) [Book reviews]. Metrika, 9(1):154–161, ???? 1965. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02614178.

Adam:1965:SVG

[172] Prof. Dr. A. Adam. Semiotische und verbandstheoretische Grundzüge einer elementaren Statistik. (German) []. Metrika, 9(1):163–194, ???? 1965. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614179.

Stange:1965:WVP

[173] Prof. Dr. K. Stange. Die wirtschaftlichste Verteilung des Prüfaufwands auf die zwei Stufen eines Modells mit unterschiedlicher Zahl und Varianz und unterschiedlichen Erhebungskosten der Einheiten zweiter Stufe. (German) [The most cost-effective distribution of test efforts on the two stages of a model with different number and different variance and collection costs of secondstage units]. Metrika, 9(1):195–211, ???? 1965. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614180.

Ghosh:1965:OAS

[174] S. P. Ghosh. Optimum allocation in stratified sampling with replacement. Metrika, 9(1):212–221, ???? 1965. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614181.

Raade:1965:SWT

[175] L. Råde. Some waiting time problems. Metrika, 9(1):222–227, ???? 1965. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614182.

Bach:1965:BCG

[176] G. Bach. Über die Bestimmung von charakteristischen Größen einer Kugelverteilung aus der unvollständigen Verteilung der Schnittkreise. (German) []. Metrika, 9(1):228–233, ???? 1965. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614183.

Subrahmanya:1965:GES

[177] M. T. Subrahmanya. On generalized estimators in sampling from finite populations. Metrika, 9(1):234–239, ???? 1965. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614184.

Roppert:1965:BGB

[178] J. Roppert and E. M. Fels. Buchbesprechungen. (German) [Book reviews]. Metrika, 9(1):240–242, ???? 1965. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02614185.

Anonymous:1965:HC

[179] Anonymous. Help & contacts. Metrika, 9(1):??, ???? 1965. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Bennett:1966:MAD

[180] Prof. Dr. B. M. Bennett. Multivariate ‘analysis of dispersion’ in the presence of intra-class correlation. Metrika, 10(1):1–5, ???? 1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613412.

Sengupta:1966:RCS

[181] J. K. Sengupta. Recursive constraints and stochastic linear programming. Metrika, 10(1):6–12, ???? 1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613413.

Kabe:1966:EDF

[182] D. G. Kabe. On the exact distribution of the Fisher–Behrens–Welch statistic. Metrika, 10(1):13–15, ???? 1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613414.

Linhart:1966:SPV

[183] H. Linhart. Streuungszerlegung für Paar–Vergleiche. (German) []. Metrika, 10(1):16–38, ???? 1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613415.

Strobel:1966:ESS

[184] D. Strobel. Estimation of the sum of statistical variables by using classes with random limits. Metrika, 10 (1):39–45, ???? 1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613416.

Beck:1966:OAT

[185] U. P. Beck. Ein optisches Analogiegerät für den tetrachorischen korrelationskoeffizienten. Metrika, 10(1):46–51, ???? 1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613417.

Bennett:1966:NCL

[186] Prof. B. M. Bennett. Note on confidence limits for a ratio of bivariate medians. Metrika, 10(1):52–54, ???? 1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613418.

Quandt:1966:ONM

[187] Prof. Dr. R. E. Quandt. Old and new methods of estimation and the Pareto distribution. Metrika, 10(1):55–82, ???? 1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613419.

Anonymous:1966:A

[188] Anonymous. Announcement. Metrika,

10(1):83, ????  1966.  CODEN

MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613420.

Wegmuller:1966:B

[189] Prof. W. Wegmüller, Dr. W. Leinfellner, and G. Fischer. Buchbesprcchungen. Metrika, 10(1):84–88, ???? 1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613421.

Rao:1966:VRE

[190] T. J. Rao. On the variance of the ratio estimator for Midzuno–Sen sampling scheme. Metrika, 10(1):89–91, ???? 1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613422.

Stange:1966:NBO

[191] Prof. Dr. K. Stange. Ein Näherungsverfahren zur Berechnung optimaler Pläne für messende Prüfung bei bekannten Kosten und bekannter Verteilung der Schlechtanteile in den vorgelegten Liefermengen. (German) [An approximate method for computing optimal plans for measuring tests at a known cost and known distribution of bad play in the deliverables submitted]. Metrika, 10(1):92–136, ???? 1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613423.

Rustagi:1966:AQT

[192] J. S. Rustagi. An application of queueing theory to a problem of environmental health. Metrika, 10(1):137–144, ???? 1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613424.

Uhlmann:1966:VHB

[193] Prof. Dr. W. Uhlmann. Vergleich der hypergeometrischen mit der Binomial– Verteilung. (German) []. Metrika, 10 (1):145–158, ???? 1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613425.

Anonymous:1966:JEM

[194] Anonymous. Joint European meeting of the Econometric Society and the Institute of Management Sciences. Metrika, 10(1):159, ???? 1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02613426.

Tichy:1966:B

[195] Gunther Tichy. Buchbesprechung. Metrika, 10(1):159–160, ???? 1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
accesspage/article/10.1007/BF02613427.

Fischer:1966:TEY

[196] Dr. G. Fischer and Dr. J. Roppert. Über ein Theorem von Eckart und Young: “Eine Verallgemeinerung eines Transformationsverfahrens von Jacobi”. (German) [On a theorem of Eckart and Young: “A generalization of a transformation method of Jacobi”]. Metrika, 10(1):161–170, ????

1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613428.

Kanpur:1966:QBP

[197] S. K. Gupta Kanpur and J. K. Goyal Kanpur. Queues with batch Poisson arrivals and hyper-exponential service time distribution. Metrika, 10 (1):171–178, ???? 1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613429.

Krickeberg:1966:CTC

[198] K. Krickeberg. On Cramér’s theorems concerning weak convergence of distributions. Metrika, 10(1):179–181, ???? 1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613430.

Kalin:1966:VUB

[199] Prof. Dr. A. Kälin. Versuchsanordnungen in unvollständigen Blöcken mit zusätzlichen Kontrollbehandlungen in jedem Block. (German) []. Metrika, 10(1):182–218, ???? 1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613431.

Anonymous:1966:HC

[200] Anonymous. Help & contacts. Metrika, 10(1):??, ???? 1966. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Dillon:1967:HM

[201] Thaddeus M. Dillon. Hypothesis modification. Metrika, 11(1):1–14, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613571.

Kramer:1967:EEV

[202] Giselbert Kramer. Entscheidungsproblem, Entscheidungskriterien bei völliger Ungewißheit und Chernoffsches Axiomensystem. (German) []. Metrika, 11(1):15–38, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613572.

Schneeberger:1967:IRG

[203] Dr. H. Schneeberger. Identitätstest von Regressionskoeffizienten. (German) []. Metrika, 11(1):39–45, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613573.

Likes:1967:DDS

[204] Ing. J. Likes. Distribution of Dixon’s statistics in the case of an exponential population. Metrika, 11(1):46–54, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613574.

Studer:1967:PAE

[205] Herbert Studer. Prüfung der Annäherung der exakten x^{2} Verteilung durch die stetige x^{2} Verteilung. (German) []. Metrika, 11(1):55–78, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613575.

Bandelow:1967:BGB

[206] Christoph Bandelow and S. Sagoroff. Buchbesprechungen. (German) [Book reviews]. Metrika, 11(1):79–80, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02613576.

Morgenstern:1967:DVW

[207] Professor Dr. Dietrich Morgenstern. Eine diskrete Verteilung für ein Warteschlangenproblem. (German) []. Metrika, 11(1):81–84, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF02613577.

Tintner:1967:MDS

[208] G. Tintner and R. Narayanan. A multi-dimensional stochastic process for the explanation of economic development. Metrika, 11(1):85–90, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613578.

Aczel:1967:RMS

[209] Prof. Dr. J. Aczel and Prof. Dr. J. Pfanzagl. Remarks on the measurement of subjective probability and information. Metrika, 11(1):91–105, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613579.

Harders:1967:SAG

[210] H. Harders. Stochastische Approximation der größten unwirksamen Dosis. (German) [Stochastic approximation of the biggest ineffective dose]. Metrika, 11(1):106–114, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613580.

Holla:1967:PIG

[211] M. S. Holla. On a Poisson-inverse Gaussian distribution. Metrika, 11 (1):115–121, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613581.

Penkov:1967:SBP

[212] Dr. Boyan Penkov and Dr. Radu Theodorescu. Über die spieltheoretische Behandlung von Prüfplänen. (German) [On the game-theoretic treatment of test plans]. Metrika, 11(1):122–126, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613582.

Mathai:1967:GHD

[213] A. M. Mathai and Prof. Dr. R. K. Saxena. On a generalized hypergeometric distribution. Metrika, 11(1):127–132, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613583.

Bhattacharya:1967:MBF

[214] Samir Kumar Bhattacharya. A modified Bessel function model in life testing. Metrika, 11(1):133–144, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613584.

Zahel:1967:NGM

[215] G. Zahel. Notiz über die Genauigkeit von Messungen im Gelände. (German) []. Metrika, 11(1):145–149, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613585.

Feichtinger:1967:BGB

[216] Dr. Feichtinger, G. Bruckmann, W. Winkler, A. Linder, W. Nörauer, S. Sagoroff, and E. M. Fels. Buchbesprechungen. (German) [Book reviews]. Metrika, 11(1):150–156, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613586.

Goyal:1967:QHP

[217] J. K. Goyal. Queues with hyper-Poisson arrivals and bulk exponential service. Metrika, 11(1):157–167, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613587.

Kashyap:1967:FRD

[218] B. R. K. Kashyap. Further results for the double ended queue. Metrika, 11(1):168–186, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613588.

Sagoroff:1967:FCM

[219] Prof. Dr. S. Sagoroff. Formale Charakteristiken der metrischen Modelle in den empirischen Wissenschaften. (German) []. Metrika, 11(1):187–198, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613589.

Singh:1967:ESS

[220] M. P. Singh. On the efficiency of systematic sampling in ratio and product estimation. Metrika, 11(1):199–205, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613590.

Anonymous:1967:HCa

[221] Anonymous. Help & contacts. Metrika, 11(1):??, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Abt:1967:ISIa

[222] Dr. Klaus Abt. On the identification of the significant independent variables in linear models. Metrika, 12(1):1–15, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613477.

Ajgaonkar:1967:NHT

[223] S. G. Prabhu Ajgaonkar. A note on Horvitz’s and Thompson’s T_{3} class of linear estimators. Metrika, 12 (1):16–21, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613478.

Bennett:1967:ERM

[224] B. M. Bennett. On estimation of a ratio of multivariate means by nonparametric methods. Metrika, 12 (1):22–28, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613479.

Morgenstern:1967:MUG

[225] Prof. Dr. Dietrich Morgenstern. Mehrdimensionale Überschreitungswahrscheinlichkeiten. (German) [Multidimensional exceedance probabilities]. Metrika, 12(1):29–33, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613480.

Singh:1967:RPM

[226] M. P. Singh. Ratio cum product method of estimation. Metrika, 12 (1):34–42, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613481.

Subrahmanya:1967:PVE

[227] M. T. Subrahmanya. On the positivity of a variance-estimator. Metrika, 12(1):43–47, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613482.

Stange:1967:WVP

[228] Prof. Dr. K. Stange. Die wirtschaftlichste Verteilung des Prüfaufwands bei einem Modell mit Zufallskomponenten für zweifache Zerlegung. (German) []. Metrika, 12(1):48–67, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613483.

Wiezorke:1967:ARG

[229] Dipl.-Math. Bernhard Wiezorke. Auswahlverfahren in der Regressionsanalyse. (German) []. Metrika, 12(1):68–79, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613484.

Anonymous:1967:PDD

[230] Anonymous. Prof. Dr. Dr. hc. Dr. Wilhelm Winkler, Ehrenmitglied des Internationalen Statistischen Instituts. (German) []. Metrika, 12(1):80, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02613486.

Rey:1967:PDR

[231] Dr. H. Rey. Prof. Dr. rer. nat. habil. Paul Lorenz 80 Jahre. (German) [Prof. Dr. rer. nat. habil. Paul Lorenz 80 years]. Metrika, 12(1):80, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02613485.

Abt:1967:ISIb

[232] Dr. Klaus Abt. On the identification of the significant independent variables in linear models. Metrika, 12(1):81–96, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613487.

Aggarwal:1967:NQC

[233] S. P. Aggarwal. A note on quasi convex programming. Metrika, 12(1):97–105, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613488.

Aggarwal:1967:PLF

[234] S. P. Aggarwal. Parametric linear fractional functionals programming. Metrika, 12(1):106–114, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613489.

Basler:1967:BKP

[235] Dr. Herbert Basler. Bestimmung kostenoptimaler Prüfpläne mittels des Mini–Max-Prinzips. (German) []. Metrika, 12(1):115–154, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF02613490.

Kabe:1967:MVQ

[236] D. G. Kabe. Minima of vector quadratic forms with applications to statistics. Metrika, 12(1):155–160, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613491.

Likes:1967:DTO

[237] Ing. J. Likes. Difference of two ordered observations divided by standard deviation in the case of negatively exponential distribution. Metrika, 12 (1):161–172, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613492.

Rao:1967:QBR

[238] S. Subba Rao. Queuing with balking and reneging in M/G/1 systems. Metrika, 12(1):173–188, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613493.

Rytz:1967:APP

[239] C. Rytz. Ausgewählte parameterfreie Prüfverfahren im 2- und k-Stichproben-Fall. (German) []. Metrika, 12(1):189–204, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613494.

Schmitz:1967:EGB

[240] Dr. Norbert Schmitz. Zur Existenz gleichmäßig bester unverfälschter Tests bei Exponentialfamilien. (German) []. Metrika, 12(1):205–212, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613495.

Stange:1967:PMP

[241] Prof. Dr. K. Stange. Pläne für messende Prüfung bei bekannter Varianz der Fertigung und einem nach oben und unten abgegrenzten Toleranzbereich für die Merkmalwerte. (German) []. Metrika, 12(1):213–233, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613496.

Anonymous:1967:HCb

[242] Anonymous. Help & contacts. Metrika, 12(1):??, ???? 1967. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Fels:1968:DAR

[243] Prof. Dr. E. M. Fels. On denoting arbitrary realizations. Metrika, 13(1):1–2, ???? 1968. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02613370.

Kellerer:1968:MFZ

[244] Prof. Dr. H. G. Kellerer. Ein Momentenproblem für Familien von Zufallsvariablen. (German) []. Metrika, 13(1):3–9, ???? 1968. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613371.

Mathai:1968:DHG

[245] A. M. Mathai and R. K. Saxena. Distributions having the gamma properties. Metrika, 13(1):10–16, ???? 1968. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613372.

Rytz:1968:APP

[246] C. Rytz. Ausgewählte parameterfreie Prüfverfahren im 2- und k-Stichproben-Fall. (German) []. Metrika, 13(1):17–71, ???? 1968. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613373.

Swarup:1968:NLF

[247] Kanti Swarup. Note on linear fractional functionals programming. Metrika, 13(1):72–77, ???? 1968. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613374.

Heubes:1968:BGB

[248] J. Heubes, A. Linder, and P. Sint. Buchbesprechungen. (German) [Book reviews]. Metrika, 13(1):78–80, ???? 1968. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613375.

Bennett:1968:NRE

[249] Prof. Dr. B. M. Bennett. Note on the relative efficiency of a bivariate sign test. Metrika, 13(1):81–85, ???? 1968. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613376.

Kabe:1968:SAA

[250] D. G. Kabe. Some aspects of analysis of variance and covariance theory for a certain multivariate complex Gaussian distribution. Metrika, 13(1):86–97, ???? 1968. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613377.

Murthy:1968:NEV

[251] Dr. M. N. Murthy. A note on estimation of variance components in a multi-stage sampling design based on interpenetrating sub-samples. Metrika, 13(1):98–103, ???? 1968. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613378.

Natarajan:1968:LRA

[252] R. Natarajan. Long-run availability of paralleled systems. Metrika, 13 (1):104–122, ???? 1968. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613379.

Nath:1968:ICT

[253] Prem Nath. Inaccuracy and coding theory. Metrika, 13(1):123–135, ???? 1968. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613380.

Nath:1968:EII

[254] Prem Nath. Entropy, inaccuracy and information. Metrika, 13(1):136–148, ???? 1968. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613381.

Sahler:1968:SDF

[255] W. Sahler. A survey on distributionfree statistics based on distances between distribution functions. Metrika, 13(1):149–169, ???? 1968. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613382.

Sarndal:1968:SPN

[256] C. E. Särndal. Some properties of the negative hypergeometric distribution and its limit distributions. Metrika, 13(1):170–190, ???? 1968. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613383.

Sidak:1968:NDF

[257] Z. Sidák. A note on distribution-free tests whether two distributions differ only by a parameter. Metrika, 13(1):191–195, ???? 1968. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613384.

Swarup:1968:VAP

[258] K. Swarup. On varying all the parameters in a linear fractional functionals programming problem. Metrika, 13(1):196–205, ???? 1968. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613385.

Uhlmann:1968:KOP

[259] Prof. Dr. Werner Uhlmann. Kostenoptimale Prüfpläne für die Gut– Schlecht–Prüfung. (German) []. Metrika, 13(1):206–224, ???? 1968. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613386.

Kaelin:1968:BGB

[260] A. Kaelin, A. Linder, F. Ferschl, and P. Sint. Buchbesprechungen. (German) [Book reviews]. Metrika, 13 (1):225–234, ???? 1968. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613387.

Anonymous:1968:HC

[261] Anonymous. Help & contacts. Metrika, 13(1):??, ???? 1968. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Bergmann:1969:SEP

[262] Prof. Dr. G. Bergmann. Seltene Ereignisse und Poisson–Verteilung. (German) []. Metrika, 14(1):1–20, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613638.

Mathai:1969:STG

[263] A. M. Mathai and R. K. Saxena. A short table of the generalized hypergeometric distribution. Metrika, 14(1):21–24, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613639.

Morgenstern:1969:MAG

[264] Prof. Dr. D. Morgenstern. Eine mehrdimensionale Ausdehnung des gewöhnlichen und des kanonischen Korrelationskoeffizienten. (German) []. Metrika, 14(1):25–43, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF02613640.

Morgenstern:1969:GEA

[265] Prof. Dr. D. Morgenstern. Grenzwertsätze für die Extrema der Abstände gewisser geordneter Stichproben. (German) []. Metrika, 14(1):44–47, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613641.

Randolph:1969:OSR

[266] P. H. Randolph. Optimal stopping rules for multinomial observations. Metrika, 14(1):48–61, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613642.

Roussas:1969:ANM

[267] G. G. Roussas. Asymptotic normality of the maximum likelihood estimate in Markov processes. Metrika, 14(1):62–70, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613643.

Stange:1969:PMP

[268] Prof. Dr. K. Stange. Pläne für messende Prüfung bei unbekannter Varianz der Fertigung und einem nach oben und unten abgegrenzten Toleranzbereich für die Merkmalwerte. (German) []. Metrika, 14(1):71–100, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613644.

Rao:1969:BRM

[269] S. Subba Rao. Balking and reneging in M/G/1 systems with postponable interruptions. Metrika, 14(1):101–116, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613645.

Vogt:1969:PPL

[270] H. Vogt. Zur Parameter- und Prozentpunktschätzung von Lebensdauerverteilungen bei kleinem Stichprobenumfang. (German) []. Metrika, 14(1):117–131, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613646.

Anonymous:1969:APD

[271] Anonymous. Auszeichnung für Prof. DDr. h. c. Wilhelm Winkler. (German) [Award for Prof. DDr. h. c. William Winkler]. Metrika, 14(1):132, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02613647.

Anonymous:1969:SSL

[272] Anonymous. Slawtscho Sagoroff, Leben und Wirken. (German) []. Metrika, 14(1):133–137, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613648.

Winkler:1969:DMB

[273] Prof. Dr. DDr. h. c. W. Winkler. Darstellung und Messung der Bevölkerungs-verteilung im geographischen Raum. (German) []. Metrika, 14(1):138–163, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF02613649.

Streissler:1969:RHB

[274] Prof. Dr. E. Streissler. A reexamination of the Haavelmo-bias in the least squares estimation of the Keynesian consumption function. Metrika, 14(1):164–182, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613650.

Bruckmann:1969:EBS

[275] Prof. Dr. G. Bruckmann. Einige Bemerkungen zur statistischen Messung der Konzentration. (German) []. Metrika, 14(1):183–213, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613651.

Derflinger:1969:EMO

[276] Prof. Dr. G. Derflinger. Efficient methods for obtaining the minres and maximum-likelihood solutions in factor analysis. Metrika, 14(1):214–231, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613652.

Heiler:1969:PDA

[277] Prof. S. Heiler and Prof. Dr. K. Weichselberger. Über den Permutationstest und ein daraus ableitbares Konfidenzintervall. (German) []. Metrika, 14(1):232–248, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613653.

Pfanzagl:1969:MCM

[278] Prof. Dr. J. Pfanzagl. On the measurability and consistency of minimum contrast estimates. Metrika, 14 (1):249–272, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613654.

Schmetterer:1969:PSA

[279] Prof. Dr. L. Schmetterer. Über ein Problem der stochastischen Approximation und ein Matrixsummierungsverfahren für Zahlenfolgen. (German) []. Metrika, 14(1):273–276, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613655.

Ferschl:1969:PMS

[280] Prof. Dr. F. Ferschl. Probleme und Möglichkeiten einer statistischen Dokumentation. (German) []. Metrika, 14 (1):277–292, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613656.

Bennett:1969:EIE

[281] Prof. J. F. Bennett. Ecometrics: An ideal for economics and ecology. Metrika, 14(1):293–301, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613657.

Anonymous:1969:HC

[282] Anonymous. Help & contacts. Metrika, 14(1):??, ???? 1969. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Aggarwal:1970:TTP

[283] S. P. Aggarwal. Transportation technique for a particular convex programming problem. Metrika, 15(1):1–5, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613548.

Bennett:1970:PFE

[284] B. M. Bennett. On the power function of the exact test for the r × c contingency table. Metrika, 15(1):6–8, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613549.

Delhi:1970:NCS

[285] M. S. Holla Delhi. On a noncentral chi-square distribution in the analysis of weapon systems effectiveness. Metrika, 15(1):9–14, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613550.

Kabe:1970:TOE

[286] D. G. Kabe. Testing outliers from an exponential population. Metrika, 15(1):15–18, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613551.

Malik:1970:DPS

[287] H. J. Malik. Distribution of product statistics from a Pareto population. Metrika, 15(1):19–22, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613552.

Meyer-Bahlburg:1970:NTR

[288] Dipl.-Psych. H. F. L. Meyer-Bahlburg. A nonparametric test for relative spread in k unpaired samples. Metrika, 15(1):23–29, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613553.

Pfanzagl:1970:MUE

[289] Prof. Dr. J. Pfanzagl. Median unbiased estimates for M.L.R.-families. Metrika, 15(1):30–39, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613554.

Ajgaonkar:1970:CTU

[290] S. G. Prabhu Ajgaonkar. A contribution to the theory of unique estimators for Horvitz and Thompson’s classes of linear estimators. Metrika, 15(1):40–47, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613555.

Schach:1970:ADC

[291] S. Schach. The asymptotic distribution of a class of nonparametric test statistics. Metrika, 15(1):48–58, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613556.

Sengupta:1970:AAS

[292] J. K. Sengupta. On the active approach of stochastic linear programming. Metrika, 15(1):59–70, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613557.

Subrahmanya:1970:MIE

[293] M. T. Subrahmanya. Modified integerestimators. Metrika, 15(1):71–80, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613558.

Swarup:1970:UBP

[294] K. Swarup. Upper bound problem in linear fractional functionals programming. Metrika, 15(1):81–85, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613559.

Tallis:1970:SED

[295] G. M. Tallis. Some extensions of discriminant function analysis. Metrika, 15(1):86–91, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613560.

Sint:1970:BGBa

[296] P. Sint and Wilhelm Winkler. Buchbesprechungen. (German) [Book reviews]. Metrika, 15(1):92–95, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613561.

Bandemer:1970:NBO

[297] Prof. Dr. H. Bandemer. Zur näherungsweisen Bestimmung des optimalen Stichprobenumfanges. (German) []. Metrika, 15(1):97–104, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613562.

Koop:1970:ECF

[298] Dr. J. C. Koop. Estimation of correlation for a finite universe. Metrika, 15(1):105–109, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613563.

Lutjohann:1970:SRA

[299] H. Lütjohann. The stepwise regression algorithm seen from the statistician’s point of view. Metrika, 15(1):110–125, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613564.

Malik:1970:EPP

[300] H. J. Malik. Estimation of the parameters of the Pareto distribution. Metrika, 15(1):126–132, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613565.

Nasr:1970:LSD

[301] S. K. Nasr. On a linear stochastic differential equation. Metrika, 15 (1):133–140, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613566.

Pfanzagl:1970:CEP

[302] Prof. Dr. J. Pfanzagl. Consistent estimation in the presence of incidental parameters. Metrika, 15(1):141–148, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613567.

Anonymous:1970:HCa

[303] Anonymous. Help & contacts. Metrika, 15(1):??, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Aczel:1970:USD

[304] J. Aczel, D. Z. Djoković, and J. Pfanzagl. On the uniqueness of scales derived from canonical representations. Metrika, 16(1):1–8, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613932.

Aggarwal:1970:ASL

[305] S. P. Aggarwal. Analyses of the solution to a linear fractional functionals programming. Metrika, 16(1):9–26, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613933.

Dubey:1970:CGB

[306] Satya D. Dubey. Compound gamma, beta and F distributions. Metrika, 16(1):27–31, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613934.

Frank:1970:SOS

[307] Ove Frank. Sampling from overlapping subpopulations. Metrika, 16(1):32–42, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613935.

Giger:1970:GSG

[308] H. Giger. Grundgleichungen der Stereologie I. (German) []. Metrika, 16 (1):43–57, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613936.

Puri:1970:ELP

[309] Madan Lal Puri and Pranab Kumar Sen. On the estimation of location parameters in the multivariate one sample and two sample problems. Metrika, 16(1):58–73, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613937.

Rana:1970:CCQ

[310] R. K. Rana. On a class of correlated queueing problems. Metrika, 16(1):74–84, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613938.

Sahler:1970:EMD

[311] W. Sahler. Estimation by minimumdiscrepancy methods. Metrika, 16 (1):85–106, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613939.

Subrahmanya:1970:NNN

[312] M. T. Subrahmanya. A note on nonnegative estimators of positive parameters. Metrika, 16(1):107–114, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613940.

Winckler:1970:SDF

[313] Klaus Winckler. Some distribution-free small samples methods for interval estimation and hypothesis testing in linear shock models. Metrika, 16(1):115–129, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613941.

Strecker:1970:BAZ

[314] Prof. Dr. H. Strecker. Ein Beitrag zur Analyse von Zeitreihen: Die Quotientenmethode, eine Variante der “Variate Difference” Methode. (German) []. Metrika, 16(1):130–187, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/

article/10.1007/BF02613942. See

remark [392].

Michel:1970:AN

[315] Dipl.-Math. R. Michel and Prof. Dr. J. Pfanzagl. Asymptotic normality. Metrika, 16(1):188–205, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613943.

Vogt:1970:DSM

[316] Dr. H. Vogt. Zur Definition und Schätzung von Mittelwerten für zufällige Variable auf der Sphäre. (German) [On the definition and estimation of average values for random variables on the sphere]. Metrika, 16(1):206–235, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613944.

Sint:1970:BGBb

[317] P. Sint, Dr. L. Reisinger, Josef Roppert, S. Sagoroff, Sint, A. Linder, K. Kreiter, and W. Winkler. Buchbesprechungen. (German) [Book reviews]. Metrika, 16(1):236–244, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613945.

Winkler:1970:NGM

[318] Wilhelm Winkler. Nachruf Giorgio Mortara. (German) [Obituary for Giorgio Mortara]. Metrika, 16(1):245, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02613946.

Bruckmann:1970:NSS

[319] Gerhart Bruckmann. Nachruf Slawtscho Sagoroff. (German) [Obituary for Slawtscho Sagoroff]. Metrika, 16(1): 246, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF02613947.

Anonymous:1970:HCb

[320] Anonymous. Help & contacts. Metrika, 16(1):??, ???? 1970. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Schaffer:1971:GGF

[321] K.-A. Schäffer and W. Wetzel. Geleitwort. (German) [Foreword]. Metrika, 17(1):1–2, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613805.

Anonymous:1971:SHM

[322] Anonymous. Schriftenverzeichnis von Hans Münzner. (German) [List of publications of Hans Münzner]. Metrika, 17(1):3–4, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613806.

Doring:1971:ADB

[323] Dr. H. Döring. Änderungen der durchschnittlichen Blutdruckwerte in den Jahren 1956–1968. (German) []. Metrika, 17(1):5–13, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF02613807.

Heubeck:1971:ZSU

[324] Dr. G. Heubeck. Zufällige oder systematische Unterschiede in den Sterbewahrscheinlichkeiten der Bevölkerung der BRD und der DDR?. (German) []. Metrika, 17(1):14–23, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613808.

Kellerer:1971:IKH

[325] Prof. Dr. H. Kellerer. Der Intraclass– Koeffizient — ein Homogenitätsmaß. (German) []. Metrika, 17(1):24–28, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613809.

Koller:1971:MAF

[326] Prof. Dr. S. Koller. Mögliche Aussagen bei Fragen der statistischen Ursachenforschung. (German) []. Metrika, 17(1):30–42, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613810.

Ludeke:1971:PLI

[327] Prof. Dr. D. Lüdeke. Produktivitätsorientierte Lohnpolitik und ihre vermutlichen Wirkungen auf die wirtschaftliche Entwicklung in der Bundesrepublik Deutschland während des Zeitraumes 1952–1966 — Eine ökonometrische Analyse. (German) [Productivity-oriented wage policy and its probable effects on economic development in the Federal Republic of Germany during the period 1952–1966 — an econometric analysis]. Metrika, 17(1):43–68, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613811.

Marfels:1971:ENE

[328] Prof. Dr. C. Marfels. Einige neuere Entwicklungen in der Messung der industriellen Konzentration. (German) []. Metrika, 17(1):69–81, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF02613812.

Pfanzagl:1971:BEB

[329] Prof. Dr. J. Pfanzagl. The Berry– Esseen bound for minimum contrast estimates. Metrika, 17(1):82–91, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613813.

Reichel:1971:EBP

[330] Dr. G. Reichel. Erweitertes Bernoulli– Prinzip und Petersburger Spiel. (German) []. Metrika, 17(1):92–97, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613814.

Schaffer:1971:OSL

[331] Prof. Dr. K.-A. Schäffer. Optimale Schichtabgrenzung bei logarithmischer Normalverteilung. (German) []. Metrika, 17(1):98–115, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613815.

Schonfeld:1971:UCL

[332] Prof. Dr. P. Schönfeld. A useful Central Limit Theorem for m-dependent variables. Metrika, 17(1):116–128, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613816.

Stange:1971:BOA

[333] Professor Dr. K. Stange. Eine Bemerkung zum optimalen Auswahlsatz bei geschichteten Stichproben. (German) []. Metrika, 17(1):129–140, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613817.

Strecker:1971:MEA

[334] Professor Dr. H. Strecker. Ein Modell zur Ermittlung von Angabefehlern in statistischen Erhebungen. (German) []. Metrika, 17(1):141–145, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613818.

Walter:1971:MVG

[335] Prof. Dr. E. Walter. Über die Methoden der Vaterschaftsbegutachtung. (German) [On methods of assessment of paternity]. Metrika, 17(1):146–158, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613819.

Weichselberger:1971:SEK

[336] Prof. Dr. K. Weichselberger. Über die statistischen Eigenschaften von Korrekturen aufgrund repräsentativer Verfahrenskontrollen. (German) []. Metrika, 17(1):159–188, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF02613820.

Wetzel:1971:BSO

[337] Prof. Dr. W. Wetzel. Bemerkungen zur Saisonbereinigung von ökonomischen Zeitreihen. (German) []. Metrika, 17 (1):189–196, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613821.

Wunsche:1971:PTL

[338] Prof. Dr. G. Wünsche. Über die projektive Transformation der logarithmischen Funktionsleiter (Zur graphischen Darstellung von Sterbewahrscheinlichkeiten). (German) []. Metrika, 17 (1):197–205, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613822.

Tuteja:1971:STS

[339] R. K. Tuteja. Solution of a transient state, limited space queueing problem with arrival and departure rates depending on queue length. Metrika, 17(1):207–214, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613823.

Kurotschka:1971:OVZ

[340] Dr. V. Kurotschka. Optimale Versuchspläne bei zweifach klassifizierten Beobachtungsmodellen. (German) []. Metrika, 17(1):215–232, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF02613824.

Anonymous:1971:HC

[341] Anonymous. Help & contacts. Metrika, 17(1):??, ???? 1971. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Bosch:1972:EBZ

[342] Dr. K. Bosch and Prof. Dr. H. Richter. Über den Erwartungswert der bis zur Zeit t angekommenen Kunden einer Warteschlange bei gleichverteilten Ankunftszeiten. (German) []. Metrika, 18(1):1–5, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02614230.

Landhausser:1972:ESE

[343] E. Landhäusser. Elliptische Streuflächen ebener Punkthaufen. (German) [Elliptical scatterers of planar point clusters]. Metrika, 18(1):6–14, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02614231.

Ajgaonkar:1972:CEV

[344] S. G. Prabhu Ajgaonkar. On classes of estimators of the variance function of a linear estimator. Metrika, 18(1):15–20, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614232.

Bector:1972:IQF

[345] C. R. Bector. Indefinite quadratic fractional functional programming. Metrika, 18(1):21–30, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF02614233.

Edlefsen:1972:FDP

[346] L. E. Edlefsen and C. B. Millham. On a formulation of discrete, N-person non-cooperative games. Metrika, 18 (1):31–34, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02614234.

Feichtinger:1972:WSL

[347] Dr. Gustav Feichtinger. ”Wahrscheinlichkeitslernen” in der statistischen Lerntheorie. (German) []. Metrika, 18(1):35–55, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02614235.

Plachky:1972:MEG

[348] Dr. D. Plachky. Eine Momentengleichung bei Exponentialfamilien. (German) []. Metrika, 18(1):56–59, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614236.

Tuteja:1972:QPC

[349] R. K. Tuteja. On the queueing problem of correlated arrivals with many servers in a finite waiting space. Metrika, 18(1):60–68, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02614237.

Rana:1972:QPA

[350] R. K. Rana. Queueing problems with arrivals in General Stream and phase type service. Metrika, 18(1):69–80, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02614238.

Bennett:1972:NST

[351] B. M. Bennett. Note on sign tests for equality of means of k bivariate populations. Metrika, 18(1):81–83, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02614239.

Giger:1972:GSI

[352] H. Giger. Grundgleichungen der Stereologie II. (German) []. Metrika, 18 (1):84–93, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02614240.

Feichtinger:1972:EMR

[353] Dr. Gustav Feichtinger. Über ein N-elementiges Modell der Reiz– Stichproben-Theorie mit stetigem Zeitparameter und zwei Antwortklassen. (German) []. Metrika, 18(1):94–109, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614241.

Murari:1972:QPC

[354] K. Murari. A queueing problem with correlated arrivals and correlated phase-type service zur theorie des rangtests. Metrika, 18(1):110–119, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614242.

Bock:1972:SMB

[355] Dr. Hans Hermann Bock. Statistische Modelle und Bayessche Verfahren zur Bestimmung einer unbekannten Klassifikation normalverteilter zufälliger Vektoren. (German) []. Metrika, 18 (1):120–132, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02614243.

Hasija:1972:TWI

[356] R. L. Hasija. A two warehouse inventory problem. Metrika, 18(1):133–136, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614244.

Witting:1972:TRG

[357] Prof. Dr. H. Witting. Zur Theorie der Rangtests. (German) []. Metrika, 18(1):137–144, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02614245.

Kulshrestha:1972:ADT

[358] Dr. D. K. Kulshrestha. Application of discrete transforms in reliability of a parallel redundant system. Metrika, 18(1):145–153, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02614246.

Pfanzagl:1972:MUE

[359] Prof. Dr. J. Pfanzagl. On median unbiased estimates. Metrika, 18 (1):154–173, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02614247.

Pfanzagl:1972:FRAa

[360] Prof. Dr. J. Pfanzagl. Further results on asymptotic normality. I. Metrika, 18(1):174–198, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02614248.

Fischer:1972:I

[361] P. Fischer. On the inequality Σp_{i}f(p_{i}) ≥ p_{i}f(q_{i}). Metrika, 18 (1):199–208, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02614249.

Anonymous:1972:HCa

[362] Anonymous. Help & contacts. Metrika, 18(1):??, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Sharma:1972:AIT

[363] B. D. Sharma. On amount of information of type-β and other measures. Metrika, 19(1):1–10, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893270.

Varma:1972:GPL

[364] G. K. Varma. General parametric linear fractional programming. Metrika, 19(1):11–17, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893271.

Chadha:1972:DNL

[365] S. S. Chadha and R. N. Kaul. A dual non-linear program. Metrika, 19(1):18–22, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893272.

Bennett:1972:TMS

[366] B. M. Bennett. Tests for marginal symmetry in contingency tables. Metrika, 19(1):23–26, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893273.

Murari:1972:QPA

[367] K. Murari. A queueing problem with arrival and service in batches of variable size. Metrika, 19(1):27–35, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893274.

Bennett:1972:NND

[368] B. M. Bennett. On the non-null distribution of Wilcoxon’s signed rank test. Metrika, 19(1):36–38, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893275.

Singh:1972:EVO

[369] N. Singh and N. Bhashyam. Effect of the variable offset distance on the kill probability. Metrika, 19(1):39–47, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893276.

Humburg:1972:WWG

[370] Dipl.-Math. J. Humburg. Wahrscheinlichkeitsund Wettbegriff. (German) [Probability and betting term]. Metrika, 19(1):48–53, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893277.

Heyer:1972:EDB

[371] Prof. Dr. H. Heyer. Zum Erschöpftheitsbegriff von D. Blackwell. (German) [On the exhaustion concept of D. Blackwell]. Metrika, 19(1):54–67, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893278.

Widdra:1972:VGS

[372] W. Widdra. Eine Verallgemeinerung des “Gesetzes seltener Ereignisse”.

(German) [A generalization of the “Law of Rare Events”]. Metrika, 19(1):68–71, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893279.

Schmitz:1972:NOS

[373] Dr. N. Schmitz. A note on the optimality of k-stage tests. Metrika, 19(1):72–75, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893280.

Raghavachari:1972:NPD

[374] M. Raghavachari. A note on the Pareto distribution. Metrika, 19(1):76, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01893281.

Bennett:1972:THS

[375] B. M. Bennett. On a test of homogeneity for samples from a negative binomial distribution. Metrika, 19(1):77, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01893282.

Riedwyl:1972:WW

[376] Hans Riedwyl. Walter Wegmüller. Metrika, 19(1):78, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01893283.

Sint:1972:BGB

[377] P. Sint, J. Gordesch, L. Reisinger, A. Kaelin, P. Ihm, F. Streit, Karl Trautmann, and Karl Trautmann. Buchbesprechungen. (German) [Book reviews]. Metrika, 19(1):79–88, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893284.

Pfanzagl:1972:FRAb

[378] Prof. Dr. J. Pfanzagl. Further results on asymptotic normality II. Metrika, 19(1):89–97, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01893285.

Strasser:1972:SUE

[379] H. Strasser. Sufficiency and unbiased estimation. Metrika, 19(1):98–114, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893286.

Ganssler:1972:NMC

[380] Prof. Dr. P. Gänssler. Note on minimum contrast estimates for Markov processes. Metrika, 19(1):115–130, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893287.

Wertz:1972:FKN

[381] Dr W. Wertz. Fehlerabschätzung für eine Klasse von nichtparametrischen Schätzfolgen. (German) []. Metrika, 19(1):131–139, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893288.

Gordesch:1972:EPM

[382] Dr. J. Gordesch. On the equality of pay-off matrices. Metrika, 19 (1):140–149, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893289.

Anonymous:1972:HCb

[383] Anonymous. Help & contacts. Metrika, 19(1):??, ???? 1972. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

VonSchneeberger:1973:OAS

[384] H. Von Schneeberger. Optimale Aufteilung des Stichprobenumfangs bei mehreren Merkmalen. (German) []. Metrika, 20(1):1–16, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF01893795.

Bhashyam:1973:SDC

[385] N. Bhashyam. Stochastic duels with correlated fire. Metrika, 20(1):17–24, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893796.

Chadha:1973:GUB

[386] S. S. Chadhà. A generalized upper bounded technique for a linear fractional program. Metrika, 20(1):25–35, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893797.

Bagga:1973:TMS

[387] P. C. Bagga. Two machine scheduling problem when passing allowed. Metrika, 20(1):36–40, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF01893798.

Fieger:1973:EHM

[388] Prof. Dr. W. Fieger. Zur Existenz der höheren Momente von Punktprozessen. (German) []. Metrika, 20(1):41–50, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893799.

Gupta:1973:MLE

[389] R. P. Gupta. Maximum likelihood estimate of the parameters of a truncated inverse Gaussian distribution. Metrika, 20(1):51–53, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893800.

Kpedekpo:1973:RAS

[390] G. M. K. Kpedekpo. Recent advances on some aspects of stratified sample design. a review of the literature. Metrika, 20(1):54–64, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893801.

Muszely:1973:CSF

[391] Gy. Muszély. On continuous solutions of a functional inequality. Metrika, 20(1):65–69, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893802.

Schips:1973:EBE

[392] B. Schips and W. Stier. Einige Bemerkungen zu einem Beitrag zur Zeitreihenanalyse von H. Strecker. (German) [Some remarks on a contribution to time series analysis by H. Strecker]. Metrika, 20(1):70–73, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF01893803. See

[314].

Sint:1973:BGB

[393] P. Sint, J. Gordesch, E. Neuwirth, L. Reisinger, F. Streit, E. Fürst, and R. Wytek. Buchbesprechungen. (German) [Book reviews]. Metrika, 20(1):74–80, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893804.

Sharda:1973:QPB

[394] Miss Sharda. A queueing problem with batch arrivals and correlated departures. Metrika, 20(1):81–92, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893805.

Sharda:1973:CTP

[395] Miss Sharda. On a certain type of priority queueing problem. Metrika, 20(1):93–100, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01893806.

Wheeler:1973:NKS

[396] Donald J. Wheeler. Note on the Kolmogorov statistic for a general discontinuous variable. Metrika, 20 (1):101–102, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF01893807.

Laube:1973:WCC

[397] G. Laube. Weak convergence and convergence in the mean of distribution functions. Metrika, 20(1):103–105, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893808.

Goldstein:1973:APT

[398] B. H. Goldstein. Über eine Anwendung der Potentialtheorie transienter Markoff–Ketten. (German) []. Metrika, 20(1):106–113, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01893809.

Vogt:1973:VZO

[399] Dr. H. Vogt. Vergleich zweier operationscharakteristiken. Metrika, 20 (1):114–121, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893810.

Rathie:1973:SCT

[400] Dr. P. N. Rathie. Some characterization theorems for generalized measures of uncertainty and information. Metrika, 20(1):122–130, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893811.

Seeger:1973:DEM

[401] P. Seeger. The determination of expected mean squares for ANOVA in balanced designs. Metrika, 20 (1):131–136, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893812.

Linhart:1973:SKS

[402] H. Linhart. Schätzen der Kovarianzdichte stationärer Punktprozesse aus Zählungen. (German) []. Metrika, 20(1):137–147, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893813.

Michel:1973:BBE

[403] R. Michel. The bound in the Berry– Esséen result for minimum contrast estimates. Metrika, 20(1):148–155, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893814.

Anonymous:1973:HC

[404] Anonymous. Help & contacts. Metrika, 20(1):??, ???? 1973. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Strasser:1974:BEE

[405] Dr. H. Strasser. Eine Bemerkung zur Existenz erwartungstreuer Schätzer. (German) []. Metrika, 21(1):1–5, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893888.

Stenger:1974:MS

[406] Prof. Dr. H. Stenger. Mehrstufige stichprobenverfahren. Metrika, 21(1):7–18, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893889.

Koop:1974:NUT

[407] Dr. J. C. Koop. Notes for a unified theory of estimation for sample surveys taking into account response errors. Metrika, 21(1):19–39, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893890.

Sharma:1974:RIF

[408] B. D. Sharma and R. Autar. Relative information functions and their type (α, β) generalizations. Metrika, 21(1):41–50, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01893891.

Gordesch:1974:BGB

[409] J. Gordesch, G. Pflug, H.-J. Lenz, H. Strasser, W. Vogel, H. Basler, W. Hazod, A. Linder, G. Feichtinger, F. Ferschl, W. Polasek, H. Drygas, D. Morgenstern, J. Gordesch, P. Sint, and et al. Buchbesprechungen. (German) [Book reviews]. Metrika, 21 (1):51–64, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893892.

Schips:1974:PSO

[410] B. Schips and W. Stier. Zum Problem der Saisonbereinigung ökonomischer Zeitreihen. (German) []. Metrika, 21(1):65–81, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893893.

Ludwig:1974:BKP

[411] Dr. R. Ludwig. Bestimmung kostenoptimaler Parameter für den kontinuierlichen Stichprobenplan von Dodge. (German) []. Metrika, 21(1):83–126, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893894.

Liu:1974:BEV

[412] T. P. Liu. Bayes estimation for the variance of a finite population. Metrika, 21(1):127–132, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893895.

Dickinson:1974:RMO

[413] Dr. J. P. Dickinson. On the resolution of a mixture of observations from two gamma distributions by the method of maximum likelihood. Metrika, 21(1):133–141, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893896.

Stenger:1974:SPS

[414] Prof. Dr. H. Stenger. Schätzbare Parameterfunktionen in der Stichprobentheorie. (German) []. Metrika, 21 (1):143–154, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01893897.

Richter:1974:KSM

[415] Dr. G. Richter. Kostenoptimale Stichprobenpläne für mehrere qualitative Merkmale. (German) []. Metrika, 21 (1):155–173, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01893898.

vCollani:1974:WMS

[416] Dipl Mathem. E. v. Collani. Zur Wahl eines mehrstufigen Stichprobenplanes. (German) []. Metrika, 21(1):175–196, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893899.

Pfanzagl:1974:CSP

[417] Prof. Dr. J. Pfanzagl. A characterization of sufficiency by power functions. Metrika, 21(1):197–199, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893900.

Padmanabhan:1974:TCP

[418] A. R. Padmanabhan. Two comments on a paper of Bahadur. Metrika, 21 (1):201–203, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893901.

Michel:1974:BEA

[419] Dipl.-Math. R. Michel. Bounds for the efficiency of approximately median unbiased estimates. Metrika, 21 (1):205–211, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893902.

Zimmermann:1974:NBT

[420] Dipl.-Math. H. Zimmermann. A note on Bennett’s “Test for marginal symmetry in contingency tables”, Metrika 19, 1972, H. 1, 23–26. Metrika, 21 (1):213–214, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01893903.

Naik:1974:BOA

[421] U. D. Naik. On Bayesian optimum allocation in some two phase sampling problems. Metrika, 21(1):215–221, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893904.

Bomsdorf:1974:CLD

[422] Dr. E. Bomsdorf. Zur Charakterisierung von Lokalisations-und Dispersionsmaßen. (German) []. Metrika, 21(1):223–229, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01893905.

Stange:1974:PMP

[423] Professor Dr. K. Stange. Pläne für messende Prüfung bei Berücksichtigung von “Vorkenntnissen” über die Verteilung der Mittelwerte. (German) []. Metrika, 21(1):231–247, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01893906.

Drygas:1974:LMH

[424] Priv.-Doz. Dr. H. Drygas. Lineare Modelle und Hochrechnung von Wahlergebnissen. (German) []. Metrika, 21(1):249–255, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF01893907.

Anonymous:1974:HC

[425] Anonymous. Help & contacts. Metrika, 21(1):??, ???? 1974. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Basler:1975:UOL

[426] Dr. H. Basler. Eine Untersuchung der Olympia-Lotterie “Glücksspirale 1971”. (German) []. Metrika, 22 (1):1–11, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01899709.

Deistler:1975:TIL

[427] Dr. M. Deistler. z-Transform and identification of linear econometric models with autocorrelated errors. Metrika, 22(1):13–25, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01899710.

Sharma:1975:II

[428] B. D. Sharma and R. Autar. On information-improvement. Metrika, 22(1):27–34, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01899711.

Mittal:1975:SFE

[429] D. P. Mittal. On some functional equations concerning entropy, directed divergence and inaccuracy. Metrika, 22(1):35–45, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01899712.

Davenport:1975:BFP

[430] J. M. Davenport and J. T. Webster. The Behrens–Fisher problem, an old solution revisited. Metrika, 22(1):47–54, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01899713.

Prabhu-Ajgaonkar:1975:UHT

[431] S. G. Prabhu-Ajgaonkar. On unordering Horvitz and Thompson’s T_{1} class of linear estimators. Metrika, 22(1):55–59, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01899714.

Morgenstern:1975:BGB

[432] D. Morgenstern, W. Uhlmann, and P. Sint. Buchbesprechungen. (German) [Book reviews]. Metrika, 22(1):61–64, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01899715.

Talukder:1975:RBD

[433] M. A. H. Talukder. Response bias and difference estimate in double sampling. Metrika, 22(1):65–75, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01899716.

Schaefer:1975:BET

[434] Dr. M. Schaefer. Berry–Esseen-type inequality for a special class of distributions. Metrika, 22(1):77–83, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01899717.

Gupta:1975:FMF

[435] B. C. Gupta. Finite moment formulae of h-statistics and their estimates. Metrika, 22(1):85–95, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01899718.

Pfaffelhuber:1975:MSA

[436] E. Pfaffelhuber. Moving shift averages for ergodic transformations. Metrika, 22(1):97–101, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01899719.

Diefenbach:1975:BVV

[437] L. Diefenbach. Eine Bemerkung über das Verhalten der Verbesserungen und ein neuer Vorschlag zur Gewichtung der Meßwerte bei Ausgleichung bedingter direkter Beobachtungen. (German) []. Metrika, 22(1):103–118, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01899720.

Mukhopadhyay:1975:OSD

[438] P. Mukhopadhyay. An optimum sampling design to base HT-method of estimating a finite population total. Metrika, 22(1):119–127, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01899721.

Kuss:1975:MPE

[439] Dr. rer. nat. U. Kuß. Maximum probability estimators in the case of exponential distribution. Metrika, 22 (1):129–146, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01899722.

Chaudhuri:1975:SMS

[440] Dr. A. Chaudhuri. A simple method of sampling without replacement with inclusion-probabilities exactly proportional to size. Metrika, 22(1):147–152, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01899723.

Bhat:1975:IMC

[441] U. N. Bhat. Imbedded Markov chains in queueing systems M/G/1 and GI/M/1 with limited waiting room. Metrika, 22(1):153–160, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01899724.

Buhler:1975:OSG

[442] Dr. W. Bühler and Dr. T. Deutler. Optimal stratification and grouping by dynamic programming. Metrika, 22(1):161–175, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01899725.

Polasek:1975:BGB

[443] W. Polasek, P. Sint, A. Leitsch, E. Neuwirth, W. Polasek, F. Ferschl, G. Feichtinger, L. Reisinger, E. Neuwirth, W. Till, L. Schmetterer, and H. Störmer. Buchbesprechungen. (German) [Book reviews]. Metrika, 22(1):177–188, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01899726.

Reiss:1975:CCC

[444] R. D. Reiss. Consistency of a certain class of empirical density functions. Metrika, 22(1):189–203, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01899727.

Sharma:1975:ETO

[445] B. D. Sharma and I. J. Taneja. Entropy of type (α, β) and other generalized measures in information theory. Metrika, 22(1):205–215, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01899728.

Anonymous:1975:HC

[446] Anonymous. Help & contacts. Metrika, 22(1):??, ???? 1975. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Rehder:1976:RRL

[447] W. Rehder. A remark on robustness of linear best estimates. Metrika, 23(1):1– 6, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902844.

Reiss:1976:MDE

[448] Dr. R.-D. Reiss. On minimum distance estimators for unimodal densities. Metrika, 23(1):7–14, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902845.

Thomsen:1976:CAO

[449] I. Thomsen. A comparison of approximately optimal stratification given proportional allocation with other methods of stratification and allocation. Metrika, 23(1):15–25, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902846.

Lingappaiah:1976:EOE

[450] G. S. Lingappaiah. Effect of outliers on the estimation of parameters. Metrika, 23(1):27–30, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902847.

Mittal:1976:CSF

[451] D. P. Mittal. On continuous solutions of a functional equation. Metrika, 23(1):31–40, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01902848.

Brown:1976:RRA

[452] Dr. C. C. Brown. The restricted regularity of admissible minimax criteria for many well known parametric statistical problems. Metrika, 23(1):41–63, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902849.

Anderson:1976:SNT

[453] O. D. Anderson. Some new time series results. Metrika, 23(1):65–76, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902850.

Schuler:1976:KSM

[454] Dr. L. Schüler. Über die Konsistenz einer Schätzung mehrdimensionaler Dichten auf der Basis trigonometrischer Reihen. (German) []. Metrika, 23(1):77–82, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902851.

Brown:1976:CSP

[455] Dr. C. C. Brown. A conventional statistical problem with irregular minimax behavior. Metrika, 23(1):83–89, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902852.

Strasser:1976:NE

[456] Prof. Dr. H. Strasser. A note on efficiency. Metrika, 23(1):91–100, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902853.

Schneeweiss:1976:CER

[457] Prof. Dr. H. Schneeweiß. Consistent estimation of a regression with errors in the variables. Metrika, 23(1):101–115, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902854.

Morgenstern:1976:ZVP

[458] Prof. Dr. D. Morgenstern. Eine zweifache Verallgemeinerung der Pólya-Verteilung und ihre verschiedenen Deutungen. (German) []. Metrika, 23 (1):117–122, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01902855.

Ferschl:1976:BGB

[459] F. Ferschl, P. Rice, W. Wetzel, J. Casti, W. Grossmann, and B. Rüger. Buchbesprechungen. (German) [Book reviews]. Metrika, 23(1):123–126, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902856.

Shukla:1976:AUP

[460] N. D. Shukla. Almost unbiased product-type estimator. Metrika, 23 (1):127–133, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01902857.

Chiu:1976:EDP

[461] W. K. Chiu. On the estimation of data parameters for economic optimum 135–1135–1135–1-charts. Metrika, 23 (1):135–147, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01902858.

Schuler:1976:SDG

[462] Dr. L. Schüler and Dr. H. Wolff. Zur Schätzung eines Dichtefunktionals. (German) []. Metrika, 23(1):149–154, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902859.

Sharma:1976:SAM

[463] B. D. Sharma and H. C. Gupta. Sub additive measures of relative information and inaccuracy. Metrika, 23 (1):155–165, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01902860.

Hilden:1976:SLT

[464] Dr. H. Hilden. Die Schätzung des linearen Todesprozesses bei fehlerbehafteten Beobachtungen. (German) []. Metrika, 23(1):167–191, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902861.

Pflug:1976:BGB

[465] G. Pflug, Yu. A. Rozanov, W. Vogel, A. Kalliauer, and W. Grossmann. Buchbesprechungen. (German) [Book reviews]. Metrika, 23(1):193–195, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902862.

Anonymous:1976:PCS

[466] Anonymous. Proceedings in computational statistics. Metrika, 23(1):196, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01902863.

Anonymous:1976:HC

[467] Anonymous. Help & contacts. Metrika, 23(1):??, ???? 1976. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Nath:1977:EPI

[468] G. B. Nath. Estimation of the parameter of the inverse Gaussian distribution from generalised censored samples. Metrika, 24(1):1–6, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893387.

Fieger:1977:TNC

[469] Prof. Dr. W. Fieger. Transformationen, die die Normalverteilung charakterisieren. (German) []. Metrika, 24(1):7–22, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01893388.

Patni:1977:ACS

[470] G. C. Patni and Dr. K. C. Jain. On axiomatic characterization of some non-additive measures of information. Metrika, 24(1):23–34, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893389.

delJunco:1977:MAE

[471] A. del Junco and J. M. Steele. Moving averages of ergodic processes. Metrika, 24(1):35–43, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893390.

Prajneshu:1977:SSS

[472] Prajneshu. Statistical study of a stochastic model of two interacting species. Metrika, 24(1):45–53, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893391.

Gordesch:1977:BGB

[473] J. Gordesch, E. Neuwirth, H. Strasser, J. Ledolter, H. Abele, F. Ferschl, W. Polasek, A. Szöllösi-Nagy, G. Pflug, G. Bonelli, E. Neuwirth, and Eberl. Buchbesprechungen. (German) [Book reviews]. Metrika, 24(1):55–63, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893392.

Bosch:1977:EWG

[474] Prof. Dr. K. Bosch. Erneuerungsprozesse mit Wartungsstrategien. (German) []. Metrika, 24(1):65–74, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF01893393.

Jensen:1977:ADF

[475] Prof. D. R. Jensen. On approximating the distributions of Friedman’s 75-175-175-1 and related statistics. Metrika, 24(1):75–85, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893394.

Mudholkar:1977:NET

[476] Prof. G. S. Mudholkar, P. Subbaiah, and S. George. A note on the estimation of two ordered Poisson parameters. Metrika, 24(1):87–98, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893395.

Dyer:1977:SAF

[477] D. Dyer. Statistical analysis of Fourier coefficients in a restricted harmonic dial. Metrika, 24(1):99–105, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893396.

Fabian:1977:MML

[478] Professor V. Fabian. A modified maximum likelihood test. Metrika, 24 (1):107–112, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01893397.

Myoken:1977:GRE

[479] Professor H. Myoken and Professor Y. Uchida. The generalized ridge estimator and improved adjustments for regression parameters. Metrika, 24 (1):113–124, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01893398.

Erber:1977:BGB

[480] G. Erber, G. Knolmayer, J. Gordesch, E. v. Collani, G. Pflug, and W. Grossmann. Buchbesprechungen.

(German) [Book reviews]. Metrika, 24(1):125–128, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893399.

Talwalker:1977:NCC

[481] Dr. S. Talwalker. A note on characterization by the conditional expectation. Metrika, 24(1):129–136, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893400.

Steinebach:1977:ECP

[482] J. Steinebach. Exponential convergence properties of linear estimators under exponential and uniform distribution. Metrika, 24(1):137–161, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893401.

Irle:1977:DTC

[483] Dr. A. Irle. Decision theory for continuous observations: Minimax solutions. Metrika, 24(1):163–168, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893402.

Dale:1977:CMN

[484] A. I. Dale. Convergence in r-mean of normalized partial sums in a separable Banach space. Metrika, 24(1):169–173, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893403.

Knusel:1977:VZT

[485] Prof. Dr. L. Knüsel. Vergleich zweier Testverfahren zur Symmetrieprüfung in 4-Felder-Tafeln. (German) []. Metrika, 24(1):175–186, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893404.

Anderson:1977:ABJ

[486] O. D. Anderson. An appraisal of the Box–Jenkins approach to univariate time series analysis. Metrika, 24 (1):187–194, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01893405.

Unbehauen:1977:KBG

[487] H. Unbehauen, B. Göring, and B. Bauer. Korrektur zu der Buchbesprechung. (German) []. Metrika, 24(1): 195, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF01893407.

Anonymous:1977:HC

[488] Anonymous. Help & contacts. Metrika, 24(1):??, ???? 1977. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Platek:1978:SDC

[489] R. Platek and Dr. M. P. Singh. A strategy for up-dating continuous surveys. Metrika, 25(1):1–7, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02204346.

Reiss:1978:ADM

[490] R.-D. Reiss. Approximate distribution of the maximum deviation of histograms. Metrika, 25(1):9–26, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02204347.

Thomsen:1978:DEP

[491] I. Thomsen. Design and estimation problems when estimating a regression coefficient from survey data. Metrika, 25(1):27–35, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02204348.

Fieger:1978:SVS

[492] Prof. Dr. W. Fieger. Suffizienz bei verschiebungsinvarianten Schätzproblemen. (German) []. Metrika, 25(1):37–48, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02204349.

Vogt:1978:VEV

[493] Dr. H. Vogt. Über eine Variante der empirischen Verteilungsfunktion. (German) []. Metrika, 25(1):49–58, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02204350.

Lorah:1978:EFP

[494] Ms. D. J. Lorah and Prof. R. M. Stark. On elementary functions of Pareto variables. Metrika, 25(1):59–63, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02204351.

Chaudhuri:1978:EVF

[495] Dr. A. Chaudhuri. On estimating the variance of a finite population. Metrika, 25(1):65–76, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02204352.

Ledolter:1978:COE

[496] Dr. J. Ledolter and Prof. G. E. P. Box. Conditions for the optimality of exponential smoothing forecast procedures. Metrika, 25(1):77–93, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02204353.

Landers:1978:NGA

[497] Prof. Dr. D. Landers and Prof. Dr. L. Rogge. On nonuniform Gaussian approximation for random summation. Metrika, 25(1):95–114, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02204354.

Mukhopadhyay:1978:EVF

[498] P. Mukhopadhyay. Estimating the variance of a finite population under a superpopulation model. Metrika, 25 (1):115–122, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02204355.

Uberhuber:1978:BGB

[499] Chr W. Überhuber, R. Schlittgen, G. Pflug, J. Ledolter, J. Gordesch, W. Grossmann, R. Schlittgen, G. Bonelli, and A. Linder. Buchbesprechungen. (German) [Book reviews]. Metrika, 25(1):123–128, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02204356.

Reiss:1978:CMC

[500] R.-D. Reiss. Consistency of minimum contrast estimators in non-standard cases. Metrika, 25(1):129–142, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02204357.

Morgenstern:1978:EVK

[501] Prof. Dr. D. Morgenstern. Erwartungswerte, Varianzen und Kovarianzen mehrerer Maxima der Abstände unabhängiger rechteckverteilter Zufallsvariabler. (German) []. Metrika, 25(1):143–148, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02204358.

Pedgaonkar:1978:CSS

[502] A. M. Pedgaonkar and S. G. Prabhu-Ajgaonkar. Comparison of sampling strategies. Metrika, 25(1):149–153, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02204359.

Patel:1978:MVUa

[503] S. R. Patel. Minimum variance unbiased estimation of multivariate modified power series distribution. Metrika, 25(1):155–161, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02204360.

Satterthwaite:1978:GGB

[504] S. P. Satterthwaite and Dr. T. P. Hutchinson. A generalisation of Gumbel’s bivariate logistic distribution. Metrika, 25(1):163–170, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02204361.

Ruger:1978:MST

[505] B. Rüger. Das maximale signifikanzniveau des Tests: “Lehne H_{o} ab, wenn k unter n gegebenen tests zur ablehnung führen”. (German) []. Metrika, 25(1):171–178, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02204362.

Hackl:1978:NSL

[506] Dipl. Ing. Dr. P. Hackl. Nichtparametrisches Schätzen in Lebensdauer — Untersuchungen bei reduzierter Information. (German) []. Metrika, 25(1):179–192, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF02204363.

El-Sayed:1978:CCE

[507] A.-B. El-Sayed. Cascaded channels and the equivocation inequality. Metrika, 25(1):193–208, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02204364.

Patel:1978:MVUb

[508] S. R. Patel and Dr. S. M. Shah. Minimum variance unbiased estimation of left truncated multivariate power series distributions. Metrika, 25(1):209–218, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02204365.

Anonymous:1978:HC

[509] Anonymous. Help & contacts. Metrika, 25(1):??, ???? 1978. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Wong:1979:EQT

[510] Prof. P. G. Wong and Prof. S. P. Wong. An extremal quotient test for exponential distributions. Metrika, 26(1):1–4, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893465.

Stormer:1979:TAV

[511] Prof. Dr. H. Störmer. Über das Testen von Ausschußanteilen mit verallgemeinerten t-Tests. (German) []. Metrika, 26(1):5–24, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF01893466.

Talwalker:1979:NCM

[512] Dr. Sh. Talwalker. A note on the characterization of the multivariate normal distribution. Metrika, 26(1):25–30, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893467.

Fieger:1979:TMP

[513] Prof. Dr. W. Fieger. r-te Momente von Punktprozessen. (German) []. Metrika, 26(1):31–41, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01893468.

Ledolter:1979:IRA

[514] J. Ledolter. Inference robustness of ARIMA models under non-normality-special application to stock price data. Metrika, 26(1):43–56, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893469.

Hutcheson:1979:MVL

[515] K. Hutcheson and N. I. Lyons. The mean and variance of the likelihood ratio criterion for the multinomial distribution. Metrika, 26(1):57–61, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893470.

Feichtinger:1979:BGB

[516] G. Feichtinger, H. Strelec, G. Pflug, K. H. Keil, and P. Fleissner. Buchbesprechungen. (German) [Book reviews]. Metrika, 26(1):63–64, ????

1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01893472.

Wertz:1979:A

[517] Prof. Dr. W. Wertz. Ankündigung. Metrika, 26(1):63, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01893471.

Anderson:1979:PRB

[518] O. D. Anderson. A proof of a relationship between the generalized variances for associated autoregressive and moving average processes. Metrika, 26(1):65–70, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893473.

Bosch:1979:FHM

[519] Prof. Dr. K. Bosch. Funktionen homogener Markoffscher Ketten als homogene bzw. inhomogene Markoffsche Ketten. (German) []. Metrika, 26(1):71–85, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893474.

Patel:1979:MVU

[520] S. R. Patel. Minimum variance unbiased estimation of left truncated multivariate modified power series distribution. Metrika, 26(1):87–94, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893475.

Schlittgen:1979:UMT

[521] R. Schlittgen. Use of a median test for a generalized Behrens–Fisher problem. Metrika, 26(1):95–103, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893476.

Dallas:1979:EL

[522] Dr. A. C. Dallas. On the exponential law. Metrika, 26(1):105–108, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893477.

Sendler:1979:SIC

[523] W. Sendler. On statistical inference in concentration measurement. Metrika, 26(1):109–122, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01893478.

Anonymous:1979:A

[524] Anonymous. Announcement. Metrika, 26(1):123–124, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF01893479.

Wilrich:1979:BGB

[525] P.-Th. Wilrich, H. Fleischner, G. Pflug, F. Landler, W. Fauni, H. Strelec, and W. Polasek. Buchbesprechungen. (German) [Book reviews]. Metrika, 26 (1):124–127, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893480.

Grossmann:1979:EBT

[526] Dr. W. Grossmann. Einige Bemerkungen zur Theorie der Maximum Probability Schätzer. (German) []. Metrika, 26(1):129–137, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893481.

Pflug:1979:SSA

[527] Dr. G. Pflug. Stetige stochastische approximation. Metrika, 26(1):139–150, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893482.

Plachky:1979:BUD

[528] Prof. Dr. D. Plachky. Bestimmung ungünstigster Dichten durch Minimierung von Abständen bei Verteilungen. (German) []. Metrika, 26(1): 151–155, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893483.

Wertz:1979:NSG

[529] Prof. Dr. W. Wertz. Über ein nichtparametrisches Schätzproblem. (German) []. Metrika, 26(1):157–167, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893484.

Anonymous:1979:HC

[530] Anonymous. Help & contacts. Metrika, 26(1):??, ???? 1979. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Oberhofer:1980:NMS

[531] Prof. Dr. W. Oberhofer. Die Nichtkonsistenz der M.-L. Schätzer im “Switching Regression” Problem. (German) [The consistency of the non-M.-L. estimator in the “switching regression” problem]. Metrika, 27(1):1–13, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893572.

Irle:1980:LBT

[532] Dr. A. Irle. Locally best tests for Gaussian processes. Metrika, 27(1):15–28, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893573.

Beg:1980:EPX

[533] M. A. Beg. On the estimation of PrY<X for the two-parameter exponential distribution. Metrika, 27 (1):29–34, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01893574.

Kohn:1980:LIA

[534] R. Kohn. Local identification of AR-MAX structures subject to nonlinear constraints. Metrika, 27(1):35–41, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893575.

Pollard:1980:MDM

[535] D. Pollard. The minimum distance method of testing. Metrika, 27 (1):43–70, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893576.

Neuwirth:1980:BR

[536] E. Neuwirth, F. Ferschl, and G. Pflug. Book reviews. Metrika, 27(1):71–72, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01893577.

Ruger:1980:AZE

[537] Priv.-Doz. Dr. B. Rüger. Alternativtests für zwei einfache Hypothesen mit endlich vielen Ergebnissen. (German) [Alternative tests for two simple hypotheses with finitely many results]. Metrika, 27(1):73–90, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893578.

Kannappan:1980:MTI

[538] Pl. Kannappan. A mixed theory of information- IV: Inset-inaccuracy and directed divergence. Metrika, 27(1):91–98, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893579.

Srivastava:1980:NEM

[539] V. K. Srivastava. A note on the estimation of mean in normal population. Metrika, 27(1):99–102, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893580.

Pukelsheim:1980:MES

[540] F. Pukelsheim. Multilinear estimation of skewness and kutosis in linear models. Metrika, 27(1):103–113, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893581.

Talwalker:1980:PDP

[541] Dr. S. Talwalker. On the property of dullness of Pareto distribution. Metrika, 27(1):115–119, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893582.

Srinath:1980:EVM

[542] Dr. K. P. Srinath and Dr. M. A. Hidiroglou. Estimation of variance in multi-stage sampling. Metrika, 27 (1):121–125, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01893583.

Schader:1980:HAC

[543] M. Schader. Hierarchical analysis: Classification with ordinal object dissimilarities. Metrika, 27(1):127–132, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893584.

Baxter:1980:MVU

[544] M. A. Baxter. Minimum variance unbiased estimation of the parameters of the Pareto distribution. Metrika, 27(1):133–138, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893585.

Anonymous:1980:A

[545] Anonymous. Announcement. Metrika, 27(1):139, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01893587.

Anonymous:1980:IN

[546] Anonymous. Important notice. Metrika, 27(1):139, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
accesspage/article/10.1007/BF01893588.

Dallas:1980:A

[547] Ag. Dallas. Acknowledgement. Metrika, 27(1):139, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
accesspage/article/10.1007/BF01893586.

Morgenstern:1980:BR

[548] D. Morgenstern, W. Grossmann, W. Polasek, G. Feichtinger, G. Pflug, A. Kalliauer, and J. Gordesch. Book reviews. Metrika, 27(1):140–143, ????

1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893589.

Anonymous:1980:F

[549] Anonymous. Fehlerhinweis. Metrika, 27(1):143, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF01893590.

Yaqub:1980:DDF

[550] M. Yaqub and A. H. Khan. On the distribution of a distance function on the sphere. Metrika, 27(1):145–151, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893591.

Anonymous:1980:HC

[551] Anonymous. Help & contacts. Metrika, 27(1):??, ???? 1980. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Chaudhuri:1981:NNV

[552] Dr. A. Chaudhuri and R. Arnab. On non-negative variance-estimation. Metrika, 28(1):1–12, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF01902872.

Gebhardt:1981:SEH

[553] Dr. D. Gebhardt. On the statistical evaluation of hit patterns. Metrika, 28(1):13–21, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01902873.

Jahn:1981:BSG

[554] J. Jahn. Zur Bewertung von Schätzungen bei der gewöhnlichen Regressionsanalyse. (German) [On the evaluation of estimates in ordinary regression analysis]. Metrika, 28 (1):23–33, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01902874.

Janssen:1981:RCC

[555] P. Janssen. Rate of convergence in the Central Limit Theorem and in the Strong law of Large Numbers for von Mises statistics. Metrika, 28(1):35–46, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902875.

Garrelfs:1981:RVB

[556] Dr. H. Garrelfs. Rangvariable bei Vorliegen von Bindungen. (German) [Rank in the presence of variable bindings]. Metrika, 28(1):47–52, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902876.

Pereira:1981:CSM

[557] B. de B. Pereira. Choice of a survival model for patients with a brain tumour. Metrika, 28(1):53–61, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902877.

vonChossy:1981:ISK

[558] R. von Chossy and U. G. Oppel. Eine Invarianzeigenschaft für Startverteilungen einer Klasse stochastischer Prozesse. (German) [An invariance property for a class of stochastic processes starting distributions]. Metrika, 28(1):63–69, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01902878.

VonRuger:1981:SUO

[559] Dr. B. Von Rüger. Scharfe untere und obere Schranken für die Wahrscheinlichkeit der Realisation von k unter n Ereignissen. (German) [Sharp lower and upper bounds for the probability of realization of k in n events]. Metrika, 28(1):71–77, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01902879.

Pfanzagl:1981:SRS

[560] Prof. Dr. J. Pfanzagl. A special representation of a sufficient randomization kernel. Metrika, 28(1):79–81, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01902880.

Charalambides:1981:BGB

[561] Ch. A. Charalambides and H. Papageorgiou. On bivariate generalized binomial and negative binomial distributions. Metrika, 28(1):83–92, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01902881.

Behl:1981:MPA

[562] Dipl.-Math. M. Behl. Die messende Prüfung bei Abweichung von der Normalverteilungsannahme. (German) [The measurement test for deviation from the normal distribution assumption]. Metrika, 28(1):93–107, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902882.

Konijn:1981:BVC

[563] H. S. Konijn. Biases, variances and covariances of raking ratio estimators for marginal and cell totals and averages of observed characteristics. Metrika, 28(1):109–121, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902883.

Sekkappan:1981:SBM

[564] R. M. Sekkappan. Subjective Bayesian multivariate stratified sampling from finite populations. Metrika, 28(1): 123–132, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01902884.

Mukhopadhyay:1981:SAS

[565] Asst. Prof. Dr. N. Mukhopadhyay. Sequential approach to simultaneous estimation of the mean and variance. Metrika, 28(1):133–136, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01902885.

Ferschl:1981:BR

[566] F. Ferschl and F. Eicker. Book reviews. Metrika, 28(1):137, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF01902886.

Kardos:1981:GBF

[567] Dr. P. Kardos. On generalized Bellman’s functional equation. Metrika, 28(1):139–149, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01902887.

Dallas:1981:CUC

[568] Dr. A. C. Dallas. A characterization using the conditional variance. Metrika, 28(1):151–153, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01902888.

Gawronski:1981:SHM

[569] W. Gawronski and U. Stadtmüller. Smoothing histograms by means of lattice-and continuous distributions. Metrika, 28(1):155–164, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF01902889.

Sen:1981:IPP

[570] P. K. Sen and Y. Tsong. An invariance principle for progressively truncated likelihood ratio statistics. Metrika, 28(1):165–177, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01902890.

Schneeberger:1981:POS

[571] Prof. Dr. H. Schneeberger. The problem of optimum stratification and allocation with q = n/N > 0. Metrika, 28(1):179–189, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01902891.

Anonymous:1981:HC

[572] Anonymous. Help & contacts. Metrika, 28(1):??, ???? 1981. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Subrahmanya:1982:CIB

[573] M. T. Subrahmanya. The combined inverse binomial sampling procedure. Metrika, 29(1):1–17, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893362.

Sendler:1982:FOS

[574] Dr. W. Sendler. On functionals of order statistics. Metrika, 29(1):19–54, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893363.

Findeisen:1982:CNN

[575] Dr. P. Findeisen. Die Charakterisierung der Normalverteilung nach Gauß. (German) [The characterization of the Gaussian normal distribution]. Metrika, 29(1):55–63, ????

1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893364.

Heyer:1982:BR

[576] H. Heyer, F. Ferschl, A. C. Harvey, P.- Th. Wilrich, W. Grossmann, G. Pflug, A. Linder, and K. D. Schmidt. Book reviews. Metrika, 29(1):65–70, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893365.

Ahsan:1982:OAM

[577] M. J. Ahsan and S. U. Khan. Optimum allocation in multivariate stratified random sampling with overhead cost. Metrika, 29(1):71–78, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893366.

Stadje:1982:SRA

[578] Dr. W. Stadje. Some remarks on the average sample number of sequential and non-sequential tests. Metrika, 29(1):79–86, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893367.

Grant:1982:SPI

[579] Dr. P. Grant. Secretary problems with inspection costs as a game. Metrika, 29(1):87–93, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893368.

Findeisen:1982:CZE

[580] Dr. P. Findeisen. Charakterisierung der zweiseitigen Exponentialverteilung. (German) [Characterization of the two-sided exponential distribution]. Metrika, 29(1):95–102, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF01893369.

Beg:1982:OTE

[581] M. A. Beg. Optimal tests and estimators for truncated exponential families. Metrika, 29(1):103–113, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893370.

Zinzius:1982:UVE

[582] Dr. E. Zinzius. Ungünstige Verteilungen bei einparametrigen Parameterschätzproblemen. (German) [Unfavorable distributions in one-parameter parameter estimation problems]. Metrika, 29(1):115–128, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893371.

Bhattacharya:1982:AII

[583] Prof. Dr. S. K. Bhattacharya and G. S. Pandey. Applications of an inequality involving conditional expectations. Metrika, 29(1):129–133, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893372.

Henze:1982:VCQ

[584] N. Henze. Verhalten des Chi– Quadrat-Tests für Prüfung der Gleichwahrscheinlichkeit der Lotto–Zahlen bei Nichtgültigkeit der Hypothese. (German) [Behavior of the chi-square test for examination of equal probability of lotto-numbers with nonvalidity of the hypothesis]. Metrika, 29 (1):135–139, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893373.

Dutter:1982:BR

[585] R. Dutter, Ch. Mandl, W. Polasek, and J. Gordesch. Book reviews. Metrika, 29(1):141–142, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01893374.

Mukhopadhyay:1982:OSE

[586] P. Mukhopadhyay. Optimum strategies for estimating the variance of a finite population under a superpopulation model. Metrika, 29(1):143–158, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893375.

Kremer:1982:LCL

[587] Dr. E. Kremer. Local comparison of linear rank tests, in the Bahadur sense. Metrika, 29(1):159–173, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893376.

Sengupta:1982:ISU

[588] S. Sengupta. On interpenetrating samples of unequal sizes. Metrika, 29 (1):175–188, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01893377.

Panaretos:1982:EDM

[589] J. Panaretos. An extension of the damage model. Metrika, 29(1):189–194, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893378.

Strauss:1982:AEF

[590] I. Strauss. On the admissibility of estimators for the finite population variance. Metrika, 29(1):195–202, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893379.

Wong:1982:CTS

[591] Dr. P. G. Wong and Dr. S. P. Wong. A curtailed test for the shape parameter of the Weibull distribution. Metrika, 29(1):203–209, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01893380.

Feichtinger:1982:BR

[592] G. Feichtinger, E. J. Hannan, K. D. Schmidt, P. Mandl, and L. E. Moses. Book reviews. Metrika, 29(1):211–214, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01893381.

Anonymous:1982:HC

[593] Anonymous. Help & contacts. Metrika, 29(1):??, ???? 1982. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

George:1983:CLR

[594] E. O. George and G. S. Mudholkar. On the convolution of logistic random variables. Metrika, 30(1):1–13, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02056895.

Kern:1983:MVU

[595] D. M. Kern. Minimum variance unbiased estimation in the Pareto distribution. Metrika, 30(1):15–19, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02056896.

Ahmad:1983:CNE

[596] Professor I. A. Ahmad and Professor P.-E. Lin. Consistency of a nonparametric estimation of a density functional. Metrika, 30(1):21–29, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02056897.

Stadje:1983:SPG

[597] Dr. W. Stadje. Summen von Produkten gemeinsam normalverteilter Zufallsvariablen. (German) [Sums of products of jointly normally distributed random variables]. Metrika, 30(1):31–36, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02056898.

Theodorescu:1983:SPD

[598] Prof. Dr. R. Theodorescu and Prof. R. L. Tweedie. Solidarity properties and a Doeblin decomposition for a class of non-Markovian stochastic processes. Metrika, 30(1):37–47, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02056899.

Bondesson:1983:UMV

[599] L. Bondesson. On uniformly minimum variance unbiased estimation when no complete sufficient statistics exist. Metrika, 30(1):49–54, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02056900.

Ruschendorf:1983:SSO

[600] Dr. L. Rüschendorf. Solution of a statistical optimization problem by rearrangement methods. Metrika, 30 (1):55–61, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02056901.

Kalin:1983:BR

[601] D. Kalin, H. Witting, H. G. Seifert, Prof. Dr. H. Schneeberger, F. Eicker, G. Rothenberger, H. Schoch, O. Richter, H. Rost, P. A. Meyer, G. Leha, E. Dettweiler, H. Heyer, E. Siebert, H. Rieder, and et al. Book reviews. Metrika, 30(1):63–72, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02056902.

Rothe:1983:BRS

[602] Dr. G. Rothe. Block rank statistics. Metrika, 30(1):73–83, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02056903.

Heyer:1983:BR

[603] H. Heyer. Book reviews. Metrika, 30 (1):84, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF02056904.

Roebruck:1983:OTL

[604] Dr. P. Roebruck. Zur Optimalität des F -Testes für lineare Hypothesen über lineare Modelle mit gemischten Effekten. (German) [On the optimality of F -test for linear hypotheses about linear mixed-effects models]. Metrika, 30(1):85–91, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02056905.

Gotze:1983:BR

[605] F. Götze and D. Plachky. Book reviews. Metrika, 30(1):92, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02056906.

Srivastava:1983:PEF

[606] S. K. Srivastava. Predictive estimation of finite population mean using product estimator. Metrika, 30(1):93–99, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02056907.

Heyer:1983:SRP

[607] H. Heyer. Statistics of random processes. I: General theory. Metrika, 30(1):100, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02056908.

Hayre:1983:SED

[608] L. S. Hayre. Sequential estimation of the difference between the means of two normal populations. Metrika, 30(1):101–107, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02056909.

Ryan:1983:DTI

[609] M. J. Ryan. Decision theory and incomplete knowledge. Metrika, 30(1): 108, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF02056910.

Glazebrook:1983:SBP

[610] K. D. Glazebrook and D. M. Jones. Some best possible results for a discounted one armed bandit. Metrika, 30(1):109–115, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02056911.

Sendler:1983:BR

[611] W. Sendler and D. Morgenstern. Book reviews. Metrika, 30(1):116, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02056912.

Kalin:1983:SRC

[612] Prof. Dr. D. Kalin and Prof. Dr. R. Theodorescu. On a stopping rule for a class of sequential decision problems. Metrika, 30(1):117–123, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02056913.

Louton:1983:BR

[613] T. Louton and R. Schassberger. Book reviews. Metrika, 30(1):124, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF02056914.

Anonymous:1983:HC

[614] Anonymous. Help & contacts. Metrika, 30(1):??, ???? 1983. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Prabhu-Ajgaonkar:1984:NNB

[615] S. G. Prabhu-Ajgaonkar. A note on necessary best estimator of order two. Metrika, 31(1):1–4, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01915176.

Katzenbeisser:1984:AVP

[616] Dr. W. Katzenbeisser. Die asymptotische Verteilung des Prognosefehlers bei Prognosen mit Hilfe eines dynamischen ökonometrischen Modells höherer als erster Ordnung. (German) [The asymptotic distribution of the forecast error in predictions using a dynamic econometric model of higher than first order]. Metrika, 31(1):5–23, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01915177.

Schaich:1984:BR

[617] E. Schaich and D. Morgenstern. Book reviews. Metrika, 31(1):24, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01915178.

Rao:1984:SAR

[618] T. J. Rao. Some aspects of random permutation models in finite population sampling theory. Metrika, 31(1):25–32, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01915179.

Ketellapper:1984:REM

[619] R. H. Ketellapper and A. E. Ronner. Are robust estimation methods useful in the structural errors-in-variables model? Metrika, 31(1):33–41, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01915180.

Krafft:1984:BR

[620] O. Krafft and H. Heyer. Book reviews. Metrika, 31(1):42, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF01915181.

Niemi:1984:ISA

[621] H. Niemi. The invertibility of sampled and aggregated ARMA models. Metrika, 31(1):43–50, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01915182.

Stormer:1984:BSJ

[622] Prof. Dr. H. Störmer. Bemerkungen zu den Sätzen von Jordan. (German) [Comments on Jordan sets]. Metrika, 31(1):51–59, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01915183.

Strasser:1984:BR

[623] H. Strasser and P. Gänssler. Book reviews. Metrika, 31(1):60–62, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01915184.

Kabe:1984:CSA

[624] D. G. Kabe. Classical statistical analysis based on a certain hypercomplex multivariate normal distribution. Metrika, 31(1):63–76, ????

1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01915185.

Saksena:1984:BUE

[625] Ass. Prof. S. K. Saksena and Prof. A. M. Johnson. Best unbiased estimators for the parameters of a two-parameter Pareto distribution. Metrika, 31(1):77–83, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF01915186.

Schneeweiss:1984:SAW

[626] H. Schneeweiß. Studien zur angewandten Wirtschaftsforschung und Statistik. (German) [Studies on applied economic research and statistics]. Metrika, 31(1):84, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01915187.

Sharma:1984:EVG

[627] D. Sharma. On estimating the variance of a generalized Laplace distribution. Metrika, 31(1):85–88, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01915188.

Padmawar:1984:ESF

[628] V. R. Padmawar and P. Mukhopadhyay. Estimation of symmetric functions of a finite population. Metrika, 31(1):89–97, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01915189.

Husler:1984:ERP

[629] J. Hüsler. Extremes and related properties of random sequences and processes. Metrika, 31(1):98, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF01915190.

Landers:1984:NQC

[630] Prof. Dr. D. Landers and Prof. Dr. L. Rogge. Natural α-quantiles and conditional α-quantiles. Metrika, 31 (1):99–113, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01915191.

Haslinger:1984:ETN

[631] F. Haslinger. Economic theory of natural resources. Metrika, 31(1):114, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF01915192.

Charalambides:1984:MVU

[632] Priv.-Doz. Dr. Ch. A. Charalambides. Minimum variance unbiased estimation for the zero class truncated bivariate Poisson and logarithmic series distributions. Metrika, 31(1):115–123, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01915193.

Georgii:1984:BR

[633] H.-O. Georgii and F. Pukelsheim. Book reviews. Metrika, 31(1):124–126, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01915194.

Kapadia:1984:MSS

[634] C. H. Kapadia and D. L. Weeks. Minimal sufficient statistics for the group divisible partially balanced incomplete block design (GD-PBIBD) with interaction under an Eisenhart Model II. Metrika, 31(1):127–144, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01915195.

Anonymous:1984:HC

[635] Anonymous. Help & contacts. Metrika, 31(1):??, ???? 1984. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Arnold:1985:KPE

[636] Dr. B. Arnold. Kostenoptimale Prüfpläne bei einem Problem mit drei möglichen Entscheidungen. (German) [Costa-optimized test plans for a problem with three possible decisions]. Metrika, 32(1):1–14, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01897793.

Holm:1985:ODS

[637] S. Holm. On the optimality of differentiated SPR tests of composite hypotheses. Metrika, 32(1):15–33, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01897794.

Uhlmann:1985:FSQ

[638] W. Uhlmann. Frontiers in statistical quality control 2. Metrika, 32(1):34, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF01897795.

Ronning:1985:NXR

[639] Prof. Dr. Gerd Ronning. On the nonnegativity of XX^{+} and its relevance in econometrics. Metrika, 32(1):35–47, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01897796.

Edgar:1985:ASF

[640] G. A. Edgar. Amarts and set function processes. Metrika, 32(1):48, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF01897797.

Saleh:1985:LSE

[641] A. K. Md. E. Saleh, M. M. Ali, and D. Umbach. Large sample estimation of Pareto quantiles using selected order statistics. Metrika, 32(1):49–56, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01897798.

Sonnemann:1985:ORC

[642] E. Sonnemann. U-optimum rowcolumns designs for the comparison of two treatments. Metrika, 32(1):57–63, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01897799.

Prabhu-Ajgaonkar:1985:NNB

[643] S. G. Prabhu-Ajgaonkar. A note on necessary best estimator of order two. Metrika, 32(1):64, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01897800.

Engel:1985:SCD

[644] J. Engel. Some characterizations of distributions by regression models for ordinal response data. Metrika, 32 (1):65–72, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01897801.

Gajek:1985:RLU

[645] L. Gajek. A remark on Lehmannunbiased estimators for scale resp. location parameters. Metrika, 32(1):73–84, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01897802.

Gupta:1985:TFP

[646] A. K. Gupta and C. F. Wong. On three and five parameter bivariate beta distributions. Metrika, 32(1):85–91, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01897803.

Reschenhofer:1985:RRE

[647] E. Reschenhofer. A remark on the recursive estimation of ARMA order. Metrika, 32(1):93–96, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01897804.

Ronner:1985:OOE

[648] A. E. Ronner and A. G. M. Steerneman. The occurrence of outliers in the explanatory variable considered in an errors-in-variables framework. Metrika, 32(1):97–107, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01897805.

Terpstra:1985:OSM

[649] T. J. Terpstra. One-sided and multisided rank tests against trend in the case of concordant observers. Metrika, 32(1):109–123, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01897806.

Hey:1985:BR

[650] John D. Hey, W. Wetzel, M. Metivier, D. M. Chibisov, and V. K. Malinovski. Book reviews. Metrika, 32(1):124–128, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01897807.

Potscher:1985:BLM

[651] B. M. Pötscher. The behaviour of the Lagrangian multiplier test in testing the orders of an ARMAmodel. Metrika, 32(1):129–150, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01897808.

Kramer:1985:AVE

[652] Dr. W. Krämer. Asymptotische Verteilungen einiger Schätzverfahren bei Trend und Fehlern in den Variablen. (German) [Asymptotic distributions of some estimations in trend and errors in variables]. Metrika, 32 (1):151–162, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01897809.

Morgenstern:1985:EBI

[653] Prof. Dr. D. Morgenstern. Einfache Begündung für die Irrtumswahrscheinlichkeitsschranke der HOMMEL’schen Mehrheitstestkombination. (German) [Simple basis for imploring the probability of error bounds of the majority Hommel test combination]. Metrika, 32(1):163–164, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01897810.

Vandemaele:1985:CTL

[654] M. Vandemaele and N. Veraverbeke. Cramér type large deviations for Studentized U-statistics. Metrika, 32 (1):165–179, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01897811.

Linhart:1985:CSM

[655] H. Linhart and P. Volkers. On a criterion for the selection of models for stationary time series. Metrika, 32 (1):181–196, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01897812.

Anonymous:1985:HC

[656] Anonymous. Help & contacts. Metrika, 32(1):??, ???? 1985. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Lee:1986:BNC

[657] P. A. Lee and S. H. Ong. The bivariate non-central negative binomial distributions. Metrika, 33(1):1–28, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01894723.

Ong:1986:BNC

[658] S. H. Ong and P. A. Lee. Bivariate non-central negative binomial distribution: Another generalisation. Metrika, 33(1):29–46, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01894724.

Lee:1986:RSP

[659] Y. J. Lee and E. J. Dudewicz. Robust selection procedures based on vector ranks. Metrika, 33(1):47–63, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01894725.

Schmitz:1986:BR

[660] N. Schmitz and D. W. Stroock. Book reviews. Metrika, 33(1):64, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01894726.

Sharma:1986:ICW

[661] D. Sharma and K. Krishnamoorthy. An identity concerning a Wishart matrix. Metrika, 33(1):65–68, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01894727.

Laake:1986:OEO

[662] P. Laake. Optimal estimates and optimal predictors of finite population characteristics in the presence of nonresponse. Metrika, 33(1):69–77, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01894728.

Priouret:1986:ITL

[663] P. Priouret. An introduction to the theory of large deviations. Metrika, 33(1): 78, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF01894729.

Arnab:1986:OPF

[664] Dr. R. Arnab. Optimal prediction for a finite population total with connected designs and related model-based results. Metrika, 33(1):79–84, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01894730.

Takeuchi:1986:NMV

[665] K. Takeuchi and M. Akahira. A note on minimum variance. Metrika, 33(1):85–91, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01894731.

Singh:1986:SUP

[666] M. P. Singh. Sampling with unequal probabilities. Metrika, 33(1):92, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF01894732.

Arnold:1986:PDO

[667] Dr. B. F. Arnold. Procedures to determine optimum two-stage sampling plans by attributes. Metrika, 33(1):93–109, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01894733.

Loynes:1986:SEA

[668] R. M. Loynes. Statistical extremes and applications. Metrika, 33(1):100, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF01894734.

Eberl:1986:CNT

[669] W. Eberl, Jr. Characterizations of normality in translation classes by properties of Bayes estimators. Metrika, 33(1):111–119, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01894735.

Hackl:1986:CMP

[670] P. Hackl. Computational methods for parsimonious data fitting. Metrika, 33(1):120, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01894736.

Bauer:1986:MTP

[671] P. Bauer, P. Hackl, G. Hommel, and E. Sonnemann. Multiple testing of pairs of one-sided hypotheses. Metrika, 33(1):121–127, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01894737.

Follmer:1986:BR

[672] H. Föllmer and K. Wittkowski. Book reviews. Metrika, 33(1):128, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01894738.

Mukhopadhyay:1986:EUL

[673] P. Mukhopadhyay. Estimation under linear regression models. Metrika, 33(1):129–134, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01894739.

Viollaz:1986:RCS

[674] Dr. A. J. Viollaz. On the reliability of the chi-square test. Metrika, 33(1):135–142, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01894740.

Mader:1986:KPQ

[675] Dr. U. Mäder. Kostenoptimale Prüfpläne für ein quantitatives Merkmal. (German) [Cost-optimized test plans for a quantitative characteristic]. Metrika, 33(1):143–163, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01894741.

Hackl:1986:BR

[676] P. Hackl and W. Hansen. Book reviews. Metrika, 33(1):164, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        BF01894742.

Anonymous:1986:HC

[677] Anonymous. Help & contacts. Metrika, 33(1):??, ???? 1986. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Akahira:1987:LMV

[678] M. Akahira and K. Takeuchi. Locally minimum variance unbiased estimator in a discontinuous density function. Metrika, 34(1):1–15, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613124.

Wiegert:1987:BR

[679] R. Wiegert. Book review. Metrika, 34 (1):16, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF02613125.

Koziol:1987:AFN

[680] Prof. J. A. Koziol, Ph. D. An alternative formulation of Neyman’s smooth goodness of fit tests under composite alternatives. Metrika, 34(1):17–24, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613126.

Goldmann:1987:REN

[681] Dr. Werner Goldmann. Risk-efficient nonparametric sequential estimators. Metrika, 34(1):25–30, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF02613127.

Knusel:1987:AEP

[682] L. Knüsel and J. Michalk. Asymptotic expansion of the power function of the two-sample binomial test with and without randomization. Metrika, 34(1):31–44, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613128.

Spruill:1987:OED

[683] M. C. Spruill. Optimal extrapolation of derivatives. Metrika, 34(1):45–60, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613129.

Gotze:1987:BR

[684] F. Götze, H. Stenger, U. Dieter, and J. W. Thompson. Book reviews. Metrika, 34(1):61–64, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613130.

Begun:1987:ERR

[685] Dr. J. M. Begun. Estimates of relative risk. Metrika, 34(1):65–82, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613131.

Mukhopadhyay:1987:TSE

[686] N. Mukhopadhyay and A. Mauromoustakos. Three-stage estimation procedures for the negative exponential distributions. Metrika, 34(1):83–93, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613132.

Borges:1987:BR

[687] R. Borges. Book review. Metrika, 34 (1):94, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF02613133.

Bagchi:1987:OCA

[688] S. Bagchi. On the E-optimality of certain asymmetrical designs under mixed effects model. Metrika, 34(1):95–105, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613134.

Ledrappier:1987:BR

[689] F. Ledrappier. Book review. Metrika, 34(1):106, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613135.

Irle:1987:LEB

[690] Prof. Dr. A. Irle. Largest excess of boundary crossings for martingales. Metrika, 34(1):107–115, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613136.

Jacod:1987:BR

[691] J. Jacod. Book review. Metrika, 34(1):116, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613137.

Dupac:1987:QIR

[692] V. Dupac. Quasi-isotonic regression and stochastic approximation. Metrika, 34(1):117–123, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613138.

Beran:1987:BR

[693] R. Beran, S. Schach, and A. Dress. Book reviews. Metrika, 34(1):124–127, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613139.

Sandstrom:1987:ANL

[694] Dr. Arne Sandström. Asymptotic normality of linear functions of concomitants of order statistics. Metrika, 34 (1):129–142, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613140.

Konecny:1987:APM

[695] Dr. Franz Konecny. The asymptotic properties of maximum likelihood estimators for marked Poisson processes with a cyclic intensity measure. Metrika, 34(1):143–155, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613141.

Takacs:1987:BR

[696] L. Takács. Book review. Metrika, 34(1):156, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613142.

Zamudio:1987:EMP

[697] F. Zamudio and Prof. Dr. Malay Ghosh. Estimation of multiple Poisson means: A compromise between maximum likelihood and empirical Bayes estimators. Metrika, 34(1):157–166, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613143.

Anonymous:1987:HC

[698] Anonymous. Help & contacts. Metrika, 34(1):??, ???? 1987. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Bagchi:1988:CNB

[699] S. Bagchi. A class of non binary unequally replicated E-optimal designs. Metrika, 35(1):1–12, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613280.

Chaubey:1988:AUE

[700] Y. P. Chaubey and B. Singh. Almost unbiased estimation in multiplicative models. Metrika, 35(1):13–28, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613281.

Padmawar:1988:CCC

[701] V. R. Padmawar. On the completeness of certain classes of two stage sampling strategies. Metrika, 35(1):29–40, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613282.

Hamedani:1988:TRC

[702] Prof. G. G. Hamedani. On two recent characterizations of multivariate normal distribution. Metrika, 35(1):41–47, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613283.

Kurotschka:1988:BR

[703] V. Kurotschka. Book review. Metrika, 35(1):48, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613284.

Grubel:1988:MCC

[704] R. Grübel. A minimal characterization of the covariance matrix. Metrika, 35(1):49–52, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613285.

Sengupta:1988:CBP

[705] Prof. Dr. S. Sengupta. A comparison between PPSWR and Chaudhuri’s IPPS procedures. Metrika, 35(1):53–57, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613286.

Wertz:1988:BR

[706] W. Wertz, V. Barnett, Sándor Csörgő, and K. V. Mardia. Book reviews. Metrika, 35(1):58–62, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF02613287.

Horalek:1988:SAT

[707] V. Horálek and V. Benes. On a special application of the truncation problem. Metrika, 35(1):63–76, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613288.

Stenger:1988:AEM

[708] Prof. Dr. H. Stenger. Asymptotic expansion of the minimax value in survey sampling. Metrika, 35(1):77–92, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613289.

Stadje:1988:GML

[709] W. Stadje. A generalized maximum likelihood characterization of the normal distribution. Metrika, 35(1):93–97, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613290.

vonCollani:1988:BR

[710] E. von Collani and W. Stute. Book reviews. Metrika, 35(1):98, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF02613291.

Arnold:1988:CCM

[711] B. C. Arnold and M. Pourahmadi. Conditional characterizations of multivariate distributions. Metrika, 35 (1):99–108, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613292.

Shanubhogue:1988:DFT

[712] A. Shanubhogue. Distribution-free test for homogeneity against stochastic ordering. Metrika, 35(1):109–119, ????

1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613293.

Schmetterer:1988:BR

[713] L. Schmetterer, P. Ressel, H. Heyer, and H. Basler. Book reviews. Metrika, 35(1):120–125, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613294.

Walter:1988:EBE

[714] Prof. G. G. Walter and Prof. G. G. Hamedani. Empiric Bayes estimation of hypergeometric probability. Metrika, 35(1):127–143, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613295.

Meyer:1988:BR

[715] P. A. Meyer. Book review. Metrika, 35(1):144, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02613296.

vonCollani:1988:UAO

[716] PD Dr. E. von Collani. A unified approach to optimal process control. Metrika, 35(1):145–159, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613297.

Rao:1988:BR

[717] C. R. Rao. Book review. Metrika, 35(1):160, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613298.

Yanagimoto:1988:CML

[718] T. Yanagimoto. The conditional maximum likelihood estimator of the shape parameter in the gamma distribution. Metrika, 35(1):161–175, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613299.

Anonymous:1988:HC

[719] Anonymous. Help & contacts. Metrika, 35(1):??, ???? 1988. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Gupta:1989:APD

[720] Sudhir Gupta and Murari Singh. Analysis of PBIB designs using association matrices. Metrika, 36(1):1–6, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614072.

PhD:1989:ITS

[721] Tatsuya Kubokawa PhD. Improving on two-stage estimators for scale families. Metrika, 36(1):7–13, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614073.

Turnwald:1989:BR

[722] G. Turnwald. Book review. Metrika, 36 (1):14, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF02614074.

Gupta:1989:TDR

[723] A. K. Gupta and W. R. Javier. On testing the dimensionality of regression coefficients. Metrika, 36(1):15–27, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02614075.

Schneeweiss:1989:BR

[724] H. Schneeweiß. Book review. Metrika, 36(1):28, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02614076.

Gaffke:1989:CPA

[725] Norbert Gaffke and Rudolf Mathar. A cyclic projection algorithm via duality. Metrika, 36(1):29–54, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02614077.

Chaudhuri:1989:ERE

[726] Dr. Arijit Chaudhuri and Dr. Arun Kumar Adhikary. On efficiency of the ratio estimator. Metrika, 36(1):55–59, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02614078.

Strecker:1989:BR

[727] H. Strecker and F. Mosteller. Book reviews. Metrika, 36(1):60–62, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614079.

Terpstra:1989:AON

[728] Prof. Dr. T. J. Terpstra. Asymptotically optimal nonparametric oneand two-way analysis of variance tests for the log linear model under censoring. Metrika, 36(1):63–90, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614080.

Stevens:1989:NMC

[729] Prof. Gary Stevens. A nonparametric multiple comparison test for differences in scale parameters. Metrika, 36(1):91–106, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02614081.

Csorgo:1989:CST

[730] Prof. Sándor Csörgő. Consistency of some tests for multivariate normality. Metrika, 36(1):107–116, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614082.

Pham:1989:PSM

[731] Tuan D. Pham and Professor Dr. Lanh T. Tran. Predicting the sample mean by extreme order statistics. Metrika, 36(1):117–125, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614083.

Kirchgassner:1989:BR

[732] G. Kirchgässner, M. Csörgő, and L. Rüschendorf. Book reviews. Metrika, 36(1):126–128, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF02614084.

Gil:1989:FMU

[733] Maria Angeles Gil, Rigoberto Perez, and Prof. Dr. Pedro Gil. A family of measures of uncertainty involving utilities: Definition, properties, applications and statistical inferences. Metrika, 36(1):129–147, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02614085.

Bandemer:1989:BR

[734] H. Bandemer. Book review. Metrika, 36(1):148, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/BF02614086.

Chen:1989:MPF

[735] Lanxiang Chen. Monotonicity of the power function and unbiasedness of some likelihood ratio tests. Metrika, 36(1):149–159, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02614087.

Denker:1989:BR

[736] M. Denker. Book review. Metrika, 36(1):160, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02614088.

Adhikary:1989:TPU

[737] Dr. Arun Kumar Adhikary and Dr. Arijit Chaudhuri. On two properties of an unequal probability sampling scheme. Metrika, 36(1):161–166, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614089.

Dikta:1989:SFW

[738] G. Dikta, B. Kurtz, and Prof. Dr. W. Stute. Sequential fixed-width confidence bands for distribution functions under random censoring. Metrika, 36 (1):167–176, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02614090.

Abdous:1989:SUC

[739] Dr. B. Abdous and Prof. Dr. R. Theodorescu. On the strong uniform consistency of a new kernel density estimator. Metrika, 36(1):177–194, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614091.

Anonymous:1989:HC

[740] Anonymous. Help & contacts. Metrika, 36(1):??, ???? 1989. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Chen:1990:GME

[741] L. Chen, J. Eichenauer-Herrmann, and J. Lehn. Gamma-minimax estimation of a multivariate normal mean. Metrika, 37(1):1–6, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613500.

Henze:1990:ALD

[742] N. Henze. An approximation to the limit distribution of the epps-pulley test statistic for normality. Metrika, 37(1):7–18, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613501.

Al-Mahmeed:1990:SEL

[743] Dr. M. Al-Mahmeed and Dr. H. I. Hamdy. Sequential estimation of linear models in three stages. Metrika, 37(1):19–36, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613502.

Beg:1990:DDC

[744] Prof. M. I. Beg and K. Balasubramanian. Distributions determined by conditioning on a single order statistic. Metrika, 37(1):37–43, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613503.

Mau:1990:BR

[745] J. Mau. Book review. Metrika, 37(1): 44, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF02613504.

Ahmad:1990:TWM

[746] I. A. Ahmad and S. C. Kochar. Testing whether F is more IFR than G. Metrika, 37(1):45–58, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613505.

Heyer:1990:BRa

[747] H. Heyer and J. de Leeuw. Book reviews. Metrika, 37(1):59–62, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613506.

Prayag:1990:DER

[748] V. R. Prayag and Dr. A. P. Gore. Density estimation for randomly distributed circular objects. Metrika, 37(1):63–69, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613507.

Heyer:1990:BRb

[749] H. Heyer. Book review. Metrika, 37 (1):70, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF02613508.

Ghorai:1990:KES

[750] Prof. Dr. J. K. Ghorai and V. Susarla. Kernel estimation of a smooth distribution function based on censored data. Metrika, 37(1):71–86, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613509.

Gabler:1990:ESD

[751] S. Gabler and R. Schweigkoffer. The existence of sampling designs with preassigned inclusion probabilities. Metrika, 37(1):87–96, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613510.

Auinger:1990:QGF

[752] Dr. K. Auinger. Quasi goodness of fit tests for lifetime distributions. Metrika, 37(1):97–116, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613511.

Gabler:1990:BR

[753] S. Gabler, J. L. Teugels, L. Afflerbach, J. Steinebach, P. L. Davies, H. O. Georgii, R. Trost, H. Heyer, and H. Strasser. Book reviews. Metrika, 37(1):117–127, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613512.

Ramallingam:1990:EBE

[754] Prof. T. Ramallingam. Exact Bahadur efficiencies of tests for superadditivity of the mean function of a non-homogeneous Poisson process. Metrika, 37(1):129–143, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613513.

Amari:1990:BR

[755] Shun ichi Amari. Book review. Metrika, 37(1):144, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
accesspage/article/10.1007/BF02613514.

Royen:1990:PIR

[756] Prof. Dr. T. Royen. A probability inequality for ranges and its application to maximum range test procedures. Metrika, 37(1):145–154, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613515.

Targhetta:1990:NMP

[757] Prof. M. Luisa Targhetta. A note on a mixing problem and the Schützenberger inequality. Metrika, 37(1):155–161, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613516.

Nakano:1990:BR

[758] Yuji Nakano. Book review. Metrika, 37(1):162, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613517.

Jacroux:1990:CTF

[759] Prof. Dr. M. Jacroux and R. SahaRay. On the construction of trend free row-column 2-level factorial experiments. Metrika, 37(1):163–180, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613518.

Mukerjee:1990:OBD

[760] Prof. Rahul Mukerjee. Optimality of balanced designs for minimum norm quadratic unbiased estimation of variance components. Metrika, 37 (1):181–188, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613519.

Anonymous:1990:HC

[761] Anonymous. Help & contacts. Metrika, 37(1):??, ???? 1990. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Gupta:1991:SND

[762] Professor A. K. Gupta and Professor D. G. Kabe. On some noncentral distribution problems for the mixture of two normal populations. Metrika, 38(1):1–10, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613593.

Naik:1991:GCE

[763] V. D. Naik and P. C. Gupta. A general class of estimators for estimating population mean using auxiliary information. Metrika, 38(1):11–17, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613594.

Pruscha:1991:BR

[764] H. Pruscha. Book review. Metrika, 38 (1):18, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF02613595.

vCollani:1991:NAS

[765] Prof. Dr. E. v. Collani. A note on acceptance sampling for variables. Metrika, 38(1):19–36, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613596.

Fahrmeir:1991:KFP

[766] L. Fahrmeir and H. Kaufmann. On Kalman filtering, posterior mode estimation and Fisher scoring in dynamic exponential family regression. Metrika, 38(1):37–60, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613597.

Kageyama:1991:BR

[767] Sanpei Kageyama, Olav Kallenberg, G. S. Watson, and M. Deistler. Book reviews. Metrika, 38(1):61–66, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613598.

Cheng:1991:LMN

[768] Prof. S. S. Cheng and B. R. Chen. The locally MIMSQE of nonnormal error variance in quadratically balanced models. Metrika, 38(1):67–70, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613599.

Sengupta:1991:AUV

[769] S. Sengupta and D. Kundu. Admissible unbiased variance estimation in finite population sampling under randomized response. Metrika, 38(1):71–82, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613600.

Pavur:1991:CIM

[770] Dr. R. Pavur and Dean M. D. M. Young. Conditions for invariance of the multivariate versions of Grubbs’ test and Bartlett’s test under a general dependency structure. Metrika, 38(1):83–97, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613601.

Heyer:1991:BR

[771] H. Heyer. Book review. Metrika, 38 (1):98, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF02613602.

Hassanein:1991:ESF

[772] Prof. K. M. Hassanein, A. K. Md E. Saleh, and E. F. Brown. Estimating survivor function using optimally selected order statistics. Metrika, 38 (1):99–115, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613603.

Kloeden:1991:BR

[773] P. E. Kloeden. Book review. Metrika, 38(1):116, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613604.

Aly:1991:TLO

[774] Dr. Emad-Eldin A. A. Aly. On testing for Lorenz ordering. Metrika, 38 (1):117–124, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613605.

Deistler:1991:BR

[775] M. Deistler, G. Koch, and G. Pap. Book reviews. Metrika, 38(1):125–128, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613606.

Draper:1991:FSO

[776] N. R. Draper, N. Gaffke, and F. Pukelsheim. First and second order rotatability of experimental designs, moment matrices, and information surfaces. Metrika, 38(1):129–161, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613607.

Kusama:1991:BR

[777] T. Kusama. Book review. Metrika, 38(1):162, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613608.

Shiraishi:1991:HTP

[778] T. Shiraishi. Hypothesis testing and parameter estimation based on Mstatistics in k samples with unequal variances. Metrika, 38(1):163–178, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613609.

Luschgy:1991:TOS

[779] H. Luschgy. Testing one-sided hypotheses for the mean of a Gaussian process. Metrika, 38(1):179–194, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613610.

Gupta:1991:TDN

[780] S. Gupta and S. Kageyama. Type S designs with nested rows and columns. Metrika, 38(1):195–202, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613611.

Jacroux:1991:NOB

[781] Prof. Dr. M. Jacroux. A note on the E-optimality of block designs. Metrika, 38(1):203–214, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613612.

Anonymous:1991:HC

[782] Anonymous. Help & contacts. Metrika, 38(1):??, ???? 1991. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Magiera:1992:BSE

[783] R. Magiera. Bayes sequential estimation for an exponential family of processes: A discrete time approach. Metrika, 39(1):1–20, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613977.

Lee:1992:MSE

[784] K. R. Lee and Prof. C. H. Kapadia. Mean squared error efficient estimators of the variance components. Metrika, 39(1):21–26, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613978.

Eichenauer-Herrmann:1992:MEU

[785] Dr. J. Eichenauer-Herrmann and Prof. Dr. W. Fieger. Minimax estimation under convex loss when the parameter interval is bounded. Metrika, 39(1):27–43, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613979.

Obata:1992:BR

[786] Nobuaki Obata. Book review. Metrika, 39(1):44, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613980.

Morgenthaler:1992:BLS

[787] Prof. S. Morgenthaler. The bias of least squares polynomial interpolants. Metrika, 39(1):45–55, ????

1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613981.

Arnold:1992:BR

[788] B. Arnold, W. Wertz, B. M. Pötscher, D. Voss, R. Shimizu, R. Dahlhaus, M. Leitner, O. Krafft, and G. Pflug. Book reviews. Metrika, 39(1):56–66, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613982.

Stoyan:1992:SEM

[789] Prof. Dr. D. Stoyan. Statistical estimation of model parameters of planar Neyman–Scott cluster processes. Metrika, 39(1):67–74, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613983.

Giri:1992:LMT

[790] N. Giri, M. Behara, and P. Banerjee. Locally minimax test of independence in elliptically symmetrical distributions with additional observations. Metrika, 39(1):75–84, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613984.

Kim:1992:EDO

[791] S. B. Kim and D. S. Bai. Economic design of one-sided screening procedures based on a correlated variable with all parameters unknown. Metrika, 39(1):85–93, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613985.

Lasser:1992:BR

[792] R. Lasser. Book review. Metrika, 39 (1):94, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF02613986.

Luschgy:1992:ATD

[793] H. Luschgy. Approximations and timediscretization in testing one-sided hypotheses for the mean of a Gaussian process. Metrika, 39(1):95–105, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613987.

Draper:1992:BR

[794] N. R. Draper. Book review. Metrika, 39(1):106, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613988.

Balasubramanian:1992:DDC

[795] K. Balasubramanian and M. I. Beg. Distributions determined by conditioning on a pair of order statistics. Metrika, 39(1):107–112, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613989.

Ramanathan:1992:RTT

[796] T. V. Ramanathan and M. B. Rajarshi. Rank tests for testing randomness of a regression coefficient in a linear regression model. Metrika, 39(1):113–124, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613990.

Hazod:1992:BR

[797] W. Hazod, N. C. Giri, H. Bergström, H. Heyer, P. K. Sen, and H. Heyer. Book reviews. Metrika, 39(1):125–130, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613991.

Selvavel:1992:UMV

[798] K. Selvavel. Uniform minimum variance unbiased estimators for bivariate families. Metrika, 39(1):131–138, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613992.

Gob:1992:TOC

[799] R. Göb. Type A operating characteristic functions of defects sampling plans. Metrika, 39(1):139–153, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613993.

Kloeden:1992:BR

[800] P. E. Kloeden. Book review. Metrika, 39(1):154, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613994.

Mukhopadhyay:1992:SCM

[801] N. Mukhopadhyay and T. N. Sriram. On sequential comparisons of means of first-order autoregressive models. Metrika, 39(1):155–164, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613995.

Bar-Lev:1992:VSN

[802] S. K. Bar-Lev and P. Enis. Variance stabilizing and normalizing transformations for the compound Poisson process. Metrika, 39(1):165–175, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613996.

Anonymous:1992:HC

[803] Anonymous. Help & contacts. Metrika, 39(1):??, ???? 1992. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Hafner:1993:CMT

[804] Robert Hafner. Construction of minimax-tests for bounded families of probability-densities. Metrika, 40 (1):1–23, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613659.

Bickel:1993:BR

[805] P. J. Bickel. Book review. Metrika, 40 (1):24, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF02613660.

Mulekar:1993:FSS

[806] Madhuri S. Mulekar and Linda J. Young. A fixed sample size selection procedure for negative binomial populations. Metrika, 40(1):25–35, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02613661.

Arnold:1993:BR

[807] B. Arnold. Book review. Metrika, 40 (1):36, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/BF02613662.

Dette:1993:NIO

[808] Holger Dette. A new interpretation of optimality for E-optimal designs in linear regression models. Metrika, 40(1):37–50, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02613663.

Takada:1993:UMA

[809] Yoshikazu Takada. Uniformly most accurate equivariant prediction limit. Metrika, 40(1):51–61, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF02613664.

Birge:1993:BR

[810] L. Birgé and H. H. Bock. Book reviews. Metrika, 40(1):62–66, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613665.

Heinrich:1993:APM

[811] Lothar Heinrich. Asymptotic properties of minimum contrast estimators for parameters of Boolean models. Metrika, 40(1):67–94, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613666.

Bai:1993:EDS

[812] Do Sun Bai and Min Koo Lee. Economic designs of single and double screening procedures. Metrika, 40 (1):95–113, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613667.

Eberle:1993:BR

[813] E. Eberle. Book review. Metrika, 40(1):114, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613668.

Eichenauer-Herrmann:1993:LSN

[814] Jürgen Eichenauer-Herrmann. The lattice structure of nonlinear congruential pseudorandom numbers. Metrika, 40(1):115–120, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613669.

Mukhopadhyay:1993:NTS

[815] N. Mukhopadhyay and A. R. Padmanabhan. A note on three-stage confidence intervals for the difference of locations: The exponential case. Metrika, 40(1):121–128, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613670.

Bartlett:1993:BR

[816] M. S. Bartlett, S. T. Rachev, E. Dettweiler, D. A. Berry, and E. Mammen. Book reviews. Metrika, 40(1):129–136, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613671.

Lee:1993:CUE

[817] Kwan R. Lee and C. H. Kapadia. Combined unbiased estimators based on Q-reduced model. Metrika, 40 (1):137–147, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613672.

Arato:1993:BR

[818] M. Arató. Book review. Metrika, 40(1):148, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613673.

Lin:1993:TSR

[819] C. H. Lin and Shashikala Sukhatme. Two-sample rank tests with truncated populations. Metrika, 40(1):149–172, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613674.

Mukhopadhyay:1993:FDE

[820] N. Mukhopadhyay, S. Chattopadhyay, and S. K. Sahu. Further developments in estimation of the largest mean of K normal populations. Metrika, 40(1):173–183, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613675.

Anger:1993:BR

[821] B. Anger. Book review. Metrika, 40(1):184, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613676.

vandeVen:1993:BMN

[822] R. van de Ven and N. C. Weber. Bounds for the median of the negative binomial distribution. Metrika, 40(1):185–189, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613677.

Pfeifer:1993:BR

[823] D. Pfeifer. Book review. Metrika, 40(1):190, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02613678.

Anonymous:1993:HC

[824] Anonymous. Help & contacts. Metrika, 40(1):??, ???? 1993. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Riedel:1994:CED

[825] Manfred Riedel and Hans-Joachim Rossberg. Characterization of the exponential distribution function by properties of the difference X_{k+s:n} − X_{k:n} of order statistics. Metrika, 41(1):1–19, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01895297.

Csaki:1994:ARW

[826] E. Csáki. Analysis of random walks, studies in probability, optimization and statistics. Metrika, 41(1):20, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01895298.

Ramchandra:1994:SMS

[827] Vidyasagar Ramchandra and R. Padmawar. Stratification, Midzuno– Sen strategy and nonnegative unbiased variance estimation. Metrika, 41(1):21–27, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01895299.

Heyer:1994:NME

[828] H. Heyer. Natur und mathematisches Erkennen. (German) [Natural and mathematical cognition]. Metrika, 41 (1):28, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/BF01895300.

Bagchi:1994:OCS

[829] Sunanda Bagchi. Optimality and construction of some rectangular designs. Metrika, 41(1):29–41, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01895301.

Rao:1994:DIF

[830] P. S. R. S. Rao. Design and inference in finite population sampling. Metrika, 41(1):42, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01895302.

Gob:1994:BMP

[831] Rainer Göb. Bounds for median and 50 percentage point of binomial and negative binomial distribution. Metrika, 41(1):43–54, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01895303.

Liski:1994:ALN

[832] Erkki P. Liski and Song-Gui Wang. Another look at the naive estimator in a regression model. Metrika, 41(1):55–64, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01895304.

Hognas:1994:BR

[833] G. Högnäs, H. Heyer, and M. Falk. Book reviews. Metrika, 41(1):65–70, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01895305.

Kunert:1994:OBD

[834] Joachim Kunert. Optimality of block designs with variable block sizes and random block effects. Metrika, 41 (1):71–81, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01895306.

Nather:1994:OED

[835] W. Näther. Optimum experimental designs. Metrika, 41(1):82, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01895307.

Seoh:1994:BER

[836] Munsup Seoh and Madan L. Puri. Berry-Esséen rate in asymptotic normality for perturbed sample quantiles. Metrika, 41(1):83–98, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01895308.

Kim:1994:AGB

[837] Byung Hwee Kim. Admissibility of generalized Bayes estimators in an one parameter nonregular family. Metrika, 41(1):99–108, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01895309.

Zografos:1994:ITA

[838] Konstantinos Zografos and Kosmas Ferentinos. An information theoretic argument for the validity of the exponential model. Metrika, 41(1):109–119, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01895310.

Woess:1994:BR

[839] W. Woess, W. Weil, W. Wertz, N. Jacob, and P. van der Laan. Book reviews. Metrika, 41(1):120–126, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01895311.

Dixit:1994:BAP

[840] Ulhas J. Dixit. Bayesian approach to prediction in the presence of outliers for Weibull distribution. Metrika, 41(1):127–136, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01895312.

Draper:1994:TOR

[841] Norman R. Draper and Friedrich Pukelsheim. On third order rotatability. Metrika, 41(1):137–161, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01895313.

Ferschl:1994:AS

[842] F. Ferschl. Analogy and structure. Metrika, 41(1):162, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01895314.

Albers:1994:STL

[843] W. Albers, W. C. M. Kallenberg, and G. D. Otten. Setting test limits under prescribed consumer loss. Metrika, 41(1):163–181, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01895315.

Burdzy:1994:SAB

[844] K. Burdzy. Some aspects of Brownian motion. Part I: Some special functionals. Metrika, 41(1):182, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01895316.

Anonymous:1994:HC

[845] Anonymous. Help & contacts. Metrika, 41(1):??, ???? 1994. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Bai:1995:EDT

[846] Do Sun Bai and Hyuck Moo Kwon. Economic design of a two-stage screening procedure with a prescribed outgoing quality. Metrika, 42(1):1–18, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01894285.

Happacher:1995:EAO

[847] Maximilian Happacher. Exact and approximate D-optimal designs in polynomial regression. Metrika, 42 (1):19–27, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01894286.

Gaffke:1995:ORI

[848] Norbert Gaffke and Berthold Heiligers. Optimal and robust invariant designs for cubic multiple regression. Metrika, 42(1):29–48, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01894287.

Heinrich:1995:PEB

[849] G. Heinrich and U. Jensen. Parameter estimation for a bivariate lifetime distribution in reliability with multivariate extensions. Metrika, 42(1):49–65, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01894288.

Silverman:1995:BR

[850] B. W. Silverman, N. M. Laird, A. Wakolbinger, M. Fukushima, P. C. Consul, D. M. Titterington, W. Fieger, K. Jacobs, and H. Drygas. Book reviews. Metrika, 42(1):66–78, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01894289.

Anonymous:1995:NAP

[851] Anonymous. Newly accepted papers. Metrika, 42(1):79–97, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01894290.

Lin:1995:SPC

[852] Dennis K. J. Lin and Norman R. Draper. Screening properties of certain two-level designs. Metrika, 42 (1):99–118, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01894291.

Dijkstra:1995:PLI

[853] Theo K. Dijkstra. Pyrrho’s lemma, or have it your way. Metrika, 42 (1):119–125, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01894292.

Famoye:1995:BGP

[854] Felix Famoye and P. C. Consul. Bivariate generalized Poisson distribution with some applications. Metrika, 42(1):127–138, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF01894293.

Heyer:1995:BR

[855] H. Heyer, K. D. Elworthy, N. Cressie, R. J. Williams, H. Büning, R. Schassberger, and H. Lütkepohl. Book reviews. Metrika, 42(1):139–148, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF01894294.

Pukelsheim:1995:E

[856] Friedrich Pukelsheim, Werner Uhlman, and Herbert Heyer. Editorial. Metrika, 42(1):149, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01894295.

Pazman:1995:P

[857] Andrej Pázman. Preface. Metrika, 42(1):151, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01894296.

LeBreton:1995:SRA

[858] Alain Le Breton and Alexander Novikov. Some results about averaging in stochastic approximation. Metrika, 42(1):153–171, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01894297.

Gaffke:1995:AOD

[859] Norbert Gaffke and Berthold Heiligers. Algorithms for optimal design with application to multiple polynomial regression. Metrika, 42(1):173–190, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01894298.

Hougaard:1995:NRC

[860] Philip Hougaard. Nonlinear regression and curved exponential families. improvement of the approximation to the asymptotic distribution. Metrika, 42(1):191–202, ????

1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF01894299.

Rao:1995:RCC

[861] C. Radhakrishna Rao and L. C. Zhao. Recent contributions to censored regression models. Metrika, 42 (1):203–213, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01894300.

Searle:1995:OVC

[862] Shayle R. Searle. An overview of variance component estimation. Metrika, 42(1):215–230, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF01894301.

Calinski:1995:BCB

[863] Tadeusz Caliński. The basic contrasts of a block design with special reference to the recovery of inter-block information. Metrika, 42(1):231–232, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
BF01894302.

Cheng:1995:MEC

[864] Ching-Shui Cheng and Don X. Sun. Maximum estimation capacity and minimum aberration. Metrika, 42 (1):232–233, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01894303.

Chuiv:1995:ELP

[865] Nóra Ní Chuív, Bimal K. Sinha, and Zhong Wu. Estimation of the location parameter of a Cauchy distribution using a ranked set sample. Metrika, 42(1):234–235, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF01894304.

Anonymous:1995:HC

[866] Anonymous. Help & contacts. Metrika, 42(1):??, ???? 1995. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Szkutnik:1996:QML

[867] Zbigniew Szkutnik. Quasi-maximum likelihood estimation of parameters in a multivariate Poisson process. Metrika, 43(1):1–16, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613893.

Ahmed:1996:SEP

[868] S. E. Ahmed and V. K. Rohatgi. Shrinkage estimation of the proportion in randomized response. Metrika, 43(1):17–30, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613894.

Flak:1996:OTL

[869] Thomas Flak and Wolfgang Schmid. An outlier test for linear processes-II. large contamination. Metrika, 43(1):31–42, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613895.

Wu:1996:LSL

[870] Yanhong Wu. A less sensitive linear detector for the change point based on kernel smoothing method. Metrika, 43(1):43–55, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613896.

Boutahar:1996:LSE

[871] Mohamed Boutahar and Claude Deniau. Least squares estimator for regression models with some deterministic time varying parameters. Metrika, 43(1):57–67, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613897.

Gaschler:1996:CAN

[872] Birgit Gaschler. Consistency and asymptotic normality of maximum likelihood estimation for Gaussian Markov processes from discrete observations. Metrika, 43(1):69–90, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613898.

Sharif:1996:BR

[873] A. H. Sharif, S. Csörgö, S. Takenaka, D. Dacunha-Castelle, and H. Basler. Book-reviews. Metrika, 43(1):91–100, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613899.

Garcia:1996:RSM

[874] M. Rueda Garcia and A. Arcos Cebrian. Repeated substitution method: The ratio estimator for the population variance. Metrika, 43(1):101–105, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613900.

Sellke:1996:GGC

[875] Thomas Sellke. Generalized Gauss– Chebyshev inequalities for unimodal distributions. Metrika, 43(1):107–121, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613901.

Chaudhuri:1996:ERC

[876] Arijit Chaudhuri and Tapabrata Maiti. Estimating regression coefficients from survey data by asymptotic design-cum-model based approach. Metrika, 43 (1):123–133, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613902.

Wu:1996:CDC

[877] Jong-Wuu Wu and L. Y. Ouyang. On characterizing distributions by conditional expectations of functions of order statistics. Metrika, 43(1):135–147, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613903.

Bathe:1996:MRS

[878] Falk Bathe and Jürgen Franz. Modelling of repairable systems with various degrees of repair. Metrika, 43 (1):149–164, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613904.

Dragalin:1996:SES

[879] Vladimir Dragalin. A simple and effective scanning rule for a multi-channel system. Metrika, 43(1):165–182, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613905.

Koeden:1996:BR

[880] P. Koeden, R. Viertl, R. Schassberger, and G. Casella. Bookreviews. Metrika, 43(1):183–187, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613906.

DelCastillo:1996:RLD

[881] Enrique Del Castillo. Run length distributions and economic design of X̄ charts with unknown process variance. Metrika, 43(1):189–201, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02613907.

Eccleston:1996:OME

[882] J. A. Eccleston and J. A. John. Orthogonal main effect plans for two and three factors in small blocks. Metrika, 43(1):203–211, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613908.

Stenger:1996:MPL

[883] H. Stenger and S. Gabler. A minimax property of Lahiri–Midzuno– Sen’s sampling scheme. Metrika, 43 (1):213–220, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613909.

Ong:1996:CDD

[884] S. H. Ong. On a class of discrete distributions arising from the birth-death-with-immigration process. Metrika, 43(1):221–235, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613910.

Gyorfi:1996:MKD

[885] L. Györfi, I. Vajda, and E. van der Meulen. Minimum Kolmogorov distance estimates of parameters and parametrized distributions. Metrika, 43(1):237–255, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613911.

Kahle:1996:BCW

[886] Waltraud Kahle. Bartlett corrections for the Weibull distribution. Metrika, 43(1):257–263, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02613912.

Anonymous:1996:HCa

[887] Anonymous. Help & contacts. Metrika,

43(1):??, ????  1996.  CODEN

MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Stute:1996:SPR

[888] W. Stute. The sequential probability ratio test under random censorship. Metrika, 44(1):1–8, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02614050.

Pazman:1996:DPE

[889] Andrej Pázman. The density of the parameter estimators when the observations are distributed exponentially. Metrika, 44(1):9–26, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02614051.

Kahle:1996:EPW

[890] Waltraud Kahle. Estimation of the parameters of the Weibull distribution for censored samples. Metrika, 44(1):27–40, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02614052.

Chiou:1996:EEV

[891] Paul Chiou and Chien-Pai Han. Estimation of error variance in one-way random model. Metrika, 44(1):41–51, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02614053.

Bai:1996:EDT

[892] Do Sun Bai and Min Koo Lee. Economic designs of two-sided single and double screening procedures. Metrika, 44(1):53–69, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02614054.

Ma:1996:MSF

[893] Chunsheng Ma. Multivariate survival functions characterized by constant product of mean remaining lives and hazard rates. Metrika, 44(1):71–83, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614055.

Heyer:1996:BR

[894] H. Heyer, H. Rieder, F. Papangelou, A. Janssen, L. Rüschendorf, R. L. Hudson, and A. Greven. Bookreviews. Metrika, 44(1):84–93, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614056.

Anonymous:1996:NAP

[895] Anonymous. Newly accepted papers. Metrika, 44(1):94, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/BF02614057.

El-Din:1996:IPM

[896] M. M. Mohie El-Din, S. E. Abu-Youssef, and K. S. Sultan. An identity for the product moments of order statistics. Metrika, 44(1):95–100, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614058.

Jensen:1996:SLM

[897] D. R. Jensen. Straight-line models in star-shaped mixtures. Metrika, 44 (1):101–117, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02614059.

Arnold:1996:AFH

[898] Bernhard F. Arnold. An approach to fuzzy hypothesis testing. Metrika, 44(1):119–126, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02614060.

Szkutnik:1996:QCQ

[899] Zbigniew Szkutnik. Quick consistency of quasi-maximum likelihood estimators in a multivariate Poisson process. Metrika, 44(1):127–134, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02614061.

Gob:1996:EMS

[900] Rainer Göb. An elementary model for statistical lot inspection and its application to sampling by variables. Metrika, 44(1):135–163, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02614062.

Schwabe:1996:SAP

[901] Rainer Schwabe and Harro Walk. On a stochastic approximation procedure based on averaging. Metrika, 44 (1):165–180, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02614063.

Gather:1996:BR

[902] U. Gather, M. Schwaiger, W. Uhlmann, H. Heyer, N. Yacob, P. Embrechto, and K. Dietz. Book reviews. Metrika, 44(1):181–187, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02614064.

Chen:1996:JCR

[903] Zhenmin Chen. Joint confidence region for the parameters of Pareto distribution. Metrika, 44(1):191–197, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614065.

Eichenauer-Herrmann:1996:EPI

[904] Jürgen Eichenauer-Herrmann. Equidistribution properties of inversive congruential pseudorandom numbers with power of two modulus. Metrika, 44 (1):199–205, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02614066.

Vogt:1996:ANE

[905] Herbert Vogt. The average number of events per time unit and its estimation. Metrika, 44(1):207–221, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02614067.

Gob:1996:TSM

[906] Rainer Göb. Tests of significance for the mean of a finite lot. Metrika, 44(1):223–238, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02614068.

Muller-Gronbach:1996:OAE

[907] Thomas Müller-Gronbach and Rainer Schwabe. On optimal allocations for estimating the surface of a random field. Metrika, 44(1):239–258, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02614069.

Anonymous:1996:HCb

[908] Anonymous. Help & contacts. Metrika, 44(1):??, ???? 1996. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Tenbusch:1997:NCE

[909] Axel Tenbusch. Nonparametric curve estimation with Bernstein estimates. Metrika, 45(1):1–30, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/BF02717090.

Targhetta:1997:TSS

[910] M. Luisa Targhetta. A two-stage sampling by unequal size samples from two normal populations with unknown variances. Metrika, 45(1):31–37, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02717091.

Rost:1997:NCD

[911] Daniel Rost. A note on compact differentiability and the δ-method. Metrika, 45(1):39–51, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02717092.

Liu:1997:EMO

[912] Shuangzhe Liu and Heinz Neudecker. Experiments with mixtures: Optimal allocations for Becker’s models. Metrika, 45(1):53–66, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02717093.

Morgan:1997:ODI

[913] J. P. Morgan. Optimal design for interacting blocks with OAVS incidence. Metrika, 45(1):67–83, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02717094.

Ronning:1997:BR

[914] G. Ronning, C. C. Heyde, O. O. Aalen, P. Huber, P. A. Loeb, D. L. Burkholder, Kh. Alam, O. E. Barndorff-Nielsen, and P. E. Kloeden. Book reviews. Metrika, 45(1):84–93, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02717095.

Mukhopadhyay:1997:EAS

[915] N. Mukhopadhyay and T. K. S. Solanky. Estimation after sequential selection and ranking. Metrika, 45 (1):95–106, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02717096.

Franco:1997:CDE

[916] Manuel Franco and Jose M. Ruiz. On characterizations of distributions by expected values of order statistics and record values with gap. Metrika, 45 (1):107–119, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02717097.

Henze:1997:DCS

[917] Norbert Henze. Do components of smooth tests of fit have diagnostic properties? Metrika, 45(1): 121–130, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02717098.

Gob:1997:TSA

[918] Rainer Göb. Tests of significance for the average square deviation from target in a finite lot. Metrika, 45 (1):131–169, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02717099.

Cebrian:1997:VEU

[919] A. Arcos Cebrián and M. Rueda García. Variance estimation using auxiliary information: An almost unbiased multivariate ratio estimator. Metrika, 45(1):171–178, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02717100.

Obata:1997:BR

[920] N. Obata, L. Arnold, J. P. Buonaccorsi, H. Witting, L. M. Bloom, J. Stark, and E. von Collani. Book reviews. Metrika, 45(1):179–188, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02717101.

Balasubramanian:1997:DCT

[921] K. Balasubramanian and Aloke Dey. Distributions characterized through conditional expectations. Metrika, 45 (1):189–196, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02717102.

LaMotte:1997:LUB

[922] Lynn Roy LaMotte. On limits of uniquely best linear estimators. Metrika, 45(1):197–211, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF02717103.

Gupta:1997:NPC

[923] Arjun K. Gupta and S. Kotz. A new process capability index. Metrika, 45 (1):213–224, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02717104.

Dewan:1997:RSL

[924] Isha Dewan and B. L. S. Prakasa Rao. Remarks on the Strong Law of Large Numbers for a triangular array of associated random variables. Metrika, 45(1):225–234, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02717105.

Mukhopadhyay:1997:ERA

[925] Nitis Mukhopadhyay, A. R. Padmanabhan, and T. K. S. Solanky. On estimating the reliability after sequentially estimating the mean: The exponential case. Metrika, 45(1):235–252, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02717106.

Schwabe:1997:RBE

[926] Rainer Schwabe and Weng Kee Wong. A relationship between efficiencies of marginal designs and the product design. Metrika, 45(1):253–257, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02717107.

Aly:1997:CPT

[927] Emad-Eldin A. A. Aly and Subhash C. Kochar. Change point tests based on U-statistics with applications in reliability. Metrika, 45(1):259–269, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02717108.

Georgii:1997:BR

[928] H. O. Georgii, H. Prodinger, Y. Chikuse, H. Rost, St. E. Fienberg, and P. Imkeller. Book reviews. Metrika, 45(1):271–280, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02717109.

Anonymous:1997:HCa

[929] Anonymous. Help & contacts. Metrika, 45(1):??, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Steinebach:1997:DCM

[930] Josef Steinebach and Vera R. Eastwood. Detecting changes in a multivariate renewal process. Metrika, 46(1):1–19, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02717162.

Joarder:1997:ESM

[931] Anwar H. Joarder and Mir M. Ali. Estimation of the scale matrix of a multivariate t-model under entropy loss. Metrika, 46(1):21–32, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02717163.

Zhang:1997:SBR

[932] Jin Zhang and Xueren Wang. Selecting the best regression equation via the P -value of F -test. Metrika, 46(1):33–40, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02717164.

Engel:1997:MH

[933] Joachim Engel. The multiresolution histogram. Metrika, 46(1):41–57, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02717165.

Bansal:1997:NSR

[934] Naveen K. Bansal and Sudhir Gupta. On the natural selection rule in general linear models. Metrika, 46 (1):59–69, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02717166.

Dette:1997:NUD

[935] Holger Dette. A note on the uniform distribution on the arcsin points. Metrika, 46(1):71–82, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02717167.

Jing:1997:BR

[936] Bing-Yi Jing, G. Kallianpur, D. W. Müller, H. Strasser, H. G. Weidner, H. J. Roßberg, and M. Akahira. Book reviews. Metrika, 46(1):83–92, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02717168.

Cramer:1997:UBT

[937] Erhard Cramer and Udo Kamps. The UMVUE of P (X < Y) based on Type-II censored samples from Weinman multivariate exponential distributions. Metrika, 46(1):93–121, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02717169.

Masse:1997:NML

[938] Jean-Claude Massé. Nonparametric maximum likelihood estimation in a non locally compact setting. Metrika, 46(1):123–145, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02717170.

Fedorov:1997:AVO

[939] Valery V. Fedorov and Werner G. Müller. Another view on optimal design for estimating the point of extremum in quadratic regression. Metrika, 46(1):147–157, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02717171.

Kaluszka:1997:DSR

[940] Marek Kaluszka and Wlodzimierz Krysicki. On decompositions of some random variables. Metrika, 46(1): 159–175, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02717172.

Schal:1997:BR

[941] M. Schäl, E. Sonnemann, P. Gänßler, M. Falk, and H. Riedwyl. Book reviews. Metrika, 46(1):177–186, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02717173.

Cramer:1997:CBT

[942] Michael Cramer. Convergence of a branching type recursion with nonstationary immigration. Metrika, 46 (1):187–211, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02717174.

Chaturvedi:1997:EPF

[943] Ajit Chaturvedi and Uma Rani. Estimation procedures for a family of density functions representing various life-testing models. Metrika, 46 (1):213–219, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02717175.

Zhang:1997:EDF

[944] Biao Zhang. Estimating a distribution function in the presence of auxiliary information. Metrika, 46 (1):221–244, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02717176.

Raqab:1997:NRP

[945] Mohammad Z. Raqab and Wael A. Amin. A note on reliability properties of k-record statistics. Metrika, 46 (1):245–251, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02717177.

Irle:1997:BEO

[946] Albrecht Irle and Vladimir Ivanovich Lotov. A bound on the expected overshoot for some concave boundaries. Metrika, 46(1):253–267, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02717178.

Dudley:1997:BR

[947] R. M. Dudley, H. Heyer, A. Greven, V. Schmidt, R. K. Getoor, and W. Wertz. Book reviews. Metrika, 46 (1):269–278, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02717179.

Anonymous:1997:HCb

[948] Anonymous. Help & contacts. Metrika, 46(1):??, January 1997. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Albers:1998:ATL

[949] Willem Albers, Wilbert C. M. Kallenberg, and Gertjan D. Otten. Accurate test limits under nonnormal measurement error. Metrika, 47(1):1–33, January 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02742862.

Ahmed:1998:LSE

[950] S. E. Ahmed. Large-sample estimation strategies for eigenvalues of a Wishart matrix. Metrika, 47(1):35–45, January 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02742863.

Krummenauer:1998:LTM

[951] Frank Krummenauer. Limit theorems for multivariate discrete distributions. Metrika, 47(1):47–69, January 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02742864.

Bhattacharya:1998:TCS

[952] Bhaskar Bhattacharya. Testing conditional symmetry against one-sided alternatives in square contingency tables. Metrika, 47(1):71–84, January 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02742865.

Dettweiler:1998:BR

[953] E. Dettweiler, H. Heyer, Hiroshi Sato, M. Duflo, R. Münnich, W. Grölz, H. Stenger, H. Luschgy, and N. Jacob. Book reviews. Metrika, 47 (1):85–94, January 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02742866.

Pakes:1998:EOS

[954] Anthony G. Pakes. Extreme order statistics on Galton–Watson trees. Metrika, 47(1):95–117, January 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF02742867.

Ryden:1998:AER

[955] Tobias Rydén. Asymptotically efficient recursive estimation for incomplete data models using the observed information. Metrika, 47(1):119–145, January 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02742868.

Kohler:1998:NRF

[956] Michael Kohler. Nonparametric regression function estimation using interaction least squares splines and comlexity regularization. Metrika, 47 (1):147–163, January 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02742869.

Lopez-Blazquez:1998:UBC

[957] Fernando López-Blázquez and Begoña Salamanca-Miño. An upper bound for the correlation ratio of records. Metrika, 47(1):165–174, January 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/BF02742870.

Stoyan:1998:BR

[958] D. Stoyan, J. W. Thompson, P. K. Sen, D. W. Stroock, M. M. A. Ananda, and J. Potthoff. Book reviews. Metrika, 47(1):175–184, January 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02742871.

Mukhopadhyay:1998:MMM

[959] Nitis Mukhopadhyay and Makoto Aoshima. Multivariate multistage methodologies for simultaneous all pairwise comparisons. Metrika, 47 (1):185–201, January 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02742872.

Garcia:1998:QIE

[960] M. Rueda García, A. Arcos Cebrián, and E. Artés Rodríguez. Quantile interval estimation in finite population using a multivariate ratio estimator. Metrika, 47(1):203–213, January 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02742873.

Mukhopadhyay:1998:POL

[961] Parimal Mukhopadhyay and Rainer Schwabe. On the performance of the ordinary least squares method under an error component model. Metrika, 47(1):215–226, January 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02742874.

Ma:1998:CPM

[962] Chunsheng Ma. Characteristic properties of multivariate survival functions in terms of residual life distributions. Metrika, 47(1):227–240, January 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02742875.

Singh:1998:EFP

[963] Sarjinder Singh and Anwar H. Joarder. Estimation of finite population variance using random non-response in survey sampling. Metrika, 47(1): 241–249, January 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/BF02742876.

Steland:1998:BRS

[964] Ansgar Steland. Bootstrapping rank statistics. Metrika, 47(1):251–264, January 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02742877.

Marion:1998:BR

[965] J. Marion, H. Heyer, R. B. Bapat, M. Stadler, R. Norberg, M. Fukushima, A. Weron, and I. Csiszár. Book reviews. Metrika, 47(1):265–278, January 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02742878.

Anonymous:1998:HCa

[966] Anonymous. Help & contacts. Metrika, 47(1):??, January 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Muller:1998:ANS

[967] Christine H. Müller and Andrej Pázman. Applications of necessary and sufficient conditions for maximin efficient designs. Metrika, 48 (1):1–19, September 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050001.

Nguyen:1998:CND

[968] Truc T. Nguyen and Khoan T. Dinh. Characterizations of normal distributions supporting goodness-of-fit tests based on sample skewness and sample kurtosis. Metrika, 48 (1):21–30, September 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050002.

Knoth:1998:QSC

[969] Sven Knoth. Quasi-stationarity of CUSUM schemes for Erlang distributions. Metrika, 48(1):31–48, September 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840050003.

Whittinghill:1998:NRB

[970] Dexter C. Whittinghill. A note on the robustness of Box–Behnken designs to the unavailability of data. Metrika, 48 (1):49–52, September 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/PL00003971.

Chattopadhyay:1998:SEN

[971] Saibal Chattopadhyay. Sequential estimation of normal mean under asymmetric loss function with a shrinkage stopping rule. Metrika, 48(1): 53–59, September 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050005.

Malov:1998:RVG

[972] Sergey V. Malov. Random variables generated by ranks in dependent schemes. Metrika, 48(1): 61–67, September 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/PL00003972.

Anonymous:1998:HCb

[973] Anonymous. Help & contacts. Metrika, 48(1):??, September 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Kundu:1998:APL

[974] Debasis Kundu and Rameshwar D. Gupta. Asymptotic properties of the least squares estimators of a two dimensional model. Metrika, 48(2): 83–97, December 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/PL00020898.

Singh:1998:AEM

[975] Sarjinder Singh and Stephen Horn. An alternative estimator for multicharacter surveys. Metrika, 48(2): 99–107, December 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/PL00020899.

Chen:1998:APM

[976] Di Chen, Jye-Chyi Lu, Jacqueline M. Hughes-Oliver, and Chin-Shang Li.

Asymptotic properties of maximum likelihood estimates for a bivariate exponential distribution and mixed censored data. Metrika, 48(2): 109–125, December 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/PL00020900.

Martin:1998:ERT

[977] R. J. Martin, G. Jones, and J. A. Eccleston. Eight-run two-level factorial designs under dependence. Metrika, 48 (2):127–139, December 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050004.

Schmidt:1998:PLM

[978] Klaus D. Schmidt. Prediction in the linear model: A direct approach. Metrika, 48(2):141–147, December 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840050011.

Joarder:1998:ESM

[979] A. H. Joarder and S. E. Ahmed. Estimation of the scale matrix of a class of elliptical distributions. Metrika, 48 (2):149–160, December 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050014.

Anonymous:1998:HCc

[980] Anonymous. Help & contacts. Metrika, 48(2):??, December 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Toutenburg:1998:ERP

[981] H. Toutenburg and V. K. Srivastava. Estimation of ratio of population means in survey sampling when some observations are missing. Metrika, 48(3):177–187, December 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00003973.

Gather:1998:ESF

[982] Ursula Gather and Jörg Pawlitschko. Estimating the survival function under a generalized Koziol–Green model with partially informative censoring. Metrika, 48(3):189–207, December 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00020902.

Maruyama:1998:MEN

[983] Yuzo Maruyama. Minimax estimators of a normal variance. Metrika, 48 (3):209–214, December 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/PL00003974.

Dembinska:1998:LRN

[984] Anna Dembińska and Jacek Wesolowski. Linearity of regression for non-adjacent order statistics. Metrika, 48(3): 215–222, December 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050016.

Liu:1998:TSP

[985] Jeng-Fu Liu. Two-stage procedures for approximating the expected reward: The negative exponential case. Metrika, 48(3):223–230, December 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840050017.

Padmawar:1998:END

[986] Vidyasagar Ramchandra Padmawar. On estimating nonnegative definite quadratic forms. Metrika, 48(3): 231–244, December 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/PL00003975.

Anonymous:1998:HCd

[987] Anonymous. Help & contacts. Metrika, 48(3):??, December 1998. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Anonymous:1999:E

[988] Anonymous. Editorial. Metrika, 49(1): 1, July 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        PL00020904.

Anonymous:1999:RR

[989] Anonymous. Recognition of referees. Metrika, 49(1):2, July 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s001840050020.

Taheri:1999:NPL

[990] S. M. Taheri and Javad Behboodian. Neyman–Pearson lemma for fuzzy hypotheses testing. Metrika, 49(1):3–17, July 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050021.

Imhof:1999:EDO

[991] Lorens Imhof and Olaf Krafft. Extending design-optimality from an initial model to augmented models. Metrika, 49(1):19–26, July 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050022.

Keseling:1999:CDG

[992] Claudia Keseling. Conditional distributions of generalized order statistics and some characterizations. Metrika, 49(1):27–40, July 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050023.

Payton:1999:SPT

[993] Mark E. Payton and Linda J. Young. A sequential procedure to test three values of a binomial parameter. Metrika, 49(1):41–52, July 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050024.

Klar:1999:GFT

[994] Bernhard Klar. Goodness-of-fit tests for discrete models based on the integrated distribution function. Metrika, 49(1):53–69, July 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050025.

Anonymous:1999:HCa

[995] Anonymous. Help & contacts. Metrika, 49(1):??, July 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Brodeau:1999:ART

[996] F. Brodeau. Asymptotic results and tests for the choice of approximative models in nonlinear two-phases regression models, heteroscedatic case. Metrika, 49(2):85–105, September 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840050026.

Jensen:1999:REN

[997] Donald R. Jensen and Donald E. Ramirez. Recovered errors and normal diagnostics in regression. Metrika, 49 (2):107–119, September 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050027.

Asadi:1999:MDC

[998] Majid Asadi. Multivariate distributions characterized by a relationship between mean residual life and hazard rate. Metrika, 49(2):121– 126, September 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050028.

Dodge:1999:MM

[999] Yadolah Dodge and Valentin Rousson. Multivariate L_{1} mean. Metrika, 49(2): 127–134, September 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050029.

Huang:1999:MFG

[1000] J. S. Huang and Samuel Kotz.

Modifications of the Farlie–Gumbel– Morgenstern distributions. A tough hill to climb. Metrika, 49(2):135– 145, September 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050030.

Nguyen:1999:SCN

[1001] Truc T. Nguyen and Khoan T. Dinh.

Stability of characterizations of normal distributions based on the conditional expected values of the sample skewness and the sample kurtosis. Metrika, 49 (2):147–158, September 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050031.

Anonymous:1999:HCb

[1002] Anonymous. Help & contacts. Metrika, 49(2):??, September 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Fahrmeir:1999:PLS

[1003] Ludwig Fahrmeir and Rita Künstler.

Penalized likelihood smoothing in robust state space models. Metrika, 49 (3):173–191, October 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050007.

Liski:1999:DOD

[1004] Erkki P. Liski, Arto Luoma, and Alexander Zaigraev. Distance optimality design criterion in linear models. Metrika, 49(3):193–211, October 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840050008.

Rousseeuw:1999:DFP

[1005] Peter J. Rousseeuw and Ida Ruts. The depth function of a population distribution. Metrika, 49(3):213–244, October 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00020903.

Snethlage:1999:BRH

[1006] Martin Snethlage. Is bootstrap really helpful in point process statistics? Metrika, 49(3):245–255, October 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840050010.

Huang:1999:RSP

[1007] Wen-Jang Huang and Jyh-Cherng Su. Reverse submartingale property arising from exchangeable random variables. Metrika, 49(3):257–262, October 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00020901.

Anonymous:1999:HCc

[1008] Anonymous. Help & contacts. Metrika, 49(3):??, October 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Munk:1999:OIC

[1009] Axel Munk. Optimal inference for circular variation diminishing experiments with applications to the von-Mises distribution and the Fisher– Efron parabola model. Metrika, 50 (1):1–17, November 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050032.

Ioannides:1999:ECM

[1010] D. A. Ioannides. Estimating the conditional mode of a stationary stochastic process from noisy observations. Metrika, 50(1):19–35, November 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840050033.

Prescott:1999:ARD

[1011] Philip Prescott and Rahim Shahlaee.

The analysis of ranked data in blocked factorial experiments. Metrika, 50 (1):37–54, November 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050034.

Romano:1999:SIM

[1012] Joseph P. Romano and Michael Wolf.

Subsampling inference for the mean in the heavy-tailed case. Metrika, 50 (1):55–69, November 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840050035.

Zou:1999:VEU

[1013] Guohua Zou. Variance estimation for unequal probability sampling. Metrika, 50(1):71–82, November 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840050036.

Anonymous:1999:HCd

[1014] Anonymous. Help & contacts. Metrika, 50(1):??, November 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Parsian:1999:EMS

[1015] Ahmad Parsian and N. Sanjari Farsipour. Estimation of the mean of the selected population under asymmetric loss function. Metrika, 50(2): 89–107, December 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840050037.

LaMotte:1999:CHD

[1016] Lynn Roy LaMotte. Collapsibility hypotheses and diagnostic bounds in regression analysis. Metrika, 50(2): 109–119, December 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840050038.

Braun:1999:RLD

[1017] W. John Braun. Run length distributions for estimated attributes charts.

Metrika, 50(2):121–129, December 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840050039.

Plachky:1999:NCE

[1018] Detlef Plachky and Andrew L. Rukhin.

Nonparametric covariance estimation in multivariate distributions. Metrika, 50(2):131–136, December 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840050040.

Gob:1999:RLA

[1019] Rainer Göb, Enrique Del Castillo, and Klaus Dräger. Run length analysis of Shewhart charts applied to drifting processes under an integrative SPC/ EPC model. Metrika, 50 (2):137–161, December 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050041.

Rudas:1999:MIF

[1020] Tamás Rudas. The mixture index of fit and minimax regression. Metrika, 50 (2):163–172, December 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050042.

Anonymous:1999:HCe

[1021] Anonymous. Help & contacts. Metrika, 50(2):??, December 1999. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Aaberge:2000:UUT

[1022] Rolf Aaberge. UMP unbiased tests for multiparameter testing problems with restricted alternatives. Metrika, 50(3):179–193, April 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840050043.

Bischoff:2000:AOT

[1023] Wolfgang Bischoff. Asymptotically optimal tests for some growth curve models under non-normal error structure. Metrika, 50(3):195–203, April 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840050044.

Wojtasik:2000:CAQ

[1024] Leszek Wojtasik and Stefan Zontek. On construction of admissible quadratic estimators. Metrika, 50 (3):205–209, April 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840050045.

Fernandez:2000:MLP

[1025] Arturo J. Fernández. On maximum likelihood prediction based on type II doubly censored exponential data. Metrika, 50(3):211–220, April 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840050046.

Wu:2000:SMR

[1026] Lang Wu, Jaxk Reeves, and Michael D. Perlman. Some model-restricted shrinkage estimators for contingency tables with missing data. Metrika, 50(3):221–245, April 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050047.

Huang:2000:SEP

[1027] Wen-Jang Huang and Wen-Ching Chang. On a study of the exponential and Poisson characteristics of the Poisson process. Metrika, 50 (3):247–254, April 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050048.

Ma:2000:NAC

[1028] Chang-Xing Ma, Kai-Tai Fang, and Erkki Liski. A new approach in constructing orthogonal and nearly orthogonal arrays. Metrika, 50 (3):255–268, April 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840050049.

Anonymous:2000:HCa

[1029] Anonymous. Help & contacts. Metrika, 50(3):??, April 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Unwin:2000:I

[1030] Antony Unwin and Claus Weihs. Introduction. Metrika, 51(1):1–4, July 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00003976.

Huber:2000:CID

[1031] Peter J. Huber. Contradictions of interactive data analysis. Metrika, 51(1):5–9, July 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840000040.

Hofmann:2000:ECD

[1032] Heike Hofmann. Exploring categorical data: interactive mosaic plots. Metrika, 51(1):11–26, July 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840000041.

Ward:2000:VTE

[1033] Matthew O. Ward and Benjamin N. Lipchak. A visualization tool for exploratory analysis of cyclic multivariate data. Metrika, 51(1):27–37, July 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840000042.

Kim:2000:IVH

[1034] Sung-Soo Kim, Sunhee Kwon, and Dianne Cook. Interactive visualization of hierarchical clusters using MDS and MST. Metrika, 51(1):39–51, July 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840000043.

Blauth:2000:IAH

[1035] Angelika Blauth, Iris Pigeot, and François Bry. Interactive analysis of high-dimensional association structures with graphical models. Metrika, 51(1):53–65, July 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840000044.

Lane:2000:SCS

[1036] David M. Lane and David W. Scott.

Simulations, case studies, and an online text: a web-based resource for teaching statistics. Metrika, 51(1):67–90, July 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840000045.

Velleman:2000:DPT

[1037] Paul F. Velleman. Design principles for technology-based statistics education. Metrika, 51(1):91–104, July 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00003977.

Anonymous:2000:HCb

[1038] Anonymous. Help & contacts.

Metrika, 51(1):??, July 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Sultan:2000:RVG

[1039] Khalaf S. Sultan and Mohamed E.

Moshref. Record values from generalized Pareto distribution and associated inference. Metrika, 51(2): 105–116, August 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840000025.

Baran:2000:CEG

[1040] Sándor Baran. A consistent estimator in general functional errorsin-variables models. Metrika, 51 (2):117–132, August 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840000030.

Dette:2000:NST

[1041] Holger Dette and Natalie Neumeyer. A note on a specification test of independence. Metrika, 51(2):133–144, August 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840000031.

Liu:2000:ECB

[1042] Shuangzhe Liu. Efficiency comparisons between two estimators based on matrix determinant Kantorovichtype inequalities. Metrika, 51(2): 145–155, August 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840000050.

Dahlhaus:2000:GIM

[1043] Rainer Dahlhaus. Graphical interaction models for multivariate time series. Metrika, 51(2):157–172, August 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840000055.

Kourouklis:2000:UAS

[1044] Stavros Kourouklis. On universal admissibility of scale parameter estimators. Metrika, 51(2):173–179, August 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840000048.

Anonymous:2000:HCc

[1045] Anonymous. Help & contacts. Metrika, 51(2):??, August 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

DasGupta:2000:BCC

[1046] Anirban DasGupta. Best constants in Chebyshev inequalities with various applications. Metrika, 51(3): 185–200, September 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s184-000-8316-9.

Nather:2000:RFV

[1047] Wolfgang Näther. On random fuzzy variables of second order and their application to linear statistical inference with fuzzy data. Metrika, 51(3): 201–221, September 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840000047.

Hu:2000:RWL

[1048] Feifang Hu, William F. Rosenberger, and James V. Zidek. Relevance weighted likelihood for dependent data. Metrika, 51(3):223–243, September 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840000051.

Ruschendorf:2000:NAR

[1049] Ludger Rüschendorf and Ludger Uckelmann. Numerical and analytical results for the transportation problem of Monge–Kantorovich. Metrika, 51(3): 245–258, September 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840000052.

Zuo:2000:NFS

[1050] Yijun Zuo. A note on finite sample breakdown points of projection based multivariate location and scatter statistics. Metrika, 51(3):259– 265, September 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840000053.

Singh:2000:CIS

[1051] Sarjinder Singh and Stephen Horn. Compromised imputation in survey sampling. Metrika, 51(3):267– 276, September 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840000054.

Anonymous:2000:HCd

[1052] Anonymous. Help & contacts. Metrika, 51(3):??, September 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Szkutnik:2000:UIF

[1053] Zbigniew Szkutnik. Unfolding intensity function of a Poisson process in models with approximately specified folding operator. Metrika, 52(1):1–26, October 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00003978.

Gajek:2000:SBM

[1054] Lestaw Gajek and Andrzej Okolewski.

Sharp bounds on moments of generalized order statistics. Metrika, 52 (1):27–43, October 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840000058.

Qian:2000:SUC

[1055] Weimin Qian and Volker Mammitzsch.

Strong uniform convergence for the estimator of the regression function under φ-mixing conditions. Metrika, 52 (1):45–61, October 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840000059.

Raqab:2000:CPR

[1056] Mohammad Z. Raqab and Adnan M.

Awad. Characterizations of the Pareto and related distributions. Metrika, 52(1):63–67, October 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840000061.

Zuo:2000:FST

[1057] Yijun Zuo. Finite sample tail behavior of the multivariate trimmed mean based on Tukey–Donoho halfspace depth. Metrika, 52(1):69–75, October 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840000062.

Nummi:2000:AMG

[1058] Tapio Nummi and Jyrki Möttönen. On the analysis of multivariate growth curves. Metrika, 52(1):77–89, October 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840000063.

Anonymous:2000:HCe

[1059] Anonymous. Help & contacts. Metrika, 52(1):??, October 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Morales:2000:EWS

[1060] D. Morales, L. Pardo, M. C. Pardo, and I. Vajda. Extension of the Wald statistic to models with dependent observations. Metrika, 52(2): 97–113, December 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840000060.

Chaudhuri:2000:MSE

[1061] Arijit Chaudhuri, Arun Kumar Adhikary, and Shankar Dihidar. Mean square error estimation in multistage sampling. Metrika, 52(2): 115–131, December 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/PL00003979.

Gombay:2000:SSC

[1062] Edit Gombay. U-statistics for sequential change detection. Metrika, 52(2):

133–145, December 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/PL00003980.

Mandal:2000:CTV

[1063] Nripes K. Mandal, Kirti R. Shah, and Bikas K. Sinha. Comparison of test vs. control treatments using distance optimality criterion. Metrika, 52(2): 147–162, December 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840000073.

Takada:2000:SPE

[1064] Yoshikazu Takada. Sequential point estimation of normal mean under LINEX loss function. Metrika, 52(2): 163–171, December 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840000074.

Zhang:2000:SFR

[1065] Bao-Xue Zhang and Bai-Sen Liu. Some further remarks on a superiority problem in misspecified restricted singular linear models. Metrika, 52(2): 173–181, December 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/PL00003981.

Anonymous:2000:BR

[1066] Anonymous. Book reviews. Metrika, 52(2):183, December 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s001840000065.

Anonymous:2000:HCf

[1067] Anonymous. Help & contacts. Metrika, 52(2):??, December 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Neubauer:2000:OWD

[1068] Michael G. Neubauer, William Watkins, and Joel Zeitlin. D-optimal weighing designs for six objects. Metrika, 52 (3):185–211, December 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/PL00003982.

Jensen:2000:USD

[1069] D. R. Jensen. The use of Studentized diagnostics in regression. Metrika, 52 (3):213–223, December 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/PL00003983.

Chattopadhyay:2000:SEL

[1070] Saibal Chattopadhyay, Ajit Chaturvedi, and Raghu Nandan Sengupta. Sequential estimation of a linear function of normal means under asymmetric loss function. Metrika, 52(3): 225–235, December 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840000086.

Klar:2000:DST

[1071] Bernhard Klar. Diagnostic smooth tests of fit. Metrika, 52(3):237– 252, December 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/PL00003984.

Childs:2000:CIP

[1072] Aaron Childs and N. Balakrishnan. Conditional inference procedures for the Laplace distribution when the observed samples are progressively censored. Metrika, 52(3): 253–265, December 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840000092.

Anonymous:2000:HCg

[1073] Anonymous. Help & contacts. Metrika, 52(3):??, December 2000. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Chauhan:2001:ONR

[1074] N. Chauhan and R. J. Martin. Optimal nested row-column designs for blocks of size 2 × 2 under dependence. Metrika, 53(1):1–14, April 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00003985.

Menendez:2001:TBD

[1075] M. L. Menéndez, J. A. Pardo, and L. Pardo. Tests based on ϕ-divergences for bivariate symmetry. Metrika, 53(1):15–29, April 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840000088.

Chen:2001:ECP

[1076] Zhenmin Chen, Dietmar Ferger, and Jie Mi. Estimation of the change point of a distribution based on the number of failed test items. Metrika, 53(1):31–38, April 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840000090.

Gupta:2001:FRM

[1077] Pushpa L. Gupta and Ramesh C. Gupta. Failure rate of the minimum and maximum of a multivariate normal distribution. Metrika, 53(1):39–49, April 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840000093.

Bourke:2001:SSB

[1078] Patrick D. Bourke. Sample size and the binomial CUSUM control chart: the case of 100 inspection. Metrika, 53(1):51–70, April 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/PL00003986.

Scherb:2001:DUM

[1079] Hagen Scherb. Determination of uniformly most powerful tests in discrete sample spaces. Metrika, 53 (1):71–84, April 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840000095.

Ma:2001:NGA

[1080] Chang-Xing Ma and Kai-Tai Fang.

A note on generalized aberration in factorial designs. Metrika, 53 (1):85–93, April 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100112.

Anonymous:2001:BRa

[1081] Anonymous. Book reviews. Metrika, 53(1):95–100, April 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840000081.

Anonymous:2001:HCa

[1082] Anonymous. Help & contacts. Metrika, 53(1):??, April 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Croux:2001:MCR

[1083] Christophe Croux and Gentiane Haesbroeck. Maxbias curves of robust scale estimators based on subranges. Metrika, 53(2):101–122, May 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00003987.

Fernholz:2001:MHO

[1084] Luisa Turrin Fernholz. On multivariate higher order von Mises expansions. Metrika, 53(2):123–140, May 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840000097.

Ziegler:2001:BMN

[1085] Klaus Ziegler. On bootstrapping the mode in the nonparametric regression model with random design. Metrika, 53(2):141–170, May 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/PL00003988.

Happacher:2001:DDS

[1086] Max Happacher. The discrepancy distribution of stationary multiplier rules for rounding probabilities. Metrika, 53(2):171–181, May 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840000099.

Gupta:2001:MOD

[1087] Sudhir Gupta and Sanpei Kageyama. More optimal designs for diallel cross experiments. Metrika, 53 (2):183–187, May 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840100111.

Anonymous:2001:HCb

[1088] Anonymous. Help & contacts. Metrika, 53(2):??, May 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Ahmadi:2001:FIR

[1089] J. Ahmadi and N. R. Arghami. On the Fisher information in record values. Metrika, 53(3):195–206, July 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840000089.

Liski:2001:SCL

[1090] Erkki P. Liski and Alexander Zaigraev.

A stochastic characterization of Loewner optimality design criterion in linear models. Metrika, 53 (3):207–222, July 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840000106.

Jensen:2001:ESC

[1091] Uwe Jensen and Jürgen Wiedmann.

Estimation of a survival curve with randomly censored data in the presence of a covariate. Metrika, 53 (3):223–244, July 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840000107.

Morgenthaler:2001:MTW

[1092] Stephan Morgenthaler and John W.

Tukey. Multipolishing and two-way plots. Metrika, 53(3):245–267, July 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840000108.

Chang:2001:EPS

[1093] Horng-Jinh Chang and Kuo-Chung Huang. Estimation of proportion and sensitivity of a qualitative character. Metrika, 53(3):269–280, July 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840100109.

Anonymous:2001:HCc

[1094] Anonymous. Help & contacts. Metrika, 53(3):??, July 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Ghosh:2001:TGC

[1095] Subir Ghosh and Colleen Burns. Two general classes of search designs for factor screening experiments with factors at three levels. Metrika, 54 (1):1–17, August 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/PL00003989.

Arts:2001:SIS

[1096] Gerda R. J. Arts, Willem Albers, and Wilbert C. M. Kallenberg. Simultaneous inspection of several product characteristics. Metrika, 54 (1):19–41, August 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840100113.

Shalabh:2001:EBS

[1097] Shalabh. Estimation of bias and standard error of an improved estimator of normal mean. Metrika, 54 (1):43–51, August 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840100114.

Joarder:2001:ETS

[1098] A. H. Joarder and Sarjinder Singh. Estimation of the trace of the scaled covariance matrix of a multivariate t-model using a known information. Metrika, 54(1):53–58, August 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00003990.

Garcia:2001:EMS

[1099] M. Rueda García and A. Arcos Cebrián. On estimating the median from survey data using multiple auxiliary information. Metrika, 54(1):59–76, August 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100116.

Finner:2001:ISS

[1100] Helmut Finner and Klaus Strassburger.

Increasing sample sizes do not always increase the power of UMPUtests for 2 × 2 tables. Metrika, 54 (1):77–91, August 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100117.

Anonymous:2001:BRb

[1101] Anonymous. Book reviews. Metrika, 54(1):93–96, August 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100119.

Anonymous:2001:HCd

[1102] Anonymous. Help & contacts. Metrika, 54(1):??, August 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Barranco-Chamorro:2001:APM

[1103] I. Barranco-Chamorro, J. L. Moreno-Rebollo, and M. T. Gómez-Gómez. Asymptotic properties of MLE and UMVUE in one-parameter truncated distributions. Metrika, 54(2):97– 110, November 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100118.

vonCollani:2001:CNM

[1104] Elart von Collani and Monica Dumitrescu. Complete Neyman measurement procedure. Metrika, 54(2): 111–130, November 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100126.

Peiris:2001:MSA

[1105] M. Shelton Peiris and A. Thavaneswaran. Multivariate stable ARMA processes with time dependent coefficients. Metrika, 54(2): 131–138, November 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100127.

Riedle:2001:VEC

[1106] M. Riedle and J. Steinebach. Variance estimation in the change analysis of a linear regression model. Metrika, 54 (2):139–157, November 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100128.

Arnab:2001:EFP

[1107] Raghunath Arnab. Estimation of a finite population total in varying probability sampling for multicharacter surveys. Metrika, 54(2): 159–177, November 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100129.

Stufken:2001:PDB

[1108] John Stufken and James H. Wright.

Polygonal designs with blocks of size k ≤ 10. Metrika, 54(2): 179–184, November 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100137.

Anonymous:2001:HCe

[1109] Anonymous. Help & contacts. Metrika, 54(2):??, November 2001. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Heiligers:2002:TPR

[1110] Berthold Heiligers. Totally positive regression: E-optimal designs. Metrika, 54(3):191–213, February 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s184-002-8364-7.

Jones:2002:MBD

[1111] M. C. Jones. Multivariate t and beta distributions associated with the multivariate F distribution. Metrika, 54 (3):215–231, February 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s184-002-8365-4.

Chatterjee:2002:TLS

[1112] K. Chatterjee, Lih-Yuan Deng, and Dennis K. J. Lin. Two-level search design for main-effect plus two plan. Metrika, 54(3):233–245, February 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s184-002-8366-1.

Toutenburg:2002:UMR

[1113] H. Toutenburg and Shalabh. Use of minimum risk approach in the estimation of regression models with missing observations. Metrika, 54(3): 247–259, February 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s184-002-8367-y.

Pearn:2002:EPC

[1114] W. L. Pearn, P. C. Lin, and K. S. Chen. Estimating process capability index c”pmk for asymmetric tolerances: Distributional properties. Metrika, 54 (3):261–279, February 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s184-002-8368-8.

Anonymous:2002:BRa

[1115] Anonymous. Book reviews. Metrika, 54 (3):281–284, February 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s184-002-8369-5.

Anonymous:2002:HCa

[1116] Anonymous. Help & contacts. Metrika, 54(3):??, February 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Anonymous:2002:P

[1117] Anonymous. Preface. Metrika, 55(1–2):

1–2, April 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s001840200181.

Agullo:2002:EAC

[1118] J. Agulló. An exchange algorithm for computing the least quartile difference estimator. Metrika, 55(1–2):3–16, April 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200182.

Bassett:2002:QME

[1119] Gilbert W. Bassett, Jr., Mo-Yin S.

Tam, and Keith Knight. Quantile models and estimators for data analysis. Metrika, 55(1–2):17–26, April 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200183.

Bednarski:2002:EGP

[1120] Tadeusz Bednarski. Estimation in the generalized Poisson model via robust testing. Metrika, 55(1–2):27–36, April 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200184.

Arslan:2002:ACE

[1121] O. Arslan, O. Edlund, and H. Ekblom. Algorithms to compute CM- and S-estimates for regression. Metrika, 55(1–2):37–51, April 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840200185.

Fernholz:2002:RIR

[1122] Luisa Turrin Fernholz. Robustness issues regarding content-corrected tolerance limits. Metrika, 55(1– 2):53–66, April 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840200186.

He:2002:MES

[1123] Xuming He and Mi-Ok Kim. On marginal estimation in a semiparametric model for longitudinal data with time-independent covariates. Metrika, 55(1–2):67–74, April 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840200187.

Koul:2002:ADS

[1124] Hira L. Koul. Asymptotic distributions of some scale estimators in nonlinear models. Metrika, 55(1–2):75–90, April 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200188.

Morgenthaler:2002:OWE

[1125] Stephan Morgenthaler and Robert Staudte. Optimal weights of evidence with bounded influence. Metrika, 55 (1–2):91–97, April 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840200189.

Muller:2002:REE

[1126] Christine H. Müller. Robust estimators for estimating discontinuous functions. Metrika, 55(1–2):99–109, April 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200190.

Pison:2002:SSC

[1127] G. Pison, S. Van Aelst, and G. Willems.

Small sample corrections for LTS and MCD. Metrika, 55(1–2):111–123, April 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200191.

Willems:2002:RHT

[1128] G. Willems, G. Pison, P. J. Rousseeuw, and S. Van Aelst. A robust Hotelling test. Metrika, 55(1–2):125–138, April 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200192.

Lauprete:2002:RPO

[1129] G. J. Lauprete, A. M. Samarov, and R. E. Welsch. Robust portfolio optimization. Metrika, 55(1– 2):139–149, April 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840200193.

Anonymous:2002:HCb

[1130] Anonymous. Help & contacts. Metrika, 55(1–2):??, April 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Rezaei:2002:RCD

[1131] A. H. Rezaei and N. R. Arghami. Right censoring in a discrete life model. Metrika, 55(3):151–160, June 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840100135.

Yue:2002:DSS

[1132] Rong-Xian Yue and Fred J. Hickernell. Designs for smoothing spline ANOVA models. Metrika, 55(3):161–176, June 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840100136.

Bohning:2002:NMD

[1133] Dankmar Böhning and Roumporn Ruangroj. A note on the maximum deviation of the scale-contaminated normal to the best normal distribution. Metrika, 55(3):177–182, June 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840100138.

Nagar:2002:DCC

[1134] Daya K. Nagar and María Eugenia Castañeda. Distribution of correlation coefficient under mixture normal model. Metrika, 55(3):183–190, June 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840100139.

Bhargava:2002:ECC

[1135] Manoj Bhargava and Ravindra Singh.

On the efficiency comparison of certain randomized response strategies. Metrika, 55(3):191–197, June 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840100140.

Varghese:2002:RWS

[1136] Cini Varghese, A. R. Rao, and V. K.

Sharma. Robustness of Williams square change-over designs. Metrika, 55(3):199–208, June 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100141.

Moreno-Rebollo:2002:IDS

[1137] J. L. Moreno-Rebollo, A. Muñoz-Reyes, M. D. Jiménez-Gamero, and J. Muñoz-Pichardo. Influence diagnostic in survey sampling: Estimating the conditional bias. Metrika, 55(3):209–214, June 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100142.

Uno:2002:SPE

[1138] Chikara Uno and Eiichi Isogai. Sequential point estimation of the powers of a normal scale parameter. Metrika, 55(3):215–232, June 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100143.

Pazman:2002:MNB

[1139] A. Pázman and J.-B. Denis. Measures of nonlinearity for a biadditive ANOVA model. Metrika, 55 (3):233–245, June 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840100144.

Castillo:2002:CMD

[1140] Enrique Castillo, Ali S. Hadi, Beatriz Lacruz, and José María Sarabia. Constrained mixture distributions. Metrika, 55(3):247–269, June 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840100149.

Anonymous:2002:BRb

[1141] Anonymous. Book reviews. Metrika, 55(3):271–278, June 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840100146.

Anonymous:2002:HCc

[1142] Anonymous. Help & contacts. Metrika, 55(3):??, June 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Ozawa:2002:OEI

[1143] Kazuhiro Ozawa, Masakazu Jimbo, Sanpei Kageyama, and Stanisław Mejza. Optimality and efficiency of incomplete split-block designs. Metrika, 56(1):1–18, July 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100150.

Yang:2002:ITB

[1144] Zhenlin Yang, Stanley P. See, and Min Xie. An investigation of transformation-based prediction interval for the Weibull median life. Metrika, 56(1):19–29, July 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840100151.

Mohammadi:2002:MIE

[1145] Leila Mohammadi and Willem R. van Zwet. Minimax invariant estimation of a continuous distribution function under entropy loss. Metrika, 56(1):31–42, July 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100152.

Bobecka:2002:TDR

[1146] Konstancja Bobecka and Jacek Wesolowski.

Three dual regression schemes for the Lukacs theorem. Metrika, 56 (1):43–54, July 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100153.

Bairamov:2002:DSS

[1147] Ismihan Bairamov and Samuel Kotz.

Dependence structure and symmetry of Huang–Kotz FGM distributions and their extensions. Metrika, 56(1):55–72, July 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840100158.

Krishnamoorthy:2002:CIP

[1148] K. Krishnamoorthy and Brett C. Moore. Combining information for prediction in linear regression. Metrika, 56(1):73–81, July 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840100159.

Bhar:2002:RRC

[1149] Lalmohan Bhar and V. K. Gupta. Robust row-column designs for complete diallel cross experiments. Metrika, 56(1):83–90, July 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840100161.

Anonymous:2002:BRc

[1150] Anonymous. Book reviews. Metrika, 56(1):91–95, July 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840100163.

Anonymous:2002:HCd

[1151] Anonymous. Help & contacts. Metrika, 56(1):??, July 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Gil:2002:LSF

[1152] María Angeles Gil, María Asunción Lubiano, Manuel Montenegro, and María Teresa López. Least squares fitting of an affine function and strength of association for interval-valued data.

Metrika, 56(2):97–111, October 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840100160.

Pazman:2002:ODN

[1153] Andrej Pázman. Optimal design of nonlinear experiments with parameter constraints. Metrika, 56(2): 113–130, October 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100162.

Singh:2002:NSR

[1154] Sarjinder Singh. A new stochastic randomized response model. Metrika, 56 (2):131–142, October 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100168.

Misra:2002:SEE

[1155] Neeraj Misra, P. K. Choudhary, I. D.

Dhariyal, and D. Kundu. Smooth estimators for estimating order restricted scale parameters of two gamma distributions. Metrika, 56 (2):143–161, October 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100169.

Huang:2002:FFD

[1156] Mong-Na Lo Huang and Kam-Fai Wong. s^{k−p} Fractional factorial designs in s^{b} blocks. Metrika, 56(2): 163–170, October 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840100170.

Bouza:2002:EMR

[1157] Carlos N. Bouza. Estimation of the mean in ranked set sampling with non responses. Metrika, 56 (2):171–179, October 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840100171.

Anonymous:2002:HCe

[1158] Anonymous. Help & contacts. Metrika, 56(2):??, October 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Ullrich:2002:OPC

[1159] Peter Ullrich. Officers, playing cards, and sheep. Metrika, 56(3): 189–204, December 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840100172.

Yamada:2002:CML

[1160] Shu Yamada and Dennis K. J. Lin. Construction of mixed-level supersaturated design. Metrika, 56(3): 205–214, December 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840100173.

Henschel:2002:SIS

[1161] Volkmar Henschel. Statistical inference in simplicially contoured sample distributions. Metrika, 56(3):

215–228, December 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100174.

Ting:2002:OER

[1162] Chao-Ping Ting. Optimal and efficient repeated measurements designs for comparing test treatments with a control. Metrika, 56(3):229–238, December 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840100175.

Srivastav:2002:OTB

[1163] Sudesh K. Srivastav and John P. Morgan. On the E-optimality of truncated block designs. Metrika, 56(3): 239–245, December 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100176.

Shin:2002:EOR

[1164] Dong Wan Shin, Dai-Gyoung Kim, and Han Joon Kim. Efficiency of the OLSE for regressions on two-dimensional grids with sinusoidal regressors and spatially correlated errors. Metrika, 56 (3):247–258, December 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840100177.

Zaigraev:2002:SOD

[1165] Alexander Zaigraev. Shape optimal design criterion in linear models. Metrika, 56(3):259–273, December 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840100179.

Anonymous:2002:BRd

[1166] Anonymous. Book reviews. Metrika, 56 (3):275–280, December 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840200212.

Anonymous:2002:HCf

[1167] Anonymous. Help & contacts. Metrika, 56(3):??, December 2002. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Mayoral:2003:EET

[1168] A. M. Mayoral, D. Morales, J. Morales, and I. Vajda. On efficiency of estimation and testing with data quantized to fixed number of cells. Metrika, 57(1):1–27, February 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840100178.

Evangelaras:2003:IPH

[1169] H. Evangelaras, S. Georgiou, and C. Koukouvinos. Inequivalent projections of Hadamard matrices of orders 16 and 20. Metrika, 57(1): 29–35, February 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840100180.

Fang:2003:CMG

[1170] Kai-Tai Fang, Gen-Nian Ge, Min-Qian Liu, and Hong Qin. Construction of minimum generalized aberration designs. Metrika, 57(1):37–50, February 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200197.

Carsoule:2003:NMT

[1171] Frédéric Carsoule and Philip Hans Franses. A note on monitoring timevarying parameters in an autoregression. Metrika, 57(1):51–62, February 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200198.

Mohammadi:2003:API

[1172] Leila Mohammadi. On an admissibility problem involving the moment generating function of the lognormal distribution. Metrika, 57 (1):63–70, February 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840200199.

Park:2003:AFI

[1173] Sangun Park. On the asymptotic Fisher information in order statistics. Metrika, 57(1):71–80, February 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200200.

Arnold:2003:TFL

[1174] Bernhard F. Arnold and Oke Gerke.

Testing fuzzy linear hypotheses in linear regression models. Metrika, 57 (1):81–95, February 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840200201.

Ickstadt:2003:BR

[1175] Katja Ickstadt, Uwe Jensen, Michael Kohler, Rolf-Dieter Reiss, G. Last, Jochen Mau, and R. Fried. Book reviews. Metrika, 57(1):97–104, February 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840200225.

Anonymous:2003:HCa

[1176] Anonymous. Help & contacts. Metrika, 57(1):??, February 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Watson:2003:LSR

[1177] G. A. Watson. On the linear statistical relationship. Metrika, 57 (2):105–114, May 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840200202.

SahaRay:2003:DOD

[1178] Rita SahaRay and Subir Kumar Bhandari. DS-optimal designs in one way ANOVA. Metrika, 57(2):115–125, May 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840200203.

Das:2003:POA

[1179] Ashish Das, Aloke Dey, and Chand K. Midha. On a property of orthogonal arrays and optimal blocking of fractional factorial plans. Metrika, 57(2):127–135, May 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840200204.

Kunert:2003:ODU

[1180] J. Kunert, R. J. Martin, and S. Pooladsaz. Optimal designs under two related models for interference. Metrika, 57(2):137–143, May 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840200205.

Stehlik:2003:DET

[1181] Milan Stehlík. Distributions of exact tests in the exponential family. Metrika, 57(2):145–164, May 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200206.

Al-Hussaini:2003:BPD

[1182] Essam K. Al-Hussaini and Abd EL-Baset A. Ahmad. On Bayesian predictive distributions of generalized order statistics. Metrika, 57(2):165–176, May 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200207.

Hofmann:2003:FIR

[1183] Glenn Hofmann and H. N. Nagaraja.

Fisher information in record data. Metrika, 57(2):177–193, May 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200208.

Christofides:2003:GRR

[1184] Tasos C. Christofides. A generalized randomized response technique.

Metrika, 57(2):195–200, May 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840200216.

Muller:2003:BR

[1185] Christine Muller. Book reviews. Metrika, 57(2):201, May 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s001840200242.

Novak:2003:BR

[1186] E. Novak. Book reviews. Metrika, 57 (2):202, May 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s001840200243.

Zhigljavsky:2003:BR

[1187] Anatoly Zhigljavsky. Book reviews. Metrika, 57(2):203–205, May 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840200245.

Konig:2003:BR

[1188] Inke R. Konig and Andreas Ziegler. Book reviews. Metrika, 57(2): 206–208, May 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840200246.

Anonymous:2003:HCb

[1189] Anonymous. Help & contacts. Metrika, 57(2):??, May 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Aly:2003:TMC

[1190] Emad-Eldin A. A. Aly, Abd-Elnaser S.

Abd-Rabou, and Noriah M. Al-Kandari. Tests for multiple change points under ordered alternatives. Metrika, 57(3):209–221, July 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200209.

Kim:2003:PTE

[1191] H. M. Kim and A. K. Md. E. Saleh.

Preliminary test estimators of the parameters of simple linear model with measurement error. Metrika, 57(3):223–251, July 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840200215.

Kukush:2003:CEB

[1192] A. Kukush, I. Markovsky, and S. Van Huffel. Consistent estimation in the bilinear multivariate errors-in-variables model. Metrika, 57(3):253–285, July 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200217.

Zaigraev:2003:ISO

[1193] Alexander Zaigraev. Integral stochastic optimal design criteria in linear models. Metrika, 57(3):287–301, July 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200218.

Pearn:2003:BLE

[1194] W. L. Pearn and G. H. Lin. A Bayesian-like estimator of the process capability index c_{pmk}. Metrika, 57(3):303–312, July 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840200219.

Cramer:2003:BR

[1195] Erhard Cramer. Book reviews. Metrika, 57(3):313–314, July 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        accesspage/article/10.1007/s001840200247.

Georgii:2003:BR

[1196] Hans-Otto Georgii. Book reviews. Metrika, 57(3):314, July 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s001840200248.

Kuhne:2003:BR

[1197] Rolf Kühne. Book reviews. Metrika, 57(3):315–316, July 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s001840200249.

Anonymous:2003:HCc

[1198] Anonymous. Help & contacts. Metrika, 57(3):??, July 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Ivanova:2003:PWT

[1199] Anastasia Ivanova. A play-the-winner-type urn design with reduced variability. Metrika, 58(1):1–13, August 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200220.

Abu-Taleb:2003:PIP

[1200] Ahmad Abu-Taleb, M. Bhaskara Rao, and Haimeng Zhang. Periodic inspection plans: The case of Weibull distribution. Metrika, 58(1):15–30, August 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200221.

Steinberg:2003:OSM

[1201] David M. Steinberg and Nitzan Rabinowitz. Optimal seismic monitoring for event location with application to on site inspection of the comprehensive nuclear test ban treaty. Metrika, 58(1):31–57, August 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840200222.

Seier:2003:TFK

[1202] Edith Seier and Douglas Bonett. Two families of kurtosis measures. Metrika, 58(1):59–70, August 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840200223.

Nather:2003:EOR

[1203] Wolfgang Näther and Jaroslav Simák.

Effective observation of random processes using derivatives. Metrika, 58 (1):71–84, August 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840200224.

vanDorp:2003:GTD

[1204] J. René van Dorp and Samuel Kotz. Generalized trapezoidal distributions. Metrika, 58(1):85–97, August 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840200230.

Pap:2003:BR

[1205] Gyula Pap. Book reviews. Metrika, 58(1):99–106, August 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840200250.

Anonymous:2003:HCd

[1206] Anonymous. Help & contacts. Metrika, 58(1):??, August 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Jacroux:2003:NDO

[1207] Mike Jacroux. A note on the determination of A-optimal type block designs under fixed and mixed effects models. Metrika, 58(2):107–115, September 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840200231.

Menendez:2003:SBP

[1208] M. L. Menéndez, L. Pardo, Ch. Tsairidis, and K. Zografos. Selection of the best population: an information theoretic approach. Metrika, 58 (2):117–147, September 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840200232.

Nguyen:2003:CND

[1209] Truc T. Nguyen and Khoan T. Dinh.

Characterizations of normal distributions and EDF goodness-of-fit tests. Metrika, 58(2):149–157, September 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200233.

Kaluszka:2003:SEE

[1210] M. Kaluszka and A. Okolewski.

Sharp exponential and entropy bounds on expectations of generalized order statistics. Metrika, 58(2):159– 171, September 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840200234.

Misra:2003:EMS

[1211] Neeraj Misra and Edward C. van der Meulen. On estimating the mean of the selected normal population under the LINEX loss function. Metrika, 58 (2):173–183, September 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840200235.

Raghavarao:2003:HAH

[1212] Damaraju Raghavarao and Stan Altan. A heuristic analysis of highly fractionated 2^{n} factorial experiments. Metrika, 58(2):185–191, September 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840200236.

Melas:2003:DEE

[1213] Viatcheslav B. Melas, Andrey Pepelyshev, and Russell C. H. Cheng. Designs for estimating an extremal point of quadratic regression models in a hyperball. Metrika, 58(2): 193–208, September 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840200237.

Becker:2003:BR

[1214] C. Becker. Book reviews. Metrika, 58 (2):209–212, September 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840300288.

Anonymous:2003:HCe

[1215] Anonymous. Help & contacts. Metrika, 58(2):??, September 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Norell:2003:ONL

[1216] Lennart Norell. On the optimum number of levels in the one-way model with random effects. Metrika, 58(3): 213–220, December 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840200238.

Piperigou:2003:TBD

[1217] Violetta E. Piperigou and H. Papageorgiou. On truncated bivariate discrete distributions: A unified treatment.

Metrika, 58(3):221–233, December 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840200239.

Jureckova:2003:GFT

[1218] Jana Jurecková, Jan Picek, and Pranab Kumar Sen. Goodness-of-fit test with nuisance regression and scale. Metrika, 58(3):235–258, December 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840300262.

Bieniek:2003:CDL

[1219] Mariusz Bieniek and Dominik Szynal. Characterizations of distributions via linearity of regression of generalized order statistics. Metrika, 58(3): 259–271, December 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300263.

Jainz:2003:NPP

[1220] M. Jainz. A note on the projection properties of Hadamard matrices obtained by the first Payley construction. Metrika, 58(3): 273–277, December 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300265.

Fang:2003:OML

[1221] Kai-Tai Fang, Dennis K. J. Lin, and Min-Qian Liu. Optimal mixed-level supersaturated design. Metrika, 58(3): 279–291, December 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840300266.

Cramer:2003:MDS

[1222] Erhard Cramer and Udo Kamps. Marginal distributions of sequential and generalized order statistics. Metrika, 58(3):293–310, December 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840300268.

Anonymous:2003:BR

[1223] Anonymous. Book reviews. Metrika, 58 (3):311–314, December 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840300291.

Anonymous:2003:HCf

[1224] Anonymous. Help & contacts. Metrika, 58(3):??, December 2003. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Huang:2004:MRS

[1225] Kuo-Chung Huang. Mixed random systematic sampling designs. Metrika, 59(1):1–11, February 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840300264.

Chinipardaz:2004:NDT

[1226] Rahim Chinipardaz and Trevor Cox. Nonparametric discrimination of time series data. Metrika, 59(1):13–20, February 2004. CODEN MTRKA8.

ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

springer.com/article/10.1007/s001840300267.

Filzmoser:2004:THF

[1227] P. Filzmoser and R. Viertl. Testing hypotheses with fuzzy data: The fuzzy pvalue. Metrika, 59(1):21–29, February 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840300269.

Montenegro:2004:ABT

[1228] Manuel Montenegro, Ana Colubi, María Rosa Casals, and María Ángeles Gil. Asymptotic and bootstrap techniques for testing the expected value of a fuzzy random variable. Metrika, 59(1):31–49, February 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300270.

Evangelaras:2004:EIP

[1229] H. Evangelaras, S. Georgiou, and C. Koukouvinos. Evaluation of inequivalent projections of Hadamard matrices of order 24. Metrika, 59 (1):51–73, February 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300271.

Kukush:2004:CEW

[1230] Alexander Kukush and Sabine Van Huffel. Consistency of elementwiseweighted total least squares estimator in a multivariate errors-in-variables model AX=B. Metrika, 59 (1):75–97, February 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840300272.

Anonymous:2004:BRa

[1231] Anonymous. Book reviews. Metrika, 59(1):99–104, February 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840300313.

Anonymous:2004:HCa

[1232] Anonymous. Help & contacts. Metrika, 59(1):??, February 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Kibria:2004:PTR

[1233] B. M. Golam Kibria and A. K. Md. E. Saleh. Preliminary test ridge regression estimators with Student’s t errors and conflicting test-statistics. Metrika, 59(2):105–124, May 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840300273.

Zhang:2004:GML

[1234] Haimeng Zhang and M. Bhaskara Rao. On generalized maximum likelihood estimation in the proportional hazards model with partially informative censoring. Metrika, 59(2):125–136, May 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840300274.

Irle:2004:NST

[1235] A. Irle and V. I. Lotov. A nonsymmetric sequential test. Metrika, 59(2):137–146, May 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300275.

Belzunce:2004:SRR

[1236] Felix Belzunce, Jorge Navarro, José M.

Ruiz, and Yolanda del Aguila. Some results on residual entropy function. Metrika, 59(2):147–161, May 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840300276.

Baklizi:2004:SEE

[1237] Ayman Baklizi and Abed El Qader El-Masri. Shrinkage estimation of P (X < Y) in the exponential case with common location parameter. Metrika, 59(2):163–171, May 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300277.

Eskandari:2004:EBA

[1238] Farzad Eskandari and Mohammad R.

Meshkani. Empirical Bayes analysis of log-linear models for a generalized finite stationary Markov chain. Metrika, 59(2):173–191, May 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300278.

Chen:2004:NSB

[1239] Kung-Yu Chen and Chien-Tai Lin. A note on strategies for bandit problems with infinitely many arms. Metrika, 59(2):193–203, May 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840300279.

Anonymous:2004:BRb

[1240] Anonymous. Book reviews. Metrika, 59 (2):205, May 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

                                        springer.com/accesspage/article/
                                        10.1007/s001840300316.

Anonymous:2004:HCb

[1241] Anonymous. Help & contacts. Metrika, 59(2):??, May 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Albers:2004:ESC

[1242] Willem Albers and Wilbert C. M. Kallenberg. Estimation in Shewhart control charts: effects and corrections. Metrika, 59(3):207–234, June 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840300280.

Lopez-Fidalgo:2004:EMD

[1243] J. López-Fidalgo and J. M. Rodríguez-Díaz. Elfving’s method for mdimensional models. Metrika, 59 (3):235–244, June 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840300281.

Wang:2004:SUC

[1244] Li Wang and Han-Ying Liang. Strong uniform convergence of the recursive regression estimator under ϕ-mixing conditions. Metrika, 59(3):245–261, June 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840300282.

Viertl:2004:GBT

[1245] Reinhard Viertl and Dietmar Hareter.

Generalized Bayes’ theorem for nonprecise a-priori distribution. Metrika, 59(3):263–273, June 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300283.

Barakat:2004:APR

[1246] H. M. Barakat, E. M. Nigm, and M. E. El-Adll. Asymptotic properties of random extremes under general normalization from nonidentical distributions. Metrika, 59(3):275–287, June 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840300284.

Tsao:2004:STP

[1247] C. Andy Tsao and Yu-Ling Tseng. A statistical treatment of the problem of division. Metrika, 59(3):289–303, June 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840300285.

Thomas:2004:BSP

[1248] Alice Thomas and K. K. Jose. Bivariate semi-Pareto minification processes. Metrika, 59(3):305–313, June 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840300287.

Anonymous:2004:HCc

[1249] Anonymous. Help & contacts. Metrika, 59(3):??, June 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Kus:2004:NDE

[1250] Václav Kus. Nonparametric density estimates consistent of the order of n^{−1/2} in the L_{1}-norm. Metrika, 60(1):1–14, July 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840300286.

Petropoulos:2004:IEE

[1251] Constantinos Petropoulos and Stavros Kourouklis. Improved estimation of extreme quantiles in the multivariate Lomax (Pareto II) distribution. Metrika, 60(1):15–24, July 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840300293.

Pewsey:2004:LSJ

[1252] Arthur Pewsey. The large-sample joint distribution of key circular statistics. Metrika, 60(1):25–32, July 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840300294.

Ferger:2004:MAP

[1253] Dietmar Ferger. Maximal asymptotic power and efficiency of twosample tests based on generalized Ustatistics. Metrika, 60(1):33–57, July 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840300295.

Qin:2004:DDF

[1254] Hong Qin and Kai-Tai Fang. Discrete discrepancy in factorial designs. Metrika, 60(1):59–72, July 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840300296.

Beyersmann:2004:ERT

[1255] Jan Beyersmann, Arnold Janssen, and Claus-Dieter Mayer. Efficient rank tests for semiparametric competing risk models. Metrika, 60(1):73–91, July 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840300297.

Wang:2004:NCU

[1256] Dabuxilatu Wang. A note on consistency and unbiasedness of point estimation with fuzzy data. Metrika, 60(1):93–104, July 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300298.

Anonymous:2004:HCd

[1257] Anonymous. Help & contacts.

Metrika, 60(1):??, July 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Stulajter:2004:EVO

[1258] Frantisek Stulajter and Viktor Witkovský.

Estimation of variances in orthogonal finite discrete spectrum linear regression models. Metrika, 60(2):

105–118, September 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300299.

Pearn:2004:IAT

[1259] W. L. Pearn, P. C. Lin, and K. S. Chen. The C′′_{pk} index for asymmetric tolerances: Implications and inference. Metrika, 60(2):119–136, September 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840300300.

Harman:2004:MED

[1260] Radoslav Harman. Minimal efficiency of designs under the class of orthogonally invariant information criteria. Metrika, 60(2):137–153, September 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840300301.

Liang:2004:ANR

[1261] Han-Ying Liang and Jong-Il Baek. Asymptotic normality of recursive density estimates under some dependence assumptions. Metrika, 60(2): 155–166, September 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300302.

Kratschmer:2004:PTF

[1262] Volker Krätschmer. Probability theory in fuzzy sample spaces. Metrika, 60 (2):167–189, September 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300303.

Schwingenschlogl:2004:SAD

[1263] Udo Schwingenschlögl and Mathias Drton. Seat allocation distributions and seat biases of stationary apportionment methods for proportional representation. Metrika, 60(2): 191–202, September 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840400347.

Anonymous:2004:BRc

[1264] Anonymous. Book reviews. Metrika, 60 (2):203–209, September 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840400337.

Anonymous:2004:HCe

[1265] Anonymous. Help & contacts. Metrika, 60(2):??, September 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Lopez-Blazquez:2004:OEG

[1266] Fernando López-Blázquez and Antonia Castaño-Martínez. Orthogonal expansions for the generalized Cramér– von Mises statistics. Metrika, 60(3): 211–221, November 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300304.

Chaudhuri:2004:CRR

[1267] Arijit Chaudhuri. Christofides’ randomized response technique in complex sample surveys. Metrika, 60(3):

223–228, November 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300305.

Steland:2004:SCT

[1268] Ansgar Steland. Sequential control of time series by functionals of kernelweighted empirical processes under local alternatives. Metrika, 60(3): 229–249, November 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300306.

Varghese:2004:SRD

[1269] Cini Varghese and V. K. Sharma. A series of resolvable PBIB(3) designs with two replicates. Metrika, 60(3): 251–254, November 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300307.

Bar-Lev:2004:NRR

[1270] Shaul K. Bar-Lev, Elizabeta Bobovitch, and Benzion Boukai. A note on randomized response models for quantitative data. Metrika, 60(3):255– 260, November 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300308.

Osius:2004:ABT

[1271] Gerhard Osius. The association between two random elements: A complete characterization and odds ratio models. Metrika, 60(3):261– 277, November 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300309.

Ai:2004:PJG

[1272] Ming-Yao Ai and Run-Chu Zhang. Projection justification of generalized minimum aberration for asymmetrical fractional factorial designs. Metrika, 60 (3):279–285, November 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300310.

Ahmad:2004:FSB

[1273] Ibrahim A. Ahmad, Wei Liu, and Ying Zhang. On the finite sample behavior of testing whether new is better than used of a specified age. Metrika, 60 (3):287–294, November 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300311.

Li:2004:NDH

[1274] Yingfu Li and Jiantian Wang. Nonregular designs from Hadamard matrices and their estimation capacity. Metrika, 60(3):295–303, November 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840300312.

Willie:2004:BR

[1275] H. Willie. Book reviews. Metrika, 60 (3):305–311, November 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840300317.

Anonymous:2004:HCf

[1276] Anonymous. Help & contacts. Metrika, 60(3):??, November 2004. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Kim:2005:TSS

[1277] Jong-Min Kim and M. E. Elam. A two-stage stratified Warner’s randomized response model using optimal allocation. Metrika, 61(1):1–7, February 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840400319.

Kuriki:2005:RSB

[1278] Shinji Kuriki, Iwona Mejza, Masakazu Jimbo, Stanisław Mejza, and Kazuhiro Ozawa. Resolvable semi-balanced incomplete split-block designs. Metrika, 61(1):9–16, February 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840400320.

Filipiak:2005:OEC

[1279] K. Filipiak and A. Markiewicz. Optimality and efficiency of circular neighbor balanced designs for correlated observations. Metrika, 61 (1):17–27, February 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840400321.

Henze:2005:RCT

[1280] Norbert Henze and Simos G. Meintanis. Recent and classical tests for exponentiality: a partial review with comparisons. Metrika, 61(1):

29–45, February 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840400322.

Kumar:2005:TIB

[1281] Usha A. Kumar and A. Subramanyam.

Tests of independence in a bivariate exponential distribution. Metrika, 61 (1):47–62, February 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840400323.

Kuroki:2005:BAC

[1282] Manabu Kuroki. Bounds on average causal effects in studies with a latent response variable. Metrika, 61 (1):63–71, February 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840400324.

Gonzalez-Barrios:2005:RAD

[1283] José M. González-Barrios and Silvia Ruiz-Velasco. Regression analysis and dependence. Metrika, 61 (1):73–87, February 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840400325.

Sliwa:2005:MCC

[1284] Przemyslaw Śliwa and Wolfgang Schmid. Monitoring the crosscovariances of a multivariate time series. Metrika, 61(1):89–115, February 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840400326.

Anonymous:2005:HCa

[1285] Anonymous. Help & contacts. Metrika, 61(1):??, February 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Hamedani:2005:CCU

[1286] G. G. Hamedani and Hans Volkmer. Certain characterizations of the uniform distribution. Metrika, 61 (2):117–125, April 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840400244.

Li:2005:COI

[1287] Peng-Fei Li, Min-Qian Liu, and Run-Chu Zhang. Choice of optimal initial designs in sequential experiments. Metrika, 61(2):127–135, April 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840400327.

Stenger:2005:CRS

[1288] Horst Stenger and Siegfried Gabler. Combining random sampling and census strategies — justification of inclusion probabilities equal to 1. Metrika, 61(2):137–156, April 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840400328.

Lin:2005:OAM

[1289] Tsai-Yu Lin and Chen-Tuo Liao. Optimal allocation of measurements in a linear calibration process. Metrika, 61(2):157–168, April 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840400329.

Morillas:2005:MON

[1290] Patricia Mariela Morillas. A method to obtain new copulas from a given one. Metrika, 61(2):169–184, April 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840400330.

Pooladsaz:2005:OEC

[1291] S. Pooladsaz and R. J. Martin. Optimal extended complete block designs for dependent observations. Metrika, 61(2):185–197, April 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840400331.

Al-Hussaini:2005:RRM

[1292] Essam K. Al-Hussaini, Abd EL-Baset A. Ahmad, and M. A. AL-Kashif. Recurrence relations for moment and conditional moment generating functions of generalized order statistics. Metrika, 61(2):199–220, April 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840400332.

Pearn:2005:PCA

[1293] W. L. Pearn and Chien-Wei Wu.

Process capability assessment for index C_{pk} based on Bayesian approach. Metrika, 61(2):221–234, April 2005.

CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840400333.

Anonymous:2005:BRa

[1294] Anonymous. Book review. Metrika, 61(2):235–236, April 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00184-005-0366-
                                        3.

Anonymous:2005:HCb

[1295] Anonymous. Help & contacts. Metrika, 61(2):??, April 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Zhang:2005:BLT

[1296] Yao Zhang and William Q. Meeker. Bayesian life test planning for the Weibull distribution with given shape parameter. Metrika, 61(3):237–249, June 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s001840400334.

Marohn:2005:EIT

[1297] Frank Marohn. Exponential inequalities for tail probabilities of generalized order statistics. Metrika, 61 (3):251–260, June 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840400335.

Yeh:2005:EOD

[1298] Hong-Gwa Yeh and Mong-Na Lo Huang. On exact D-optimal designs with 2 two-level factors and n autocorrelated observations. Metrika, 61(3):261–275, June 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840400336.

Klimczak:2005:RPF

[1299] Monika Klimczak and Tomasz Rychlik. Reconstruction of previous failure times and records. Metrika, 61 (3):277–290, June 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840400344.

Kundu:2005:EGE

[1300] Debasis Kundu and Rameshwar D.

Gupta. Estimation of P [Y < X] for generalized exponential distribution. Metrika, 61(3):291–308, June 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840400345.

Nadarajah:2005:TSP

[1301] Saralees Nadarajah. On the twosided power distribution. Metrika, 61(3):309–321, June 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840400346.

Iwasaki:2005:NBD

[1302] Masakazu Iwasaki and Hiroe Tsubaki. A new bivariate distribution in natural exponential family. Metrika, 61(3):323–336, June 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840400348.

Adell:2005:MPD

[1303] J. A. Adell and P. Jodrá. The median of the Poisson distribution. Metrika, 61(3):337–346, June 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840400350.

Anonymous:2005:BRb

[1304] Anonymous. Book review. Metrika, 61(3):347–350, June 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-005-0371-6.

Anonymous:2005:HCc

[1305] Anonymous. Help & contacts. Metrika, 61(3):??, June 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Baran:2005:CEN

[1306] S. Baran. A consistent estimator for nonlinear regression models. Metrika, 62(1):1–15, September 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840400349.

Choudhury:2005:SDM

[1307] Amit Choudhury. A simple derivation of moments of the exponentiated Weibull distribution. Metrika, 62 (1):17–22, September 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840400351.

Drton:2005:ASB

[1308] Mathias Drton and Udo Schwingenschlögl. Asymptotic seat bias formulas. Metrika, 62(1):23–31, September 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s001840400352.

Das:2005:SAF

[1309] Ashish Das, Aloke Dey, and Paramita Saha. Small asymmetric fractional factorial plans for main effects and specified two-factor interactions. Metrika, 62(1):33–52, September 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840400353.

Christofides:2005:RRT

[1310] Tasos C. Christofides. Randomized response technique for two sensitive characteristics at the same time. Metrika, 62(1):53–63, September 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840400355.

Damaraju:2005:SIA

[1311] Lakshmi Damaraju and Damaraju Raghavarao. Statistical inferences accounting for human behavior. Metrika, 62(1):65–72, September 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s001840400356.

Ai:2005:OCE

[1312] Ming-Yao Ai, Peng-Fei Li, and Run-Chu Zhang. Optimal criteria and equivalence for nonregular fractional factorial designs. Metrika, 62(1): 73–83, September 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840400357.

Bischoff:2005:ACP

[1313] Wolfgang Bischoff, Enkelejd Hashorva, Jürg Hüsler, and Frank Miller. Analysis of a change-point regression problem in quality control by partial sums processes and Kolmogorov type tests. Metrika, 62(1):85–98, September 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001840400354.

Epps:2005:TLS

[1314] T. W. Epps. Tests for location-scale families based on the empirical characteristic function. Metrika, 62(1): 99–114, September 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001840400358.

Anonymous:2005:BRc

[1315] Anonymous. Book review. Metrika, 62 (1):115–124, September 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-005-0368-1.

Anonymous:2005:HCd

[1316] Anonymous. Help & contacts. Metrika, 62(1):??, September 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Anonymous:2005:P

[1317] Anonymous. Preface. Metrika, 62(2– 3):125–126, November 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00184-005-0402-
3.

Atkinson:2005:EOD

[1318] Anthony C. Atkinson. Efficiencies for optimum designs when transforming the response in nonlinear models with nonconstant variance. Metrika, 62(2– 3):127–138, November 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-005-0406-z.

Naiman:2005:ABB

[1319] D. Q. Naiman and H. P. Wynn. The algebra of Bonferroni bounds: discrete tubes and extensions. Metrika, 62(2– 3):139–147, November 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-005-0403-2.

Charalambides:2005:DJO

[1320] Ch. A. Charalambides. Derivation of a joint occupancy distribution via a bivariate inclusion and exclusion formula. Metrika, 62(2–3):149–160, November 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-005-
0413-0.

Bailey:2005:SFE

[1321] R. A. Bailey. Six families of efficient resolvable designs in three replicates. Metrika, 62(2–3):161–173, November 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-005-
0405-0.

Crnkovic:2005:CBD

[1322] Dean Crnković and Sanja Rukavina. Construction of block designs admitting an Abelian automorphism group. Metrika, 62(2–3):175–183, November 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-005-
0407-y.

Craigen:2005:RMO

[1323] R. Craigen and H. Kharaghani. A recursive method for orthogonal designs. Metrika, 62(2–3):185–193, November 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-005-
0404-1.

Orrick:2005:MDO

[1324] William P. Orrick. The maximal {−1, 1}-determinant of order 15. Metrika, 62(2–3):195–219, November 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.

com/article/10.1007/s00184-005-
0410-3.

Seberry:2005:SAH

[1325] Jennifer Seberry, Beata J. Wysocki, and Tadeusz A. Wysocki. On some applications of Hadamard matrices. Metrika, 62(2–3):221–239, November 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-005-
0415-y.

Evangelaras:2005:PPC

[1326] H. Evangelaras, C. Koukouvinos, A. M.

Dean, and C. A. Dingus. Projection properties of certain three level orthogonal arrays. Metrika, 62(2–3): 241–257, November 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-005-0409-9.

Xu:2005:CTL

[1327] Hongquan Xu. A catalogue of three-level regular fractional factorial designs. Metrika, 62(2–3):259– 281, November 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-005-0408-x.

Karagiannis:2005:COF

[1328] Vassilis Karagiannis and Chronis Moyssiadis. Construction of D-optimal s_{1} × s_{2} × s_{3} factorial designs using graph theory. Metrika, 62(2–3): 283–307, November 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-005-0414-z.

Antognini:2005:SCS

[1329] Alessandro Baldi Antognini. On the speed of convergence of some urn designs for the balanced allocation of two treatments. Metrika, 62(2–3): 309–322, November 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-005-0412-1.

Katsis:2005:SSD

[1330] Athanassios Katsis. Sample size determination of binomial data with the presence of misclassification. Metrika, 62(2–3):323–329, November 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-005-
                                        0411-2.

Salmaso:2005:MAT

[1331] Luigi Salmaso and Aldo Solari. Multiple aspect testing for casecontrol designs. Metrika, 62(2–3): 331–340, November 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-005-0416-x.

Anonymous:2005:HCe

[1332] Anonymous. Help & contacts. Metrika, 62(2–3):??, November 2005. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Singh:2006:SSC

[1333] Sarjinder Singh. Survey statisticians celebrate golden jubilee year 2003 of the linear regression estimator. Metrika, 63(1):1–18, February 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-005-
0002-2.

Takada:2006:ASO

[1334] Yoshikazu Takada. Asymptotic secondorder efficiency of three-stage procedure with a warranted confidence level. Metrika, 63(1):19–31, February 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-005-
0003-1.

Chaudhuri:2006:ETS

[1335] Arijit Chaudhuri and Amitava Saha.

Estimation from two-stage unequal probability sampling with missing units. Metrika, 63(1):33–41, February 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-005-
0004-0.

Auer:2006:HAH

[1336] C. Auer and J. Kunert. On a heuristic analysis of highly fractionated 2^{n} factorial experiments. Metrika, 63 (1):43–54, February 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-005-0005-z.

Berrendero:2006:NUA

[1337] José R. Berrendero and Ruben H. Zamar. A note on the uniform asymptotic normality of location M-estimates.

Metrika, 63(1):55–69, February 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00184-005-0006-
                                        y.

Hillebrand:2006:CRK

[1338] Martin Hillebrand and Christine H. Müller. On consistency of redescending M-kernel smoothers. Metrika, 63 (1):71–90, February 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-005-0007-x.

Madi:2006:DTE

[1339] Mohamed T. Madi. Decision theoretic estimation using record statistics. Metrika, 63(1):91–97, February 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-005-
                                        0009-8.

Jeffrey:2006:FTS

[1340] Andrew Jeffrey, Oliver Linton, and Thong Nguyen. Flexible term structure estimation: Which method is preferred? Metrika, 63(1):99–122, February 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-005-
                                        0017-8.

Anonymous:2006:HCa

[1341] Anonymous. Help & contacts. Metrika, 63(1):??, February 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Nie:2006:SCM

[1342] L. Nie. Strong consistency of the maximum likelihood estimator in generalized linear and nonlinear mixed– effects models. Metrika, 63(2):123–143, April 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-005-
0001-3.

Huskova:2006:CPA

[1343] Marie Husková and Simos G. Meintanis. Change point analysis based on empirical characteristic functions. Metrika, 63(2):145–168, April 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s00184-005-0008-
9.

daFonseca:2006:DSI

[1344] V. Grunert da Fonseca and N. R. J.

Fieller. Distortion in statistical inference: the distinction between data contamination and model deviation. Metrika, 63(2):169–190, April 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-005-
0010-2.

Racckauskas:2006:HPP

[1345] Alfredas Racckauskas. Hölderian properties of partial sums of regression residuals. Metrika, 63(2):191–205, April 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-005-
0011-1.

Trautmann:2006:DDI

[1346] Heike Trautmann and Claus Weihs. On the distribution of the desirability index using Harrington’s desirability function. Metrika, 63(2):207–213, April 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-005-
                                        0012-0.

Atienza:2006:NCI

[1347] N. Atienza, J. Garcia-Heras, and J. M. Muñoz-Pichardo. A new condition for identifiability of finite mixture distributions. Metrika, 63(2):215–221, April 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-005-
                                        0013-z.

Gupta:2006:MRH

[1348] Ramesh C. Gupta, Rameshwar D. Gupta, and Pushpa L. Gupta. Monotonicity of the (reversed) hazard rate of the (maximum) minimum in bivariate distributions. Metrika, 63(2):223–241, April 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-005-
                                        0014-y.

Chatterjee:2006:LBC

[1349] Kashinath Chatterjee, Kai-Tai Fang, and Hong Qin. A lower bound for the centered L_{2}-discrepancy on asymmetric factorials and its application. Metrika, 63(2):243–255, April 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00184-005-
0015-x.

Anonymous:2006:HCb

[1350] Anonymous. Help & contacts. Metrika, 63(2):??, April 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Adekeye:2006:CRO

[1351] Kayode S. Adekeye and Joachim Kunert. On the comparison of run orders of unreplicated 2^{k−p} designs in the presence of a time trend. Metrika, 63(3):257–269, June 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-005-0016-9.

Evangelaras:2006:SOA

[1352] H. Evangelaras, E. Kolaiti, and C. Koukouvinos. Some orthogonal arrays with 32 runs and their projection properties. Metrika, 63(3):271–281, June 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-005-
0018-7.

Salibian-Barrera:2006:AME

[1353] Matias Salibian-Barrera. The asymptotics of MM-estimators for linear regression with fixed designs. Metrika, 63(3):283–294, June 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-005-0019-6.

Ruckdeschel:2006:MSN

[1354] Peter Ruckdeschel. A motivation for 1/\sqrt{n}-shrinking neighborhoods.

Metrika, 63(3):295–307, June 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s00184-005-0020-
0.

Jandhyala:2006:ICP

[1355] Venkata K. Jandhyala, Stergios B. Fotopoulos, and Douglas M. Hawkins. On the inconsistency of the change–point estimator for the NE family. Metrika, 63(3):309–315, June 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-005-0021-z.

Liu:2006:RBD

[1356] Min-Qian Liu and Fred J. Hickernell. The relationship between discrepancies defined on a domain and on its subset. Metrika, 63(3):317–327, June 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-005-
0022-y.

Karlsson:2006:ERP

[1357] Maria Karlsson. Estimators of regression parameters for truncated and censored data. Metrika, 63(3):329–341, June 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-005-
0023-x.

Garcia-Perez:2006:CST

[1358] A. García-Pérez. Chi-square tests under models close to the normal distribution. Metrika, 63(3):343–354, June 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-005-
0024-9.

Perez:2006:DSC

[1359] T. Pérez and J. A. Pardo. The K_{φ}- divergence statistic for categorical data problems. Metrika, 63(3):355–369, June 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-005-
0025-8.

Balakrishnan:2006:EES

[1360] N. Balakrishnan and T. Rychlik. Evaluating expectations of L-statistics by the Steffensen inequality. Metrika, 63(3):371–384, June 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-005-0026-7.

Anonymous:2006:HCc

[1361] Anonymous. Help & contacts.

Metrika, 63(3):??, June 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Pham-Gia:2006:BAN

[1362] T. Pham-Gia, N. Turkkan, and A. Bekker. Bayesian analysis in the L^{1}-norm of the mixing proportion using discriminant analysis. Metrika, 64(1):1–22, August 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0027-1.

Perlman:2006:SIT

[1363] Michael D. Perlman and Lang Wu. Some improved tests for multivariate one–sided hypotheses. Metrika, 64(1):23–39, August 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0028-0.

Cheng:2006:NEF

[1364] Chi-Lun Cheng and Alexander Kukush. Non-existence of the first moment of the adjusted least squares estimator in multivariate errors-in-variables model. Metrika, 64(1):41–46, August 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0029-z.

Nanda:2006:SPP

[1365] Asok K. Nanda and Prasanta Paul. Some properties of past entropy and their applications. Metrika, 64 (1):47–61, August 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0030-6.

Pardo:2006:ADL

[1366] L. Pardo and M. L. Menéndez. Analysis of divergence in loglinear models when expected frequencies are subject to linear constraints. Metrika, 64(1):63–76, August 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0034-2.

Gonzalez-Barrios:2006:GFD

[1367] José M. González-Barrios, Federico O’Reilly, and Raúl Rueda. Goodness of fit for discrete random variables using the conditional density. Metrika, 64(1):77–94, August 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0035-1.

Ai:2006:GWP

[1368] Ming-Yao Ai and Shu-Yuan He. Generalized wordtype pattern for nonregular factorial designs with multiple groups of factors. Metrika, 64(1):95–108, August 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0037-z.

Grossmann:2006:ODA

[1369] Heiko Großmann, Heinz Holling, Ulrike Graßhoff, and Rainer Schwabe. Optimal designs for asymmetric linear paired comparisons with a profile strength constraint. Metrika, 64 (1):109–119, August 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0038-y.

Fried:2006:BRH

[1370] R. Fried. Book review: Härdle, W., Simar, L.: Applied multivariate statistical analysis, Springer Verlag, Berlin–Heidelberg-New York 2003, pp. XVIII/486, Eur69.95, ISBN 3-540-03079-4. Metrika, 64(1):121–122, August 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00184-006-0031-5.

Repsilber:2006:BRS

[1371] Dirk Repsilber and Andreas Ziegler. Book review: Simon, R. M., Korn, E. L., McShane, L. M., Radmacher, M. D., Wright, G. W. and Zhao, Y., Design and Analysis of DNA Microarray Investigations, Springer New York 2004, ISBN 0-387-00135-2. Metrika, 64 (1):123–124, August 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00184-006-0032-
4.

Falk:2006:BRB

[1372] Michael Falk. Book review: Introduction to Rare Event Simulation, by James Antonio Bucklew Springer Berlin, ISBN 0-387-20078-9. Metrika, 64(1):125–126, August 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00184-006-0033-
3.

Drees:2006:BRM

[1373] Holger Drees. Book review: Michael Falk, Jürg Hüsler, Rolf-Dieter Reiss, Laws of Small Numbers: Extremes and Rare Events. Metrika, 64(1): 127–129, August 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0036-0.

Anonymous:2006:HCd

[1374] Anonymous. Help & contacts. Metrika, 64(1):??, August 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Castano-Martinez:2006:HAC

[1375] Antonia Castaño-Martínez and Fernando López-Blázquez. Heuristic approximation to Cramér–von Mises type statistics. Metrika, 64(2):131–138, October 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0039-x.

Stoyan:2006:ENN

[1376] Dietrich Stoyan. On estimators of the nearest neighbour distance distribution function for stationary point processes. Metrika, 64(2):139–150, October 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0040-4.

Hesse:2006:IDE

[1377] Christian H. Hesse. Iterative density estimation from contaminated observations. Metrika, 64(2):151–165, October 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0041-3.

Kulkarni:2006:CMM

[1378] H. V. Kulkarni. Characterizations and modelling of multivariate lack of memory property. Metrika, 64 (2):167–180, October 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0042-2.

Zaigraev:2006:DOD

[1379] Alexander Zaigraev. On DS-optimal design matrices with restrictions on rows or columns. Metrika, 64(2): 181–189, October 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0043-1.

Heinrich:2006:GFC

[1380] Lothar Heinrich and Udo Schwingenschlögl. Goodness-of-fit criteria for the Adams and Jefferson rounding methods and their limiting laws. Metrika, 64(2):191–207, October 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0044-0.

Georgiou:2006:MLC

[1381] S. Georgiou and C. Koukouvinos. Multi-level k-circulant supersaturated designs. Metrika, 64(2):209–220, October 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0045-z.

Batsidis:2006:DOO

[1382] A. Batsidis and K. Zografos. Discrimination of observations into one of two elliptic populations based on monotone training samples. Metrika, 64 (2):221–241, October 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0046-y.

Wakolbinger:2006:BRR

[1383] Anton Wakolbinger. Book review: Rasmus Nielsen (ed.): Statistical methods in molecular evolution. Metrika, 64 (2):243–245, October 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0086-3.

Creutzig:2006:BRW

[1384] Jakob Creutzig. Book review: W.

Hörmann, J. Leydold, G. Derflinger: Automatic nonuniform random variate generation. Metrika, 64(2):247– 248, October 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s00184-006-0087-2.

Droge:2006:BRP

[1385] Bernd Droge. Book review: Phillip Good: Permutation, parametric, and bootstrap tests of hypotheses. Metrika, 64(2):249–250, October 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00184-006-0088-
1.

Schmidt:2006:BRJ

[1386] Volker Schmidt. Book review: James E. Gentle: Random number generation and Monte Carlo methods. Metrika, 64 (2):251–252, October 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00184-006-0089-
                                        0.

Nather:2006:BRA

[1387] Wolfgang Näther. Book review: Antony C. Atkinson, Marco Riani, Andrea Cerioli: Exploring multivariate data with the forward search. Metrika, 64(2):253–254, October 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00184-006-0090-
                                        7.

Steinebach:2006:BRL

[1388] J. Steinebach. Book review: E. L. Lehmann, J. P. Romano: Testing statistical hypotheses. Metrika, 64 (2):255–256, October 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00184-006-0091-
                                        6.

Kohler:2006:BRW

[1389] Michael Kohler. Book review: Wallace, C. S: Statistical and inductive inference by minimum message length. Metrika, 64(2):257, October 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00184-006-0093-
                                        4.

Anonymous:2006:HCe

[1390] Anonymous. Help & contacts. Metrika, 64(2):??, October 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Jokiel-Rokita:2006:GMP

[1391] Alicja Jokiel-Rokita. Gamma-minimax prediction for the multinomial distribution. Metrika, 64(3):259–269, December 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0047-x.

Wang:2006:EVC

[1392] Qihua Wang and Lili Yao. Estimation in varying-coefficient proportional hazard regression model. Metrika, 64 (3):271–288, December 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0048-9.

Torabi:2006:NPL

[1393] Hamzeh Torabi, Javad Behboodian, and S. Mahmoud Taheri. Neyman– Pearson lemma for fuzzy hypotheses testing with vague data. Metrika, 64 (3):289–304, December 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0049-8.

Muralidharan:2006:AIE

[1394] K. Muralidharan and P. Lathika. Analysis of instantaneous and early failures in Weibull distribution. Metrika, 64 (3):305–316, December 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0050-2.

Chacko:2006:CRV

[1395] Manoj Chacko and P. Yageen Thomas. Concomitants of record values arising from Morgenstern type bivariate logistic distribution and some of their applications in parameter estimation. Metrika, 64(3):317–331, December 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0051-1.

Tsou:2006:PRT

[1396] Tsung-Shan Tsou. Parametric robust test for several variances with unknown underlying distributions. Metrika, 64 (3):333–349, December 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0052-0.

Evangelaras:2006:PPH

[1397] H. Evangelaras, C. Koukouvinos, and K. Mylona. Projection properties of Hadamard matrices of order 36 obtained from Paley’s constructions. Metrika, 64(3):351–359, December 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0053-z.

Wang:2006:NLI

[1398] Jian-Feng Wang and Li-Xin Zhang. A nonclassical Law of the Iterated Logarithm for functions of positively associated random variables. Metrika, 64 (3):361–378, December 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0054-y.

Pandit:2006:LMP

[1399] Parameshwar V. Pandit. Locally most powerful and other rank tests for independence — with a contaminated weighted alternative. Metrika, 64(3): 379–387, December 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0056-9.

Anonymous:2006:HCf

[1400] Anonymous. Help & contacts. Metrika, 64(3):??, December 2006. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Kamps:2007:E

[1401] Udo Kamps, Friedrich Pukelsheim, and Jürgen Lehn. Editorial. Metrika, 65(1):1, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00184-006-0115-
2.

Kossler:2007:AER

[1402] Wolfgang Kössler and Egmar Rödel.

The asymptotic efficacies and relative efficiencies of various linear rank tests for independence. Metrika, 65 (1):3–28, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0055-x.

Stenger:2007:OSS

[1403] Horst Stenger and Siegfried Gabler.

Optimal strategies in 2-stage sampling. Metrika, 65(1):29–42, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0057-8.

Kasturiratna:2007:CND

[1404] Dhanuja Kasturiratna, Truc T. Nguyen, and Arjun K. Gupta. Characterization of normal distribution related to two samples based on regression. Metrika, 65(1):43–52, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0058-7.

Goos:2007:OMS

[1405] Peter Goos and Alexander N. Donev. D-optimal minimum support mixture designs in blocks. Metrika, 65 (1):53–68, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0059-6.

Lopez-Blazquez:2007:TL

[1406] Fernando López-Blázquez and Jacek Wesolowski. Top- k-lists. Metrika, 65(1):69–82, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0060-0.

Dasgupta:2007:OAC

[1407] Nairanjana Dasgupta and Rita SahaRay. Optimal allocation for comparing k test treatments to positive and negative control with unequal weighting under A-optimality and MV-optimality. Metrika, 65 (1):83–92, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0061-z.

Liu:2007:LRT

[1408] Xinsheng Liu. Likelihood ratio test for and against nonlinear inequality constraints. Metrika, 65(1):93–108, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0062-y.

Nather:2007:CVF

[1409] Wolfgang Näther and Andreas Wünsche.

On the conditional variance of fuzzy random variables. Metrika, 65(1): 109–122, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0063-x.

Zi:2007:AFD

[1410] Xue-Min Zi, Min-Qian Liu, and Run-Chu Zhang. Asymmetrical factorial designs containing clear effects. Metrika, 65(1):123–131, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0064-9.

Anonymous:2007:HCa

[1411] Anonymous. Help & contacts. Metrika, 65(1):??, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Eichler:2007:FDB

[1412] Michael Eichler. A frequency-domain based test for non-correlation between stationary time series. Metrika, 65 (2):133–157, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-006-0065-8.

Kundu:2007:AHL

[1413] Debasis Kundu and Rameshwar D. Gupta. Analysis of hybrid lifetests in presence of competing risks. Metrika, 65(2):159–170, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00184-006-0066-
                                        7.

Druilhet:2007:IMN

[1414] Pierre Druilhet and Augustyn Markiewicz. Information matrices for non full rank subsystems. Metrika, 65(2): 171–182, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-006-0067-6.

Schmelter:2007:OSG

[1415] Thomas Schmelter. The optimality of single-group designs for certain mixed models. Metrika, 65(2):183–193, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-006-
                                        0068-5.

Zhang:2007:LDB

[1416] Li-Xin Zhang and Xiao-Rong Yang.

The limit distribution of the bootstrap for the unit root test statistic when the residuals are dependent. Metrika, 65 (2):195–206, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0070-y.

Savani:2007:EPE

[1417] V. Savani and A. A. Zhigljavsky. Efficient parameter estimation for independent and INAR(1) negative binomial samples. Metrika, 65(2): 207–225, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0071-x.

Chang:2007:MSO

[1418] Fu-Chuen Chang and Hung-Ming Lin. On minimally-supported Doptimal designs for polynomial regression with log-concave weight function. Metrika, 65(2):227–233, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0072-9.

Jacroux:2007:MRM

[1419] Mike Jacroux. Maximal rank minimum aberration foldover plans for 2^{m−k} fractional factorial designs. Metrika, 65 (2):235–242, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0073-8.

Fischer:2007:CGF

[1420] Matthias Fischer and Ingo Klein. Constructing generalized FGM copulas by means of certain univariate distributions. Metrika, 65(2):243–260, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-006-
                                        0075-6.

Anonymous:2007:HCb

[1421] Anonymous. Help & contacts. Metrika, 65(2):??, February 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Krishnamoorthy:2007:IRT

[1422] K. Krishnamoorthy, Shubhabrata Mukherjee, and Huizhen Guo. Inference on reliability in two-parameter exponential stress-strength model. Metrika, 65(3):261–273, May 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00184-006-0074-
                                        7.

Shklyar:2007:QSM

[1423] Sergiy Shklyar, Hans Schneeweiß, and Alexander Kukush. Quasi score is more efficient than corrected score in a polynomial measurement error model. Metrika, 65(3):275–295, May 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-006-
                                        0076-5.

Bieniek:2007:VDP

[1424] Mariusz Bieniek. Variation diminishing property of densities of uniform generalized order statistics. Metrika, 65(3):297–309, May 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0077-4.

Tian:2007:EOD

[1425] Yongge Tian and Agnes M. Herzberg.

Estimation and optimal designs for linear Haar-wavelet models. Metrika, 65(3):311–324, May 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0078-3.

Hu:2007:IFR

[1426] Taizhong Hu and Ying Li. Increasing failure rate and decreasing reversed hazard rate properties of the minimum and maximum of multivariate distributions with log-concave densities. Metrika, 65(3):325–330, May 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0079-2.

Stulajter:2007:MSE

[1427] Frantisek Stulajter. Mean squared error of the empirical best linear unbiased predictor in an orthogonal finite discrete spectrum linear regression model. Metrika, 65(3):331–348, May 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00184-006-
                                        0080-9.

Rueda:2007:PEL

[1428] M. Rueda, J. F. Muñoz, Y. G. Berger, A. Arcos, and S. Martínez. Pseudo empirical likelihood method in the presence of missing data. Metrika, 65(3):349–367, May 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-006-0081-8.

Filipiak:2007:ODM

[1429] K. Filipiak and A. Markiewicz. Optimal designs for a mixed interference model. Metrika, 65(3):369–386, May 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-006-
                                        0082-7.

Anonymous:2007:HCc

[1430] Anonymous. Help & contacts. Metrika, 65(3):??, May 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Wu:2007:GLM

[1431] Kunling Wu and Lang Wu. Generalized linear mixed models with informative dropouts and missing covariates. Metrika, 66(1):1–18, July 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-006-
                                        0083-6.

Dominguez-Menchero:2007:AEI

[1432] J. Santos Domínguez-Menchero and Gil González-Rodríguez. Analyzing an extension of the isotonic regression problem. Metrika, 66(1):19–30, July 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0084-5.

Evangelaras:2007:RNT

[1433] H. Evangelaras, C. Koukouvinos, and E. Lappas. 18-run nonisomorphic three level orthogonal arrays. Metrika, 66(1):31–37, July 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0085-4.

Muller:2007:WLS

[1434] Ursula U. Müller. Weighted least squares estimators in possibly misspecified nonlinear regression. Metrika, 66(1):39–59, July 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0092-5.

Jokiel-Rokita:2007:MEC

[1435] Alicja Jokiel-Rokita and Ryszard Magiera. Minimax estimation of a cumulative distribution function by converting to a parametric problem. Metrika, 66(1):61–73, July 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0094-3.

Li:2007:ADP

[1436] Yun-Xia Li and Jian-Feng Wang.

Asymptotic distribution for products of sums under dependence. Metrika, 66(1):75–87, July 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-006-0095-2.

Li:2007:SCN

[1437] Yun Li and Quanxi Shao. Slow convergence of the number of near-maxima for Burr XII distributions. Metrika, 66(1):89–104, July 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-006-0096-1.

Li:2007:NMM

[1438] Yuling Li, Wei Gao, and Ning-Zhong Shi. A note on multinomial maximum likelihood estimation under ordered restrictions and the EM algorithm. Metrika, 66(1):105–114, July 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-006-
                                        0098-z.

Sreekumar:2007:EPL

[1439] N. V. Sreekumar and P. Yageen Thomas. Estimation of the parameters of log-gamma distribution using order statistics. Metrika, 66(1):115–127, July 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-006-
                                        0099-y.

Anonymous:2007:HCd

[1440] Anonymous. Help & contacts. Metrika, 66(1):??, July 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Danielak:2007:CDD

[1441] Katarzyna Danielak and Anna Dembińska.

On characterizing discrete distributions via conditional expectations of weak record values. Metrika, 66(2): 129–138, September 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0100-9.

Angelopoulos:2007:ESA

[1442] P. Angelopoulos, H. Evangelaras, C. Koukouvinos, and E. Lappas. An effective step-down algorithm for the construction and the identification of nonisomorphic orthogonal arrays. Metrika, 66(2):139–149, September 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0101-8.

Jozani:2007:MEC

[1443] Mohammad Jafari Jozani and Éric Marchand. Minimax estimation of constrained parametric functions for discrete families of distributions. Metrika, 66(2):151–160, September 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0102-7.

Rodriguez-Torreblanca:2007:LOD

[1444] C. Rodríguez-Torreblanca and J. M.

Rodríguez-Díaz. Locally D- and c-optimal designs for Poisson and negative binomial regression models. Metrika, 66(2):161–172, September 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-006-
                                        0103-6.

Chen:2007:PSF

[1445] Zhenlong Chen. Polar sets of fractional Brownian sheets. Metrika, 66 (2):173–196, September 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-006-0104-5.

Belzunce:2007:RTC

[1446] Félix Belzunce, Eva-María Ortega, Franco Pellerey, and José M. Ruiz. On rankings and top choices in random utility models with dependent utilities. Metrika, 66(2):197–212, September 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-006-
                                        0105-4.

Abdel-Hamid:2007:PSA

[1447] Alaa H. Abdel-Hamid and Essam K. Al-Hussaini. Progressive stress accelerated life tests under finite mixture models. Metrika, 66(2):213–231, September 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-006-
                                        0106-3.

Bieniek:2007:CDR

[1448] Mariusz Bieniek. On characterizations of distributions by regression of adjacent generalized order statistics. Metrika, 66(2):233–242, September 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0107-2.

Gjestvang:2007:FQR

[1449] Christopher R. Gjestvang and Sarjinder Singh. Forced quantitative randomized response model: a new device. Metrika, 66(2):243–257, September 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0108-1.

Anonymous:2007:HCe

[1450] Anonymous. Help & contacts. Metrika, 66(2):??, September 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Mostofi:2007:MSB

[1451] Amin Ghalamfarsa Mostofi and Javad Behboodian. On model selection in Bayesian regression. Metrika, 66(3): 259–268, November 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0109-0.

Gottard:2007:IBM

[1452] Anna Gottard. On the inclusion of bivariate marked point processes in graphical models. Metrika, 66(3): 269–287, November 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0110-7.

Zhengyan:2007:ABE

[1453] Lin Zhengyan, Li Degui, and Chen Jia. Asymptotic behavior for Sestimators in random design linear model with long-range-dependent errors. Metrika, 66(3):289–303, November 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0111-6.

Campbell:2007:GCM

[1454] Paul J. Campbell. Games of chance with multiple objectives. Metrika, 66 (3):305–313, November 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0112-5.

Nguyen:2007:CNM

[1455] Truc T. Nguyen, Arjun K. Gupta, Diem M. Nguyen, and Yining Wang. Characterizations of negative multinomial distributions based on conditional distributions. Metrika, 66(3): 315–322, November 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0113-4.

Schmid:2007:NIM

[1456] Friedrich Schmid and Rafael Schmidt. Nonparametric inference on multivariate versions of Blomqvist’s beta and related measures of tail dependence. Metrika, 66(3):323–354, November 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.

com/article/10.1007/s00184-006-
0114-3.

Rueda:2007:CME

[1457] M. Rueda, S. Martínez-Puertas, H. Martínez-Puertas, and A. Arcos. Calibration methods for estimating quantiles. Metrika, 66(3): 355–371, November 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0116-1.

Niaki:2007:MMA

[1458] Seyed Taghi Akhavan Niaki and Babak Abbasi. On the monitoring of multiattributes high-quality production processes. Metrika, 66(3):373–388, November 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0117-0.

Anonymous:2007:HCf

[1459] Anonymous. Help & contacts. Metrika, 66(3):??, November 2007. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Li:2008:ASM

[1460] Yun-Xia Li and Jian-Feng Wang.

An application of Stein’s method to limit theorems for pairwise negative quadrant dependent random variables. Metrika, 67(1):1–10, January 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-006-
0118-z.

Jia:2008:SLE

[1461] Chen Jia, Zhang Lixin, and Li Degui. Spatial local M-estimation under association. Metrika, 67(1):11–29, January 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-006-
                                        0119-y.

Lin:2008:BBW

[1462] Lu Lin, Yunzheng Fan, and Lin Tan. Blockwise bootstrap wavelet in nonparametric regression model with weakly dependent processes. Metrika, 67(1):31–48, January 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-006-0120-5.

Xie:2008:IAP

[1463] Feng-Chang Xie and Bo-Cheng Wei. Influence analysis for Poisson inverse Gaussian regression models based on the EM algorithm. Metrika, 67 (1):49–62, January 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-006-0121-4.

Szucs:2008:PBT

[1464] Gábor Szűcs. Parametric bootstrap tests for continuous and discrete distributions. Metrika, 67(1):63–81, January 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-006-
                                        0122-3.

Kurian:2008:GCI

[1465] K. M. Kurian, Thomas Mathew, and G. Sebastian. Generalized confidence intervals for process capability indices in the one-way random model. Metrika, 67(1):83–92, January 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-006-0123-2.

Lu:2008:ELA

[1466] Xuewen Lu, Jie Sun, and Yongcheng Qi. Empirical likelihood for average derivatives of hazard regression functions. Metrika, 67(1):93–112, January 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-007-
0124-9.

Kumar:2008:UAB

[1467] C. Satheesh Kumar. A unified approach to bivariate discrete distributions. Metrika, 67(1):113–123, January 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-007-
0125-8.

Anonymous:2008:HCa

[1468] Anonymous. Help & contacts. Metrika, 67(1):??, January 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Dette:2008:E

[1469] Holger Dette, Udo Kamps, and Jürgen Lehn. Editorial. Metrika, 67(2):125, March 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00184-008-0169-4.

Bodnar:2008:TWG

[1470] Taras Bodnar and Wolfgang Schmid. A test for the weights of the global minimum variance portfolio in an elliptical model. Metrika, 67(2):127–143, March 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-007-
0126-7.

Comas:2008:GRP

[1471] C. Comas and J. Mateu. Growing and reproducing particles evolving through space and time. Metrika, 67(2):145–169, March 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-007-0127-6.

Patel:2008:MBV

[1472] P. A. Patel and R. D. Chaudhari. Model-based variance estimation under unequal probability sampling. Metrika, 67(2):171–187, March 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-007-0128-5.

Boente:2008:ECP

[1473] Graciela Boente, Ana M. Pires, and Isabel M. Rodrigues. Estimators for the common principal components model based on reweighting: influence functions and Monte Carlo study. Metrika, 67(2):189–218, March 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-007-0129-4.

Hafner:2008:AQM

[1474] Christian M. Hafner and Helmut Herwartz. Analytical quasi maximum likelihood inference in multivariate volatility models. Metrika, 67(2):219–239, March 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-007-
0130-y.

Karlsson:2008:SST

[1475] Andreas Karlsson. Survey sampling:

theory and methods. Metrika, 67 (2):241–242, March 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00184-007-0161-
4.

Sturtz:2008:BRH

[1476] Sibylle Sturtz. Book review: H. Rue, L.

Held: Gaussian Markov random fields. Theory and applications. Metrika, 67 (2):243–244, March 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00184-007-0162-
3.

Unwin:2008:BRG

[1477] Antony Unwin. Book review: Gentle, J., Härdle, W., Mori, Y.: Handbook of Computational Statistics: Concepts and Methods. Metrika, 67(2):245–246, March 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.

                                        com/accesspage/article/10.1007/
                                        s00184-007-0163-2.

Kainhofer:2008:BRH

[1478] Reinhold Kainhofer. Book review: H. Föllmer, A. Schied: Stochastic finance: an introduction in discrete time. de Gruyter Studies in Mathematics 27. Metrika, 67(2):247–249, March 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-007-
                                        0164-1.

Anonymous:2008:HCb

[1479] Anonymous. Help & contacts. Metrika, 67(2):??, March 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Yu:2008:TNM

[1480] Jun-Wu Yu, Guo-Liang Tian, and Man-Lai Tang. Two new models for survey sampling with sensitive characteristic: design and analysis. Metrika, 67(3):251–263, April 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-007-0131-x.

Hancova:2008:NEV

[1481] Martina Hancová. Natural estimation of variances in a general finite discrete spectrum linear regression model. Metrika, 67(3):265–276, April 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-007-
                                        0132-9.

Navarro:2008:MRL

[1482] Jorge Navarro and Pedro J. Hernandez.

Mean residual life functions of finite mixtures, order statistics and coherent systems. Metrika, 67(3):277–298, April 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-007-
0133-8.

Eryilmaz:2008:ROS

[1483] Serkan Eryilmaz and Alexei Stepanov.

Runs in an ordered sequence of random variables. Metrika, 67(3):299–313, April 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-007-
0134-7.

Shin:2008:URT

[1484] Dong Wan Shin and Oesook Lee. Unit root tests for panel MTAR model with cross-sectionally dependent error. Metrika, 67(3):315–326, April 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-007-
0135-6.

Nagaraja:2008:FIO

[1485] H. N. Nagaraja and Z. A. Abo-Eleneen.

Fisher information in order statistics and their concomitants in bivariate censored samples. Metrika, 67 (3):327–347, April 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-007-0136-5.

Giovagnoli:2008:NAI

[1486] Alessandra Giovagnoli, Johnny Marzialetti, and Henry P. Wynn. A new approach to inter-rater agreement through stochastic orderings: the discrete case. Metrika, 67(3):349–370, April 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-007-
                                        0137-4.

Choudhury:2008:BIP

[1487] Amit Choudhury and Arun C. Borthakur. Bayesian inference and prediction in the single server Markovian queue. Metrika, 67(3):371–383, April 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-007-
                                        0138-3.

Anonymous:2008:HCc

[1488] Anonymous. Help & contacts. Metrika, 67(3):??, April 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Antoch:2008:DDR

[1489] Jaromír Antoch, Marie Husková, Alicja Janic, and Teresa Ledwina. Data driven rank test for the change point problem. Metrika, 68(1):1–15, June 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-007-
                                        0139-2.

Chen:2008:MLR

[1490] Huaihou Chen and Taizhong Hu. Multivariate likelihood ratio orderings between spacings of heterogeneous exponential random variables. Metrika, 68(1):17–29, June 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-007-0140-9.

Sanhueza:2008:NCI

[1491] Antonio Sanhueza, Víctor Leiva, and N. Balakrishnan. A new class of inverse Gaussian type distributions. Metrika, 68(1):31–49, June 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-007-0141-8.

Yanev:2008:CPD

[1492] George P. Yanev, M. Ahsanullah, and M. I. Beg. Characterizations of probability distributions via bivariate regression of record values. Metrika, 68(1):51–64, June 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-007-0142-7.

Hsu:2008:GCI

[1493] Bi-Min Hsu, Chien-Wei Wu, and Ming-Hung Shu. Generalized confidence intervals for the process capability index C_{pm}. Metrika, 68(1):65–82, June 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-007-
0143-6.

Liang:2008:MGU

[1494] Jiajuan Liang and Kai Wang Ng. A method for generating uniformly scattered points on the L_{p}-norm unit sphere and its applications. Metrika, 68(1):83–98, June 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-007-0144-5.

Koukouvinos:2008:GCO

[1495] C. Koukouvinos, P. Mantas, and K. Mylona. A general construction of E(s^{2})-optimal large supersaturated designs. Metrika, 68(1):99–110, June 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-007-
                                        0145-4.

Munoz-Pichardo:2008:IMP

[1496] J. M. Muñoz-Pichardo, J. L. Moreno-Rebollo, A. Enguix-González, and A. Pascual-Acosta. Influence measures on profile analysis with elliptical data through Frèchet’s metric. Metrika, 68(1):111–127, June 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-007-0146-3.

Anonymous:2008:HCd

[1497] Anonymous. Help & contacts. Metrika, 68(1):??, June 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Xu:2008:GFT

[1498] Wang-Li Xu and Li-Xing Zhu. Goodness-of-fit testing for varyingcoefficient models. Metrika, 68(2): 129–146, September 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-007-0147-2.

Chaudhuri:2008:ESP

[1499] Arijit Chaudhuri and Sanghamitra Pal.

Estimating sensitive proportions from Warner’s randomized responses in alternative ways restricting to only distinct units sampled. Metrika, 68(2): 147–156, September 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-007-0148-1.

Synowka-Bejenka:2008:CAL

[1500] Ewa Synówka-Bejenka and Stefan Zontek. A characterization of admissible linear estimators of fixed and random effects in linear models. Metrika, 68 (2):157–172, September 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-007-0149-0.

Gupta:2008:SGR

[1501] Ramesh C. Gupta, Mohammad Tajdari, and Henrik Bresinsky. Some general results for moments in bivariate distributions. Metrika, 68(2): 173–187, September 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-007-0150-7.

Georgiou:2008:COS

[1502] Stelios D. Georgiou. On the construction of E(s^{2})-optimal supersaturated designs. Metrika, 68(2): 189–198, September 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-007-0151-6.

Nadarajah:2008:PPD

[1503] Saralees Nadarajah and Arjun K. Gupta. A product Pareto distribution. Metrika, 68(2):199–208, September 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-007-
                                        0152-5.

Goliforushani:2008:DMP

[1504] S. Goliforushani and M. Asadi. On the discrete mean past lifetime. Metrika, 68 (2):209–217, September 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-007-0153-4.

Seshadri:2008:MCB

[1505] V. Seshadri and J. Wesolowski. More on connections between Wishart and matrix GIG distributions. Metrika, 68 (2):219–232, September 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-007-0154-3.

Balakrishnan:2008:PTT

[1506] N. Balakrishnan, Anna Dembińska, and Alexei Stepanov. Precedencetype tests based on record values. Metrika, 68(2):233–255, September 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-007-
                                        0155-2.

Anonymous:2008:HCe

[1507] Anonymous. Help & contacts. Metrika, 68(2):??, September 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Santos:2008:LML

[1508] José António Santos and M. Manuela Neves. A local maximum likelihood estimator for Poisson regression. Metrika, 68(3):257–270, November 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-007-
0156-1.

Kumar:2008:ERN

[1509] Somesh Kumar and Yogesh Mani Tripathi. Estimating a restricted normal mean. Metrika, 68(3):271–288, November 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-007-
0157-0.

Hashorva:2008:NFB

[1510] Enkelejd Hashorva. A new family of bivariate max-infinitely divisible distributions. Metrika, 68(3): 289–304, November 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-007-0158-z.

Liang:2008:BEB

[1511] Han-Ying Liang and Jong-Il Baek.

Berry–Esseen bounds for density estimates under NA assumption. Metrika, 68(3):305–322, November 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-007-
0159-y.

Cai:2008:SLW

[1512] Guang hui Cai. Strong laws for weighted sums of NA random variables. Metrika, 68(3):323–331, November 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00184-
007-0160-5.pdf.

Sankaran:2008:PRH

[1513] P. G. Sankaran and V. L. Gleeja. Proportional reversed hazard and frailty models. Metrika, 68(3): 333–342, November 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-007-0165-0.

Uddin:2008:OES

[1514] Nizam Uddin. Optimality and efficiency of small chessboard designs for correlated errors. Metrika, 68(3): 343–350, November 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-007-0166-z.

Nordman:2008:ELM

[1515] Daniel J. Nordman. An empirical likelihood method for spatial regression. Metrika, 68(3):351–363, November 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-007-
0167-y.

Lee:2008:REO

[1516] Sangyeol Lee and Taewook Lee. Robust estimation for the order of finite mixture models. Metrika, 68(3): 365–390, November 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-007-0168-x.

Anonymous:2008:HCf

[1517] Anonymous. Help & contacts. Metrika, 68(3):??, November 2008. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Bandyopadhyay:2009:DLD

[1518] Uttam Bandyopadhyay, Atanu Biswas, and Rahul Bhattacharya. Drop-the-loser design in the presence of covariates. Metrika, 69(1):1–15, January 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-008-
0170-y.

Bathke:2009:UAN

[1519] Arne C. Bathke. A unified approach to nonparametric trend tests for dependent and independent samples. Metrika, 69(1):17–29, January 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-008-
0171-x.

Sarr:2009:EPM

[1520] Amadou Sarr, Arjun K. Gupta, and Anwar H. Joarder. Estimation of the precision matrix of multivariate Pearson type II model. Metrika, 69 (1):31–44, January 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-008-0172-9.

Ou:2009:MRD

[1521] Beiyan Ou and Julie Zhou. Minimax robust designs for field experiments. Metrika, 69(1):45–54, January 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-008-
                                        0173-8.

Qingguo:2009:ANE

[1522] Tang Qingguo. Asymptotic normality of M-estimators in a semiparametric model with longitudinal data. Metrika, 69(1):55–67, January 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-008-0175-6.

Doostparast:2009:NEB

[1523] Mahdi Doostparast. A note on estimation based on record data. Metrika, 69(1):69–80, January 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-008-0178-3.

Anonymous:2009:HCa

[1524] Anonymous. Help & contacts. Metrika, 69(1):??, January 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Gantert:2009:P

[1525] Nina Gantert and Christine H. Müller. Preface. Metrika, 69(2–3):81–84, March 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00184-
008-0208-1.pdf.

Evans:2009:HPP

[1526] Steven N. Evans and Alex Gottlieb. Hyperdeterminantal point processes. Metrika, 69(2–3):85–99, March 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-008-
0209-0.

Flandoli:2009:RUS

[1527] Franco Flandoli. Remarks on uniqueness and strong solutions to deterministic and stochastic differential equations. Metrika, 69(2–3):101–123, March 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-008-
0210-7.

Baraka:2009:HPL

[1528] D. Baraka, T. Mountford, and Y. Xiao.

Hölder properties of local times for fractional Brownian motions. Metrika, 69(2–3):125–152, March 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-008-0211-6.

Lytova:2009:CLT

[1529] Anna Lytova and Leonid Pastur. Central limit theorem for linear eigenvalue statistics of the Wigner and the sample covariance random matrices. Metrika, 69(2–3):153–172, March 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-008-0212-5.

Cariboni:2009:JIM

[1530] Jessica Cariboni and Wim Schoutens. Jumps in intensity models: investigating the performance of Ornstein– Uhlenbeck processes in credit risk modeling. Metrika, 69(2–3):173–198, March 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-008-
                                        0213-4.

Doukhan:2009:WDM

[1531] Paul Doukhan, Nathanaël Mayo, and Lionel Truquet. Weak dependence, models and some applications. Metrika, 69(2–3):199–225, March 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-008-0216-1.

Robins:2009:QSM

[1532] James Robins, Lingling Li, Eric Tchetgen, and Aad W. van der Vaart. Quadratic semiparametric von Mises calculus. Metrika, 69(2–3):227–247, March 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00184-
                                        008-0214-3.pdf.

Schick:2009:CRD

[1533] Anton Schick and Wolfgang Wefelmeyer. Convergence rates of density estimators for sums of powers of observations. Metrika, 69(2–3):249–264, March 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-008-
0215-2.

Jongbloed:2009:CLB

[1534] Geurt Jongbloed. Consistent likelihoodbased estimation of a star-shaped distribution. Metrika, 69(2–3):265–282, March 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00184-
008-0217-0.pdf.

Schafer:2009:SLM

[1535] Chad M. Schafer and Kjell A. Doksum. Selecting local models in multiple regression by maximizing power. Metrika, 69(2–3):283–304, March 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s00184-008-0218-
z.

Sotto:2009:MPM

[1536] Cristina Sotto, Caroline Beunckens, Geert Molenberghs, Ivy Jansen, and Geert Verbeke. Marginalizing pattern-mixture models for categorical data subject to monotone missingness. Metrika, 69(2–3):305–336, March 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-008-
0219-y.

Banjevic:2009:RUL

[1537] Dragan Banjevic. Remaining useful life in theory and practice. Metrika, 69 (2–3):337–349, March 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-008-0220-5.

Balakrishnan:2009:SEI

[1538] N. Balakrishnan. A synthesis of exact inferential results for exponential step-stress models and associated optimal accelerated life-tests. Metrika, 69 (2–3):351–396, March 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-008-0221-4.

Riccomagno:2009:SHA

[1539] Eva Riccomagno. A short history of algebraic statistics. Metrika, 69 (2–3):397–418, March 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-008-0222-3.

Anonymous:2009:HCb

[1540] Anonymous. Help & contacts. Metrika, 69(2–3):??, March 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Amblard:2009:NEB

[1541] Cécile Amblard and Stéphane Girard. A new extension of bivariate FGM copulas. Metrika, 70(1):1–17, June 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-008-
                                        0174-7.

Ahmadi:2009:PRG

[1542] Jafar Ahmadi, Mohammad Jafari Jozani, Éric Marchand, and Ahmad Parsian. Prediction of k-records from a general class of distributions under balanced type loss functions. Metrika, 70(1):19–33, June 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-008-0176-5.

Dodge:2009:PQR

[1543] Yadolah Dodge and Joe Whittaker.

Partial quantile regression. Metrika, 70(1):35–57, June 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-008-0177-4.

Lin:2009:HDL

[1544] Jin-Guan Lin, Li-Xing Zhu, and Feng-Chang Xie. Heteroscedasticity diagnostics for t linear regression models. Metrika, 70(1):59–77, June 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-008-
0179-2.

Qin:2009:CBU

[1545] Hong Qin, Na Zou, and Kashinath Chatterjee. Connection between uniformity and minimum moment aberration. Metrika, 70(1):79–88, June 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-008-
0180-9.

Balakrishnan:2009:RPR

[1546] N. Balakrishnan, M. Doostparast, and J. Ahmadi. Reconstruction of past records. Metrika, 70(1):89–109, June 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-008-
                                        0181-8.

Albers:2009:CC

[1547] Willem Albers and Wilbert C. M. Kallenberg. CUMIN charts. Metrika, 70(1):111–130, June 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s00184-008-0184-5.pdf.

Anonymous:2009:HCc

[1548] Anonymous. Help & contacts. Metrika, 70(1):??, June 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Dukes:2009:EBS

[1549] Peter J. Dukes and Alan C. H. Ling. Existence of balanced sampling plans avoiding cyclic distances. Metrika, 70 (2):131–140, September 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-008-0182-7.

Zou:2009:CCS

[1550] Changliang Zou, Yukun Liu, and Zhaojun Wang. Comparisons of control schemes for monitoring the means of processes subject to drifts. Metrika, 70 (2):141–163, September 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-008-0183-6.

Kharrati-Kopaei:2009:CMS

[1551] Mahmood Kharrati-Kopaei. Comparing means of several treatments with a control when one observation exists per treatment. Metrika, 70(2): 165–176, September 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-008-0185-4.

Roelant:2009:MWC

[1552] Ella Roelant, Stefan Van Aelst, and Gert Willems. The minimum weighted covariance determinant estimator. Metrika, 70(2):177–204, September 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-008-
0186-3.

Steinebach:2009:MRR

[1553] Josef G. Steinebach. Monitoring risk in a ruin model perturbed by diffusion. Metrika, 70(2):205–224, September 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-008-
0187-2.

Ortiz:2009:ODG

[1554] I. Ortiz, I. Martínez, C. Rodríguez, and Y. del Águila. Optimal designs for generalized linear models with biased response. Metrika, 70(2): 225–237, September 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-008-0188-1.

Huang:2009:ODL

[1555] Mong-Na Lo Huang and Miao-Kuan Huang. ϕ_{p}-Optimal designs for a linear log contrast model for experiments with mixtures. Metrika, 70(2): 239–256, September 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-008-0190-7.

Anonymous:2009:HCd

[1556] Anonymous. Help & contacts. Metrika, 70(2):??, September 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Koukouvinos:2009:GCO

[1557] C. Koukouvinos and K. Mylona. A general construction of E(s^{2})-optimal supersaturated designs via supplementary difference sets. Metrika, 70(3): 257–265, November 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-008-0189-0.

Bouza:2009:RSS

[1558] Carlos N. Bouza. Ranked set sampling and randomized response procedures for estimating the mean of a sensitive quantitative character. Metrika, 70 (3):267–277, November 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-008-0191-6.

Hu:2009:JBP

[1559] Aiping Hu, Zuoxiang Peng, and Yongcheng Qi. Joint behavior of point process of exceedances and partial sum from a Gaussian sequence. Metrika, 70 (3):279–295, November 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-008-0192-5.

Reyen:2009:SMT

[1560] Salem S. Reyen, John J. Miller, and Edward J. Wegman. Separating a mixture of two normals with proportional covariances. Metrika, 70(3): 297–314, November 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-008-0193-4.

Quessy:2009:TEC

[1561] Jean-François Quessy. Theoretical efficiency comparisons of independence tests based on multivariate versions of Spearman’s rho. Metrika, 70(3): 315–338, November 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-008-0194-3.

Chang:2009:ODP

[1562] Fu-Chuen Chang, Hsiu-Ching Chang, and Sheng-Shian Wang. D-optimal designs for polynomial regression with exponential weight function. Metrika, 70 (3):339–354, November 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-008-0195-2.

Tsai:2009:CRO

[1563] Min-Hsiao Tsai. Criterion-robust optimal product designs for discrimination between nested response surface models. Metrika, 70(3):355–367, November 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-008-
                                        0196-1.

Rychlik:2009:BES

[1564] Tomasz Rychlik. Bounds on expectations of small order statistics from decreasing density populations. Metrika, 70(3):369–381, November 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-008-
                                        0200-9.

Anonymous:2009:HCe

[1565] Anonymous. Help & contacts. Metrika, 70(3):??, November 2009. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Yue:2010:BRD

[1566] Rong-Xian Yue and Xiao-Dong Zhou. Bayesian robust designs for linear models with possible bias and correlated errors. Metrika, 71(1):1–15, January 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-008-
                                        0197-0.

Takagi:2010:ALR

[1567] Y. Takagi. Asymptotics of likelihood ratio statistic for non-inferiority testing under exponential distribution with fixed censoring times. Metrika, 71 (1):17–31, January 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-008-0198-z.

Rueda:2010:MCE

[1568] M. Rueda, I. Sánchez-Borrego, A. Arcos, and S. Martínez. Model-calibration estimation of the distribution function using nonparametric regression. Metrika, 71(1):33–44, January 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s00184-008-0199-
y.

Chen:2010:BCS

[1569] Qiang Chen, Lu Lin, and Lixing Zhu.

Bias-corrected smoothed score function for single-index models. Metrika, 71(1):45–58, January 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-008-0201-8.

Gupta:2010:GCB

[1570] Ramesh C. Gupta and Hui Tao. A generalized correlated binomial distribution with application in multiple testing problems. Metrika, 71 (1):59–77, January 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-008-0202-7.

Arashi:2010:EPP

[1571] M. Arashi, A. K. Md. E. Saleh, and S. M. M. Tabatabaey. Estimation of parameters of parallelism model with elliptically distributed errors. Metrika, 71(1):79–100, January 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-008-0203-6.

Tsou:2010:RLI

[1572] Tsung-Shan Tsou. Robust likelihood inferences about regression parameters for general bivariate continuous data. Metrika, 71(1):101–115, January 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-008-
                                        0204-5.

Sreehari:2010:EMD

[1573] M. Sreehari and S. Ravi. On extremes of mixtures of distributions. Metrika, 71(1):117–123, January 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-008-0206-3.

Anonymous:2010:HCa

[1574] Anonymous. Help & contacts. Metrika, 71(1):??, January 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Vexler:2010:OHT

[1575] Albert Vexler, Chengqing Wu, and Kai Fun Yu. Optimal hypothesis testing: from semi to fully Bayes factors. Metrika, 71(2):125–138, March 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-008-
                                        0205-4.

Durante:2010:NEN

[1576] Fabrizio Durante and Pier Luigi Papini. Non-exchangeability of negatively dependent random variables. Metrika, 71(2):139–149, March 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-008-0207-2.

Li:2010:ACN

[1577] Qizhai Li, Guoying Li, and Shifeng Xiong. Assessment of cell number for a multinomial distribution with application to genomic data. Metrika, 71 (2):151–164, March 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-008-0223-2.

Presno:2010:CLD

[1578] María José Presno and Manuel Landajo. Computation of limiting distributions in stationarity testing with a generic trend. Metrika, 71(2):165–183, March 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-008-
0224-1.

Ramos-Guajardo:2010:OST

[1579] Ana Belén Ramos-Guajardo, Ana Colubi, Gil González-Rodríguez, and María Ángeles Gil. One-sample tests for a generalized Fréchet variance of a fuzzy random variable. Metrika, 71 (2):185–202, March 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-008-0225-0.

Carleos:2010:NFD

[1580] Carlos Carleos and Miguel López-Díaz. A new family of dispersive orderings. Metrika, 71(2):203–217, March 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-008-
                                        0226-z.

Pronzato:2010:OSA

[1581] Luc Pronzato. One-step ahead adaptive D-optimal design on a finite design space is asymptotically optimal. Metrika, 71(2):219–238, March 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-008-
                                        0227-y.

Siburg:2010:MMC

[1582] Karl Friedrich Siburg and Pavel A. Stoimenov. A measure of mutual complete dependence. Metrika, 71 (2):239–251, March 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-008-0229-9.

Anonymous:2010:HCb

[1583] Anonymous. Help & contacts. Metrika, 71(2):??, March 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Leskow:2010:BPR

[1584] Jacek Leśkow and Rafał Synowiecki. On bootstrapping periodic random arrays with increasing period. Metrika, 71(3):253–279, May 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-008-0228-x.

Kuhnt:2010:BCC

[1585] Sonja Kuhnt. Breakdown concepts for contingency tables. Metrika, 71 (3):281–294, May 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-008-0230-3.

Dasgupta:2010:PRF

[1586] Nairanjana Dasgupta, Mike Jacroux, and Rita SahaRay. Partially replicated fractional factorial designs. Metrika, 71(3):295–311, May 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0231-x.

Navarro:2010:SPM

[1587] Jorge Navarro and Moshe Shaked.

Some properties of the minimum and the maximum of random variables with joint logconcave distributions. Metrika, 71(3):313–317, May 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0232-9.

Chen:2010:LLE

[1588] Jia Chen and Li-Xin Zhang. Local linear M-estimation for spatial processes in fixed-design models. Metrika, 71(3):319–340, May 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0233-8.

Huang:2010:UEM

[1589] Kuo-Chung Huang. Unbiased estimators of mean, variance and sensitivity level for quantitative characteristics in finite population sampling. Metrika, 71(3):341–352, May 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-009-0234-7.

Mi:2010:EQD

[1590] Jie Mi. Estimation of quantities determined as implicit functions of unknown parameters. Metrika, 71(3):353–359, May 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-009-
                                        0235-6.

Sharma:2010:OBD

[1591] M. K. Sharma and Sileshi Fanta. Optimal block designs for diallel crosses. Metrika, 71(3):361–372, May 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/content/pdf/10.1007/s00184-
                                        009-0236-5.pdf.

Anonymous:2010:HCc

[1592] Anonymous. Help & contacts. Metrika, 71(3):??, May 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Liang:2010:ANE

[1593] Han-Ying Liang and Jacobo de Uña-Álvarez. Asymptotic normality for estimator of conditional mode under left-truncated and dependent observations. Metrika, 72(1):1–19, July 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0237-4.

Miyaoka:2010:NCA

[1594] Etsuo Miyaoka and Kazuo Noda. Necessary conditions for admissibility of matrix linear estimators in a multivariate linear model. Metrika, 72(1):21–35, July 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0238-3.

Stepien-Baran:2010:MIE

[1595] Agnieszka Stepień-Baran. Minimax invariant estimator of a continuous distribution function under a general loss function. Metrika, 72(1):37–49, July 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0239-2.

Arashi:2010:ELP

[1596] M. Arashi and S. M. M. Tabatabaey.

Estimation of the location parameter under LINEX loss function: multivariate case. Metrika, 72(1):51–57, July 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0240-9.

Tavangar:2010:SMP

[1597] Mahdi Tavangar and Majid Asadi. A study on the mean past lifetime of the components of (n − k + 1)-out-of-n system at the system level. Metrika, 72(1):59–73, July 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-009-0241-8.

Barabesi:2010:BEP

[1598] Lucio Barabesi and Marzia Marcheselli. Bayesian estimation of proportion and sensitivity level in randomized response procedures. Metrika, 72(1):75–88, July 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-009-
                                        0242-7.

Balakrishnan:2010:SMM

[1599] N. Balakrishnan and G. Iliopoulos. Stochastic monotonicity of the MLEs of parameters in exponential simple stepstress models under Type-I and Type-II censoring. Metrika, 72(1):89–109, July 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-009-
                                        0243-6.

Harms:2010:KNR

[1600] Torsten Harms and Pierre Duchesne. On kernel nonparametric regression designed for complex survey data. Metrika, 72(1):111–138, July 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-009-
                                        0244-5.

Dodge:2010:DD

[1601] Yadolah Dodge and Iraj Yadegari. On direction of dependence. Metrika, 72(1):139–150, July 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0273-0.

Anonymous:2010:HCd

[1602] Anonymous. Help & contacts.

Metrika, 72(1):??, July 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Borzadaran:2010:VBB

[1603] G. R. Mohtashami Borzadaran, A. H. Rezaei Roknabadi, and M. Khorashadizadeh. A view on Bhattacharyya bounds for inverse Gaussian distributions. Metrika, 72(2):151–161, September 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0245-4.

Mielniczuk:2010:EFI

[1604] Jan Mielniczuk and Małgorzata Wojtyś. Estimation of Fisher information using model selection. Metrika, 72(2): 163–187, September 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0246-3.

Irle:2010:PMS

[1605] Albrecht Irle and Vladimir Lotov. On the properties of a multiple sequential test. Metrika, 72(2):189–198, September 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0247-2.

Wen:2010:SML

[1606] Chi-Chung Wen. Semiparametric maximum likelihood estimation in Cox proportional hazards model with covariate measurement errors. Metrika, 72 (2):199–217, September 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0248-1.

Yue:2010:BTD

[1607] Rong-Xian Yue and Kashinath Chatterjee. Bayesian U-type design for nonparametric response surface prediction. Metrika, 72(2):219–231, September 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0249-0.

Withers:2010:MEM

[1608] Christopher S. Withers and Saralees Nadarajah. The misclassification error of the maximum likelihood procedure when deciding among a finite choice of distributions. Metrika, 72 (2):233–250, September 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0250-7.

Asadi:2010:MRR

[1609] Majid Asadi and Mohammad Z. Raqab. The mean residual of record values at the level of previous records. Metrika, 72(2):251–264, September 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.

com/article/10.1007/s00184-009-
0251-6.

Bakouch:2010:ZTP

[1610] Hassan S. Bakouch and Miroslav M.

Ristić. Zero truncated Poisson integervalued AR(1) model. Metrika, 72(2): 265–280, September 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0252-5.

Harman:2010:OPD

[1611] Radoslav Harman and Frantisek Stulajter. Optimal prediction designs in finite discrete spectrum linear regression models. Metrika, 72(2): 281–294, September 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0253-4.

Anonymous:2010:HCe

[1612] Anonymous. Help & contacts. Metrika, 72(2):??, September 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

McGree:2010:IDS

[1613] J. M. McGree and J. A. Eccleston.

Investigating design for survival models. Metrika, 72(3):295–311, November 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0254-3.

Tian:2010:EEP

[1614] Yongge Tian. On equalities of estimations of parametric functions under a general linear model and its restricted models. Metrika, 72(3): 313–330, November 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0255-2.

Rodriguez-Lallena:2010:MCQ

[1615] José Antonio Rodríguez-Lallena and Manuel Úbeda-Flores. Multivariate copulas with quadratic sections in one variable. Metrika, 72(3): 331–349, November 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0256-1.

Keating:2010:PEL

[1616] J. P. Keating, R. L. Mason, and N. Balakrishnan. Percentile estimators in location-scale parameter families under absolute loss. Metrika, 72(3): 351–367, November 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0257-0.

Gonzalez-Barrios:2010:DRS

[1617] José M. González-Barrios, Federico O’Reilly, and Raúl Rueda. Durbin’s random substitution and conditional Monte Carlo. Metrika, 72(3):369– 383, November 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0258-z.

Mikusinski:2010:SAC

[1618] Piotr Mikusiński and Michael D. Taylor. Some approximations of n-copulas.

Metrika, 72(3):385–414, November 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0259-y.

Zhou:2010:MCE

[1619] Xingcai Zhou, Xinsheng Liu, and Shuhe Hu. Moment consistency of estimators in partially linear models under NA samples. Metrika, 72(3): 415–432, November 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0260-5.

Yu:2010:GBB

[1620] Qiqing Yu, Cuixian Chen, and G. Y. C.

Wong. The GMLE based Buckley– James estimator with modified casecohort data. Metrika, 72(3):433– 464, November 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0261-4.

Anonymous:2010:HCf

[1621] Anonymous. Help & contacts. Metrika, 72(3):??, November 2010. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Carlini:2011:CSM

[1622] Enrico Carlini and Fabio Rapallo. A class of statistical models to weaken independence in two-way contingency tables. Metrika, 73(1):1–22, January 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.

com/article/10.1007/s00184-009-
0262-3.

Todem:2011:EML

[1623] David Todem and KyungMann Kim. On the existence of maximum likelihood estimates in random effects models for clustered multivariate binary data. Metrika, 73(1):23–41, January 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0263-2.

Mattner:2011:OOO

[1624] Lutz Mattner. One optional observation inflates α by 100/\sqrt{n} per cent. Metrika, 73(1):43–59, January 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0264-1.

Peng:2011:ASC

[1625] Zuoxiang Peng, Zhongquan Tan, and Saralees Nadarajah. Almost sure Central Limit Theorem for the products of U-statistics. Metrika, 73(1):61–76, January 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0265-0.

Garcia-Perez:2011:ALT

[1626] A. García-Pérez. Another look at the tail area influence function. Metrika, 73(1):77–92, January 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0266-z.

Chang:2011:SEG

[1627] Yuan chin Ivan Chang. Sequential estimation in generalized linear models when covariates are subject to errors. Metrika, 73(1):93–120, January 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0267-y.

Zhou:2011:EEE

[1628] Bin Zhou, Qinfeng Xu, and Jinhong You. Efficient estimation for error component seemingly unrelated nonparametric regression models. Metrika, 73 (1):121–138, January 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0268-x.

Anonymous:2011:HCa

[1629] Anonymous. Help & contacts. Metrika, 73(1):??, January 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Saha:2011:OSR

[1630] Amitava Saha. An optional scrambled randomized response technique for practical surveys. Metrika, 73 (2):139–149, March 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0269-9.

Chiang:2011:MFJ

[1631] Chin-Tsang Chiang. A more flexible joint latent model for longitudinal and survival time data. Metrika, 73(2):151–170, March 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0270-3.

Wang:2011:ELI

[1632] Xiuli Wang, Gaorong Li, and Lu Lin. Empirical likelihood inference for semiparametric varying-coefficient partially linear EV models. Metrika, 73 (2):171–185, March 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0271-2.

Boudt:2011:REE

[1633] Kris Boudt, Derya Caliskan, and Christophe Croux. Robust explicit estimators of Weibull parameters. Metrika, 73(2):187–209, March 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0272-1.

Zhu:2011:ETP

[1634] Fukang Zhu and Dehui Wang. Estimation and testing for a Poisson autoregressive model. Metrika, 73 (2):211–230, March 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0274-z.

Pardo:2011:TER

[1635] M. C. Pardo. Testing equality restrictions in generalized linear models for multinomial data. Metrika, 73 (2):231–253, March 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0275-y.

Grane:2011:DTE

[1636] Aurea Grané and Josep Fortiana. A directional test of exponentiality based on maximum correlations. Metrika, 73(2):255–274, March 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0276-x.

Anonymous:2011:HCb

[1637] Anonymous. Help & contacts. Metrika, 73(2):??, March 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Yang:2011:PTL

[1638] Hu Yang and Jianwen Xu. Preliminary test Liu estimators based on the conflicting W, LR and LM tests in a regression model with multivariate Student-t error. Metrika, 73(3):275–292, May 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0277-9.

Eryilmaz:2011:DRS

[1639] Serkan Eryilmaz and Femin Yalçin.

Distribution of run statistics in partially exchangeable processes. Metrika, 73(3):293–304, May 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0278-8.

Sun:2011:CBU

[1640] Fasheng Sun, Jie Chen, and Min-Qian Liu. Connections between uniformity and aberration in general multi-level factorials. Metrika, 73(3):305–315, May 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0279-7.

Diaz-Garcia:2011:NBV

[1641] José A. Díaz-García and Ramón Gutiérrez-Jáimez. Noncentral bimatrix variate generalised beta distributions. Metrika, 73(3):317–333, May 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0280-1.

Balakrishnan:2011:EIP

[1642] N. Balakrishnan, Donghoon Han, and G. Iliopoulos. Exact inference for progressively Type-I censored exponential failure data. Metrika, 73(3):335–358, May 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0281-0.

Coiffard:2011:HDE

[1643] Claire Coiffard. On the Hausdorff dimension of exceptional random sets generated by multivariate spacings. Metrika, 73(3):359–371, May 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0282-z.

Chernov:2011:FCS

[1644] N. Chernov. Fitting circles to scattered data: parameter estimates have no moments. Metrika, 73(3):373–384, May 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0283-y.

Chatzopoulos:2011:ORO

[1645] Stavros A. Chatzopoulos, Fotini Kolyva-Machera, and Kashinath Chatterjee. Optimality results on orthogonal arrays plus p runs for s^{m} factorial experiments. Metrika, 73(3):385–394, May 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0284-x.

Rosadi:2011:ECF

[1646] Dedi Rosadi and Manfred Deistler. Estimating the codifference function of linear time series models with infinite variance. Metrika, 73(3):395–429, May 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0285-9.

Anonymous:2011:HCc

[1647] Anonymous. Help & contacts.

Metrika, 73(3):??, May 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Kumar:2011:DMI

[1648] Vikas Kumar, H. C. Taneja, and R. Srivastava. A dynamic measure of inaccuracy between two past lifetime distributions. Metrika, 74(1):1–10, July 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0286-8.

Madadi:2011:SIR

[1649] Mohsen Madadi and Mahbanoo Tata. Shannon information in record data. Metrika, 74(1):11–31, July 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s00184-009-0287-
7.

Mirakhmedov:2011:CTL

[1650] Sherzod M. Mirakhmedov, Syed Ikram A. Tirmizi, and Muhammad Naeem. A Cramér-type large deviation theorem for sums of functions of higher order non-overlapping spacings. Metrika, 74(1):33–54, July 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0288-6.

Baran:2011:AIO

[1651] Sándor Baran and Gyula Pap. Asymptotic inference for a one-dimensional simultaneous autoregressive model. Metrika, 74(1):55–66, July 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s00184-009-0289-
5.

Albers:2011:CCH

[1652] Willem Albers. Control charts for health care monitoring under overdispersion. Metrika, 74(1):67–83, July 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00184-
009-0290-z.pdf.

Anido:2011:APA

[1653] Carmen Anido, Carlos Rivero, and Teofilo Valdes. An algorithm for panel ANOVA with grouped data. Metrika, 74(1):85–107, July 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-009-0291-y.

Ou:2011:LBV

[1654] Zujun Ou, Kashinath Chatterjee, and Hong Qin. Lower bounds of various discrepancies on combined designs. Metrika, 74(1):109–119, July 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0292-x.

Hwang:2011:RAO

[1655] Leng-Cheng Hwang. A robust asymptotically optimal sequential estimation procedure for the Poisson process. Metrika, 74(1):121–133, July 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-009-
0293-9.

Anonymous:2011:HCd

[1656] Anonymous. Help & contacts.

Metrika, 74(1):??, July 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Muller:2011:DDS

[1657] Christine H. Müller. Data depth for simple orthogonal regression with application to crack orientation. Metrika, 74(2):135–165, September 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-009-
                                        0294-8.

DeOliveira:2011:MLR

[1658] Victor De Oliveira and Marco A. R. Ferreira. Maximum likelihood and restricted maximum likelihood estimation for a class of Gaussian Markov random fields. Metrika, 74(2):167– 183, September 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-009-0295-7.

Zhou:2011:VSA

[1659] Zhangong Zhou, Rong Jiang, and Weimin Qian. Variable selection for additive partially linear models with measurement error. Metrika, 74(2): 185–202, September 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-009-0296-6.

Zhao:2011:DOF

[1660] Peng Zhao and N. Balakrishnan. Dispersive ordering of fail-safe systems with heterogeneous exponential components. Metrika, 74(2): 203–210, September 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-010-0297-5.

Jokiel-Rokita:2011:BSE

[1661] Alicja Jokiel-Rokita. Bayes sequential estimation for a particular exponential family of distributions under LINEX loss. Metrika, 74(2):211–219, September 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-010-
0298-4.

Hron:2011:SPT

[1662] Karel Hron and Lubomír Kubácek.

Statistical properties of the total variation estimator for compositional data. Metrika, 74(2):221–230, September 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-010-
0299-3.

Zhao:2011:VSV

[1663] Peixin Zhao and Liugen Xue. Variable selection for varying coefficient models with measurement errors. Metrika, 74 (2):231–245, September 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-010-0300-1.

Withers:2011:DMF

[1664] Christopher S. Withers and Saralees Nadarajah. The distribution of the maximum of a first order autoregressive process: the continuous case. Metrika, 74(2):247–266, September 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00184-010-
                                        0301-0.

Zahle:2011:RAS

[1665] Henryk Zähle. Rates of almost sure convergence of plug-in estimates for distortion risk measures. Metrika, 74 (2):267–285, September 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-010-0302-z.

Volkmer:2011:DOS

[1666] Hans Volkmer and G. G. Hamedani. On distributions of order statistics and their applications to uniform distribution. Metrika, 74(2):287–295, September 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-010-
                                        0303-y.

Anonymous:2011:HCe

[1667] Anonymous. Help & contacts. Metrika, 74(2):??, September 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Kim:2011:CPT

[1668] Moosup Kim and Sangyeol Lee. Change point test for tail index for dependent data. Metrika, 74(3): 297–311, November 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-010-0304-x.

Plancade:2011:MSH

[1669] Sandra Plancade. Model selection for hazard rate estimation in presence of censoring. Metrika, 74(3):

313–347, November 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-010-0305-9.

Lin:2011:APM

[1670] Chien-Tai Lin and N. Balakrishnan.

Asymptotic properties of maximum likelihood estimators based on progressive Type-II censoring. Metrika, 74 (3):349–360, November 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-010-0306-8.

Raqab:2011:BMR

[1671] Mohammad Z. Raqab and Tomasz Rychlik. Bounds for the mean residual life function of a k-out-of-n system. Metrika, 74(3):361–380, November 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-010-
0307-7.

Haslett:2011:EBU

[1672] Stephen J. Haslett and Simo Puntanen.

On the equality of the BLUPs under two linear mixed models. Metrika, 74 (3):381–395, November 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-010-0308-6.

Arnold:2011:UPM

[1673] Bernhard F. Arnold and Peter Stahlecker. An unexpected property of minimax estimation in the relative squared error approach to linear regression analysis. Metrika, 74(3):

397–407, November 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-010-0309-5.

Li:2011:TCV

[1674] Na Li, Xingzhong Xu, and Xuhua Liu. Testing the constancy in varying-coefficient regression models. Metrika, 74(3):409–438, November 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-010-
                                        0310-z.

Anonymous:2011:HCf

[1675] Anonymous. Help & contacts. Metrika, 74(3):??, November 2011. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Anonymous:2012:WU

[1676] Anonymous. Werner Uhlmann (1928– 2011). Metrika, 75(1):1–3, January 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-011-
                                        0370-8.

Arefi:2012:FDE

[1677] Mohsen Arefi, Reinhard Viertl, and S. Mahmoud Taheri. Fuzzy density estimation. Metrika, 75(1):5–22, January 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-010-
                                        0311-y.

Pang:2012:NCA

[1678] Fang Pang and Min-Qian Liu. A note on connections among criteria for asymmetrical factorials. Metrika, 75(1):23–32, January 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-010-0312-x.

Liu:2012:CEW

[1679] Yan Liu and Min-Qian Liu. Construction of equidistant and weak equidistant supersaturated designs. Metrika, 75(1):33–53, January 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-010-0313-9.

Huang:2012:ELV

[1680] Zhensheng Huang. Empirical likelihood for varying-coefficient singleindex model with right-censored data. Metrika, 75(1):55–71, January 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s00184-010-0314-
8.

Kumar:2012:LBD

[1681] Vikas Kumar and H. C. Taneja. On length biased dynamic measure of past inaccuracy. Metrika, 75(1):73–84, January 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-010-
0315-7.

Singh:2012:CDW

[1682] Sarjinder Singh. On the calibration of design weights using a displacement function. Metrika, 75(1):85–107, January 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-010-
                                        0316-6.

Maiboroda:2012:NDE

[1683] Rostyslav Maiboroda and Olena Sugakova. Nonparametric density estimation for symmetric distributions by contaminated data. Metrika, 75 (1):109–126, January 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-010-0317-5.

Koul:2012:GFT

[1684] Hira L. Koul and Nao Mimoto. A goodness-of-fit test for GARCH innovation density. Metrika, 75(1): 127–149, January 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-010-0318-4.

Anonymous:2012:HCa

[1685] Anonymous. Help & contacts. Metrika, 75(1):??, January 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Gabler:2012:OAS

[1686] Siegfried Gabler, Matthias Ganninger, and Ralf Münnich. Optimal allocation of the sample size to strata under box constraints. Metrika, 75(2): 151–161, February 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-010-0319-3.

Hashorva:2012:LTS

[1687] Enkelejd Hashorva and Alexei Stepanov.

Limit theorems for the spacings of weak records. Metrika, 75(2): 163–180, February 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-010-0320-x.

Gupta:2012:LCM

[1688] Ramesh C. Gupta and N. Balakrishnan. Log-concavity and monotonicity of hazard and reversed hazard functions of univariate and multivariate skew-normal distributions. Metrika, 75 (2):181–191, February 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-010-0321-9.

Abtahi:2012:GCU

[1689] A. Abtahi, J. Behboodian, and M. Sharafi. A general class of univariate skew distributions considering Stein’s lemma and infinite divisibility. Metrika, 75(2):193–206, February 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-010-
0322-8.

Withers:2012:UEL

[1690] Christopher S. Withers and Saralees Nadarajah. Unbiased estimates for a lognormal regression problem and a nonparametric alternative. Metrika, 75 (2):207–227, February 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-010-0323-7.

Khardani:2012:SUC

[1691] Salah Khardani, Mohamed Lemdani, and Elias Ould Saïd. On the strong uniform consistency of the mode estimator for censored time series. Metrika, 75 (2):229–241, February 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-010-0324-6.

Papadatos:2012:LEL

[1692] Nickos Papadatos. Linear estimation of location and scale parameters using partial maxima. Metrika, 75(2): 243–270, February 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-010-0325-5.

Sreehari:2012:CMG

[1693] M. Sreehari and R. Vasudeva. Characterizations of multivariate geometric distributions in terms of conditional distributions. Metrika, 75(2): 271–286, February 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-010-0326-4.

Anonymous:2012:HCb

[1694] Anonymous. Help & contacts. Metrika, 75(2):??, February 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Fontanella:2012:MKL

[1695] Lara Fontanella, Luigi Ippoliti, and Arcangelo Merla. Multiresolution Karhunen Loéve analysis of galvanic skin response for psycho-physiological studies. Metrika, 75(3):287–309, April 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-010-
0327-3.

Saleh:2012:EPM

[1696] A. K. Md. Ehsanes Saleh, Jan Picek, and Jan Kalina. R-estimation of the parameters of a multiple regression model with measurement errors. Metrika, 75(3):311–328, April 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-010-
0328-2.

Fang:2012:ASB

[1697] Yun Fang and Li-Xing Zhu. Asymptotics of SIMEX-based variance estimation. Metrika, 75(3):329–345, April 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-010-
0329-1.

Liu:2012:DDF

[1698] Juxin Liu and Paul Gustafson. On the detectability of different forms of interaction in regression models. Metrika, 75(3):347–365, April 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-010-0330-8.

Ng:2012:PIS

[1699] Hon Keung Tony Ng, Jorge Navarro, and Narayanaswamy Balakrishnan. Parametric inference from system lifetime data under a proportional hazard rate model. Metrika, 75(3):367–388, April 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-010-
                                        0331-7.

Nair:2012:SRA

[1700] N. Unnikrishnan Nair and P. G. Sankaran. Some results on an additive hazards model. Metrika, 75 (3):389–402, April 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-010-0332-6.

Malyarenko:2012:PEP

[1701] A. Malyarenko and V. Vasiliev. On parameter estimation of partly observed bilinear discrete-time stochastic systems. Metrika, 75(3):403–424, April 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-010-
                                        0333-5.

Jahufer:2012:ILI

[1702] Aboobacker Jahufer and Jianbao Chen. Identifying local influential observations in Liu estimator. Metrika, 75 (3):425–438, April 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-010-0334-4.

Anonymous:2012:HCc

[1703] Anonymous. Help & contacts. Metrika, 75(3):??, April 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Salehi:2012:RPL

[1704] E. T. Salehi and M. Asadi. Results on the past lifetime of (n − k + 1)- out-of-n structures with nonidentical components. Metrika, 75(4):439–454, May 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-010-
0335-3.

Zheng:2012:TPC

[1705] Xu Zheng. Testing parametric conditional distributions using the nonparametric smoothing method. Metrika, 75(4):455–469, May 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-010-0336-2.

Wu:2012:SOE

[1706] Mi-Xia Wu, Kai-Fun Yu, Aiyi Liu, and Tie-Feng Ma. Simultaneous optimal estimation in linear mixed models. Metrika, 75(4):471–489, May 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-010-
0337-1.

Petropoulos:2012:NCI

[1707] Constantinos Petropoulos and Stavros Kourouklis. New classes of improved confidence intervals for the variance of a normal distribution. Metrika, 75(4):491–506, May 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-010-0338-0.

Misiewicz:2012:WPP

[1708] Jolanta Misiewicz and Jacek Wesolowski. Winding planar probabilities. Metrika, 75(4):507–519, May 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00184-010-0339-z.pdf.

Mojirsheibani:2012:SRC

[1709] Majid Mojirsheibani. Some results on classifier selection with missing covariates. Metrika, 75(4):521–539, May 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-010-
0340-6.

Diaz-Garcia:2012:GST

[1710] José A. Díaz-García and Francisco J. Caro-Lopera. Generalised shape theory via SV decomposition I. Metrika, 75(4):541–565, May 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-010-0341-5.

Mahmoudi:2012:AND

[1711] Eisa Mahmoudi. Asymptotic nondeficiency of the Bayes sequential estimation in a family of transformed chi-square distributions. Metrika, 75 (4):567–580, May 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-011-0342-z.

Anonymous:2012:HCd

[1712] Anonymous. Help & contacts.

Metrika, 75(4):??, May 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Kang:2012:FII

[1713] Shin-Soo Kang, Kenneth J. Koehler, and Michael D. Larsen. Fractional imputation for incomplete two-way contingency tables. Metrika, 75 (5):581–599, July 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-011-0343-y.

Hering:2012:MBE

[1714] Christian Hering and Jan-Frederik Mai. Moment-based estimation of extendible Marshall–Olkin copulas. Metrika, 75(5):601–620, July 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s00184-011-0344-
x.

Lee:2012:SEL

[1715] Shen-Ming Lee, Chin-Shang Li, Shu-Hui Hsieh, and Li-Hui Huang. Semiparametric estimation of logistic regression model with missing covariates and outcome. Metrika, 75(5):621–653, July 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00184-
011-0345-9.pdf.

Chang:2012:EPP

[1716] Horng-Jinh Chang and Mei-Pei Kuo. Estimation of population proportion in randomized response sampling using weighted confidence interval construction. Metrika, 75(5):655–672, July 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-011-
                                        0346-8.

Huskova:2012:BSC

[1717] Marie Husková and Claudia Kirch. Bootstrapping sequential change-point tests for linear regression. Metrika, 75(5):673–708, July 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-011-0347-7.

Filova:2012:CME

[1718] Lenka Filová, Mária Trnovská, and Radoslav Harman. Computing maximin efficient experimental designs using the methods of semidefinite programming. Metrika, 75(5):709–719, July 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-011-
                                        0348-6.

Anonymous:2012:HCe

[1719] Anonymous. Help & contacts. Metrika, 75(5):??, July 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Schneeweiss:2012:PQE

[1720] Hans Schneeweiß, Daniel Rost, and Matthias Schmid. Probability and quantile estimation from individually micro-aggregated data. Metrika, 75 (6):721–742, August 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-011-0349-5.

Yanev:2012:CED

[1721] George P. Yanev. Characterization of exponential distribution via regression of one record value on two nonadjacent record values. Metrika, 75 (6):743–760, August 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-011-0350-z.

Ceyhan:2012:DRA

[1722] Elvan Ceyhan. The distribution of the relative arc density of a family of interval catch digraph based on uniform data. Metrika, 75(6):761–793, August 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-011-
0351-y.

Wu:2012:VSJ

[1723] Liucang Wu and Huiqiong Li. Variable selection for joint mean and dispersion models of the inverse Gaussian distribution. Metrika, 75(6):795–808, August 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-011-
0352-x.

Tsai:2012:EDD

[1724] Min-Hsiao Tsai. Efficient discriminating design for a class of nested polynomial regression models. Metrika, 75 (6):809–817, August 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-011-0353-9.

Frick:2012:MMA

[1725] Melanie Frick. Measures of multivariate asymptotic dependence and their relation to spectral expansions. Metrika, 75(6):819–831, August 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-011-0354-8.

Li:2012:DFT

[1726] Erning Li, Johan Lim, Kyunga Kim, and Shin-Jae Lee. Distribution-free tests of mean vectors and covariance matrices for multivariate paired data. Metrika, 75(6):833–854, August 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-011-
                                        0355-7.

Anonymous:2012:HCf

[1727] Anonymous. Help & contacts. Metrika, 75(6):??, August 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Garcia-Perez:2012:LAP

[1728] A. García-Pérez. A linear approximation to the power function of a test. Metrika, 75(7):855–875, October 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-011-
                                        0356-6.

Beska:2012:MSM

[1729] Marek Beśka, Krzysztof Jasiński, Tomasz Rychlik, and Marcin Spryszyński. Mixed systems with minimal and maximal lifetime variances. Metrika, 75 (7):877–894, October 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00184-011-0357-5.pdf.

Burch:2012:ODD

[1730] Nathanial Burch, Jennifer A. Hoeting, and Donald Estep. Optimal design and directional leverage with applications in differential equation models. Metrika, 75(7):895–911, October 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-011-
0358-4.

Sinha:2012:RAL

[1731] Sanjoy K. Sinha. Robust analysis of longitudinal data with nonignorable missing responses. Metrika, 75 (7):913–938, October 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-011-0359-3.

Lacko:2012:CDA

[1732] Vladimír Lacko and Radoslav Harman. A conditional distribution approach to uniform sampling on spheres and balls in l_{p} spaces. Metrika, 75 (7):939–951, October 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-011-0360-x.

Zhao:2012:MLF

[1733] Shengli Zhao and Xiangfei Chen. Mixed-level fractional factorial splitplot designs containing clear effects. Metrika, 75(7):953–962, October 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00184-011-0361-
                                        9.

Diaz-Garcia:2012:MMM

[1734] José A. Díaz-García and Ramón Gutiérrez-Jáimez. Matricvariate and matrix multivariate T distributions and associated distributions. Metrika, 75 (7):963–976, October 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-011-0362-8.

Coeurjolly:2012:GND

[1735] Jean-François Coeurjolly and Nicolas Le Bihan. Geodesic normal distribution on the circle. Metrika, 75 (7):977–995, October 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-011-0363-7.

Tavangar:2012:SUC

[1736] Mahdi Tavangar and Majid Asadi. Some unified characterization results on the generalized Pareto distributions based on generalized order statistics. Metrika, 75(7):997–1007, October 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-011-
                                        0364-6.

Anonymous:2012:HCg

[1737] Anonymous. Help & contacts. Metrika, 75(7):??, October 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Arevalillo:2012:ERB

[1738] Jorge M. Arevalillo. Exploring the relation between the r∗ approximation and the Edgeworth expansion. Metrika, 75(8):1009–1024, November 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-011-
0365-5.

Ruckdeschel:2012:YAB

[1739] Peter Ruckdeschel and Nataliya Horbenko. Yet another breakdown point notion: EFSBP. Metrika, 75(8): 1025–1047, November 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-011-0366-4.

Ferraro:2012:MLR

[1740] Maria Brigida Ferraro and Paolo Giordani. A multiple linear regression model for imprecise information. Metrika, 75(8):1049–1068, November 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-011-
0367-3.

Filipiak:2012:MEM

[1741] Katarzyna Filipiak and Dietrich von Rosen. On MLEs in an extended multivariate linear growth curve model. Metrika, 75(8):1069–1092, November 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00184-
011-0368-2.pdf.

Langel:2012:ILZ

[1742] Matti Langel and Yves Tillé. Inference by linearization for Zenga’s new inequality index: a comparison with the Gini index. Metrika, 75(8): 1093–1110, November 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-011-0369-1.

Wied:2012:NFT

[1743] Dominik Wied, Matthias Arnold, Nicolai Bissantz, and Daniel Ziggel. A new fluctuation test for constant variances with applications to finance. Metrika, 75(8):1111–1127, November 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-011-
0371-7.

Wang:2012:RWF

[1744] Yunyan Wang, Lixin Zhang, and Mingtian Tang. Re-weighted functional estimation of second-order diffusion processes. Metrika, 75(8): 1129–1151, November 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-011-0372-6.

Anonymous:2012:HCh

[1745] Anonymous. Help & contacts. Metrika, 75(8):??, November 2012. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Balakrishnan:2013:ITA

[1746] N. Balakrishnan, C. Koukouvinos, and C. Parpoula. An information theoretical algorithm for analyzing supersaturated designs for a binary response. Metrika, 76(1):1–18, January 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-011-
0373-5.

Dabo-Niang:2013:KSD

[1747] Sophie Dabo-Niang and Anne-Françoise Yao. Kernel spatial density estimation in infinite dimension space. Metrika, 76(1):19–52, January 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-011-0374-4.

Xu:2013:TAV

[1748] Wangli Xu and Lixing Zhu. Testing the adequacy of varying coefficient models with missing responses at random. Metrika, 76(1):53–69, January 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-011-
0375-3.

Popovic:2013:NNL

[1749] Bozidar V. Popović, Saralees Nadarajah, and Miroslav M. Ristić. A new non-linear AR(1) time series model having approximate beta marginals. Metrika, 76(1):71–92, January 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00184-011-0376-
                                        2.

Cirillo:2013:UBB

[1750] Pasquale Cirillo and Pietro Muliere. An urn-based Bayesian block bootstrap. Metrika, 76(1):93–106, January 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-011-
                                        0377-1.

Skulj:2013:CEM

[1751] D. Skulj and R. Hable. Coefficients of ergodicity for Markov chains with uncertain parameters. Metrika, 76 (1):107–133, January 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-011-0378-0.

Zhang:2013:MBT

[1752] Jin-Ting Zhang and Xuefeng Liu. A modified Bartlett test for heteroscedastic one-way MANOVA. Metrika, 76 (1):135–152, January 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-011-0379-z.

Anonymous:2013:HCa

[1753] Anonymous. Help & contacts. Metrika, 76(1):??, January 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Stoyan:2013:WRE

[1754] Dietrich Stoyan. Weibull, RRSB or extreme-value theorists? Metrika, 76 (2):153–159, February 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-011-0380-6.

Balakrishnan:2013:GFP

[1755] N. Balakrishnan, M. R. Brito, and A. J. Quiroz. On the goodness-of-fit procedure for normality based on the empirical characteristic function for ranked set sampling data. Metrika, 76 (2):161–177, February 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-012-0381-0.

Bernhart:2013:DMB

[1756] German Bernhart, Marcos Escobar Anel, Jan-Frederik Mai, and Matthias Scherer. Default models based on scale mixtures of Marshall–Olkin copulas: properties and applications. Metrika, 76(2):179–203, February 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s00184-012-0382-
z.

Koul:2013:GFT

[1757] Hira L. Koul, Nao Mimoto, and Donatas Surgailis. Goodness-of-fit tests for long memory moving average marginal density. Metrika, 76 (2):205–224, February 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-012-0383-y.

Li:2013:VSP

[1758] Feng Li, Lu Lin, and Yuxia Su. Variable selection and parameter estimation for partially linear models via Dantzig selector. Metrika, 76(2): 225–238, February 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-012-0384-x.

Mattner:2013:CBS

[1759] Lutz Mattner and Frauke Mattner. Confidence bounds for the sensitivity lack of a less specific diagnostic test, without gold standard. Metrika, 76 (2):239–263, February 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-012-0385-9.

Giniewicz:2013:BTT

[1760] Andrzej Giniewicz and Alicja Jokiel-Rokita. Burn-in for a time-transformed exponential model. Metrika, 76(2): 265–285, February 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s00184-012-0386-8.pdf.

Zhou:2013:LVS

[1761] Zhangong Zhou, Rong Jiang, and Weimin Qian. LAD variable selection for linear models with randomly censored data. Metrika, 76 (2):287–300, February 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-012-0387-7.

Anonymous:2013:HCb

[1762] Anonymous. Help & contacts. Metrika, 76(2):??, February 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Henze:2013:E

[1763] Norbert Henze and Udo Kamps. Editorial. Metrika, 76(3):301, April 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00184-
013-0433-0.pdf.

Li:2013:TKV

[1764] Zaixing Li. Two kinds of variance/ covariance estimates in linear mixed models. Metrika, 76(3):303–324, April 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-012-
0388-6.

Tang:2013:OTL

[1765] Boxin Tang and Julie Zhou. D-optimal two-level orthogonal arrays for estimating main effects and some specified two-factor interactions. Metrika, 76(3):325–337, April 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-012-0389-5.

Sharma:2013:UOB

[1766] V. K. Sharma. Universally optimal balanced changeover designs with first residuals. Metrika, 76(3):339–346, April 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.

com/article/10.1007/s00184-012-
0390-z.

Schmidt:2013:EPG

[1767] Karl Michael Schmidt and Anatoly Zhigljavsky. An extremal property of the generalized arcsine distribution. Metrika, 76(3):347–355, April 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-012-
0391-y.

Withers:2013:DEL

[1768] Christopher S. Withers and Saralees Nadarajah. Density estimates of low bias. Metrika, 76(3):357–379, April 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-012-
0392-x.

Lin:2013:HOL

[1769] S. H. Lin. The higher order likelihood method for the common mean of several log-normal distributions. Metrika, 76(3):381–392, April 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-012-0393-9.

Katulska:2013:OCB

[1770] Krystyna Katulska and Lukasz Smaga. D-optimal chemical balance weighing designs with autoregressive errors. Metrika, 76(3):393–407, April 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00184-
012-0394-8.pdf.

Ma:2013:EOR

[1771] Tiefeng Ma and Shuangzhe Liu. Estimation of order-restricted means of two normal populations under the LINEX loss function. Metrika, 76(3):409–425, April 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-012-
0396-6.

Lesitha:2013:ESP

[1772] G. Lesitha and P. Yageen Thomas.

Estimation of the scale parameter of a log-logistic distribution. Metrika, 76(3):427–448, April 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-012-0397-5.

Anonymous:2013:HCc

[1773] Anonymous. Help & contacts. Metrika, 76(3):??, April 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Castano-Martinez:2013:APW

[1774] A. Castaño-Martínez, F. López-Blázquez, and B. Salamanca-Miño. An additive property of weak records from geometric distributions. Metrika, 76(4):449–458, May 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-012-0398-4.

Xu:2013:NCV

[1775] Wangli Xu and Xu Guo. Nonparametric checks for varying coefficient models with missing response at random. Metrika, 76(4):459–482, May 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-012-
0399-3.

Liu:2013:NOD

[1776] Xin Liu and Rong-Xian Yue. A note on R-optimal designs for multiresponse models. Metrika, 76(4):483–493, May 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-012-
0400-1.

Niwitpong:2013:CRE

[1777] Sa aat Niwitpong, Dankmar Böhning, Peter G. M. van der Heijden, and Heinz Holling. Capture-recapture estimation based upon the geometric distribution allowing for heterogeneity. Metrika, 76(4):495–519, May 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-012-0401-0.

Ahmadi:2013:NRV

[1778] Jafar Ahmadi and N. Balakrishnan. On the nearness of record values to order statistics from Pitman’s measure of closeness. Metrika, 76(4):521–541, May 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-012-
0402-z.

Minkova:2013:CWP

[1779] Leda D. Minkova and N. Balakrishnan. Compound weighted Poisson distributions. Metrika, 76(4):543–558, May 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-012-
0403-y.

Amiripour:2013:SOC

[1780] Fariba Amiripour, Baha-Eldin Khaledi, and Moshe Shaked. Stochastic orderings of convolution residuals. Metrika, 76(4):559–576, May 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-012-0404-x.

Zaigraev:2013:OCO

[1781] Alexander Zaigraev and Magdalena Alama-Bućko. On optimal choice of order statistics in large samples for the construction of confidence regions for the location and scale. Metrika, 76(4):577–593, May 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00184-012-0405-9.pdf.

Anonymous:2013:HCd

[1782] Anonymous. Help & contacts.

Metrika, 76(4):??, May 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Golparver:2013:ORR

[1783] Leila Golparver, Ali Karimnezhad, and Ahmad Parsian. Optimal rules and robust Bayes estimation of a gamma scale parameter. Metrika, 76(5):595–622, July 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-012-
0407-7.

Psarrakos:2013:GCR

[1784] Georgios Psarrakos and Jorge Navarro. Generalized cumulative residual entropy and record values. Metrika, 76(5):623–640, July 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-012-0408-6.

Huang:2013:TSP

[1785] Chao Huang, Jin-Guan Lin, and Yan-Yan Ren. Testing for the shape parameter of generalized extreme value distribution based on the L_{q}-likelihood ratio statistic. Metrika, 76(5):641–671, July 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-012-
                                        0409-5.

Du:2013:MBT

[1786] Yeting Du and Johanna Neslehová. A moment-based test for extreme-value dependence. Metrika, 76(5):673–695, July 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-012-
                                        0410-z.

Ghorbani:2013:CCP

[1787] Mohammad Ghorbani. Cauchy cluster process. Metrika, 76(5):697–706, July 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-012-
                                        0411-y.

Lu:2013:RBW

[1788] Changli Lu, Yuqin Sun, and Yongge Tian. On relations between weighted least-squares estimators of parametric functions under a general partitioned linear model and its small models. Metrika, 76(5):707–722, July 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-012-
0412-x.

Wei:2013:SCL

[1789] Jia-Lin Wei and Jian-Feng Yang. On simplifying the calculations leading to designs with general minimum lower-order confounding. Metrika, 76(5):723–732, July 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-013-0442-z.

Anonymous:2013:HCe

[1790] Anonymous. Help & contacts.

Metrika, 76(5):??, July 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Dowla:2013:LBB

[1791] Arif Dowla, Efstathios Paparoditis, and Dimitris N. Politis. Local block bootstrap inference for trending time series. Metrika, 76(6):733–764, August 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-012-
0413-9.

Benhenni:2013:ERE

[1792] Karim Benhenni, Mustapha Rachdi, and Yingcai Su. The effect of the regularity of the error process on the performance of kernel regression estimators. Metrika, 76(6):765–781, August 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-012-
0414-8.

Gietl:2013:API

[1793] Christoph Gietl and Fabian P. Reffel. Accumulation points of the iterative proportional fitting procedure. Metrika, 76(6):783–798, August 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-012-
0415-7.

Hron:2013:EPS

[1794] Karel Hron, Matthias Templ, and Peter Filzmoser. Estimation of a proportion in survey sampling using the logratio approach. Metrika, 76(6):799–818, August 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-012-
0416-6.

Gu:2013:CNO

[1795] Li Gu and Jian-Feng Yang. Construction of nearly orthogonal Latin hypercube designs. Metrika, 76(6): 819–830, August 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-012-0417-5.

Yang:2013:EMC

[1796] Xiao-Rong Yang. Estimation of the mean for critical branching process and its bootstrap approximation. Metrika, 76(6):831–846, August 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-012-0418-4.

Bodnar:2013:ETC

[1797] Taras Bodnar and Arjun K. Gupta. An exact test for a column of the covariance matrix based on a single observation. Metrika, 76(6):847–855, August 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-012-
0419-3.

Chang:2013:EOD

[1798] Fu-Chuen Chang, Lorens Imhof, and Yi-Ying Sun. Exact D-optimal designs for first-order trigonometric regression models on a partial circle. Metrika, 76(6):857–872, August 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-012-0420-x.

Anonymous:2013:HCf

[1799] Anonymous. Help & contacts. Metrika, 76(6):??, August 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Colombi:2013:MDG

[1800] Roberto Colombi and Sabrina Giordano. Monotone dependence in graphical models for multivariate Markov chains. Metrika, 76(7):873–885, October 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-012-
0421-9.

Hong:2013:VSH

[1801] Zhaoping Hong, Yuao Hu, and Heng Lian. Variable selection for highdimensional varying coefficient partially linear models via nonconcave penalty. Metrika, 76(7):887–908, October 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-012-
0422-8.

Rezapour:2013:PDP

[1802] M. Rezapour, M. H. Alamatsaz, and N. Balakrishnan. On properties of dependent progressively Type-II censored order statistics. Metrika, 76 (7):909–917, October 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-012-0423-7.

Barabesi:2013:DBD

[1803] Lucio Barabesi, Giancarlo Diana, and Pier Francesco Perri. Design-based distribution function estimation for stigmatized populations. Metrika, 76 (7):919–935, October 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-012-0424-6.

Nkurunziza:2013:ESI

[1804] Sévérien Nkurunziza. Extension of some important identities in shrinkagepretest strategies. Metrika, 76(7): 937–947, October 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-012-0425-5.

Wang:2013:SEL

[1805] Zhanfeng Wang and Yuan chin Ivan Chang. Sequential estimate for linear regression models with uncertain number of effective variables. Metrika, 76(7):949–978, October 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-012-0426-4.

Nama:2013:RLR

[1806] M. Kelkin Nama, M. Asadi, and Z. Zhang. On the residual lifelengths of the remaining components in a coherent system. Metrika, 76 (7):979–996, October 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-012-0427-3.

Anonymous:2013:HCg

[1807] Anonymous. Help & contacts. Metrika, 76(7):??, October 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Romeo:2013:LSP

[1808] José S. Romeo, Nelson I. Tanaka, Antonio C. Pedroso de Lima, and Victor H. Salinas-Torres. Large sample properties for a class of copulas in bivariate survival analysis. Metrika, 76(8): 997–1015, November 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-012-0428-2.

Ling:2013:SCN

[1809] Xiaoliang Ling and Ping Li. Stochastic comparisons for the number of working components of a system in random environment. Metrika, 76(8): 1017–1030, November 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-012-0429-1.

ElKolei:2013:PEH

[1810] Salima El Kolei. Parametric estimation of hidden stochastic model by contrast minimization and deconvolution. Metrika, 76(8):1031–1081, November 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-013-
0430-3.

Wang:2013:PSE

[1811] Lichun Wang, Peng Lai, and Heng Lian. Polynomial spline estimation for generalized varying coefficient partially linear models with a diverging number of components. Metrika, 76(8): 1083–1103, November 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-013-0431-2.

Bodnar:2013:ABE

[1812] Taras Bodnar, Wolfgang Schmid, and Taras Zabolotskyy. Asymptotic behavior of the estimated weights and of the estimated performance measures of the minimum VaR and the minimum CVaR optimal portfolios for dependent data. Metrika, 76(8):1105–1134, November 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-013-
0432-1.

Anonymous:2013:HCh

[1813] Anonymous. Help & contacts. Metrika, 76(8):??, November 2013. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Bauerle:2014:PGS

[1814] Nicole Bäuerle, Herold Dehling, Achim Klenke, Götz Kersting, and Jens-Peter Kreiss. Preface to the GPSD 2012 special issue of Metrika. Metrika, 77(1):1–4, January 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00184-013-0470-8.pdf.

Alvisio:2014:NEM

[1815] Marcelo Alvisio and Gregory F. Lawler.

Note on the existence and modulus of continuity of the SLE_{8} curve. Metrika, 77(1):5–22, January 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-013-0471-7.

vanderHofstad:2014:UHP

[1816] Remco van der Hofstad and Asaf Nachmias. Unlacing hypercube percolation: a survey. Metrika, 77(1):23–50, January 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-013-
0473-5.

DiGirolami:2014:CBS

[1817] Cristina Di Girolami, Giorgio Fabbri, and Francesco Russo. The covariation for Banach space valued processes and applications. Metrika, 77 (1):51–104, January 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-013-0472-6.

Levy-Leduc:2014:HRS

[1818] C. Lévy-Leduc and M. S. Taqqu. Hermite ranks and U-statistics. Metrika, 77(1):105–136, January 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-013-0474-4.

Stockbridge:2014:DDP

[1819] R. H. Stockbridge. Discussion of dynamic programming and linear programming approaches to stochastic control and optimal stopping in continuous time. Metrika, 77(1): 137–162, January 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-013-0476-2.

Comte:2014:NDE

[1820] Fabienne Comte, Céline Duval, and Valentine Genon-Catalot. Nonparametric density estimation in compound Poisson processes using convolution power estimators. Metrika, 77 (1):163–183, January 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-013-0475-3.

Bagdonavicius:2014:GFT

[1821] V. Bagdonavicius and M. S. Nikulin.

Goodness-of-fit tests for parametric nonhomogeneous Markov processes. Metrika, 77(1):185–209, January 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s00184-013-0477-
1.

Anonymous:2014:HCa

[1822] Anonymous. Help & contacts. Metrika, 77(1):??, January 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Groenitz:2014:NPP

[1823] Heiko Groenitz. A new privacyprotecting survey design for multichotomous sensitive variables. Metrika, 77(2):211–224, February 2014. CO-DEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s00184-012-0406-
8.

Zou:2014:ESI

[1824] Qingming Zou and Zhongyi Zhu.

M-estimators for single-index model using B-spline. Metrika, 77(2): 225–246, February 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-013-0434-z.

Riemer:2014:CDG

[1825] Wolfgang Riemer, Dietrich Stoyan, and Danail Obreschkow. Cuboidal dice and Gibbs distributions. Metrika, 77 (2):247–256, February 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-013-0435-y.

Mielniczuk:2014:BVM

[1826] Jan Mielniczuk and Małgorzata Wojtyś. P -value model selection criteria for exponential families of increasing dimension. Metrika, 77(2): 257–284, February 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00184-013-0436-x.pdf.

Fourdrinier:2014:GBM

[1827] Dominique Fourdrinier, Othmane Kortbi, and William E. Strawderman. Generalized Bayes minimax estimators of location vectors for spherically symmetric distributions with residual vector. Metrika, 77(2):285–296, February 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-013-
0437-9.

Withers:2014:SAB

[1828] Christopher S. Withers and Saralees Nadarajah. Simple alternatives for Box–Cox transformations. Metrika, 77 (2):297–315, February 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-013-0438-8.

Lu:2014:FPL

[1829] Ying Lu, Jiang Du, and Zhimeng Sun. Functional partially linear quantile regression model. Metrika, 77 (2):317–332, February 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-013-0439-7.

Anonymous:2014:HCb

[1830] Anonymous. Help & contacts. Metrika, 77(2):??, February 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Dolati:2014:DPB

[1831] A. Dolati, M. Amini, and S. M. Mirhosseini. Dependence properties of bivariate distributions with proportional (reversed) hazards marginals. Metrika, 77(3):333–347, April 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-013-0440-1.

Eryilmaz:2014:SRC

[1832] Serkan Eryilmaz. A study on reliability of coherent systems equipped with a cold standby component. Metrika, 77(3):349–359, April 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-013-0441-0.

daSilva:2014:SCL

[1833] João Lita da Silva. Strong consistency of least squares estimates in multiple regression models with random regressors. Metrika, 77(3):361–375, April 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-013-
0443-y.

Yang:2014:NSC

[1834] Lei Yang and Xianyi Wu. A new sufficient condition for identifiability of countably infinite mixtures. Metrika, 77(3):377–387, April 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-013-0444-x.

Gob:2014:DAS

[1835] Rainer Göb and Kristina Lurz. Design and analysis of shortest two-sided confidence intervals for a probability under prior information. Metrika, 77(3):389–413, April 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-013-0445-9.

Sun:2014:FVI

[1836] Zhimeng Sun, Zhi Su, and Jingyi Ma. Focused vector information criterion model selection and model averaging regression with missing response. Metrika, 77(3):415–432, April 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-013-
0446-8.

Jureckova:2014:RTH

[1837] Jana Jurecková and Radim Navrátil. Rank tests in heteroscedastic linear model with nuisance parameters. Metrika, 77(3):433–450, April 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-013-
0447-7.

Anonymous:2014:HCc

[1838] Anonymous. Help & contacts. Metrika, 77(3):??, April 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Hu:2014:RWE

[1839] Gaoge Hu, Shesheng Gao, Yongmin Zhong, and Chengfan Gu. Random weighting estimation of stable exponent. Metrika, 77(4):451–468, May 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-013-
0448-6.

Diaz-Garcia:2014:RD

[1840] José A. Díaz-García. On Riesz distribution. Metrika, 77(4):469–481, May 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-013-
0449-5.

Edwards:2014:FET

[1841] David J. Edwards. Follow-up experiments for two-level fractional factorial designs via double semifoldover. Metrika, 77(4):483–507, May 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s00184-013-0450-
z.

Lacaux:2014:CPP

[1842] Céline Lacaux, Aurélie Muller-Gueudin, Radu Ranta, and Samy Tindel. Convergence and performance of the peeling wavelet denoising algorithm. Metrika, 77(4):509–537, May 2014.

CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s00184-013-0451-
y.

Rychlik:2014:NPU

[1843] Tomasz Rychlik. Non-positive upper bounds on expectations of small order statistics from DDA and DFRA populations. Metrika, 77(4):539–557, May 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00184-
013-0452-x.pdf.

Arellano-Valle:2014:SMU

[1844] R. B. Arellano-Valle, Ahad Jamalizadeh, H. Mahmoodian, and N. Balakrishnan. L-statistics from multivariate unified skew-elliptical distributions. Metrika, 77(4):559–583, May 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-013-
0453-9.

Anonymous:2014:HCd

[1845] Anonymous. Help & contacts. Metrika, 77(4):??, May 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Denecke:2014:NRT

[1846] Liesa Denecke and Christine H. Müller. New robust tests for the parameters of the Weibull distribution for complete and censored data. Metrika, 77(5):585–607, July 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-013-0454-8.

Park:2014:KLI

[1847] Sangun Park. On Kullback–Leibler information of order statistics in terms of the relative risk. Metrika, 77 (5):609–616, July 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-013-0455-7.

Jokiel-Rokita:2014:DST

[1848] Alicja Jokiel-Rokita and Ryszard Magiera. Distributions of stopping times in some sequential estimation procedures. Metrika, 77(5):617–634, July 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00184-
013-0456-6.pdf.

Wang:2014:AIC

[1849] Weizhen Wang and Zhongzhan Zhang.

Asymptotic infimum coverage probability for interval estimation of proportions. Metrika, 77(5):635–646, July 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-013-
0457-5.

Withers:2014:APE

[1850] Christopher S. Withers and Saralees Nadarajah. Asymptotic properties of M-estimators in linear and nonlinear multivariate regression models. Metrika, 77(5):647–673, July 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-013-
                                        0458-4.

Yu:2014:ELI

[1851] Wei Yu, Cuizhen Niu, and Wangli Xu. An empirical likelihood inference for the coefficient difference of a twosample linear model with missing response data. Metrika, 77(5):675–693, July 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-013-
                                        0459-3.

Wang:2014:NBL

[1852] HaiYing Wang, Nancy Flournoy, and Eloi Kpamegan. A new bounded log-linear regression model. Metrika, 77(5):695–720, July 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-013-0460-x.

Anonymous:2014:HCe

[1853] Anonymous. Help & contacts. Metrika, 77(5):??, July 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Guo:2014:SRC

[1854] Bing Guo, Qi Zhou, and Runchu Zhang. Some results on constructing general minimum lower order confounding 2^{n−m} designs for n ≤ 2^{n−m−2}. Metrika, 77(6):721–732, August 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-013-
                                        0461-9.

Ma:2014:SEM

[1855] Tiefeng Ma, Shuangzhe Liu, and S. Ejaz Ahmed. Shrinkage estimation for the mean of the inverse Gaussian population. Metrika, 77 (6):733–752, August 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-013-0462-8.

Xu:2014:ORD

[1856] Xiaojian Xu and Xiaoli Shang. Optimal and robust designs for trigonometric regression models. Metrika, 77 (6):753–769, August 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-013-0463-7.

Novoa-Munoz:2014:TBP

[1857] F. Novoa-Muñoz and M. D. Jiménez-Gamero. Testing for the bivariate Poisson distribution. Metrika, 77 (6):771–793, August 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-013-0464-6.

Niu:2014:TES

[1858] Cuizhen Niu, Xu Guo, Wangli Xu, and Lixing Zhu. Testing equality of shape parameters in several inverse Gaussian populations. Metrika, 77 (6):795–809, August 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-013-0465-5.

Zhao:2014:MHG

[1859] Peng Zhao and Yiying Zhang. On the maxima of heterogeneous gamma variables with different shape and scale parameters. Metrika, 77(6):811–836, August 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-013-
                                        0466-4.

Anonymous:2014:HCf

[1860] Anonymous. Help & contacts. Metrika, 77(6):??, August 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Zhang:2014:SOL

[1861] Chen Zhang and Alwell Julius Oyet. Second order longitudinal dynamic models with covariates: estimation and forecasting. Metrika, 77(7): 837–859, October 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-013-0467-3.

Vasudeva:2014:ABN

[1862] Rasbagh Vasudeva and J. Vasantha Kumari. Asymptotic behaviour of near-maxima of Gaussian sequences. Metrika, 77(7):861–866, October 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00184-013-0468-
                                        2.

Huang:2014:MMS

[1863] Chao Huang and Jin-Guan Lin. Modified maximum spacings method for generalized extreme value distribution and applications in real data analysis. Metrika, 77(7):867–894, October 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-013-
0469-1.

Inoue:2014:SLW

[1864] Kiyoshi Inoue and Sigeo Aki. On sooner and later waiting time distributions associated with simple patterns in a sequence of bivariate trials. Metrika, 77 (7):895–920, October 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-013-0478-0.

Guo:2014:ELH

[1865] Hong Guo, Changliang Zou, Zhaojun Wang, and Bin Chen. Empirical likelihood for high-dimensional linear regression models. Metrika, 77 (7):921–945, October 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-013-0479-z.

Thomas:2014:CBD

[1866] P. Yageen Thomas and T. G. Veena.

Characterizations of bivariate distributions using concomitants of record values. Metrika, 77(7):947–963, October 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-013-
0480-6.

Anonymous:2014:HCg

[1867] Anonymous. Help & contacts. Metrika, 77(7):??, October 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Burkschat:2014:ABH

[1868] M. Burkschat and J. Navarro. Asymptotic behavior of the hazard rate in systems based on sequential order statistics. Metrika, 77(8):965–994, November 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-013-
                                        0481-5.

Yu:2014:NSC

[1869] Qiqing Yu, Yuting Hsu, and Kai Yu. A necessary and sufficient condition for justifying non-parametric likelihood with censored data. Metrika, 77(8): 995–1011, November 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-014-0482-z.

Zhao:2014:MAB

[1870] Shangwei Zhao. Model averaging based on James–Stein estimators. Metrika, 77(8):1013–1022, November 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-014-
                                        0483-y.

Jokiel-Rokita:2014:BPD

[1871] Alicja Jokiel-Rokita, Daniel Lazar, and Ryszard Magiera. Bayesian prediction in doubly stochastic Poisson process. Metrika, 77(8):1023–1039, November 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.

com/content/pdf/10.1007/s00184-
014-0484-x.pdf.

You:2014:EBR

[1872] Yinping You, Xiaohu Li, and Narayanaswamy Balakrishnan. On extremes of bivariate residual lifetimes from generalized Marshall–Olkin and time transformed exponential models. Metrika, 77(8): 1041–1056, November 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-014-0485-9.

Georgiou:2014:UTC

[1873] Stelios D. Georgiou, Christos Koukouvinos, and Min-Qian Liu. U-type and column-orthogonal designs for computer experiments. Metrika, 77(8): 1057–1073, November 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-014-0486-8.

Anonymous:2014:HCh

[1874] Anonymous. Help & contacts. Metrika, 77(8):??, November 2014. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Torgovitski:2015:DET

[1875] Leonid Torgovitski. A Darling–Erdőstype CUSUM-procedure for functional data. Metrika, 78(1):1–27, January 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-014-
0487-7.

Arashi:2015:SEM

[1876] M. Arashi, B. M. Golam Kibria, and A. Tajadod. On shrinkage estimators in matrix variate elliptical models. Metrika, 78(1):29–44, January 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-014-
0488-6.

Efthimiou:2015:CNO

[1877] Ifigenia Efthimiou, Stelios D. Georgiou, and Min-Qian Liu. Construction of nearly orthogonal Latin hypercube designs. Metrika, 78(1):45–57, January 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-014-
0489-5.

Goldmann:2015:DTG

[1878] Christian Goldmann, Bernhard Klar, and Simos G. Meintanis. Data transformations and goodness-of-fit tests for Type-II right censored samples. Metrika, 78(1):59–83, January 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-014-
0490-z.

Feng:2015:RSB

[1879] Long Feng, Changliang Zou, Zhaojun Wang, Xianwu Wei, and Bin Chen. Robust spline-based variable selection in varying coefficient model. Metrika, 78(1):85–118, January 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-014-0491-y.

Koul:2015:MDL

[1880] Hira L. Koul, Donatas Surgailis, and Nao Mimoto. Minimum distance lackof-fit tests under long memory errors. Metrika, 78(2):119–143, February 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-014-
0492-x.

Hwang:2015:RTS

[1881] Leng-Cheng Hwang and Chia-Chen Yang. A robust two-stage procedure in Bayes sequential estimation of a particular exponential family. Metrika, 78 (2):145–159, February 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-014-0493-9.

Wilk:2015:OCD

[1882] Adrian Wilk and Joachim Kunert. Optimal crossover designs in a model with self and mixed carryover effects with correlated errors. Metrika, 78 (2):161–174, February 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-014-0494-8.

Goroncy:2015:OBE

[1883] Agnieszka Goroncy and Tomasz Rychlik. Optimal bounds on expectations of order statistics and spacings from nonparametric families of distributions generated by convex transform order. Metrika, 78(2):175–204, February 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-014-
                                        0495-7.

Dolegowski:2015:LRO

[1884] Adam Dolegowski and Jacek Wesolowski. Linearity of regression for overlapping order statistics. Metrika, 78(2): 205–218, February 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s00184-014-0496-6.pdf.

Morina:2015:CIF

[1885] D. Moriña, P. Puig, and J. Valero. A characterization of the innovations of first order autoregressive models. Metrika, 78(2):219–225, February 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-014-
                                        0497-5.

Zakerzadeh:2015:ANR

[1886] H. Zakerzadeh and S. H. Moradi Zahraie. Admissibility in non-regular family under squared-log error loss. Metrika, 78(2):227–236, February 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00184-014-0499-
                                        3.

Schmidt:2015:ODC

[1887] Dennis Schmidt and Rainer Schwabe. On optimal designs for censored data. Metrika, 78(3):237–257, April 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00184-014-

0500-1. See [1888].

Schmidt:2015:EOD

[1888] Dennis Schmidt and Rainer Schwabe.

Erratum to: On optimal designs for censored data. Metrika, 78(3):259, April 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00184-

014-0504-x.pdf. See erratum [1887].

Ashrafi:2015:SDP

[1889] S. Ashrafi and M. Asadi. On the stochastic and dependence properties of the three-state systems. Metrika, 78(3):261–281, April 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-014-0501-0.

Berred:2015:APN

[1890] Alexandre Berred and Alexei Stepanov.

Asymptotic properties of the number of near minimum-concomitant observations in the case of progressive type-II censoring. Metrika, 78(3):283–294, April 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-014-
0502-z.

Shen:2015:ART

[1891] Aiting Shen, Ying Zhang, and Andrei Volodin. Applications of the Rosenthal-type inequality for negatively superadditive dependent random variables. Metrika, 78(3):295–311, April 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-014-
0503-y.

Dudek:2015:CBB

[1892] A. E. Dudek. Circular block bootstrap for coefficients of autocovariance function of almost periodically correlated time series. Metrika, 78(3):313–335, April 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-014-
0505-9.

Carvalho:2015:ITS

[1893] Francisco Carvalho, João T. Mexia, Carla Santos, and Célia Nunes. Inference for types and structured families of commutative orthogonal block structures. Metrika, 78(3):337–372, April 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-014-
0506-8.

Lin:2015:CSO

[1894] Chang-Yun Lin. Construction and selection of the optimal balanced blocked definitive screening design. Metrika, 78(4):373–383, May 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-014-0507-7.

Balakrishnan:2015:LRC

[1895] N. Balakrishnan and Alexei Stepanov. Limit results for concomitants of order statistics. Metrika, 78(4):385–397, May 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.

com/article/10.1007/s00184-014-
0508-6.

Pellerey:2015:NRB

[1896] Franco Pellerey and Saeed Zalzadeh.

A note on relationships between some univariate stochastic orders and the corresponding joint stochastic orders. Metrika, 78(4):399–414, May 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-014-
0509-5.

Kundu:2015:GMI

[1897] Chanchal Kundu. Generalized measures of information for truncated random variables. Metrika, 78(4):415–435, May 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-014-
0510-z.

Bieniek:2015:OEB

[1898] Mariusz Bieniek. Optimal evaluations for the bias of trimmed means of k th record values. Metrika, 78(4):437–460, May 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00184-
014-0511-y.pdf.

Fourdrinier:2015:RMS

[1899] Dominique Fourdrinier and William Strawderman. Robust minimax Stein estimation under invariant data-based loss for spherically and elliptically symmetric distributions. Metrika, 78 (4):461–484, May 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-014-0512-x.

Lee:2015:TWS

[1900] Myoung jae Lee and Maria Karlsson. Trimmed and Winsorized semiparametric estimator for left-truncated and right-censored regression models. Metrika, 78(4):485–495, May 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-014-
0513-9.

DiBernardino:2015:ECF

[1901] Elena Di Bernardino, Thomas Laloë, and Rémi Servien. Estimating covariate functions associated to multivariate risks: a level set approach. Metrika, 78(5):497–526, July 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-014-0498-4. See erra-

tum [1902].

DiBernardino:2015:EEC

[1902] Elena Di Bernardino, Thomas Laloë, and Rémi Servien. Erratum to: Estimating covariate functions associated to multivariate risks: a level set approach. Metrika, 78(5):527, July 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-015-
0548-6;   http://link.springer.
com/content/pdf/10.1007/s00184-
015-0548-6.pdf. See [1901].

Yang:2015:BST

[1903] Po Yang, Chang-Yun Lin, and William Li. Blocked semifoldovers of twolevel orthogonal designs. Metrika, 78(5):529–548, July 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-014-0514-8.

Mohammadi:2015:ETI

[1904] Mohammad Mohammadi, Adel Mohammadpour, and Hiroaki Ogata. On estimating the tail index and the spectral measure of multivariate α-stable distributions. Metrika, 78(5):549–561, July 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-014-
0515-7.

Pena:2015:CMD

[1905] Edsel A. Peña, Joshua D. Habiger, and Wensong Wu. Classes of multiple decision functions strongly controlling FWER and FDR. Metrika, 78(5):563–595, July 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-014-0516-6.

Zuo:2015:BRB

[1906] Yijun Zuo. Bahadur representations for bootstrap quantiles. Metrika, 78 (5):597–610, July 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-014-0517-5.

Basu:2015:RTE

[1907] A. Basu, A. Mandal, N. Martin, and L. Pardo. Robust tests for the equality of two normal means based on the density power divergence. Metrika, 78(5):611–634, July 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-014-0518-4.

Tsou:2015:USL

[1908] Tsung-Shan Tsou and Chi-Chuan Yang. Universal surrogate likelihood functions for nonnegative continuous data. Metrika, 78(6):635–646, August 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-014-
0519-3.

Weerahandi:2015:IEB

[1909] Samaradasa Weerahandi and Malwane M. A. Ananda. Improving the EBLUPs of balanced mixed-effects models. Metrika, 78(6):647–662, August 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-014-
0520-x.

Weerahandi:2015:EIE

[1910] Samaradasa Weerahandi and Malwane M. A. Ananda. Erratum to: Improving the EBLUPs of balanced mixedeffects models. Metrika, 78(6):663, August 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-015-
0539-7;   http://link.springer.

com/content/pdf/10.1007/s00184-
015-0539-7.pdf.

Pestova:2015:TSC

[1911] Barbora Pestová and Michal Pesta.

Testing structural changes in panel data with small fixed panel size and bootstrap. Metrika, 78(6):665–689, August 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-014-

0522-8. See erratum [1940].

Navarro:2015:BDC

[1912] J. Navarro, M. Esna-Ashari, M. Asadi, and J. M. Sarabia. Bivariate distributions with conditionals satisfying the proportional generalized odds rate model. Metrika, 78(6):691–709, August 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-014-
0523-7.

Furer:2015:SSR

[1913] Dmytro Furer and Michael Kohler.

Smoothing spline regression estimation based on real and artificial data. Metrika, 78(6):711–746, August 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-014-
0524-6.

Chan:2015:ELI

[1914] Ping Shing Chan, Hon Keung Tony Ng, and Feng Su. Exact likelihood inference for the two-parameter exponential distribution under Type-II progressively hybrid censoring. Metrika, 78 (6):747–770, August 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-014-0525-5.

Al-Hussaini:2015:OSB

[1915] Essam K. Al-Hussaini, Alaa H. Abdel-Hamid, and Atef F. Hashem. Onesample Bayesian prediction intervals based on progressively type-II censored data from the half-logistic distribution under progressive stress model. Metrika, 78(7):771–783, October 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-014-
0526-4.

Pi:2015:FIC

[1916] Lira Pi and H. N. Nagaraja. Fisher information in censored samples from folded and unfolded populations. Metrika, 78(7):785–806, October 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s00184-014-0527-
3.

Cha:2015:DSS

[1917] Ji Hwan Cha and Maxim Finkelstein. A dynamic stress-strength model with stochastically decreasing strength. Metrika, 78(7):807–827, October 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-015-
0528-x.

Weiss:2015:PIM

[1918] Christian H. Weiß. A Poisson INAR(1)

model with serially dependent innovations. Metrika, 78(7):829–851, October 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-015-
0529-9.

Balakrishnan:2015:IBB

[1919] N. Balakrishnan and Xiaojun Zhu. Inference for the bivariate Birnbaum– Saunders lifetime regression model and associated inference. Metrika, 78(7): 853–872, October 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-015-0530-3.

Lin:2015:GVR

[1920] Jin-Guan Lin, Xue-Ping Chen, Jian-Feng Yang, Xing-Fang Huang, and Ying-Shan Zhang. Generalized variable resolution designs. Metrika, 78 (7):873–884, October 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-015-0531-2.

Ye:2015:TVC

[1921] Rendao Ye, Tonghui Wang, Saowanit Sukparungsee, and Arjun K. Gupta. Tests in variance components models under skew-normal settings. Metrika, 78(7):885–904, October 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-015-0532-1.

Tian:2015:NDB

[1922] Yongge Tian. A new derivation of BLUPs under random-effects model. Metrika, 78(8):905–918, November 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-015-
0533-0.

Ciuperca:2015:ELT

[1923] Gabriela Ciuperca and Zahraa Salloum. Empirical likelihood test in a posteriori change-point nonlinear model. Metrika, 78(8):919– 952, November 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-015-0534-z.

Zhao:2015:MLD

[1924] Qianqian Zhao and Shengli Zhao. Mixed-level designs with resolution III or IV containing clear two-factor interaction components. Metrika, 78(8): 953–965, November 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-015-0535-y.

Zou:2015:NWD

[1925] Yu-Ye Zou, Han-Ying Liang, and Jing-Jing Zhang. Nonlinear wavelet density estimation with data missing at random when covariates are present. Metrika, 78(8):967–995, November 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-015-
0536-x.

Roosta-Khorasani:2015:SPC

[1926] Farbod Roosta-Khorasani and Gábor J.

Székely. Schur properties of convolutions of gamma random variables. Metrika, 78(8):997–1014, November 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-015-
0537-9.

Amini-Seresht:2015:MSC

[1927] Ebrahim Amini-Seresht and Baha-Eldin Khaledi. Multivariate stochastic comparisons of mixture models. Metrika, 78(8):1015–1034, November 2015. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-015-
0538-8.

Gupta:2016:PAF

[1928] Ramesh C. Gupta. Properties of additive frailty model in survival analysis. Metrika, 79(1):1–17, January 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-015-
0540-1.

Balakrishnan:2016:GPD

[1929] Narayanaswamy Balakrishnan, Hong Qin, and Kashinath Chatterjee. Generalized projection discrepancy and its applications in experimental designs. Metrika, 79(1):19–35, January 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-015-
0541-0.

Kuhlisch:2016:SAC

[1930] Wiltrud Kuhlisch, Magnus Roos, Jörg Rothe, Joachim Rudolph, Björn Scheuermann, and Dietrich Stoyan. A statistical approach to calibrating the scores of biased reviewers of scientific papers. Metrika, 79(1):37–57, January 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-015-
                                        0542-z.

Ma:2016:NVS

[1931] Xue-Jun Ma and Jing-Xiao Zhang. A new variable selection approach for varying coefficient models. Metrika, 79(1):59–72, January 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-015-0543-y.

Colombi:2016:TOR

[1932] R. Colombi and A. Forcina. Testing order restrictions in contingency tables. Metrika, 79(1):73–90, January 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-015-
                                        0544-x.

Tripathi:2016:ESP

[1933] Yogesh Mani Tripathi, Somesh Kumar, and Constantinos Petropoulos. Estimating the shape parameter of a Pareto distribution under restrictions. Metrika, 79(1):91–111, January 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00184-015-
0545-9.

Zhang:2016:RIT

[1934] Zhengcheng Zhang and N. Balakrishnan. Representations of the inactivity time for coherent systems with heterogeneous components and some ordered properties. Metrika, 79(1): 113–126, January 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-015-0546-8.

Wang:2016:EPI

[1935] Xuejun Wang, Yi Wu, and Shuhe Hu.

Exponential probability inequality for m-END random variables and its applications. Metrika, 79(2):127–147, February 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-015-
0547-7.

Rusnacko:2016:CBU

[1936] Rastislav Rusnacko and Ivan Zezula.

Connection between uniform and serial correlation structure in the growth curve model. Metrika, 79(2): 149–164, February 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-015-0549-5.

Koul:2016:GFT

[1937] Hira L. Koul, Nao Mimoto, and Donatas Surgailis. A goodness-of-fit test for marginal distribution of linear random fields with long memory. Metrika, 79(2):165–193, February 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-015-
                                        0550-z.

Janssen:2016:IBN

[1938] Arnold Janssen and Andreas Knoch. Information bounds for nonparametric estimators of L-functionals and survival functionals under censored data. Metrika, 79(2):195–220, February 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-015-
                                        0551-y.

Milosevic:2016:AEN

[1939] Bojana Milosević. Asymptotic efficiency of new exponentiality tests based on a characterization. Metrika, 79(2):221–236, February 2016. CO-DEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00184-015-0552-
                                        x.

Pestova:2016:ETS

[1940] Barbora Pestová and Michal Pesta. Erratum to: Testing structural changes in panel data with small fixed panel size and bootstrap. Metrika, 79(2):237–238, February 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-015-
                                        0562-8;   http://link.springer.
                                        com/content/pdf/10.1007/s00184-
                                        015-0562-8.pdf. See [1911].

Dutta:2016:DFE

[1941] Santanu Dutta. Distribution function estimation via Bernstein polynomial of random degree. Metrika, 79(3):239–263, April 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-015-0553-9.

Wang:2016:NSC

[1942] Xinghui Wang and Shuhe Hu. A note on the strong consistency of Mestimates in linear models. Metrika, 79(3):265–266, April 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00184-015-0554-
8;   http://link.springer.com/article/
10.1007/s00184-015-0554-8.

Rabyk:2016:ECC

[1943] Liubov Rabyk and Wolfgang Schmid.

EWMA control charts for detecting changes in the mean of a long-memory process. Metrika, 79(3):267–301, April 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-015-
0555-7.

Avlogiaris:2016:LDB

[1944] G. Avlogiaris, A. Micheas, and K. Zografos. On local divergences between two probability measures. Metrika, 79(3):303–333, April 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-015-0556-6.

Kundu:2016:CRP

[1945] Chanchal Kundu, Antonio Di Crescenzo, and Maria Longobardi. On cumulative residual (past) inaccuracy for truncated random variables. Metrika, 79(3):335–356, April 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-015-0557-5.

Eryilmaz:2016:GWT

[1946] Serkan Eryilmaz. Generalized waiting time distributions associated with runs. Metrika, 79(3):357–368, April 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-015-
                                        0558-4.

Garcia-Perez:2016:MAS

[1947] A. García-Pérez. A von Mises approximation to the small sample distribution of the trimmed mean. Metrika, 79(4):369–388, May 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-015-0559-3.

Francq:2016:FTE

[1948] Christian Francq and Simos G. Meintanis. Fourier-type estimation of the power GARCH model with stable-Paretian innovations. Metrika, 79 (4):389–424, May 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-015-0560-x.

Prause:2016:MHD

[1949] Annabel Prause, Ansgar Steland, and Mohammed Abujarad. Minimum Hellinger distance estimation for bivariate samples and time series with applications to nonlinear regression and copula-based models. Metrika, 79(4):425–455, May 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-015-0561-9.

Lukusa:2016:SEZ

[1950] T. Martin Lukusa, Shen-Ming Lee, and Chin-Shang Li. Semiparametric estimation of a zero-inflated Poisson regression model with missing covariates. Metrika, 79(4):457–483, May 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-015-
0563-7.

Attaoui:2016:ARN

[1951] Said Attaoui and Nengxiang Ling.

Asymptotic results of a nonparametric conditional cumulative distribution estimator in the single functional index modeling for time series data with applications. Metrika, 79(5):485–511, July 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-015-
0564-6.

Kayid:2016:MMS

[1952] M. Kayid, S. Izadkhah, J. Jarrahiferiz, and P. Asghari. A mixture model of size-biased distributions. Metrika, 79(5):513–529, July 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-015-0565-5.

Balakrishnan:2016:SCP

[1953] Narayanaswamy Balakrishnan, Laurent Bordes, Christian Paroissin, and Jean-Christophe Turlot. Single changepoint detection methods for small lifetime samples. Metrika, 79(5):531–551, July 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-015-
                                        0566-4.

Li:2016:SSO

[1954] Chen Li, Rui Fang, and Xiaohu Li. Stochastic comparisons of order statistics from scaled and interdependent random variables. Metrika, 79(5):553–578, July 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-015-0567-3. See pub-

lisher correction [2038].

Makri:2016:ROD

[1955] Frosso S. Makri and Zaharias M. Psillakis. On runs of ones defined on a q-sequence of binary trials. Metrika, 79(5):579–602, July 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-015-0568-2.

Pulit:2016:NMK

[1956] Michal Pulit. A new method of kernelsmoothing estimation of the ROC curve. Metrika, 79(5):603–634, July 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/content/pdf/10.1007/s00184-
015-0569-1.pdf.

Goroncy:2016:EEO

[1957] Agnieszka Goroncy and Tomasz Rychlik. Evaluations of expectations of order statistics and spacings based on IFR distributions. Metrika, 79 (6):635–657, August 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00184-015-0570-8.pdf.

Xie:2016:VOM

[1958] Hongmei Xie, Keshe Ni, and Wenyu Liu. Variability ordering of multiplicative frailty models. Metrika, 79 (6):659–670, August 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00184-015-0571-
7.

Ahmadi:2016:RBP

[1959] Jafar Ahmadi, Elham Mirfarah, and Ahmad Parsian. Robust Bayesian Pitman closeness. Metrika, 79(6):671– 691, August 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s00184-015-0572-6.

Fang:2016:LRO

[1960] Longxiang Fang and N. Balakrishnan.

Likelihood ratio order of parallel systems with heterogeneous Weibull components. Metrika, 79(6):693–703, August 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00184-015-0573-5.

Kutoyants:2016:MSM

[1961] Yu. A. Kutoyants and A. Motrunich. On multi-step MLE-process for Markov sequences. Metrika, 79(6):705–724, August 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00184-015-0574-4.

Bayramoglu:2016:RMR

[1962] Ismihan Bayramoglu. On the records of multivariate random sequences. Metrika, 79(6):725–747, August 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00184-016-0575-y.

Brau:2016:MPD

[1963] Christian Bräu and Lothar Heinrich. Multivariate Poisson distributions associated with Boolean models. Metrika, 79(6):749–761, August 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00184-016-0576-
                                        x.

Che:2016:NJM

[1964] Xiaoyu Che and John Angus. A new joint model of recurrent event data with the additive hazards model for the terminal event time. Metrika, 79 (7):763–787, October 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00184-016-0577-
9.

Barczy:2016:SIC

[1965] Mátyás Barczy, Kristóf Körmendi, and Gyula Pap. Statistical inference for critical continuous state and continuous time branching processes with immigration. Metrika, 79(7):789–816, October 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00184-016-0578-8.

Amini-Seresht:2016:SOS

[1966] Ebrahim Amini-Seresht, Jianfei Qiao, Yiying Zhang, and Peng Zhao. On the skewness of order statistics in multipleoutlier PHR models. Metrika, 79 (7):817–836, October 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00184-016-0579-
7.

Castano-Martinez:2016:ER

[1967] A. Castaño-Martínez, F. López-Blázquez, and B. Salamanca-Miño. Exceedances of records. Metrika, 79 (7):837–866, October 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00184-016-0580-
1.

Ozturk:2016:QIB

[1968] Omer Ozturk and Asuman Turkmen.

Quantile inference based on clustered data. Metrika, 79(7):867–893, October 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00184-016-0581-0.

Strohriegl:2016:QRE

[1969] Katharina Strohriegl and Robert Hable. Qualitative robustness of estimators on stochastic processes. Metrika, 79(8):895–917, November 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00184-016-0582-z.

Dion:2016:NEM

[1970] Charlotte Dion. Nonparametric estimation in a mixed-effect Ornstein– Uhlenbeck model. Metrika, 79(8): 919–951, November 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00184-016-0583-
y.

Yu:2016:TLP

[1971] Ping Yu, Zhongzhan Zhang, and Jiang Du. A test of linearity in partial functional linear regression. Metrika, 79 (8):953–969, November 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00184-016-0584-
x.

Ducharme:2016:TBD

[1972] Gilles R. Ducharme and Walid Al Akhras. Tree based diagnostic procedures following a smooth test of goodness-of-fit. Metrika, 79(8):971– 989, November 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00184-016-0585-
9.

Zhao:2016:IBS

[1973] Peixin Zhao and Xinrong Tang. Imputation based statistical inference for partially linear quantile regression models with missing responses. Metrika, 79(8):991–1009, November 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00184-016-0586-8.

Sharafi:2016:AMS

[1974] M. Sharafi and A. R. Nematollahi.

AR(1) model with skew-normal innovations. Metrika, 79(8):1011–1029, November 2016. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http://link.

springer.com/accesspage/article/
10.1007/s00184-016-0587-7.

Wang:2017:CEL

[1975] Wu Wang and Zhongyi Zhu. Conditional empirical likelihood for quantile regression models. Metrika, 80 (1):1–16, January 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00184-016-0588-
6.

Li:2017:RFS

[1976] Xiang-Jie Li, Xue-Jun Ma, and Jing-Xiao Zhang. Robust feature screening for varying coefficient models via quantile partial correlation. Metrika, 80(1):17–49, January 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00184-016-0589-
                                        5.

Wu:2017:SEL

[1977] Yueqin Wu and Yan Sun. Shrinkage estimation of the linear model with spatial interaction. Metrika, 80 (1):51–68, January 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00184-016-0590-
                                        z.

Samadi:2017:RLC

[1978] P. Samadi, M. Rezaei, and M. Chahkandi. On the residual lifetime of coherent systems with heterogeneous components. Metrika, 80(1):69–82, January 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00184-016-0591-y.

Wang:2017:ASW

[1979] Xinghui Wang and Shuhe Hu. Asymptotics of self-weighted M-estimators for autoregressive models. Metrika, 80(1):83–92, January 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00184-016-0592-

Yan:2017:FIC

[1980] Liang Yan, Rui Wang, and Xingzhong Xu. Fiducial inference in the classical errors-in-variables model. Metrika, 80(1):93–114, January 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00184-016-0593-
9.

Reisen:2017:RDF

[1981] V. A. Reisen, C. Lévy-Leduc, M. Bourguignon, and H. Boistard. Robust Dickey–Fuller tests based on ranks for time series with additive outliers. Metrika, 80(1):115–131, January 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00184-016-0594-8.

Wang:2017:BFA

[1982] Dongying Wang, Shili Ye, Qi Zhou, and Runchu Zhang. Blocked factor aliased effect-number pattern and column rank of blocked regular designs. Metrika, 80 (2):133–152, February 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00184-016-0595-
7.

Nguyen:2017:NAD

[1983] Thuong T. M. Nguyen. A new approach to distribution free tests in contingency tables. Metrika, 80(2): 153–170, February 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/

                                        article/10.1007/s00184-016-0596-
                                        6.

Tzavelas:2017:MME

[1984] George Tzavelas, Maria Douli, and Polychronis Economou. Model misspecification effects for biased samples. Metrika, 80(2):171–185, February 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00184-016-0597-5.

Zhou:2017:OSH

[1985] Sanyu Zhou. One-sided hyperbolic simultaneous confidence bands for multiple and polynomial regression models. Metrika, 80(2):187–200, February 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00184-016-0598-4.

Harman:2017:BAC

[1986] Radoslav Harman and Eva Benková. Barycentric algorithm for computing D-optimal size- and cost-constrained designs of experiments. Metrika, 80 (2):201–225, February 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00184-016-0599-
                                        3.

Goli:2017:SCI

[1987] S. Goli and M. Asadi. A study on the conditional inactivity time of coherent systems. Metrika, 80(2): 227–241, February 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/accesspage/
article/10.1007/s00184-016-0600-
1.

Hrabakova:2017:NCS

[1988] Jitka Hrabáková and Václav Kus.

Notes on consistency of some minimum distance estimators with simulation results. Metrika, 80(2):243–257, February 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00184-016-0601-0.

Eryilmaz:2017:CWE

[1989] Serkan Eryilmaz. The concept of weak exchangeability and its applications. Metrika, 80(3):259–271, April 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00184-016-0602-z.

Ghnimi:2017:RPE

[1990] Soumaya Ghnimi, Soufiane Gasmi, and Arwa Nasr. Reliability parameters estimation for parallel systems under imperfect repair. Metrika, 80(3):273–288, April 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00184-016-0603-y.

Huang:2017:ASS

[1991] Hengzhen Huang, Shuangshuang Zhou, Min-Qian Liu, and Zong-Feng Qi. Acceleration of the stochastic search variable selection via componentwise Gibbs sampling. Metrika, 80(3):289–308, April 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00184-016-0604-x.

Dey:2017:EPC

[1992] Aloke Dey, Rakhi Singh, and Ashish Das. Efficient paired choice designs with fewer choice pairs. Metrika, 80 (3):309–317, April 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/accesspage/
                                        article/10.1007/s00184-016-0605-
                                        9.

Dembinska:2017:ETP

[1993] Anna Dembińska. An ergodic theorem for proportions of observations that fall into random sets determined by sample quantiles. Metrika, 80 (3):319–332, April 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/content/pdf/
                                        10.1007/s00184-016-0606-8.pdf.

Chacko:2017:BEB

[1994] Manoj Chacko. Bayesian estimation based on ranked set sample from Morgenstern type bivariate exponential distribution when ranking is imperfect. Metrika, 80(3):333–349, April 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/accesspage/article/10.1007/
                                        s00184-016-0607-7.

Unkel:2017:SCR

[1995] Steffen Unkel. On the shape of the cross-ratio function in bivariate survival models induced by truncated and folded normal frailty distributions. Metrika, 80(3):351–362, April 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00184-016-0608-6.

Heinrich:2017:VDD

[1996] Lothar Heinrich, Friedrich Pukelsheim, and Vitali Wachtel. The variance of the discrepancy distribution of rounding procedures, and sums of uniform random variables. Metrika, 80(3):363–375, April 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/accesspage/article/10.1007/
s00184-017-0609-0.

Angelov:2017:NES

[1997] Angel G. Angelov and Magnus Ekström. Nonparametric estimation for self-selected interval data collected through a two-stage approach. Metrika, 80(4):377–399, May 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/content/pdf/
10.1007/s00184-017-0610-7.pdf.

Koley:2017:GPH

[1998] Arnab Koley and Debasis Kundu. On generalized progressive hybrid censoring in presence of competing risks. Metrika, 80(4):401–426, May 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Saulo:2017:EGB

[1999] Helton Saulo, N. Balakrishnan, Xiaojun Zhu, Jhon F. B. Gonzales, and Jeremias Leão. Estimation in generalized bivariate Birnbaum–Saunders models. Metrika, 80(4):427–453, May 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Badia:2017:BPR

[2000] F. G. Badía and Ji Hwan Cha. On bending (down and up) property of reliability measures in mixtures. Metrika, 80(4):455–482, May 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Petropoulos:2017:EOR

[2001] Constantinos Petropoulos. Estimation of the order restricted scale parameters for two populations from the Lomax distribution. Metrika, 80(4):483–502, May 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Jimenez-Gamero:2017:MDE

[2002] M. D. Jiménez-Gamero and A. Batsidis. Minimum distance estimators for count data based on the probability generating function with applications. Metrika, 80(5):503–545, July 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Du:2017:FIC

[2003] Jiang Du, Zhongzhan Zhang, and Tianfa Xie. Focused information criterion and model averaging in censored quantile regression. Metrika, 80(5):547–570, July 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Weiss:2017:TCS

[2004] Christian H. Weiß, Esmeralda Gonçalves, and Nazaré Mendes Lopes. Testing the compounding structure of the CP– INARCH model. Metrika, 80(5):571– 603, July 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Shen:2017:WSL

[2005] Aiting Shen and Andrei Volodin. Weak and strong laws of large numbers for arrays of rowwise END random variables and their applications. Metrika, 80(6– 8):605–625, November 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-017-0618-z.

Navarro:2017:SCD

[2006] Jorge Navarro and Yolanda del Águila.

Stochastic comparisons of distorted distributions, coherent systems and mixtures with ordered components. Metrika, 80(6–8):627–648, November 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-017-
0619-y.

Asadi:2017:NMA

[2007] Majid Asadi. A new measure of association between random variables. Metrika, 80(6–8):649–661, November 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-017-
0620-5.

Joutard:2017:MSL

[2008] Cyrille Joutard. Multidimensional strong large deviation results. Metrika, 80(6–8):663–683, November 2017.

CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s00184-017-0621-
4.

Bohning:2017:SGP

[2009] Dankmar Böhning, Rattana Lerdsuwansri, and Heinz Holling. Some general points on the I^{2}-measure of heterogeneity in meta-analysis. Metrika, 80(6–8):685–695, November 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
content/pdf/10.1007/s00184-017-
0622-3.pdf;   http://link.springer.
com/article/10.1007/s00184-017-
0622-3.

Li:2017:EMA

[2010] Zaixing Li, Fei Chen, and Lixing Zhu. Estimating moments in ANOVAtype mixed models. Metrika, 80(6– 8):697–715, November 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-017-0623-2.

He:2017:ODM

[2011] Lei He and Rong-Xian Yue. R-optimal designs for multi-factor models with heteroscedastic errors. Metrika, 80(6– 8):717–732, November 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-017-0624-1.

Gatto:2017:MST

[2012] R. Gatto. Multivariate saddlepoint tests on the mean direction of the von Mises–Fisher distribution.

Metrika, 80(6–8):733–747, November 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-017-
0625-0.

Mesfioui:2017:SCO

[2013] M. Mesfioui, M. Kayid, and S. Izadkhah. Stochastic comparisons of order statistics from heterogeneous random variables with Archimedean copula. Metrika, 80(6–8):749–766, November 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-017-
0626-z.

Hwang:2017:SBC

[2014] Eunju Hwang and Dong Wan Shin.

Stationary bootstrapping for common mean change detection in crosssectionally dependent panels. Metrika, 80(6–8):767–787, November 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s00184-017-0627-
y.

Filipiak:2017:UOD

[2015] Katarzyna Filipiak and Augustyn Markiewicz. Universally optimal designs under mixed interference models with and without block effects. Metrika, 80(6–8):789–804, November 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
content/pdf/10.1007/s00184-017-
0628-x.pdf;   http://link.springer.
com/article/10.1007/s00184-017-
0628-x.

Ni:2017:APC

[2016] Lyu Ni, Fang Fang, and Fangjiao Wan. Adjusted Pearson chi-square feature screening for multi-classification with ultrahigh dimensional data. Metrika, 80(6–8):805–828, November 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-017-
                                        0629-9.

Tafiadi:2017:EID

[2017] Maria Tafiadi and George Iliopoulos. Exact inference for the difference of Laplace location parameters. Metrika, 80(6–8):829–861, November 2017. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-017-
                                        0630-3.

Lu:2018:SBQ

[2018] Minggen Lu. Spline-based quasilikelihood estimation of mixed Poisson regression with single-index models. Metrika, 81(1):1–17, January 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-017-
                                        0631-2.

Fang:2018:OPS

[2019] Longxiang Fang and N. Balakrishnan. Ordering properties of the smallest order statistics from generalized Birnbaum–Saunders models with associated random shocks. Metrika, 81 (1):19–35, January 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-017-0632-1.

Psarrakos:2018:RIM

[2020] Georgios Psarrakos and Antonio Di Crescenzo. A residual inaccuracy measure based on the relevation transform. Metrika, 81(1):37–59, January 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-017-
0633-0.

Chowdhury:2018:NEC

[2021] Mohammed Chowdhury, Colin Wu, and Reza Modarres. Nonparametric estimation of conditional distribution functions with longitudinal data and time-varying parametric models. Metrika, 81(1):61–83, January 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-017-
0634-z.

Isogai:2018:TSC

[2022] Eiichi Isogai and Chikara Uno. Threestage confidence intervals for a linear combination of locations of two negative exponential distributions. Metrika, 81(1):85–103, January 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-017-0635-y.

Hazra:2018:SCM

[2023] Nil Kamal Hazra, Mithu Rani Kuiti, Maxim Finkelstein, and Asok K. Nanda. On stochastic comparisons of minimum order statistics from the location-scale family of distributions. Metrika, 81(2):105–123, February 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-017-
0636-x.

Jasinski:2018:RPM

[2024] Krzysztof Jasiński. Relations for product moments and covariances of k-th records from discrete distributions. Metrika, 81(2):125–141, February 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-017-
0637-9;   http://link.springer.
com/content/pdf/10.1007/s00184-
017-0637-9.pdf.

Pal:2018:EAD

[2025] Suvra Pal, Jacob Majakwara, and N. Balakrishnan. An EM algorithm for the destructive COM–Poisson regression cure rate model. Metrika, 81 (2):143–171, February 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-017-0638-8.

Gorny:2018:MHC

[2026] Julian Górny and Erhard Cramer. Modularization of hybrid censoring schemes and its application to unified progressive hybrid censoring. Metrika, 81(2):173–210, February 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s00184-017-0639-
7.

Liu:2018:EIL

[2027] Kai Liu, Xiaojun Zhu, and N. Balakrishnan. Exact inference for Laplace distribution under progressive Type–II censoring based on BLUEs. Metrika, 81 (2):211–227, February 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-017-0640-1.

Raqab:2018:ITP

[2028] Mohammad Z. Raqab, Omar M. Bdair, and Fahad M. Al-Aboud. Inference for the two-parameter bathtubshaped distribution based on record data. Metrika, 81(3):229–253, April 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-017-
0641-0.

Fang:2018:ELH

[2029] Jianglin Fang, Wanrong Liu, and Xuewen Lu. Empirical likelihood for heteroscedastic partially linear singleindex models with growing dimensional data. Metrika, 81(3):255–281, April 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-018-
0642-7.

Zaigraev:2018:OCO

[2030] Alexander Zaigraev and Magdalena Alama-Bućko. Optimal choice of order statistics under confidence region estimation in case of large samples. Metrika, 81(3):283–305, April 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-018-
                                        0643-6.

Elsawah:2018:NRQ

[2031] A. M. Elsawah and Kai-Tai Fang. New results on quaternary codes and their Gray map images for constructing uniform designs. Metrika, 81(3):307–336, April 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-018-
                                        0644-5.

Chatterjee:2018:MLM

[2032] K. Chatterjee, K. Drosou, S. D. Georgiou, and C. Koukouvinos. Multi-level and mixed-level k-circulant supersaturated designs. Metrika, 81(3):337–355, April 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-018-
                                        0645-4.

Hu:2018:CFS

[2033] Qinqin Hu and Lu Lin. Conditional feature screening for mean and variance functions in models with multiple-index structure. Metrika, 81(4):357–393, May 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-018-0646-3.

Jia:2018:GFT

[2034] Weijia Jia and Weixing Song. Goodnessof-fit tests in linear EV regression with replications. Metrika, 81(4):395–421, May 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00184-018-
0648-1.

Beh:2018:AGS

[2035] Eric J. Beh and Rosaria Lombardo. An algebraic generalisation of some variants of simple correspondence analysis. Metrika, 81(4):423–443, May 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-018-
0649-0.

Li:2018:PIC

[2036] Chen Li and Xiaohu Li. Preservation of increasing convex/concave order under the formation of parallel/series system of dependent components. Metrika, 81(4):445–464, May 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0651-6.

Navarro:2018:SCC

[2037] Jorge Navarro. Stochastic comparisons of coherent systems. Metrika, 81(4):465–482, May 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0650-7.

Li:2018:PCS

[2038] Chen Li, Rui Fang, and Xiaohu Li. Publisher correction: Stochastic comparisons of order statistics from scaled and interdependent random variables. Metrika, 81(4):483, May 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-018-

                                        0647-2;   http://link.springer.
                                        com/content/pdf/10.1007/s00184-
                                        018-0647-2.pdf. See [1954].

Ghosh:2018:CCB

[2039] Indranil Ghosh. A complete characterization of bivariate densities using the conditional percentile function. Metrika, 81(5):485–492, July 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-018-
                                        0652-5.

Basu:2018:RWT

[2040] Ayanendranath Basu, Abhik Ghosh, Nirian Martin, and Leandro Pardo. Robust Wald-type tests for nonhomogeneous observations based on the minimum density power divergence estimator. Metrika, 81(5):493–522, July 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-018-
                                        0653-4.

Han:2018:JAR

[2041] Miao Han, Liuquan Sun, Yutao Liu, and Jun Zhu. Joint analysis of recurrent event data with additivemultiplicative hazards model for the terminal event time. Metrika, 81 (5):523–547, July 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-018-0654-3.

Liu:2018:LRC

[2042] Yang Liu, Yukun Liu, Yan Fan, and Han Geng. Likelihood ratio confidence interval for the abundance under binomial detectability models. Metrika, 81(5):549–568, July 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0655-2.

Hossain:2018:SEL

[2043] Shakhawat Hossain, Trevor Thomson, and Ejaz Ahmed. Shrinkage estimation in linear mixed models for longitudinal data. Metrika, 81(5):569–586, July 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-018-
0656-1.

Henze:2018:SIP

[2044] N. Henze, C. Kirch, and S. G. Meintanis. Special issue with papers from the “3rd workshop on Goodness-of-fit and change-point problems”. Metrika, 81(6):587–588, August 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0677-9;   http:
//link.springer.com/content/pdf/
10.1007/s00184-018-0677-9.pdf.

Messer:2018:MFT

[2045] Michael Messer, Stefan Albert, and Gaby Schneider. The multiple filter test for change point detection in time series. Metrika, 81(6):589–607, August 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-018-
0672-1.

Nikitin:2018:LEB

[2046] Ya. Yu. Nikitin. Local exact Bahadur efficiencies of two scale-free tests of normality based on a recent characterization. Metrika, 81(6):609–618, August 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-018-
                                        0673-0.

Weiss:2018:GFT

[2047] Christian H. Weiß. Goodness-of-fit testing of a count time series’ marginal distribution. Metrika, 81 (6):619–651, August 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-018-0674-z.

Duan:2018:RBM

[2048] Fang Duan and Dominik Wied. A residual-based multivariate constant correlation test. Metrika, 81(6): 653–687, August 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-018-0675-y.

Ciuperca:2018:TAL

[2049] Gabriela Ciuperca. Test by adaptive LASSO quantile method for real-time detection of a change-point. Metrika, 81(6):689–720, August 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-018-0676-x.

Jaruskova:2018:ENS

[2050] Daniela Jarusková. Estimating nonsimultaneous changes in the mean of vectors. Metrika, 81(6):721–743, August 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-018-
0671-2.

Kaino:2018:HES

[2051] Yusuke Kaino and Masayuki Uchida.

Hybrid estimators for small diffusion processes based on reduced data. Metrika, 81(7):745–773, October 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL http://link.springer.com/
article/10.1007/s00184-018-0657-
0.

Krishnamoorthy:2018:HPM

[2052] K. Krishnamoorthy and Shanshan Lv.

Highest posterior mass prediction intervals for binomial and Poisson distributions. Metrika, 81(7):775–796, October 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-018-
0658-z.

Wang:2018:CWL

[2053] Xuejun Wang, Xin Deng, and Shuhe Hu. On consistency of the weighted least squares estimators in a semiparametric regression model. Metrika, 81 (7):797–820, October 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0659-y.

Pan:2018:NIF

[2054] Jing Pan, Yuan Yu, and Yong Zhou. Nonparametric independence feature screening for ultrahigh-dimensional survival data. Metrika, 81(7): 821–847, October 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0660-5.

Worms:2018:EVS

[2055] Julien Worms and Rym Worms. Extreme value statistics for censored data with heavy tails under competing risks. Metrika, 81(7):849–889, October 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-018-
0662-3.

Baringhaus:2018:ECW

[2056] Ludwig Baringhaus and Daniel Gaigall. Efficiency comparison of the Wilcoxon tests in paired and independent survey samples. Metrika, 81(8):891– 930, November 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0661-4.

Sun:2018:STS

[2057] Ruili Sun, Tiefeng Ma, and Shuangzhe Liu. A Stein-type shrinkage estimator of the covariance matrix for portfolio selections. Metrika, 81(8): 931–952, November 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0663-2.

Delattre:2018:AML

[2058] Maud Delattre, Valentine Genon-Catalot, and Catherine Larédo. Approximate maximum likelihood estimation for stochastic differential equations with random effects in the drift and the diffusion. Metrika, 81(8): 953–983, November 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0666-z.

Weems:2018:ARE

[2059] Kimberly S. Weems and Paul J.

Smith. Assessing the robustness of estimators when fitting Poisson inverse Gaussian models. Metrika, 81(8): 985–1004, November 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0664-1.

Bapat:2018:PSE

[2060] Sudeep R. Bapat. On purely sequential estimation of an inverse Gaussian mean. Metrika, 81(8):1005– 1024, November 2018. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0665-0.

Mori:2019:FSA

[2061] Tamás F. Móri and Gábor J. Székely.

Four simple axioms of dependence measures. Metrika, 82(1):1–16, January 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-018-
0670-3.

Krishnan:2019:QBR

[2062] Aswathy S. Krishnan, S. M. Sunoj, and P. G. Sankaran. Quantile-based reliability aspects of cumulative Tsallis entropy in past lifetime. Metrika, 82(1):17–38, January 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-018-0678-8.

Ling:2019:SFP

[2063] Nengxiang Ling, Rui Kan, Philippe Vieu, and Shuyu Meng. Semifunctional partially linear regression model with responses missing at random. Metrika, 82(1):39–70, January 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-018-
                                        0688-6.

Sanchez-Borrego:2019:KBM

[2064] I. Sánchez-Borrego, A. Arcos, and M. Rueda. Kernel-based methods for combining information of several frame surveys. Metrika, 82(1):71–86, January 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-018-
                                        0686-8.

Liu:2019:ODM

[2065] Xin Liu, Rong-Xian Yue, and Weng Kee Wong. D-optimal designs for multi-response linear mixed models. Metrika, 82(1):87–98, January 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00184-018-
0679-7.

Arriaza:2019:SMB

[2066] A. Arriaza, A. Di Crescenzo, M. A.

Sordo, and A. Suárez-Llorens. Shape measures based on the convex transform order. Metrika, 82(1):99– 124, January 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0667-y.

Zhu:2019:TLH

[2067] Shuzhi Zhu and Peixin Zhao. Tests for the linear hypothesis in semifunctional partial linear regression models. Metrika, 82(2):125–148, March 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-018-
0680-1.

Anan:2019:TEC

[2068] Orasa Anan, Dankmar Böhning, and Antonello Maruotti. On the Turing estimator in capture–recapture count data under the geometric distribution. Metrika, 82(2):149–172, March 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-018-
0695-7.

Lu:2019:EEV

[2069] Shanshan Lu, Jingjing Wu, and Xuewen Lu. Efficient estimation of the varying-coefficient partially linear proportional odds model with current status data. Metrika, 82(2):173–194, March 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-018-
                                        0698-4.

Alnosaier:2019:IWK

[2070] Waseem Alnosaier and David Birkes. Inner workings of the Kenward–Roger test. Metrika, 82(2):195–223, March 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-018-
                                        0669-9.

Pakdaman:2019:STS

[2071] Zohreh Pakdaman and Jafar Ahmadi. Switching time of the standby component to the k-out-of-n:G system in the stress-strength setup. Metrika, 82 (2):225–248, March 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-018-0694-8.

Kumar:2019:GDR

[2072] A. N. Kumar and N. S. Upadhye. Generalizations of distributions related to (k_{1}, k_{2})-runs. Metrika, 82(2):249–268, March 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-018-
                                        0668-x.

Ye:2019:PAE

[2073] Shili Ye, Dongying Wang, and Runchu Zhang. Partial aliased effect number pattern and selection of optimal compromise designs. Metrika, 82(3):269–293, April 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-00705-2.

Mostafa:2019:KDE

[2074] Sayed A. Mostafa and Ibrahim A. Ahmad. Kernel density estimation from complex surveys in the presence of complete auxiliary information. Metrika, 82(3):295–338, April 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0703-y.

Sheena:2019:ECD

[2075] Yo Sheena. Estimation of a continuous distribution on the real line by discretization methods. Metrika, 82(3):339–360, April 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0683-y.

Bohning:2019:MCL

[2076] Dankmar Böhning, Panicha Kaskasamkul, and Peter G. M. van der Heijden. A modification of Chao’s lower bound estimator in the case of oneinflation. Metrika, 82(3):361–384, April 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-018-
0689-5;   http://link.springer.
com/content/pdf/10.1007/s00184-
018-0689-5.pdf.

Nikolov:2019:APL

[2077] Nikolay I. Nikolov and Eugenia Stoimenova. Asymptotic properties of Lee distance. Metrika, 82(3):385–408, April 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-018-

0687-7. See correction [2172].

Kharrati-Kopaei:2019:EDL

[2078] Mahmood Kharrati-Kopaei and Ahad Malekzadeh. On the exact distribution of the likelihood ratio test for testing the homogeneity of scale parameters of several two-parameter exponential distributions: complete and censored samples. Metrika, 82(4):409–427, May 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-018-
00704-3.

Tzavelas:2019:CPD

[2079] George Tzavelas. A characterization of the Pareto distribution based on the Fisher information for censored data under non-regularity conditions. Metrika, 82(4):429–440, May 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-018-
0697-5.

Sankaran:2019:RPP

[2080] P. G. Sankaran and M. Dileep Kumar. Reliability properties of proportional hazards relevation transform. Metrika, 82(4):441–456, May 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0681-0.

Kim:2019:THH

[2081] Jeonghwan Kim and Woojoo Lee.

On testing the hidden heterogeneity in negative binomial regression models. Metrika, 82(4):457–470, May 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-018-
0684-x.

Zhang:2019:AMG

[2082] Hongbin Zhang and Lang Wu. An approximate method for generalized linear and nonlinear mixed effects models with a mechanistic nonlinear covariate measurement error model. Metrika, 82(4):471–499, May 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0690-z.

Negarestani:2019:MMN

[2083] Hossein Negarestani, Ahad Jamalizadeh, Sobhan Shafiei, and Narayanaswamy Balakrishnan. Mean mixtures of normal distributions: properties, inference and application. Metrika, 82 (4):501–528, May 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0692-x.

Kazemi:2019:IAS

[2084] Mohammad Reza Kazemi and Ali Akbar Jafari. Inference about the shape parameters of several inverse Gaussian distributions: testing equality and confidence interval for a common value. Metrika, 82(5):529–545, July 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-018-
                                        0693-9.

Moran-Vasquez:2019:BCE

[2085] Raúl Alejandro Morán-Vásquez and Silvia L. P. Ferrari. Box–Cox elliptical distributions with application. Metrika, 82(5):547–571, July 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-018-0682-z.

Hu:2019:COF

[2086] Liuping Hu, Hongyi Li, and Zujun Ou. Constructing optimal four-level designs via Gray map code. Metrika, 82(5):573–587, July 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-018-0685-9.

Salehi:2019:RMW

[2087] M. Salehi, Z. Shishebor, and M. Asadi. On the reliability modeling of weighted k-out-of-n systems with randomly chosen components. Metrika, 82 (5):589–605, July 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-018-0691-y.

Crescenzo:2019:PIM

[2088] Antonio Di Crescenzo, Suchandan Kayal, and Abdolsaeed Toomaj. A past inaccuracy measure based on the reversed relevation transform. Metrika, 82(5):607–631, July 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0696-6.

Yu:2019:CQE

[2089] Ping Yu, Ting Li, Zhongyi Zhu, and Zhongzhan Zhang. Composite quantile estimation in partial functional linear regression model with dependent errors. Metrika, 82(6):633–656, August 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-018-
0699-3.

Mukherjee:2019:CPM

[2090] Amitava Mukherjee and Marco Marozzi.

A class of percentile modified Lepagetype tests. Metrika, 82(6):657–689, August 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-018-
0700-1.

Patra:2019:GOA

[2091] Arijit Patra and Chanchal Kundu.

On generalized orderings and ageing classes for residual life and inactivity time at random time. Metrika, 82 (6):691–704, August 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0701-0.

Yang:2019:SPF

[2092] Po Yang and William Li. Some properties of foldover designs with column permutations. Metrika, 82 (6):705–717, August 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-018-0702-z.

Rosa:2019:EWP

[2093] Samuel Rosa. Equivalence of weighted and partial optimality of experimental designs. Metrika, 82(6):719–732, August 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-019-
00706-9.

Li:2019:LLR

[2094] Xu Li, Juxia Xiao, Weixing Song, and Jianhong Shi. Local linear regression with reciprocal inverse Gaussian kernel. Metrika, 82(6):733–758, August 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-019-
00717-6.

Zhu:2019:EUM

[2095] Xiaojun Zhu, N. Balakrishnan, and Helton Saulo. On the existence and uniqueness of the maximum likelihood estimates of parameters of Laplace Birnbaum–Saunders distribution based on Type-I, Type-II and hybrid censored samples. Metrika, 82(7):759–778, October 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-019-
00707-8.

Betsch:2019:NCG

[2096] Steffen Betsch and Bruno Ebner. A new characterization of the Gamma distribution and associated goodnessof-fit tests. Metrika, 82(7):779–806, October 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-019-
00708-7.

Forcina:2019:ETM

[2097] Antonoio Forcina. Estimation and testing of multiplicative models for frequency data. Metrika, 82(7): 807–822, October 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-019-00709-6.

Bouezmarni:2019:LSB

[2098] Taoufik Bouezmarni, Félix Camirand Lemyre, and Jean-François Quessy. On the large-sample behavior of two estimators of the conditional copula under serially dependent data. Metrika, 82 (7):823–841, October 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-019-00711-y.

Lin:2019:MEC

[2099] Yuxuan Lin and Kai-Tai Fang. The main effect confounding pattern for saturated orthogonal designs. Metrika, 82 (7):843–861, October 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-019-00713-w.

Yang:2019:MOU

[2100] Kai Yang, Yao Kang, Dehui Wang, Han Li, and Yajing Diao. Modeling overdispersed or underdispersed count data with generalized Poisson integer-valued autoregressive processes. Metrika, 82 (7):863–889, October 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-019-00714-9.

Letac:2019:CDF

[2101] Gérard Letac. Cumulative distribution functions for the five simplest natural exponential families. Metrika, 82 (8):891–902, November 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-019-00710-z.

Nishino:2019:CST

[2102] Takuya Nishino and Hidetoshi Murakami. The Cucconi statistic for Type-I censored data. Metrika, 82(8): 903–929, November 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-019-00712-x.

Scherer:2019:ESM

[2103] Matthias Scherer and Henrik Sloot. Exogenous shock models: analytical characterization and probabilistic construction. Metrika, 82(8):931–959, November 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-019-
00715-8.

Kroll:2019:NIE

[2104] Martin Kroll. Nonparametric intensity estimation from noisy observations of a Poisson process under unknown error distribution. Metrika, 82 (8):961–990, November 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-019-00716-7.

Balakrishnan:2019:REO

[2105] N. Balakrishnan, E. Castilla, N. Martín, and L. Pardo. Robust estimators for one-shot device testing data under gamma lifetime model with an application to a tumor toxicological data. Metrika, 82(8):991–1019, November 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-019-
00718-5.

Mi:2019:SLR

[2106] Jie Mi. Some limit results in estimation of proportion based on group testing. Metrika, 82(8):1021–1038, November 2019. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-019-
00719-4.

Yang:2020:CQR

[2107] Weiming Yang and Yiping Yang. Composite quantile regression estimation of linear error-in-variable models using instrumental variables. Metrika, 83(1):1–16, January 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-019-00734-5.

Ma:2020:NEC

[2108] Yaolan Ma, Bo Wei, and Wei Huang.

A nonparametric estimator for the conditional tail index of Pareto-type distributions. Metrika, 83(1):17–44, January 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-019-
                                        00723-8.

Coelho:2020:THD

[2109] Carlos A. Coelho and Anuradha Roy. Testing the hypothesis of a doubly exchangeable covariance matrix. Metrika, 83(1):45–68, January 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-019-00724-7.

Garcia-Perez:2020:SAD

[2110] A. García-Pérez. Saddlepoint approximations for the distribution of some robust estimators of the variogram. Metrika, 83(1):69–91, January 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-019-
                                        00725-6.

Comte:2020:RFE

[2111] F. Comte and V. Genon-Catalot. Regression function estimation on non compact support in an heteroscesdastic model. Metrika, 83(1):93–128, January 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-019-
                                        00727-4.

Hu:2020:CFL

[2112] Liuping Hu, Zujun Ou, and Hongyi Li. Construction of four-level and mixedlevel designs with zero Lee discrepancy. Metrika, 83(1):129–139, January 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-019-
00720-x.

Kohler:2020:NQE

[2113] Michael Kohler and Reinhard Tent.

Nonparametric quantile estimation using surrogate models and importance sampling. Metrika, 83(2): 141–169, February 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-019-00736-3.

Misra:2020:RAF

[2114] Neeraj Misra and Jisha Francis. Relative ageing in frailty and resilience models. Metrika, 83(2):171–196, February 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-019-
00726-5.

Cao:2020:BFP

[2115] Cong Cao, Markus Pauly, and Frank Konietschke. The Behrens–Fisher problem with covariates and baseline adjustments. Metrika, 83(2): 197–215, February 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-019-00729-2.

Xu:2020:GSD

[2116] Xinyi Xu and Jingxiao Zhang. Groupwise sufficient dimension reduction via conditional distance clustering. Metrika, 83(2):217–242, February 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-019-
00732-7.

Lindholm:2020:VPE

[2117] Mathias Lindholm and Felix Wahl. On the variance parameter estimator in general linear models. Metrika, 83 (2):243–254, February 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-019-00751-4;   http:
//link.springer.com/content/pdf/
10.1007/s00184-019-00751-4.pdf.

Huang:2020:ODS

[2118] Yimin Huang, Xiangshun Kong, and Mingyao Ai. Optimal designs in sparse linear models. Metrika, 83 (2):255–273, February 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-019-00722-9.

Bagdonavicius:2020:NMO

[2119] Vilijandas Bagdonavicius and Linas Petkevicius. A new multiple outliers identification method in linear regression. Metrika, 83(3):275–296, April 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-019-
00731-8.

Batsidis:2020:GFT

[2120] Apostolos Batsidis, María Dolores Jiménez-Gamero, and Artur J. Lemonte. On goodness-of-fit tests for the Bell distribution. Metrika, 83(3):297–319, April 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-019-
00733-6.

Karimnezhad:2020:GGB

[2121] Ali Karimnezhad and Mahmoud Zarepour. A general guide in Bayesian and robust Bayesian estimation using Dirichlet processes. Metrika, 83 (3):321–346, April 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-019-00737-2.

Dash:2020:CNO

[2122] Sukanta Dash, Baidya Nath Mandal, and Rajender Parsad. On the construction of nested orthogonal Latin hypercube designs. Metrika, 83(3):347–353, April 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-019-
00721-w.

Badia:2020:SCA

[2123] Francisco Germán Badía and Hyunju Lee. On stochastic comparisons and ageing properties of multivariate proportional hazard rate mixtures. Metrika, 83(3):355–375, April 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-019-
00730-9.

Liu:2020:EVC

[2124] Hua Liu, Youquan Pei, and Qunfang Xu. Estimation for varying coefficient panel data model with cross-sectional dependence. Metrika, 83(3):377–410, April 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-019-
                                        00739-0.

Chernoyarov:2020:PSL

[2125] O. V. Chernoyarov and Yu. A. Kutoyants. Poisson source localization on the plane: the smooth case. Metrika, 83(4):411–435, May 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-019-00738-1.

Gaigall:2020:TMH

[2126] Daniel Gaigall. Testing marginal homogeneity of a continuous bivariate distribution with possibly incomplete paired data. Metrika, 83(4):437–465, May 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-019-
                                        00742-5.

Li:2020:ADM

[2127] Yang Li and Yongcheng Qi. Asymptotic distribution of modularity in networks. Metrika, 83(4):467–484, May 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-019-
                                        00740-7.

Liu:2020:ETO

[2128] Xin Liu and Rong-Xian Yue. Elfving’s theorem for R-optimality of experimental designs. Metrika, 83(4):485–498, May 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00184-019-
00728-3.

Koul:2020:CTN

[2129] Hira L. Koul and Fang Li. Comparing two nonparametric regression curves in the presence of long memory in covariates and errors. Metrika, 83(4):499–517, May 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-019-00735-4.

Johnen:2020:NMR

[2130] Marcus Johnen, Stefan Bedbur, and Udo Kamps. A note on multiple roots of a likelihood equation for Weibull sequential order statistics. Metrika, 83(4):519–525, May 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-019-00743-4.

Li:2020:QCU

[2131] Hongyi Li and Hong Qin. Quadrupling:

construction of uniform designs with large run sizes. Metrika, 83(5):527–544, July 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-019-
00741-6.

Ding:2020:RQR

[2132] Xianwen Ding, Jiandong Chen, and Xueping Chen. Regularized quantile regression for ultrahigh-dimensional data with nonignorable missing responses. Metrika, 83(5):545–568, July 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00184-019-
                                        00744-3.

Goto:2020:DAB

[2133] Yuichi Goto and Masanobu Taniguchi. Discriminant analysis based on binary time series. Metrika, 83(5):569–595, July 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-019-
                                        00746-1.

Poloczanski:2020:MEB

[2134] Rafał Połoczański and Maciej Wilczyński. Minimax estimation of a bivariate cumulative distribution function. Metrika, 83(5):597–615, July 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-019-
                                        00747-0;   http://link.springer.
                                        com/content/pdf/10.1007/s00184-
                                        019-00747-0.pdf.

Fang:2020:IGR

[2135] Longxiang Fang and Wenyu Huang. Inequalities for Gaussian random variables under Archimedean copula dependence. Metrika, 83(5):617–625, July 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-019-
                                        00748-z.

Weissbach:2020:CNB

[2136] Rafael Weißbach and Lucas Radloff. Consistency for the negative binomial regression with fixed covariate. Metrika, 83(5):627–641, July 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-019-
00750-5.

Zhang:2020:LOD

[2137] Min-Jue Zhang and Rong-Xian Yue.

Locally D-optimal designs for heteroscedastic polynomial measurement error models. Metrika, 83(6): 643–656, August 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-019-00745-2.

Asadi:2020:UAC

[2138] Majid Asadi and Somayeh Zarezadeh.

A unified approach to constructing correlation coefficients between random variables. Metrika, 83(6):657–676, August 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-019-
00759-w.

Roy:2020:RCS

[2139] Achintya Roy and Nitin Gupta.

Reliability of a coherent system equipped with two cold standby components. Metrika, 83(6):677–697, August 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-019-
00752-3.

Castano-Martinez:2020:ICO

[2140] Antonia Castaño-Martínez, Gema Pigueiras, Georgios Psarrakos, and Miguel A. Sordo. Increasing concave orderings of linear combinations of order statistics with applications to social welfare. Metrika, 83(6):699–712, August 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-019-
                                        00754-1.

Ling:2020:CRK

[2141] Nengxiang Ling, Lingyu Wang, and Philippe Vieu. Convergence rate of kernel regression estimation for time series data when both response and covariate are functional. Metrika, 83 (6):713–732, August 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-019-00757-y.

Hadjicosta:2020:ITM

[2142] Elena Hadjicosta and Donald Richards. Integral transform methods in goodnessof-fit testing, I: the gamma distributions. Metrika, 83(7):733–777, October 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-019-
                                        00749-y.

Han:2020:ASP

[2143] Xiaoxue Han, Jianbin Chen, Min-Qian Liu, and Shengli Zhao. Asymmetrical split-plot designs with clear effects. Metrika, 83(7):779–798, October 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-019-
                                        00755-0.

Song:2020:RCW

[2144] Fengli Song, Peng Lai, and Baohua Shen. Robust composite weighted quantile screening for ultrahigh dimensional discriminant analysis. Metrika, 83(7):799–820, October 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-019-00758-x.

Yin:2020:VSS

[2145] Zanhua Yin. Variable selection for sparse logistic regression. Metrika, 83 (7):821–836, October 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-020-00764-4.

Coeurjolly:2020:MJP

[2146] Jean-François Coeurjolly and Joëlle Rousseau Trépanier. The median of a jittered Poisson distribution. Metrika, 83 (7):837–851, October 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-020-00765-3.

Volovskiy:2020:MPS

[2147] Grigoriy Volovskiy and Udo Kamps.

Maximum product of spacings prediction of future record values. Metrika, 83(7):853–868, October 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-020-00767-1;   http:
//link.springer.com/content/pdf/
10.1007/s00184-020-00767-1.pdf.

Das:2020:OEE

[2148] Sangita Das and Suchandan Kayal.

Ordering extremes of exponentiated location-scale models with dependent and heterogeneous random samples. Metrika, 83(8):869–893, November 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-019-
00753-2.

Miyata:2020:EFM

[2149] Yoichi Miyata, Takayuki Shiohama, and Toshihiro Abe. Estimation of finite mixture models of skew-symmetric circular distributions. Metrika, 83(8): 895–922, November 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-019-00756-z.

Kumar:2020:FOR

[2150] Jitendra Kumar, Seema Jaggi, Eldho Varghese, Arpan Bhowmik, and Cini Varghese. First order rotatable designs incorporating differential neighbour effects from experimental units up to distance 2. Metrika, 83(8): 923–935, November 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-020-00762-6.

Shi:2020:SIF

[2151] Gongming Shi, Tianfa Xie, and Zhongzhan Zhang. Statistical inference for the functional quadratic quantile regression model. Metrika, 83(8): 937–960, November 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-020-00763-5.

Balakrishnan:2020:NEB

[2152] N. Balakrishnan, A. Stepanov, and V. B. Nevzorov. North-east bivariate records. Metrika, 83(8): 961–976, November 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-020-00766-2.

Xu:2020:NQR

[2153] Dengke Xu and Jiang Du. Nonparametric quantile regression estimation for functional data with responses missing at random. Metrika, 83(8): 977–990, November 2020. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-020-00769-z.

Idais:2021:ASL

[2154] Osama Idais and Rainer Schwabe. Analytic solutions for locally optimal designs for gamma models having linear predictors without intercept. Metrika, 84(1):1–26, January 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-019-00760-3.

Lee:2021:GMN

[2155] Hyunju Lee and Ji Hwan Cha. A general multivariate new better than used (MNBU) distribution and its properties. Metrika, 84(1):27–46, January 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-020-
00773-3.

Guney:2021:CML

[2156] Yesim Güney, Y. Tuaç, S. Özdemir, and O. Arslan. Conditional maximum L_{q}-likelihood estimation for regression model with autoregressive error terms. Metrika, 84(1):47–74, January 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-020-
                                        00774-2.

Gao:2021:ELM

[2157] Qibing Gao, Xiuqing Zhou, Yanqin Feng, Xiuli Du, and XiaoXiao Liu. An empirical likelihood method for quantile regression models with censored data. Metrika, 84(1):75–96, January 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-020-
                                        00775-1.

Majumder:2021:SIB

[2158] Suman Majumder, Adhidev Biswas, Tania Roy, Subir Kumar Bhandari, and Ayanendranath Basu. Statistical inference based on a new weighted likelihood approach. Metrika, 84 (1):97–120, January 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-020-00778-y.

Zhao:2021:MAB

[2159] Shengli Zhao and Qianqian Zhao. Minimum aberration blocked designs with multiple block variables. Metrika, 84 (2):121–140, February 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-020-00761-7.

Borisov:2021:UWK

[2160] Igor S. Borisov, Yuliana Yu. Linke, and Pavel S. Ruzankin. Universal weighted kernel-type estimators for some class of regression models. Metrika, 84 (2):141–166, February 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-020-00768-0.

Jiang:2021:OMA

[2161] Rongjie Jiang, Liming Wang, and Yang Bai. Optimal model averaging estimator for semi-functional partially linear models. Metrika, 84(2):167–194, February 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-020-
00772-4.

Voutilainen:2021:AMS

[2162] Marko Voutilainen, Pauliina Ilmonen, Soledad Torres, Ciprian Tudor, and Lauri Viitasaari. On the ARCH model with stationary liquidity. Metrika, 84 (2):195–224, February 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-020-00779-x;   http:
//link.springer.com/content/pdf/
10.1007/s00184-020-00779-x.pdf.

Guo:2021:PIB

[2163] Guangbao Guo, Guoqi Qian, Lu Lin, and Wei Shao. Parallel inference for big data with the group Bayesian method.

Metrika, 84(2):225–243, February 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-020-
                                        00784-0.

Bohning:2021:GFM

[2164] Dankmar Böhning and Helen E. Ogden. General flation models for count data. Metrika, 84(2):245–261, February 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-020-
                                        00786-y;   http://link.springer.
                                        com/content/pdf/10.1007/s00184-
                                        020-00786-y.pdf.

Gayathri:2021:PGV

[2165] Ramanathan Gayathri and Enchakudiyil Ibrahim Abdul Sathar. On past geometric vitality function of order statistics. Metrika, 84(2):263–280, February 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-020-
                                        00789-9.

Kone:2021:NNB

[2166] Mamadou Koné and Annick Valibouze. Nearest neighbor balanced block designs for autoregressive errors. Metrika, 84(3):281–312, April 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-020-00770-6.

Qeadan:2021:EBL

[2167] Fares Qeadan and Ronald Christensen. On the equivalence between the LRT and F -test for testing variance components in a class of linear mixed models. Metrika, 84(3):313–338, April 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-020-
00777-z.

Hu:2021:SAJ

[2168] Zongliang Hu, Zhishui Hu, Kai Dong, Tiejun Tong, and Yuedong Wang. A shrinkage approach to joint estimation of multiple covariance matrices. Metrika, 84(3):339–374, April 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-020-
00781-3.

Philippe:2021:RDS

[2169] Anne Philippe, Caroline Robet, and Marie-Claude Viano. Random discretization of stationary continuous time processes. Metrika, 84(3):375–400, April 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-020-
00783-1.

Dorr:2021:NTM

[2170] Philip Dörr, Bruno Ebner, and Norbert Henze. A new test of multivariate normality by a double estimation in a characterizing PDE. Metrika, 84(3):401–427, April 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-020-00795-x;   http:
//link.springer.com/content/pdf/
10.1007/s00184-020-00795-x.pdf.

Tian:2021:AFR

[2171] Yuzhu Tian, Hongmei Lin, Heng Lian, and Zengyan Fan. Additive functional regression in reproducing kernel Hilbert spaces under smoothness condition. Metrika, 84(3):429–442, April 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-020-
                                        00797-9.

Nikolov:2021:CAP

[2172] Nikolay I. Nikolov and Eugenia Stoimenova. Correction to: Asymptotic properties of Lee distance. Metrika, 84(3):443–444, April 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00184-021-00810-
                                        9;   http://link.springer.com/content/
                                        pdf/10.1007/s00184-021-00810-9.

pdf. See [2077].

Singh:2021:ODU

[2173] Rakhi Singh, Angela Dean, Ashish Das, and Fangfang Sun. A-optimal designs under a linearized model for discrete choice experiments. Metrika, 84(4):445–465, May 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-020-00771-5.

Cai:2021:EEV

[2174] Xiong Cai, Liugen Xue, Xiaolong Pu, and Xingyu Yan. Efficient estimation for varying-coefficient mixed effects models with functional response data. Metrika, 84(4):467–495, May 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-020-
00776-0;   http://link.springer.
com/content/pdf/10.1007/s00184-
020-00776-0.pdf.

Singh:2021:ECD

[2175] Rakhi Singh and Joachim Kunert. Efficient crossover designs for non-regular settings. Metrika, 84(4):497–510, May 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-020-
00780-4.

Hirukawa:2021:APM

[2176] Junichi Hirukawa and Sangyeol Lee.

Asymptotic properties of mildly explosive processes with locally stationary disturbance. Metrika, 84(4):511–534, May 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-020-
00782-2.

Stephanou:2021:PHS

[2177] Michael Stephanou and Melvin Varughese. On the properties of Hermite series based distribution function estimators. Metrika, 84(4):535–559, May 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-020-
00785-z.

Auder:2021:LSM

[2178] Benjamin Auder, Elisabeth Gassiat, and Mor Absa Loum. Least squares moment identification of binary regression mixture models. Metrika, 84(4):561–593, May 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-020-00787-x.

Furmanczyk:2021:EAM

[2179] Konrad Furmańczyk. Estimation of autocovariance matrices for high dimensional linear processes. Metrika, 84(4):595–613, May 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-020-00790-2;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00184-020-00790-2.pdf.

Yang:2021:MDD

[2180] Liuqing Yang, Yongdao Zhou, and Min-Qian Liu. Maximin distance designs based on densest packings. Metrika, 84(5):615–634, July 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-020-
                                        00788-w.

Han:2021:HBR

[2181] Eric Han and Majid Mojirsheibani. On histogram-based regression and classification with incomplete data. Metrika, 84(5):635–662, July 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-020-00794-y.

Su:2021:BMW

[2182] Chun-Lung Su. Bayesian multi-way balanced nested MANOVA models with random effects and a large number of the main factor levels. Metrika, 84(5):663–692, July 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-020-00796-w.

Jin:2021:NES

[2183] Jun Jin, Tiefeng Ma, and Jiajia Dai.

New efficient spline estimation for varying-coefficient models with twostep knot number selection. Metrika, 84(5):693–712, July 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-020-00798-8.

Wang:2021:IVT

[2184] Xinyang Wang, Dehui Wang, and Kai Yang. Integer-valued time series model order shrinkage and selection via penalized quasi-likelihood approach. Metrika, 84(5):713–750, July 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-020-
00799-7.

Oliveira:2021:CIM

[2185] Gabriela Oliveira, Wagner Barreto-Souza, and Roger W. C. Silva. Convergence and inference for mixed Poisson random sums. Metrika, 84(5):751–777, July 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-020-
00800-3.

Zhang:2021:URM

[2186] Mengli Zhang and Yang Bai. On the use of repeated measurement errors in linear regression models. Metrika, 84(5):779–803, July 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-020-00801-2.

Zhao:2021:NMF

[2187] Yuna Zhao, Zhiwei Li, and Shengli Zhao. A new method of finding component orthogonal arrays for orderof-addition experiments. Metrika, 84 (6):805–824, August 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-020-00791-1.

Galarza:2021:MFT

[2188] Christian E. Galarza, Tsung-I Lin, Wan-Lun Wang, and Víctor H. Lachos. On moments of folded and truncated multivariate Student-t distributions based on recurrence relations. Metrika, 84(6):825–850, August 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-020-
00802-1.

Freise:2021:CLS

[2189] Fritjof Freise, Norbert Gaffke, and Rainer Schwabe. Convergence of least squares estimators in the adaptive Wynn algorithm for some classes of nonlinear regression models. Metrika, 84(6):851–874, August 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-020-00803-0;   http:
//link.springer.com/content/pdf/
10.1007/s00184-020-00803-0.pdf.

Tang:2021:NVA

[2190] Yiming Tang, Yang Bai, and Tao Huang. Network vector autoregression with individual effects. Metrika, 84(6):875–893, August 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-020-00805-y.

Laksaci:2021:UCN

[2191] Ali Laksaci, Elias Ould Saïd, and Mustapha Rachdi. Uniform consistency in number of neighbors of the k NN estimator of the conditional quantile model. Metrika, 84(6):895–911, August 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-021-
00806-5.

Coolen:2021:IIB

[2192] Frank P. A. Coolen, Abdullah A. H.

Ahmadini, and Tahani Coolen-Maturi. Imprecise inference based on the logrank test for accelerated life testing. Metrika, 84(6):913–925, August 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-021-
00807-4;   http://link.springer.
com/content/pdf/10.1007/s00184-
021-00807-4.pdf.

Li:2021:DTM

[2193] Mengke Li, Yan Fan, Yang Liu, and Yukun Liu. Diagnostic test metaanalysis by empirical likelihood under a Copas-like selection model. Metrika, 84(6):927–947, August 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-021-00809-2.

Chatterjee:2021:CML

[2194] K. Chatterjee and C. Koukouvinos. Construction of mixed-level supersaturated split-plot designs. Metrika, 84 (7):949–967, October 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-020-00792-0.

Hu:2021:UAL

[2195] Zongyi Hu, Jiaqi Liu, Yi Li, and Hongyi Li. Uniform augmented q-level designs. Metrika, 84(7):969–995, October 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-020-
                                        00793-z.

DeBrabanter:2021:ATR

[2196] Kris De Brabanter and Farzad Sabzikar. Asymptotic theory for regression models with fractional local to unity root errors. Metrika, 84(7): 997–1024, October 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-021-00812-7.

Sang:2021:JEL

[2197] Yongli Sang. A jackknife empirical likelihood approach for testing the homogeneity of K variances. Metrika, 84 (7):1025–1048, October 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-021-00813-6.

Maehara:2021:RBS

[2198] Rocío Maehara, Heleno Bolfarine, Filidor Vilca, and N. Balakrishnan. A robust Birnbaum–Saunders regression model based on asymmetric heavytailed distributions. Metrika, 84(7): 1049–1080, October 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-021-00815-4.

Jasinski:2021:NFC

[2199] Krzysztof Jasiński. The number of failed components in a coherent working system when the lifetimes are discretely distributed. Metrika, 84(7): 1081–1094, October 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-021-00817-2;   http:
//link.springer.com/content/pdf/
10.1007/s00184-021-00817-2.pdf.

Evangelaras:2021:PPT

[2200] H. Evangelaras and S. D. Georgiou.

Projection properties of two-level supersaturated designs constructed from Hadamard designs using Lin’s method. Metrika, 84(8):1095–1108, November 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-020-
00804-z.

Li:2021:TSE

[2201] Rui Li and Yuanyuan Zhang. Twostage estimation and simultaneous confidence band in partially nonlinear additive model. Metrika, 84(8): 1109–1140, November 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-021-00808-3.

Englert:2021:CMF

[2202] Berthold-Georg Englert, Michael Evans, Gun Ho Jang, Hui Khoon Ng, David Nott, and Yi-Lin Seah. Checking for model failure and for prior-data conflict with the constrained multinomial model. Metrika, 84(8): 1141–1168, November 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-021-00811-8;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00184-021-00811-8.pdf.

Wang:2021:HDS

[2203] Zhendong Wang and Xingzhong Xu. High-dimensional sphericity test by extended likelihood ratio. Metrika, 84(8): 1169–1212, November 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-021-00816-3.

Shojaee:2021:SPA

[2204] Omid Shojaee, Majid Asadi, and Maxim Finkelstein. On some properties of α-mixtures. Metrika, 84(8): 1213–1240, November 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-021-00818-1.

Jasinski:2021:SCR

[2205] Krzysztof Jasiński. Some conditional reliability properties of k-out-of-n system composed of different types of components with discrete independent lifetimes. Metrika, 84(8): 1241–1251, November 2021. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-021-00826-1;   http:
//link.springer.com/content/pdf/
10.1007/s00184-021-00826-1.pdf.

Zhao:2022:HDQ

[2206] Weihua Zhao, Rui Li, and Heng Lian.

High-dimensional quantile varyingcoefficient models with dimension reduction. Metrika, 85(1):1–19, January 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-021-
00814-5.

Yuan:2022:HDI

[2207] Panxu Yuan and Xiao Guo. Highdimensional inference for linear model with correlated errors. Metrika, 85 (1):21–52, January 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-021-00820-7.

He:2022:ODR

[2208] Lei He and Rong-Xian Yue. I_{L}-optimal designs for regression models under the second-order least squares estimator. Metrika, 85(1):53–66, January 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-021-
00819-0.

Beknazaryan:2022:SOI

[2209] Aleksandr Beknazaryan and Peter Adamic. On a stochastic order induced by an extension of Panjer’s family of discrete distributions. Metrika, 85(1):67–91, January 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-021-00822-5;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00184-021-00822-5.pdf.

Sun:2022:QML

[2210] Yan Sun and Wei Huang. Quasimaximum likelihood estimation of short panel data models with timevarying individual effects. Metrika, 85 (1):93–114, January 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-021-00825-2.

Bera:2022:ECM

[2211] Samadrita Bera and Nabakumar Jana. On estimating common mean of several inverse Gaussian distributions. Metrika, 85(1):115–139, January 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00184-021-
                                        00829-y.

Anastasiou:2022:DMG

[2212] Andreas Anastasiou and Piotr Fryzlewicz. Detecting multiple generalized change-points by isolating single ones. Metrika, 85(2):141–174, February 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00184-021-
00821-6;   http://link.springer.
com/content/pdf/10.1007/s00184-
021-00821-6.pdf.

Huang:2022:EPV

[2213] Zhensheng Huang, Xing Sun, and Riquan Zhang. Estimation for partially varying-coefficient single-index models with distorted measurement errors. Metrika, 85(2):175–201, February 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-021-
00823-4.

Yzenbrandt:2022:MRD

[2214] Kai Yzenbrandt and Julie Zhou.

Minimax robust designs for regression models with heteroscedastic errors. Metrika, 85(2):203–222, February 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00184-021-
00827-0.

Shao:2022:GPL

[2215] Yujing Shao and Lei Wang. Generalized partial linear models with nonignorable dropouts. Metrika, 85(2): 223–252, February 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

//link.springer.com/article/10.
1007/s00184-021-00828-z.

Balakrishnan:2022:OJB

[2216] Narayanaswamy Balakrishnan and Ritwik Bhattacharya. D-optimal joint best linear unbiased prediction of order statistics. Metrika, 85(2): 253–267, February 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00184-021-00835-0.

Bose:2022:KBE

[2217] Arup Bose and Santanu Dutta. Kernel based estimation of the distribution function for length biased data. Metrika, 85(3):269–287, April 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00184-021-00824-
                                        3.

Phuong:2022:NEC

[2218] Cao Xuan Phuong, Le Thi Hong Thuy, and Vo Nguyen Tuyet Doan. Nonparametric estimation of cumulative distribution function from noisy data in the presence of Berkson and classical errors. Metrika, 85 (3):289–322, April 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-021-00830-5.

Hansmann:2022:ECD

[2219] Matthias Hansmann, Benjamin M. Horn, and Stefan Ulbrich. Estimation of conditional distribution functions from data with additional errors applied to shape optimization. Metrika, 85(3):323–343, April 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-021-00831-4.

Wang:2022:MDT

[2220] Haoyu Wang and Chongqi Zhang. The mixture design threshold accepting algorithm for generating D⃗-optimal designs of the mixture models. Metrika, 85(3):345–371, April 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00832-3.

Diop:2022:PQC

[2221] Mamadou Lamine Diop and William Kengne. Poisson QMLE for changepoint detection in general integervalued time series models. Metrika, 85(3):373–403, April 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00834-1.

Pang:2022:AOM

[2222] Tonghui Pang, Yan Wang, and Jian-Feng Yang. Asymptotically optimal maximin distance Latin hypercube designs. Metrika, 85(4): 405–418, May 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00833-2.

Baringhaus:2022:QM

[2223] Ludwig Baringhaus and Rudolf Grübel.

The quarter median. Metrika, 85 (4):419–458, May 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00836-z.

Krause:2022:PAL

[2224] Joscha Krause, Jan Pablo Burgard, and Domingo Morales. ℓ_{2}- penalized approximate likelihood inference in logit mixed models for regional prevalence estimation under covariate rank-deficiency. Metrika, 85 (4):459–489, May 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00837-y.

Tarima:2022:MPT

[2225] Sergey Tarima and Nancy Flournoy. Most powerful test sequences with early stopping options. Metrika, 85 (4):491–513, May 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00839-w.

Hamura:2022:BPD

[2226] Yasuyuki Hamura and Tatsuya Kubokawa. Bayesian predictive density estimation with parametric constraints for the exponential distribution with unknown location. Metrika, 85 (4):515–536, May 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00840-3.

Yan:2022:MTF

[2227] Zhaohui Yan and Shengli Zhao. Mixed two- and four-level splitplot designs with combined minimum aberration. Metrika, 85 (5):537–555, July 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00838-x.

Cao:2022:HDT

[2228] Mingxiang Cao and Yuanjing He.

A high-dimensional test on linear hypothesis of means under a lowdimensional factor model. Metrika, 85(5):557–572, July 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00841-2.

Dafnis:2022:WRS

[2229] Spiros D. Dafnis and Frosso S.

Makri. Weak runs in sequences of binary trials. Metrika, 85 (5):573–603, July 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00842-1.

Ding:2022:RBT

[2230] Yanling Ding, Binghui Liu, and Long Feng. Rank-based test for slope homogeneity in high-dimensional panel data models. Metrika, 85 (5):605–626, July 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00845-y.

Rachdi:2022:ERS

[2231] Mustapha Rachdi, Ali Laksaci, and Noriah M. Al-Kandari. Expectile regression for spatial functional data analysis (sFDA). Metrika, 85 (5):627–655, July 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00846-x.

Liu:2022:CGS

[2232] Mengmeng Liu, Min-Qian Liu, and Jinyu Yang. Construction of group strong orthogonal arrays of strength two plus. Metrika, 85(6):657–674, August 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https://link.

springer.com/article/10.1007/s00184-
021-00843-0.

Navarro:2022:PRV

[2233] Jorge Navarro. Prediction of record values by using quantile regression curves and distortion functions. Metrika, 85 (6):675–706, August 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00847-w.

Li:2022:RCE

[2234] Kangqiang Li, Han Bao, and Lixin Zhang. Robust covariance estimation for distributed principal component analysis. Metrika, 85(6): 707–732, August 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00848-9.

Cali:2022:ARL

[2235] Camilla Calì, Maria Longobardi, and Barbara Pacchiarotti. Asymptotic results for linear combinations of spacings generated by i.i.d. exponential random variables. Metrika, 85(6): 733–752, August 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00849-8.

Jiang:2022:CUM

[2236] Bochuan Jiang, Fei Wang, and Yaping Wang. Construction of uniform mixed-level designs through level permutations. Metrika, 85(6): 753–770, August 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00850-1.

Huskova:2022:EGP

[2237] Marie Husková, Zuzana Prásková, and Josef G. Steinebach. Estimating a gradual parameter change in an AR(1)-process. Metrika, 85(7): 771–808, October 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00844-z.

Wang:2022:NCB

[2238] Chunlin Wang, Paul Marriott, and Pengfei Li. A note on the coverage behaviour of bootstrap percentile confidence intervals for constrained parameters. Metrika, 85 (7):809–831, October 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00851-0.

Bdair:2022:PFC

[2239] Omar M. Bdair and Mohammad Z.

Raqab. Prediction of future censored lifetimes from mixture exponential distribution. Metrika, 85(7): 833–857, October 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-021-00852-z.

Albertus:2022:AEZ

[2240] Mickael Albertus. Asymptotic Z and chi-squared tests with auxiliary information. Metrika, 85(7): 859–883, October 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-021-00853-y.

Belfadli:2022:SAN

[2241] Rachid Belfadli, Khalifa Es-Sebaiy, and Fatima-Ezzahra Farah. Statistical analysis of the non-ergodic fractional Ornstein–Uhlenbeck process with periodic mean. Metrika, 85 (7):885–911, October 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-021-00854-x.

Diaz:2022:EGB

[2242] Walter Diaz and Carles M. Cuadras. An extension of the Gumbel–Barnett family of copulas. Metrika, 85(7): 913–926, October 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-022-00859-0.

Chernoyarov:2022:NAE

[2243] O. V. Chernoyarov, A. S. Dabye, and Y. A. Kutoyants. Non asymptotic expansions of the MME in the case of Poisson observations. Metrika, 85(8): 927–950, November 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00855-w.

He:2022:CMU

[2244] Mingxing He and Jiahua Chen.

Consistency of the MLE under a two-parameter Gamma mixture model with a structural shape parameter. Metrika, 85(8):951– 975, November 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00856-9.

Tang:2022:DFS

[2245] Yinfen Tang, Tao Su, and Zhiyuan Zhang. Distribution-free specification test for volatility function based on high-frequency data with microstructure noise. Metrika, 85(8): 977–1022, November 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-021-00857-8.

Chang:2022:BEI

[2246] Yu-Wei Chang and Jyun-Ye Tu.

Bayesian estimation for an item response tree model for nonresponse modeling. Metrika, 85(8):1023– 1047, November 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-022-00858-1.

Fakhari-Esferizi:2022:UOB

[2247] Akram Fakhari-Esferizi and Saeid Pooladsaz. Universally optimal balanced block designs for interference model. Metrika, 85(8):1049– 1061, November 2022. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-022-00870-5.

Peng:2023:LRS

[2248] Ling Peng, Yan Zhu, and Wenxuan Zhong. Lasso regression in sparse linear model with φ-mixing errors. Metrika, 86(1):1–26, January 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-022-00860-7.

Cheng:2023:ABR

[2249] Fuxia Cheng and Hira L. Koul. An analog of Bickel–Rosenblatt test for fitting an error density in the two phase linear regression model. Metrika, 86 (1):27–56, January 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-022-00861-6.

Kelkinnama:2023:SRM

[2250] Maryam Kelkinnama and Serkan Eryilmaz. Some reliability measures and maintenance policies for a coherent system composed of different types of components. Metrika, 86 (1):57–82, January 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-022-00862-5.

Bar-Lev:2023:NDN

[2251] Shaul K. Bar-Lev and Gérard Letac. A note on the discretization of natural exponential families on the real line. Metrika, 86(1):83–90, January 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL https://link.springer.com/
article/10.1007/s00184-022-00863-
4.

Mercier:2023:GML

[2252] Sophie Mercier and Carmen Sangüesa.

A general multivariate lifetime model with a multivariate additive process as conditional hazard rate increment process. Metrika, 86(1): 91–129, January 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-022-00864-3.

Esna-Ashari:2023:HRO

[2253] Maryam Esna-Ashari, Narayanaswamy Balakrishnan, and Mahdi Alimohammadi. HR and RHR orderings of generalized order statistics. Metrika, 86 (1):131–148, January 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-022-00865-2.

Tang:2023:ELB

[2254] Songqiao Tang, Huiyu Wang, Guanao Yan, and Lixin Zhang. Empirical likelihood based tests for detecting the presence of significant predictors in marginal quantile regression. Metrika, 86(2): 149–179, February 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-022-00866-1.

Ng:2023:PML

[2255] Tin Lok James Ng. Penalized maximum likelihood estimator for mixture of von Mises–Fisher distributions. Metrika, 86(2):181– 203, February 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-022-00867-0.

Jasinski:2023:CRL

[2256] Krzysztof Jasiński. On conditional residual lifetimes of coherent systems consisting of components with discrete lifetimes. Metrika, 86(2): 205–218, February 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-022-00871-4.

Chen:2023:ODN

[2257] Xueping Chen, Yujie Gai, and Xiaodi Wang. A-optimal designs for nonparametric symmetrical global sensitivity analysis. Metrika, 86(2): 219–237, February 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-022-00872-3.

Huang:2023:LOC

[2258] Zhiyun Huang, Zhiming Li, Ge Zhang, and Tao Chen. Lower-order confounding information of inverse yates-order designs with three levels. Metrika, 86 (2):239–259, February 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-022-00876-z.

Yang:2023:CES

[2259] Yaohong Yang and Lei Wang. Communicationefficient sparse composite quantile regression for distributed data. Metrika, 86(3):261–283, April 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-022-00868-z.

Liu:2023:BEL

[2260] Chang-Sheng Liu and Han-Ying Liang. Bayesian empirical likelihood of quantile regression with missing observations. Metrika, 86 (3):285–313, April 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-022-00869-y.

Dette:2023:EIP

[2261] Holger Dette, Vasyl Golosnoy, and Janosch Kellermann. The effect of intraday periodicity on realized volatility measures. Metrika, 86 (3):315–342, April 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-022-00875-0.

Younso:2023:CME

[2262] Ahmad Younso. On the consistency of mode estimate for spatially dependent data. Metrika, 86 (3):343–372, April 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-022-00879-w.

Wang:2023:FFD

[2263] Lin Wang, Hongquan Xu, and Min-Qian Liu. Fractional factorial designs for Fourier-cosine models. Metrika, 86(3):373–390, April 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-022-00881-2.

Jin:2023:CML

[2264] Libin Jin, Sung Nok Chiu, Jianhua Zhao, and Lixing Zhu. A constrained maximum likelihood estimation for skew normal mixtures. Metrika, 86 (4):391–419, May 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-022-00873-2.

Kawashima:2023:RRA

[2265] Takayuki Kawashima and Hironori Fujisawa. Robust regression against heavy heterogeneous contamination. Metrika, 86(4):421–442, May 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-022-00874-1.

Jureckova:2023:NTL

[2266] Jana Jurečková, Olcay Arslan, Yeşim Güney, Jan Picek, Martin Schindler, and Yetkin Tuaç. Nonparametric tests in linear model with autoregressive errors. Metrika, 86 (4):443–453, May 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-022-00877-y.

Liu:2023:CED

[2267] Zhan Liu, Xiaoluo Zhao, and Yingli Pan. Communication-efficient distributed estimation for high-dimensional large-scale linear regression. Metrika, 86(4):455–485, May 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-022-00878-x.

McVittie:2023:NPL

[2268] James H. McVittie, David B. Wolfson, and David A. Stephens. A note on the partial likelihood estimator of the proportional hazards model for combined incident and prevalent cohort data. Metrika, 86 (4):487–497, May 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-022-00882-1.

Shojaee:2023:SSC

[2269] Omid Shojaee and Manoochehr Babanezhad. On some stochastic comparisons of arithmetic and geometric mixture models. Metrika, 86 (5):499–515, July 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-022-00880-3.

Gochanour:2023:SMR

[2270] Benjamin Gochanour, Sixia Chen, Laura Beebe, and David Haziza. A semiparametric multiply robust multiple imputation method for causal inference. Metrika, 86 (5):517–542, July 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-022-00883-0.

Navarro:2023:PFF

[2271] Jorge Navarro and Francesco Buono. Predicting future failure times by using quantile regression. Metrika, 86 (5):543–576, July 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-022-00884-z.

Bai:2023:DEM

[2272] Qiming Bai, Hongyi Li, Xingyou Huang, and Huili Xue. Design efficiency for minimum projection uniform designs with q levels. Metrika, 86(5):577–594, July 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-022-00885-y.

Jiang:2023:FSI

[2273] Zhiqiang Jiang, Zhensheng Huang, and Jing Zhang. Functional single-index composite quantile regression. Metrika, 86(5):595–603, July 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-022-00887-w.

Etminan:2023:NLR

[2274] Jalal Etminan, Mohammad Khanjari Sadegh, and Majid Chahkandi. A new light on reliability equivalence factors. Metrika, 86(6): 605–625, August 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-022-00886-x.

Bapat:2023:NSA

[2275] Sudeep R. Bapat. A novel sequential approach to estimate functions of parameters of two gamma populations. Metrika, 86(6):627–641, August 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL https://link.springer.com/
article/10.1007/s00184-022-00888-
9.

Kawai:2023:AIS

[2276] Tetsuya Kawai and Masayuki Uchida.

Adaptive inference for small diffusion processes based on sampled data. Metrika, 86(6):643–696, August 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL https://link.springer.com/
article/10.1007/s00184-022-00889-
8.

Schmidt:2023:SME

[2277] Marius Schmidt. Standardized maximin D — and c-optimal designs for the Poisson–Gamma model. Metrika, 86(6):697–721, August 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-022-00890-1.

Meintanis:2023:BEC

[2278] Simos Meintanis, Bojana Milosević, and Marko Obradović. Bahadur efficiency for certain goodness-of-fit tests based on the empirical characteristic function. Metrika, 86(7): 723–751, October 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-022-00891-0.

Yan:2023:CTO

[2279] Chongya Yan, Zhiming Li, and Mingyao Ai. Characterization of three-order confounding via consulting sets. Metrika, 86(7):753– 779, October 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00892-7.

Tsirpitzi:2023:ROD

[2280] Renata Eirini Tsirpitzi, Frank Miller, and Carl-Fredrik Burman. Robust optimal designs using a model misspecification term. Metrika, 86(7): 781–804, October 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00893-6.

Chen:2023:CDB

[2281] Huaping Chen, Qi Li, and Fukang Zhu. A covariate-driven betabinomial integer-valued GARCH model for bounded counts with an application. Metrika, 86(7):805– 826, October 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00894-5.

Ouimet:2023:RCC

[2282] Frédéric Ouimet. A refined continuity correction for the negative binomial distribution and asymptotics of the median. Metrika, 86(7):

827–849, October 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00897-2.

Canu:2023:DBL

[2283] Stéphane Canu and Dominique Fourdrinier. Data based loss estimation of the mean of a spherical distribution with a residual vector. Metrika, 86 (8):851–878, November 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00895-4.

Niaparast:2023:ODR

[2284] Mehrdad Niaparast, Sahar MehrMansour, and Rainer Schwabe. V - optimality of designs in random effects Poisson regression models. Metrika, 86 (8):879–897, November 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00896-3.

Fetitah:2023:RNE

[2285] Omar Fetitah, Mohammed Kadi Attouch, Salah Khardani, and Ali Righi. Robust nonparametric equivariant regression for functional data with responses missing at random. Metrika, 86 (8):899–929, November 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00898-1.

Yang:2023:BTP

[2286] Kai Yang, Yiwei Zhao, Han Li, and Dehui Wang. On bivariate threshold Poisson integer-valued autoregressive processes. Metrika, 86(8): 931–963, November 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00899-0.

Ding:2023:NEF

[2287] Hui Ding, Mei Yao, and Riquan Zhang. A new estimation in functional linear concurrent model with covariate dependent and noise contamination. Metrika, 86(8):965– 989, November 2023. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00900-w.

Sisheng:2024:KRE

[2288] Liu Sisheng and Yang Jing. Kernel regression for estimating regression function and its derivatives with unknown error correlations. Metrika, 87(1):1–20, January 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00901-9.

Liu:2024:DEF

[2289] Jiamin Liu, Rui Li, and Heng Lian. Distributed estimation of functional linear regression with functional responses. Metrika, 87(1): 21–30, January 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00902-8.

Neumann:2024:EBS

[2290] Michael H. Neumann. Estimation and bootstrap for stochastically monotone Markov processes. Metrika, 87 (1):31–59, January 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00903-7.

Cha:2024:PGL

[2291] Ji Hwan Cha and F. G. Badía.

Poisson generalized Lindley process and its properties. Metrika, 87 (1):61–74, January 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00906-4.

Jiang:2024:OTL

[2292] Haosheng Jiang and Chongqi Zhang.

Optimal two-level regular fractional factorial split-plot designs when the effects of subplot factors are more important. Metrika, 87(1): 75–103, January 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00907-3.

Ciuperca:2024:RTC

[2293] Gabriela Ciuperca, Matús Maciak, and Michal Pesta. Real-time changepoint detection in a nonlinear expectile model. Metrika, 87(2): 105–131, February 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00904-6.

Amini-Seresht:2024:CEO

[2294] Ebrahim Amini-Seresht, Ebrahim Nasiroleslami, and Narayanaswamy Balakrishnan. Comparison of extreme order statistics from two sets of heterogeneous dependent random variables under random shocks. Metrika, 87 (2):133–153, February 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00905-5.

Grubel:2024:RCP

[2295] R. Grübel. Ranks, copulas, and permutons. Metrika, 87(2):155– 182, February 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00908-2.

Jasinski:2024:NFC

[2296] Krzysztof Jasiński. On the number of failed components in a seriesparallel system upon system failure when the lifetimes are DNID discrete random variables. Metrika, 87(2): 183–200, February 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00909-1.

Zaigraev:2024:NAR

[2297] Alexander Zaigraev. A note on asymptotics of the risk function under confidence region estimation in case of large samples of random size. Metrika, 87 (2):201–209, February 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00910-8.

Janssen:2024:NEU

[2298] Paul Janssen and Noël Veraverbeke.

Nonparametric estimation of univariate and bivariate survival functions under right censoring: a survey. Metrika, 87(3):211–245, April 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00911-7.

Guo:2024:ADL

[2299] Wenchuan Guo and Yongcheng Qi.

Asymptotic distributions for likelihood ratio tests for the equality of covariance matrices. Metrika, 87 (3):247–279, April 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00912-6.

Yin:2024:MDP

[2300] Qing Yin, Yu Miao, Zhen Wang, and Guangyu Yang. Moderate deviation principle of modularity in network. Metrika, 87(3):281–298, April 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL https://link.springer.com/
article/10.1007/s00184-023-00914-
4.

Sun:2024:RDM

[2301] Gaoming Sun, Xiaozhou Wang, Yibo Yan, and Riquan Zhang. Robust distributed multicategory angle-based classification for massive data. Metrika, 87(3):299–323, April 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00915-3.

Shojaee:2024:ORS

[2302] Omid Shojaee, Seyed Morteza Mohammadi, and Reza Momeni. Ordering results for the smallest (largest) and the second smallest (second largest) order statistics of dependent and heterogeneous random variables. Metrika, 87(3):325–347, April 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00917-1.

Huillet:2024:BTU

[2303] Thierry Huillet and Martin Möhle. On Bernoulli trials with unequal harmonic success probabilities. Metrika, 87(4):349–378, May 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00913-5.

Chao:2024:OSM

[2304] Yue Chao, Lei Huang, Xuejun Ma, and Jiajun Sun. Optimal subsampling for modal regression in massive data. Metrika, 87 (4):379–409, May 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00916-2.

Rodenburger:2024:RAA

[2305] Daniel Rodenburger. Refining analytic approximation based estimation of mixed multinomial probit models by parameter selection. Metrika, 87(4):411–425, May 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00920-6.

Arnold:2024:BDE

[2306] Barry C. Arnold and B. G. Manjunath. Bivariate distributions with equi-dispersed normal conditionals and related models. Metrika, 87 (4):427–448, May 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00921-5.

Nair:2024:AHQ

[2307] N. Unnikrishnan Nair, S. M. Sunoj, and Namitha Suresh. Additive hazards quantile model. Metrika, 87 (4):449–469, May 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00924-2.

Ogasawara:2024:DSC

[2308] Haruhiko Ogasawara. The distribution of the sample correlation coefficient under variance-truncated normality. Metrika, 87(5):471–497, July 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL https://link.springer.com/
article/10.1007/s00184-023-00918-
0.

Beh:2024:GAA

[2309] Eric J. Beh, Duy Tran, and Irene L.

Hudson. A generalisation of the aggregate association index (AAI): incorporating a linear transformation of the cells of a 2 × 2 table. Metrika, 87(5):1–33, July 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00919-z.

Sijuwade:2024:ILT

[2310] Adebowale J. Sijuwade, Swarnita Chakraborty, and Nairanjana Dasgupta. An inverse Laplace transform oracle estimator for the normal means problem. Metrika, 87 (5):533–550, July 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00922-4.

Wu:2024:LDR

[2311] Yi Wu, Wei Yu, and Xuejun Wang. Large deviations for randomly weighted least squares estimator in a nonlinear regression model. Metrika, 87(5):551–570, July 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00926-0.

Vasudeva:2024:ABJa

[2312] R. Vasudeva. On the asymptotic behaviour of the joint distribution of the maxima and minima of observations, when the sample size is a random variable. Metrika, 87 (5):571–584, July 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00928-y.

Wang:2024:BMN

[2313] Wan-Lun Wang, Luis M. Castro, and Tsung-I Lin. Bayesian multivariate nonlinear mixed models for censored longitudinal trajectories with non-monotone missing values. Metrika, 87(5):585–605, July 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00929-x.

Mojirsheibani:2024:RCP

[2314] Majid Mojirsheibani, William Pouliot, and Andre Shakhbandaryan. On regression and classification with possibly missing response variables in the data. Metrika, 87(6): 607–648, August 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00923-3.

Mukhopadhyay:2024:SOM

[2315] Nitis Mukhopadhyay and Anhar Aloufi. Second-order (s.o.) multi-stage fixed-width confidence interval (FWCI) estimation strategies for comparing location parameters from two negative exponential (NE) populations: illustrations with cancer data. Metrika, 87 (6):649–680, August 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00925-1.

Bhakta:2024:SCT

[2316] Raju Bhakta, Priyanka Majumder, Suchandan Kayal, and Narayanaswamy Balakrishnan. Stochastic comparisons of two finite mixtures of general family of distributions. Metrika, 87 (6):681–712, August 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00930-4.

Yu:2024:PSA

[2317] Qiqing Yu. A proper selection among multiple Buckley–James estimates. Metrika, 87(6):713–728, August 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00184-023-00933-
                                        1.

Barczy:2024:MCL

[2318] Mátyás Barczy, Fanni Nedényi, and Gyula Pap. Mixing convergence of LSE for supercritical AR(2) processes with Gaussian innovations using random scaling. Metrika, 87 (6):729–756, August 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00936-y.

Kuran:2024:RPE

[2319] Özge Kuran and M. Revan Özkale. The ridge prediction error sum of squares statistic in linear mixed models. Metrika, 87(7):757–773, October 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00184-023-00927-
                                        z.

Capaldo:2024:CIG

[2320] Marco Capaldo, Antonio Di Crescenzo, and Alessandra Meoli. Cumulative information generating function and generalized Gini functions. Metrika, 87 (7):775–803, October 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00931-3.

Sun:2024:JMA

[2321] Xianwen Sun and Lixin Zhang. Jackknife model averaging for mixeddata kernel-weighted spline quantile regressions. Metrika, 87(7): 805–842, October 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00932-2.

Aleksandrov:2024:MDT

[2322] Boris Aleksandrov, Christian H. Weiß, Simon Nik, Maxime Faymonville, and Carsten Jentsch. Modelling and diagnostic tests for Poisson and negativebinomial count time series. Metrika, 87 (7):843–887, October 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00934-0.

Li:2024:BMA

[2323] Hui Li, Min-Qian Liu, and Jinyu Yang.

Bayesian minimum aberration mixedlevel split-plot designs. Metrika, 87 (7):889–906, October 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00937-x.

Vasudeva:2024:ABJb

[2324] R. Vasudeva. On the asymptotic behaviour of the joint distribution of the maxima and minima of observations, when the sample size is a random variable. Metrika, 87(7): 907–920, October 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00940-2.

He:2024:CSP

[2325] Lidan He, Qiang Liu, Zhi Liu, and Andrea Bucci. Correcting spot power variation estimator via Edgeworth expansion. Metrika, 87(8): 921–945, November 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00935-z.

Wang:2024:TIE

[2326] Xiao Wang and Lihong Wang. A tail index estimation for long memory processes. Metrika, 87(8):947– 971, November 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00938-w.

Phuong:2024:NEM

[2327] Cao Xuan Phuong and Le Thi Hong Thuy. Nonparametric estimation of P (X < Y) from noisy data samples with non-standard error distributions. Metrika, 87(8):973– 1006, November 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-023-00941-1.

Almanjahie:2024:LLF

[2328] Ibrahim M. Almanjahie, Salim Bouzebda, Zoulikha Kaid, and Ali Laksaci.

The local linear functional k NN estimator of the conditional expectile: uniform consistency in number of neighbors. Metrika, 87(8): 1007–1035, November 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00942-0.

Ando:2024:MDM

[2329] Shuji Ando. Measure of deviancy from marginal mean equality based on cumulative marginal probabilities in square contingency tables. Metrika, 87(8):1037–1048, November 2024. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL https://link.springer.com/
article/10.1007/s00184-023-00945-
x.

Hu:2025:PLL

[2330] Hongchang Hu, Mingqiu Liu, and Zhen Zeng. Penalized Lq-likelihood estimator and its influence function in generalized linear models. Metrika, 88(1):1–18, January 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-023-00943-z.

Wu:2025:BEB

[2331] Yi Wu and Xuejun Wang. On Berry–Esséen bound of frequency polygon estimation under ρ-mixing samples. Metrika, 88(1):19–41, January 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL https://link.springer.com/
article/10.1007/s00184-023-00944-
y.

DiCrescenzo:2025:SCD

[2332] Antonio Di Crescenzo, Luca Paolillo, and Alfonso Suárez-Llorens. Stochastic comparisons, differential entropy and varentropy for distributions induced by probability density functions. Metrika, 88(1):43–59, January 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-024-00947-3.

Maluf:2025:RBR

[2333] Yuri S. Maluf, Silvia L. P. Ferrari, and Francisco F. Queiroz. Robust beta regression through the logit transformation. Metrika, 88 (1):61–81, January 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-024-00949-1.

Arnroth:2025:BCQ

[2334] Lukas Arnroth. Bayesian composite L^{p}-quantile regression. Metrika, 88(1):83–97, January 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-024-00950-8.

Qi:2025:MAC

[2335] Liangwei Qi and Yongdao Zhou. Minimum θ-aberration criterion for designs with qualitative and quantitative factors. Metrika, 88(1): 99–117, January 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-024-00951-7.

Cleanthous:2025:PDE

[2336] G. Cleanthous, Athanasios G. Georgiadis, and P. A. White. Pointwise density estimation on metric spaces and applications in seismology. Metrika, 88 (2):119–148, February 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-024-00948-2.

Fung:2025:RED

[2337] Tsz Chai Fung. Robust estimation and diagnostic of generalized linear model for insurance losses: a weighted likelihood approach. Metrika, 88(2): 149–182, February 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-024-00952-6.

Aydin:2025:RCN

[2338] Dursun Aydın, Ersin Yılmaz, Nur Chamidah, Budi Lestari, and I. Nyoman Budiantara. Right-censored nonparametric regression with measurement error. Metrika, 88(2): 183–214, February 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-024-00953-5.

Babilua:2025:SPO

[2339] P. Babilua and E. Nadaraya. On some properties of one nonparametric estimate of Poisson regression function. Metrika, 88(2): 215–229, February 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-024-00955-3.

Drew:2025:ESL

[2340] Courtney Drew and Éric Marchand. Estimating the suspected larger of two normal means. Metrika, 88(2): 231–245, February 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-024-00961-5.

Chisaki:2025:OCS

[2341] Shoko Chisaki, Ryoh Fuji-Hara, and Nobuko Miyamoto. Optimality and constructions of spanning bipartite block designs. Metrika, 88(2): 247–265, February 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-024-00963-3.

Anonymous:2025:JNa

[2342] Anonymous. Journal navigation. Metrika, 88(2):??, February 2025. CO-DEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Jebalia:2025:MJP

[2343] Mohamed Jebalia and Abderrazek Karoui. A multivariate Jacobi polynomials regression estimator associated with an ANOVA decomposition model. Metrika, 88(3): 267–310, April 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-024-00954-4.

Garcia-Angulo:2025:BIA

[2344] Andrea C. Garcia-Angulo and Gerda Claeskens. Bootstrap for inference after model selection and model averaging for likelihood models. Metrika, 88(3):311–340, April 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-024-00956-2.

Liu:2025:BRB

[2345] Qing Liu, Xiaohui Liu, and Zihao Hu. Bahadur representations for the bootstrap median absolute deviation and the application to projection depth weighted mean. Metrika, 88(3):341–364, April 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-024-00958-0.

Eberl:2025:MKI

[2346] Andreas Eberl and Bernhard Klar.

Measures of kurtosis: inadmissible for asymmetric distributions? Metrika, 88(3):365–391, April 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-024-00959-z.

Cramer:2025:SHC

[2347] Erhard Cramer. Structure of hybrid censoring schemes and its implications. Metrika, 88(3):393–417, April 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL https://link.springer.com/
article/10.1007/s00184-024-00960-
6.

Anonymous:2025:JNb

[2348] Anonymous. Journal navigation. Metrika, 88(3):??, April 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Hansen:2025:VEA

[2349] Stefan Nygaard Hansen and Morten Overgaard. Variance estimation for average treatment effects estimated by g-computation. Metrika, 88(4):??, May 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-024-00962-4.

Zhang:2025:TAS

[2350] Xinqi Zhang, Yaping Wang, and Fasheng Sun. Theory and applications of stratification criteria based on space-filling pattern and projection pattern. Metrika, 88 (4):419–443, May 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00184-024-00964-2.

Kengne:2025:SPD

[2351] William Kengne and Modou Wade. Sparse-penalized deep neural networks estimator under weak dependence. Metrika, 88(4):445–468, May 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00184-024-00965-
                                        1.

Hu:2025:MXK

[2352] Ran Hu, Di Xia, and Yingli Pan. Model-X knockoffs for high-dimensional controlled variable selection under the proportional hazards model with heterogeneity parameter. Metrika, 88 (4):469–500, May 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-024-00966-0.

Liang:2025:SIL

[2353] Xiaowen Liang and Boping Tian.

Statistical inference for linear quantile regression with measurement error in covariates and nonignorable missing responses. Metrika, 88 (4):501–521, May 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-024-00967-z.

Goroncy:2025:BEO

[2354] Agnieszka Goroncy and Tomasz Rychlik. Bounds of expectations of order statistics for distributions possessing monotone reversed failure rates. Metrika, 88(4): 523–549, May 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-024-00968-y.

Anonymous:2025:JNc

[2355] Anonymous. Journal navigation.

Metrika, 88(4):551–575, May 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

Kappara:2025:AMS

[2356] Divya Kappara, Arup Bose, and Madhuchhanda Bhattacharjee. An association measure for spatio-temporal time series. Metrika, 88(5):577–599, July 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).

URL https://link.springer.com/
article/10.1007/s00184-023-00939-
9.

Tonaki:2025:PEL

[2357] Yozo Tonaki, Yusuke Kaino, and Masayuki Uchida. Parametric estimation for linear parabolic SPDEs in two space dimensions based on temporal and spatial increments. Metrika, 88(5):601–656, July 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-024-00969-x.

Wang:2025:MTH

[2358] Guanpeng Wang, Yuyuan Wang, and Hengjian Cui. Mean test for highdimensional data based on covariance matrix with linear structures. Metrika, 88(5):657–678, July 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-024-00971-3.

Zhang:2025:CTL

[2359] Tian fang Zhang, Yingxing Duan, and Zhiming Li. Construction of three-level factorial designs with general minimum lower-order confounding via resolution IV designs. Metrika, 88(5):679–687, July 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-024-00972-2.

Joshi:2025:TSP

[2360] Neeraj Joshi, Sudeep R. Bapat, and Raghu Nandan Sengupta. Twostage and purely sequential minimum risk point estimation of the scale parameter of a family of distributions under modified LINEX loss plus sampling cost. Metrika, 88 (5):689–707, July 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-024-00973-1.

Zhu:2025:SSC

[2361] Xiaorui Zhu, Yichen Qin, and Peng Wang. Sparsified simultaneous confidence intervals for highdimensional linear models. Metrika, 88(5):709–733, July 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic). URL https:

//link.springer.com/article/10.
1007/s00184-024-00975-z.

Anonymous:2025:JNd

[2362] Anonymous. Journal navigation.

Metrika, 88(5):709–733, July 2025. CODEN MTRKA8. ISSN 0026-1335 (print), 1435-926X (electronic).